A Complete Bibliography of Publications in Communications of the ACM : 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 2.93

Title word cross-reference

2^{W} [2673]. 3 [1619, 1514, 1621, 750, 1371, 1615, 1620]. = [959, 1014]. ^{st} [1531, 1840]. ^{th} [1531]. D [574].

.NET [1017, 1016].

’02 [916].

10 [1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1939, 1965, 1986, 2005, 2025, 2044]. 10th [2566]. 11th [2313]. 12 [2667, 5, 1485, 2988].

2.0 [613, 2673]. 20 [480]. 20/20 [2040]. 2002 [864]. 21st [2090, 371, 347, 2466, 2499, 1678].

3 [2865, 574]. 3083s [40]. 3D [257, 847]. 3D-enabled [847]. 3G [2191].

50-year-old [2458]. 500 [2751].

802.11-based [988]. 802.11b [990].

9/11 [792, 791].

A-Head [1503]. Abbreviation [127]. Abduction [838]. Abilities [2432]. ability [2171, 2826]. absolute [350]. Abstract [84, 2865]. Abstraction [1586, 2908, 2138]. abstractions [2909]. abundant [1786].

Abuse [604, 601]. academia [2337, 1105, 1155]. Academic [1919, 15, 1405, 1714, 2639, 248, 1972, 1295, 1945, 2990]. academics [378]. Accelerating [682, 2310]. Acceptable [607]. acceptance [934, 2590, 1448]. Access [2484, 258, 397, 376, 541, 588, 1409, 2454, 183, 699, 2596, 990, 1273, 2, 898, 1977, 705, 2718, 2819, 1722, 1871]. accessibility [2141, 1472, 2917, 2167]. Accessible [685]. Accessing [2169]. Accompaniment [2041, 1969]. Accomplishing [614]. Accomplishment [1030]. accomplishments [31]. According [2432]. account [2706]. Accountability [1655, 351, 935, 2452]. Accreditation [481, 781]. Accuracy [1028]. achieve [2010, 2366, 2315]. achievements [250]. achieving [2593]. ACM [2159, 1915, 2240, 2596, 2163, 2161, 466, 729, 921, 1303, 1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1936, 1962, 1779, 1812, 1858, 1866, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2044, 2218, 2081, 2101, 2126, 2147, 2173, 2190, 2213, 2233, 2259, 2282, 2303, 2322, 2152, 2288, 2367, 2640, 515, 1424, 2160, 2503, 2922, 2162, 2158, 66, 1659, 1505, 1781, 1913, 424, 1319, 803, 623, 520, 635, 600, 1380, 1218, 1469, 1260, 1530, 1849, 2043, 1839, 1861, 2072, 2110, 2134, 2106, 2135, 2312, 2424, 2384, 2403, 2421, 2402, 2697]. ACM [2829, 2786, 455, 858, 1282, 2566, 2157]. ACM-W [2786]. Acquire [887, 2324]. Acquisition [917, 1201]. across [628, 1924, 868, 1037, 289, 1522, 1930, 2074]. Act [220, 1629, 149, 960, 2093, 21, 620, 2125]. Action [1896, 1895, 1489, 1898]. actions [2029]. activated [2788]. Active [2652, 385, 2142, 391, 828, 2725]. activity [488, 666, 2142, 1765, 1691]. actors [747]. actual [1579]. Ad [1053, 1545]. Ad-hoc [1545]. Ada [1835]. adapt [1219]. adaptation [2096]. Adaptive [722, 699, 706, 1592, 1901, 1376, 176, 950, 702, 707, 698, 703, 1847, 523, 1595, 1848, 818, 1593]. Adaptivity [645]. Add [127, 2299]. addiction [2959]. Addictive [1753]. Additional [767]. Address [497, 1913]. addressable [1907]. addressing [2064]. adhocracies [2115]. Adjective [481]. Admediation [583]. admissions [2127]. Ado [88]. adopters [974, 2653]. Adopting [1449, 1226]. Adoption [481, 682, 2850, 2018, 2886, 1999, 2047, 2184, 1092, 1103, 249, 1669, 1298, 1605, 1729, 2847, 1774, 2388, 1170, 1111, 70]. Ads [640]. Adults [685, 1525]. Advance [300]. Advanced [866, 2062, 2184, 1839, 2135, 2402]. advances [2898]. Advancing [2812]. Advantage [1912, 588, 1090, 2658]. advantages [2818]. Advertising [583, 2573, 2123, 2958, 1907]. Advice [564, 1982, 2605]. Advising [2701, 2605, 1700]. advocate [1423]. AFF [1823]. affair [2417]. affect [151, 2875, 1103, 1263]. Affecting [328, 1669]. affective [50, 1702]. Affinity [996]. Affordable [481]. affordances [157, 2595]. Afghanistan [1808]. Afoot [1961]. Africa [467, 868, 1349]. African [1777]. After [1141, 350, 792, 1133, 2669]. after-math [2669]. aftermath [18, 791]. again [2336, 1518, 87, 744, 2800]. Against [1982, 2585]. Age [1028, 1236, 1096, 2611, 2118, 880, 2894, 456, 2520, 310, 2705, 55, 3006]. age-old [2894]. agencies [289]. Agency [560]. agenda [2857]. Agendas [1280]. Agent [381, 2114, 383, 387, 1395, 2431, 382, 2058, 2170]. Agent-based [2114, 383, 1395, 382]. Agent-oriented [381, 387]. Agents [786, 1626, 58, 78, 1046, 491, 1049, 946, 1286, 123, 710, 2228, 654]. aggregation [2525]. Agile [2069, 1769, 2920, 613, 1598, 2122, 2009, 2624, 2324, 2011]. Agility [994, 2532, 2010, 2593]. aging [2981, 1717]. agricultural [126]. Ahead [1163, 982, 2050, 1938, 2087, 359, 2341]. AI [1215, 2130, 595]. aid [1100, 727]. aided [1951]. aids [1717]. Air [147, 1961, 445]. airjacking [1458]. Airline [692]. Airport [838, 1961]. AJAX [2769]. Alaska [433]. Alerts [838, 1585]. algorithm [1581, 2680]. Algorithmic [2772, 2974, 1944]. algorithmics [2291]. algorithmization [1893]. Algorithms [2041, 2358, 2556, 770, 2987, 769, 2941, 2756]. Alice [2865]. Aligning [2010, 606]. alignment [1969, 2851]. alive [1304]. All-Seeing [618]. Allure [1141]. Alone [1835, 1925]. along [974]. Already [740, 1835, 1962]. Alter [1777, 1351]. Alternate [2368, 1091, 1284]. Alternative [2371, 1310]. alternatives [1879]. always [2513]. Ambidextrous [2008]. ambiguity [2785]. America [2921, 1472, 1236]. American [463, 1439]. Americans [1861, 1452, 2885]. amidst [2488]. Amoeba [2251]. Amoeba-based [2251]. among [2777]. amorphous [98]. analog [2161]. analogous [57]. analogy [2014, 315]. analyses [548]. Analysis [2142, 570, 2751, 735, 2050, 1816, 1266, 1187, 2171, 1619, 1875, 1341, 508, 2911, 1405, 903, 2877, 1829, 1513, 3002, 2037, 1675, 2941, 1982, 2584, 255, 28, 1820, 1402, 26, 1625]. analysts [1957, 1100]. analytic [1520]. analytics [771]. Analyzing [2571, 1422, 1004, 527, 1070, 1212]. anatomy [1514]. Ancient [1325]. ancyent [2155]. Animating [137]. Animation [574, 136, 134, 135]. animations [850]. anniversary [2329, 2566]. annotate [1876]. annotation [1443]. Announce [2097, 2363, 1508, 1551, 2109]. Annual [2072, 2820, 2367, 2640, 1781, 600, 1218, 1469]. Anonymity [740, 1586, 664]. Anonymizing [961]. Anonymous [1347, 2684, 1299]. answer [2205, 2794]. Answering [2794, 2519]. Anthropomorphic [574]. anti [2132, 1669, 2781, 2366]. anti-circumvention [2366]. anti-forensics [2132]. anti-spam [2781]. anti-spyware [1669]. Anticipating [930, 1558]. antiphishing [2886]. Antiterror [982]. Anton [2476]. Any [1632, 1964, 2446, 2252, 82]. Anymore [258]. Anyone [2032]. anyplace [877]. Anything [1778, 1674]. Anytime [877, 1674]. Anytime/anyplace [877]. anywhere [1674]. AOL [180]. AOP [529, 640, 522]. apart [1246]. API [2768]. Appeal [496]. Apple [2511, 480, 2719, 2598]. Appliances [717, 1708]. Application [1237, 812, 1032, 241, 1929, 1080, 2053, 1535, 1643, 1846, 447, 879, 1429, 77, 818, 1070]. Applications [1479, 575, 1854, 690, 772, 882, 304, 707, 185, 2141, 509, 2677, 1221, 2620, 383, 1548, 819, 899, 826, 625, 773, 947, 210, 306, 2277, 188, 189, 835, 999, 1510, 1747, 1521, 2208, 875, 1644, 904, 1175, 1766, 1059, 1036, 3016, 305]. Applying [819, 1643, 2014]. Apprenticeship [2841]. Approach [690, 591, 2098, 174, 309, 667, 2720, 2472, 626, 1931, 382, 1928, 156, 2253, 816, 530, 1191, 1763, 912, 2851, 492]. approaches [2808, 2623, 603, 2824]. Approval [1237]. approximate [2358]. apps [2822]. apt [315]. archaeological [2164]. Arches [982]. architect [2166, 2636]. Architecting [1128]. architectural [1375, 1764, 2450]. architecture [2565, 215, 2349, 1253, 954, 226, 2034, 2209, 2206, 1310, 1595, 1129, 343, 3005, 1748]. architectures [2431, 727, 1307, 2739]. Archival [416]. archives [26]. Archiving [983]. area [2031, 1434]. Areas [276]. Arena [599]. Argentina [838]. Argument [258]. Argumentation [1853]. Ariadne [417]. ARQuake [596]. array [2741]. ’art [2468, 938, 1367, 2003, 866, 1480, 1677, 2458, 824, 2451, 1268, 442, 2505]. Article [1053]. articles [2630, 1254]. artifacts [1544].

Artificial [1323]. Asia [2665]. Asian [1999, 1608]. ASIMO [717]. ASP [1070]. Aspect [521, 523, 532, 528, 530]. Aspect-oriented [521, 523, 532, 528, 530]. ASPECTJ [526]. Aspects [640, 529, 522, 527]. assembled [2247]. assembly [2993, 1154]. Assessing [1500, 3017, 1746, 1251]. Assessment [689, 1604, 2279, 151, 213, 2588, 1108, 2880, 1447, 1670]. asset [1495]. assets [854]. assimilation [2534]. assistance [1877]. Assisting [1957]. Assurance [1479, 783]. Asteroids [179]. Astronauts [330]. astronomer [2581]. Asturian [505]. asymmetric [3004]. asymmetry [714]. Asynchronous [1020, 2649]. ATM [242]. Atoms [480]. attached [226]. attack [1658, 637, 492]. Attacks [515, 739, 2774, 1743, 2585, 1624, 85, 274]. attain [2816]. Attaining [2949]. attempt [2941]. attend [1534]. attention [948, 1648, 947, 949, 2459]. Attentive [946, 943, 944]. Attract [1962]. Attracting [497, 1733, 2885]. attraction [1085]. Attractions [1434]. attributes [2270]. Attuning [949]. auction [2062, 2433, 1211, 2483]. auctions [548, 1185, 2753, 2782, 2568]. Audience [671, 1504]. Audience-specific [671]. audiences [390]. audio [1968, 29, 26]. audit [889]. Auditing [1420, 889]. augmented [752, 596, 1372]. Australia [425]. Australian [464]. Authentication [1186, 1529, 1476, 1583]. authenticity [644]. Author [1935, 1983, 2684]. authoring [1646]. Authorization [413]. Auto [1674]. Auto-ID [1674]. Automated [2050, 2940, 250, 2519, 1877, 1290]. Automatic [175, 2189, 1844, 2677, 811, 2531]. Automatically [2684, 2617, 1272, 1246]. Automating [2651, 2741]. automation [2995, 239]. automobiles [1222]. automotive [2686]. Autonomous [2792, 2247, 2267]. Autopsies [1236]. availability [1423]. available [2950]. avatar [1770]. Avatars [574]. Aviation [430]. Avoid [2022, 1351, 2139, 434]. award [2829, 2106, 2421, 2761]. awards [2288, 1380]. aware [599, 703, 636, 2837, 366, 1989]. awareness [333]. away [890]. Awkshun [1280]. awkward [613]. axiomatic [2935].

B2B [100, 2185, 1764, 1206, 1880]. B2C [1825]. babies [2241]. Baboons [559]. Babylon [2785]. baccalaureate [743]. Back [365, 497, 300, 1778, 559, 916, 2330, 1140, 1753, 1585, 1230, 2597, 1796, 1518, 2309, 2157]. backbone [2]. backdoors [2038]. Backward [1053, 985, 2823, 2793]. Bad [109, 2685, 2799, 1720, 2271]. bag [1819]. Bahgdad [1325]. balance [2779, 939, 1713, 2366]. Balancing [1155, 2948, 1845]. bald [1867]. Ballot [2919]. ban [2546]. bandwagon [1650]. Bandwidth [333, 1479, 1148, 53]. bandwidth-rich [1148]. bandwith [1425]. Bane [374]. Banished [39]. Bankers [425]. banking [438, 435, 992, 440, 1309, 1197, 2407, 1311, 1430]. Bankruptcy [1935]. banks [2886]. ’banner [2371]. banquet [1380]. Bar [1560]. Barriers [913]. base [942]. Based [1028, 2251, 2400, 883, 1013, 230, 1999, 550, 2394, 1266, 1308, 1749, 1063, 2114, 211, 383, 1619, 1131, 1875, 364, 2145, 1650, 988, 2808, 1489, 2059, 382, 1700, 306, 2391, 1276, 713, 2532, 1134, 205, 2818, 1373, 188, 2847, 161, 175, 1398, 2128, 2915, 1176, 1395, 2211, 556, 384, 778, 1691, 1129, 209, 820, 1059, 1154, 1064, 2851, 100]. Basics [40, 2309, 2558]. basis [2935, 1898]. Batteries [1885]. Battery [376]. battle [2335, 313, 2089, 2334, 2836, 2082]. Bay [838, 1434]. Be [2041, 1835, 350, 855, 851, 1302, 286, 2543, 664, 857, 613, 854, 2009, 2265, 991, 449, 1336, 1434, 1169, 2689, 428]. Beam [397]. beat [2752]. Beautiful [2465, 1708]. beauty [366, 2708]. became [789]. become [1185, 1353]. becoming [635]. bed [598]. beds [1594]. Bedside [299]. been [1304]. Before [1258, 1862, 500, 1650]. Beginning [960, 2480]. behavior [1184, 1579, 234, 834, 294, 975, 2711, 2941]. behaviors [1666, 1794]. behind [2335, 2341]. Being [202, 2468, 1452]. beings [323]. Belarus [474]. Believing [1234, 41]. Bell [2352, 2082]. Benchmarks [686]. beneficial [399]. benefit [672, 2568]. Benefits [2041, 235, 1578, 2813, 2980, 2883, 2058, 2388, 253]. Berkeley [2159]. best [2565, 1994, 251, 1731, 1855]. best-laid [1731]. Better [2219, 180, 481, 2041, 219, 2703, 2320]. Better-than-nothing [2219]. Betting [2693, 1608]. Between [786, 1182, 1651, 1316, 2657, 2562]. Beware [1261, 1609, 1029, 717, 697]. Beyond [502, 1586, 2796, 1893, 1883, 147, 1445, 2243, 760, 2519, 2471, 64, 1955, 2558, 1073, 2673, 2785, 820, 986, 2315, 979]. Bias [454, 1962, 796, 2145, 380]. biased [2370, 2528]. biases [2529]. Bible [1628]. bidding [2782, 1211]. Big [516, 480, 496, 588, 1503, 1861, 2868, 1407, 1460, 7]. bigger [2482]. Bill [2638]. Billboard [1805]. billing [232]. Billion [559, 785, 2730, 2970]. billions [2884]. Bills [1456]. biogrids [1444]. bioimaging [341]. Bioinformatics [1479, 1440, 1549, 1445]. biological [2937, 405]. Biologically [1847, 2026]. biology [1441, 342, 2422, 2710, 2772]. Biomedical [1446, 1440]. Biometric [33, 1973, 2530]. biometrics [1772, 191]. biomolecular [2936, 2247]. Birth [541, 2352]. Birthday [838]. Bistro [899]. Bite [515]. bits [2642]. Black [430, 1601, 2655]. blackboards [1966]. Blame [917, 1237]. blank [202]. Blanking [543]. Blasted [1861]. blend [732]. Blending [1243]. Blessing [618]. blind [2871]. Blinking [717, 785]. Block [1529]. Blockbusters [1966]. Blog [1561, 1503, 1733, 2002, 1464, 2822, 2788, 2925, 2958, 2756, 2990]. Blog-Free [2002]. Blogged [1834]. Blogging [1325, 1466, 1465, 2241]. blogosphere [2854, 1462]. Blogs [1504]. blogspace [1463]. Blood [1052, 193]. Blowing [389]. blue [741]. Blueprint [1145]. Blueprints [2993]. Blues [588]. Bluetooth [2721]. Blurred [1257]. BMW [1861]. boards [1857]. boasts [1710]. Body [182, 1214]. Bonds [1027]. book [2830]. Books [1710, 500, 1841, 89, 358]. Boolean [2873]. Boom [659, 1192]. Booms [1754]. Boon [374]. Boosting [639]. boosts [845]. boot [2774]. Boots [276]. Border [1708]. borrowing [1678]. Bos [496]. Boss [1861, 18]. Bosses [2627]. Bot [1934]. both [2481]. Bots [1002, 1162, 1456]. ’bottom [2730, 391]. Bottom-up [391]. boulders [1780]. Boundaryless [577]. bounds [2591]. Bounty [942]. bowls [2007]. Box [1835, 1601, 223, 2655, 1740, 2473, 2561, 2919]. Braille [257]. Brain [1709, 618, 959, 1140, 1911, 1777, 2440, 2202, 950]. Brainy [559]. branding [1644]. BRAP [2439]. breaches [1815]. Break [959]. Breakdown [147]. Breaking [2615, 504, 276, 2387]. Breakthrough [2354, 2860]. bricks [1188, 1205]. Bridging [1541]. bright [2662, 202]. Bringing [974, 1796, 1105]. Bristles [1911]. Britannica [1911]. Broadband [839, 720, 2653, 2966, 2374, 518]. broadcast [30, 29, 2032, 26, 1603]. Broadening [2997, 2049]. broader [174]. Broken [3000]. brokerage [777, 1338]. Brother [7]. brothers [2970]. browse [1876]. Browser [2869, 904, 307, 2767]. browsers [2806]. browsing [1050, 1671]. Brush [453]. Brush-on [453]. ’bubble [2258, 1557]. bubbles [986]. Buckets [412]. Budget [1503]. Budgeting [1800]. buffer [1743]. bugging [4]. Buggy [1887]. bugs [1262, 1199]. build [2292, 920, 2201, 2882]. Building [1096, 411, 2677, 3, 579, 976, 2014, 1339, 1473, 1539, 270, 977, 382, 2558, 1233, 419, 530, 2412, 665, 7, 166, 1748]. Builds [1052, 3000, 2598]. Built [1387, 2735]. Bulk [3005]. bullet [1656, 1629, 1069]. bulletin [1857]. bullets [2488]. Bullies [1585]. Bumper [785]. Bundling [438, 1116]. Buoys [1214]. Burdens [1633]. bureaucracies [285]. Burning [866]. burst [1229]. Business [772, 150, 319, 840, 1217, 1685, 1836, 2328, 425, 1127, 1003, 1655, 147, 2277, 1599, 2276, 1760, 1745, 1902, 704, 278, 457, 517, 589, 1080, 919, 984, 1261, 1411, 1327, 1532, 1610, 1480, 1780, 1914, 1984, 2069, 2242, 2107, 2179, 2381, 2832, 2930, 2637, 2666, 757, 591, 72, 1471, 2816, 1949, 1362, 714, 101, 2763, 776, 1121, 2004, 1187, 2458, 1947, 1897, 2864, 2620, 880, 976, 1092, 36, 1535, 2573, 1189, 633, 2966, 2320, 817, 271, 1724, 771, 1767, 1338, 216, 154, 4, 1765, 387, 2036, 709, 1188, 2321, 155, 1010, 2965, 933, 606, 1313, 1763, 1766]. business [1210, 2406, 153, 114, 1761, 2181]. Business-oriented [1127]. business-to-customer [704]. businesses [2168, 1212]. Bust [1754]. Busting [1672]. Button [639]. buy [2073]. Buzz [1805]. Bye [588, 2890]. Byte [2618, 2680, 2527, 2713, 3013, 2479, 2557, 2650, 2844, 2745, 2502, 2526, 2587, 2712, 2777, 2979, 3012]. Byzantine [2586].

C&C [1531]. C. [2702]. C2C [873]. caching [787]. CACM} [1783, 2328, 2822, 2788, 2925, 2353, 2958, 2506, 2538, 2569, 2459, 2599, 2630, 2992, 2959, 2789, 2823, 2724, 2757, 2329, 2338, 2535, 2990]. Café [1560]. Caging [886]. Calculating [39, 2631, 1401]. calculus [2651]. Calendar [238, 394]. Call [1778, 1983, 2218, 64, 318, 639, 1742, 642, 1839, 2135, 2402, 753, 1953]. Calling [1774, 1214]. Calls [785]. Cam [430, 458].

Camera [1325]. Cameras [937]. campaign [1892]. Campaigns [1887]. Campus [376, 692, 1434, 1887]. Can [300, 1119, 2022, 1806, 2009, 358, 1318, 1185, 2816, 2984, 554, 1419, 486, 2061, 2032, 2689, 2093, 1629, 1925]. CAN-SPAM [2093, 1629]. Canaan [129]. Canada [2063, 861]. Canadian [2551]. Canary [1052]. cancellations [1576]. Cancer [1632]. Candidates [2097, 2363, 1508, 1551, 2109]. can’t [193, 2752, 2239]. Capability [688]. Capacity [735]. Capital [1609, 2933, 1201, 665, 268]. Capitalism [1281]. captures [2719]. capturing [170]. carbon [2245]. Card [2308, 509, 1741, 2820, 755]. cardiac [2711]. Cards [586, 1346, 594, 3015]. care [2454, 2090, 2187, 2656, 2191, 2221, 115, 1798, 1773, 1350, 875, 1440, 1593, 2812, 1020]. Career [237, 393, 788, 1683, 551, 471]. cargo [445]. Carnegie [2202]. Carrier [1410, 1642]. Cart [1162]. carving [2647]. cascades [1298]. Case [683, 150, 425, 2268, 2875, 2316, 723, 2719, 291, 2914, 1712, 254, 14]. cash [2060, 1097]. casinos [200]. Cat [480]. catalog [100]. categories [1874]. cause [1223]. causes [1277]. Caustic [400]. cautionary [1731]. Caveman [1982]. CBD [822]. CD [739]. Celebrate [1586, 1380]. celebrates [2786]. Celestial [2]. Cell [87, 515, 496, 1777, 2191]. Cell-phone-free [87]. Cells [1934]. cellular [2266]. censorship [1155]. center [2355, 2202]. Centered [1529, 1208, 1445, 1643, 156, 1555]. centrality [2353]. Centralization [1518, 447]. Centre [1710, 1282]. centric [2031, 2592]. Century [371, 2313, 2090, 347, 103, 1531, 1840, 2466, 2499, 1678]. CEO [2566]. CEOs [739, 2002]. Certificate [1754]. certification [1377, 1431]. certifications [1416]. Certified [1378]. chain [1674, 444, 446, 441, 2227, 2208, 445, 2561, 1113, 443]. chains [1626, 2076]. challanges [208]. challenge [2249, 371, 974, 115, 2814, 1378, 674, 437, 2267]. Challenges [1772, 1226, 1639, 1197, 1756, 1598, 445, 724, 1168, 882, 2050, 1640, 1011, 2049, 2185, 2845, 2645, 2023, 1487, 876, 2116, 815, 1695, 2058, 1536, 123, 623, 1064]. Challenging [2485]. Champ [1302]. Champs [1528]. Chance [481, 515]. Change [431, 692, 2055, 1683, 693, 856, 2223, 2747, 2429, 2606, 229]. change-management [2429]. changed [2423]. changeovers [498]. Changes [781, 493, 2835, 1928, 2414, 471, 2482]. Changing [1974, 2196, 1045, 2451, 490, 1600, 116]. channel [2427, 756, 1048, 874]. channels [295, 276, 2715, 1360]. Chaos [1367, 1327, 924]. chaotic [2251]. chapter [2961]. chapters [1516, 1517]. Character [454, 1121, 1628, 135]. Characteristics [826, 688, 208]. Characterizing [254]. characters [6]. charging [232]. charter [1516, 1517, 1827, 2086]. charts [2905]. chasing [2130]. chasm [379]. chatterbox [204]. Cheap [916]. Check [1528, 845]. checkable [2707]. Checked [1002, 1214]. checking [1929, 2974, 1647, 2479]. checks [2185]. chemical [2246]. Chess [1528]. Chic [1777]. Chicken [147]. child [2803]. Children [1489, 2691, 1486, 1488]. Chilling [640]. chilly [2773]. China [1574, 2534, 1609, 1935, 1572, 2852, 1953, 2, 64, 515, 717, 916, 1359, 1358, 1568, 2658, 2266, 1573, 473, 1366, 1528, 1560, 1777, 1934, 1566, 1575, 1179, 1556, 1569]. Chinese [1570, 1571, 2021]. Chip [147, 480, 659, 1052, 87, 219]. Chipped [785]. chips [2876]. Choice [1754, 2286]. choiceboards [1471]. choices [2185]. Choking [937]. choosing [999]. Chord [1077]. Chrome [2869]. Church [2041, 2003]. Churchman [2846]. Cinematic [1027]. CIO [1633, 71]. CIOs [1065]. circle [322]. circuit [230, 2877]. circuits [956]. Circular [2587].

Circumvention [542, 2366]. citation [2880, 1746]. citations [2311]. cite [2239]. Cities [1241, 1244, 1242]. Citing [1609]. citizen [2652]. citizens [1355, 900, 1623]. city [754]. Civil [660]. Claremont [2802]. Clarion [1953]. Class [2383, 87, 84, 2239]. Classes [88, 2352]. Classical [2003]. classics [1710]. classification [2228]. Classifying [731]. Classroom [640, 1303, 562, 1987, 2752, 1318]. Clean [560, 2541, 1248, 855, 2311]. Cleaning [838]. clear [855]. cliché [2832]. Click [453, 2910]. Clicking [659]. clicks [1188, 1205, 1191]. Climate [127, 79]. climbing [2788]. climbs [18]. Clocking [1585]. Clogging [276]. Cloned [147]. clopen [2819]. Close [453, 1384]. Closed [180, 2819]. closer [1648, 2549]. Closing [1080, 322, 931, 373]. clothes [537]. Cloud [2460, 2870, 2736]. Clouds [807, 1528]. Clustering [1874]. clusters [2270, 2356]. CMM [1072]. CMU [1834]. Coat [1887]. Code [560, 1410, 1367, 1303, 276, 2547, 1217, 1062, 2171, 1063, 2346, 1421, 2520, 2465, 2036, 1427, 1991, 2866, 561]. Code-breaking [276]. code-review [2171]. codes [2708]. Coding [481, 1862]. coexistence [399]. Coffee [106, 430, 2499]. Coffee-Cam [430]. Coffers [106]. Coffin [219]. Cognitive [1237, 134, 2529]. Coincidental [1300]. Coke [1386]. Cold [1062, 2774]. cold-boot [2774]. collaborate [1902]. collaborate/integrate [1902]. Collaboration [582, 581, 1362, 895, 733]. Collaborative [475, 752, 575, 573, 2033, 825, 570, 571, 2115, 567, 2207, 846, 194, 2649, 626, 844, 2412, 2496, 443, 952]. Collect [1586]. collection [422, 897, 2970]. collections [2553, 2376]. collective [666, 1413]. college [2576]. Collegiate [803]. collision [314]. colonoscopy [2727, 1512]. Color [1912, 1881]. colors [1461]. column [277]. Com [318, 937, 179]. Combat [866]. combined [2720]. Combining [65, 2202].

Comcast [2850]. come [920, 2354, 2614, 2457]. Comes [516, 1983, 1198, 49, 55]. Coming [1118, 1236]. Comma [1322]. Comma-free [1322]. commander [2745]. Commendation [516]. comments [232]. Commerce [39, 682, 566, 1198, 1168, 1929, 591, 1186, 1579, 482, 696, 295, 1651, 37, 1208, 2645, 1092, 211, 756, 308, 388, 1171, 1173, 1428, 189, 1200, 1552, 264, 1203, 941, 1207, 1519, 1744, 977, 1169, 143, 1172, 1175, 386, 1174, 710, 1475, 1111, 624]. commercial [1548]. Commercialization [247]. commercializing [933]. commission [2682]. Commitment [2021, 398, 443]. committee [2766, 2312]. Common [1410, 2313, 1826, 1189, 1157, 1223]. Commons [1862, 1826, 2351]. commonsense [161, 2651, 368]. commonsense-based [161]. Communicating [960]. Communication [147, 659, 687, 578, 2116, 386, 2806, 947, 1931, 477, 1770, 2158, 675, 2849, 155, 1898, 1020, 2919, 1556, 2531]. Communications [2969, 2820, 2853, 2536, 2538, 2630, 2268, 1229, 615, 536, 2668, 655, 1120, 1179, 2690]. Communism [1281]. communities [670, 1090, 1798, 790, 2095, 672, 2919]. Community [579, 384, 1719, 671, 1466, 920, 1243, 419, 665, 1485]. Community-based [384]. compaction [1060, 1058]. companies [1570, 2751, 70]. companions [702]. Companionship [148]. Company [1121, 602, 2087, 1771, 1521, 3021]. compared [2264]. Comparison [680, 1206, 1550, 2228]. compatible [2781]. compelling [1034]. compete [2267]. Competing [2321]. Competition [1912, 1733, 2589]. Competitive [1912, 1106, 1090]. Competitor [1675, 508]. competitors [2947]. Compiler [2674, 2843, 2842, 2755]. complaint [2949]. complaints [2411]. Complementary [26]. completeness [1013]. completion [2554].

Complex [917, 1592, 2023, 2557, 382, 1594, 844, 2317, 1595, 1596, 2977, 818, 1593]. complexity [2679, 2225, 841, 2678, 1110, 1317]. Compliance [2093, 2256]. complications [759]. Component [883, 211, 206, 205, 1398, 819, 2532, 816, 2317, 1071, 209, 820, 1154, 1064]. Component-based [883, 211, 205, 1398, 2532, 209, 820, 1154, 1064]. componentization [1124]. componentizing [77]. Components [813, 1163, 425, 814, 2223, 1554, 207, 2010, 923, 2244]. composability [726]. Composable [2501]. compose [1135]. Composing [525]. composites [1361]. composition [525, 1545]. compositions [1959]. Comprehension [1410]. comprehensive [2210]. compression [1063]. computability [2572]. Computable [2041]. Computation [2245, 979, 656, 96, 252, 343, 766]. Computational [1441, 1983, 2645, 341, 2553, 553, 2796, 2717, 2489, 1740, 1842]. Computer [656, 367, 1370, 1549, 786, 1141, 1028, 1435, 1457, 1734, 1806, 2544, 843, 808, 2958, 748, 685, 287, 2497, 2888, 3007, 1575, 1225, 3006, 2292, 2352, 2045, 2112, 2394, 1614, 427, 1577, 1534, 398, 1927, 1969, 1564, 1735, 1082, 2804, 2290, 2833, 2867, 136, 257, 18, 2576, 2142, 133, 2935, 3009, 781, 2632, 1738, 844, 595, 721, 2925, 1263, 1838, 2734, 1286, 1287, 2350, 760, 1856, 434, 1687, 1563, 2355, 2498, 2467, 372, 360, 2180, 743, 2967, 2821, 2723, 1485, 2988, 343, 1537, 1820, 2026, 349, 1831, 1882, 2127, 1343, 1951, 2918, 2475, 2895, 1840, 2899]. computer-aided [1951]. computer-mediated [349]. Computer-related [1575, 2350]. Computer-Supported [843, 2394, 844]. Computers [342, 107, 1303, 257, 318, 374, 2248, 126, 1544, 9, 2446, 1542, 337, 2815, 1539, 945, 1246].

Computing [2347, 2733, 2931, 960, 1215, 1586, 1561, 1709, 1835, 2195, 2897, 2999, 886, 2544, 2771, 2722, 2080, 1122, 2244, 2982, 2249, 98, 1230, 2274, 1085, 849, 1868, 2939, 552, 1441, 882, 634, 1340, 213, 2748, 371, 2161, 622, 2779, 2870, 1507, 1786, 2679, 877, 1142, 2487, 348, 1442, 636, 966, 880, 2667, 1873, 1548, 536, 1341, 819, 222, 1079, 878, 2765, 2460, 947, 934, 1035, 964, 881, 1446, 2678, 2552, 1136, 535, 790, 2675, 2138, 2539, 1545, 2490, 923, 2013, 2158, 1086, 2314, 876, 2252, 2250, 1756, 2710, 2872, 2243, 89, 368, 2461, 623, 2246, 2043, 2923, 820]. computing [1997, 94, 2754, 2887, 2885, 2950, 2786, 2342, 2857, 2477, 1854, 1808]. concentration [2751]. Concept [2720, 1887, 1070]. conception [2381]. Conceptual [1128, 1957, 1132, 1361, 143, 3006]. Conceptualization [573]. Conceptually [2591]. Concerns [3, 2040, 525, 842, 524]. Concerts [1302]. concise [855]. concrete [2865]. concurrency [2577]. Concurrent [244, 1093, 2775, 2704]. Concurrent-convergent [1093]. conductor [28]. Conferences [2754, 1534, 1460, 1659, 2212]. conferencing [780]. Conficker [2932]. confidence [1978, 845]. confidentiality [902]. Configurable [652]. configure [929]. conflict [2268, 1227, 1048]. Confusions [2913]. congress [406, 564]. Connected [182, 1589, 1686, 1456, 126, 68, 302, 499, 720, 963, 888, 1078, 1304, 1216, 1388, 1482, 653, 1985, 1782, 1891, 2154, 2070, 2241, 2329, 2423]. Connecting [257]. connection [2789]. connections [2828]. cons [276]. Consciousness [1303]. Consequences [1609, 1346]. conservation [2242]. Consider [1410, 1156]. considerations [872, 1019]. considered [1524, 1958, 782, 538, 1137, 863]. considering [948, 518]. Consistent [431, 2641]. consortium [299].

Conspicuous [918]. constraints

[1772, 2591, 531, 2902, 254]. Constructive [603, 2915]. consulting [1771, 1093]. Consumer [793, 839, 2140, 1207, 1718, 1184, 707, 975, 1697, 1696, 1700, 2818, 2407, 1233, 1973, 2254, 1616, 2562]. Consumer-Created [839]. Consumer-perceived [1207]. Consumers [1753, 834, 1552, 1664]. contact [1037]. ContactMap [675]. Contagious [2930]. Contemporary [2824, 1199]. Content [867, 839, 1028, 2791, 973, 1203, 3016, 2141, 962, 890, 1621, 1797, 845, 1979, 2096]. contents [2720]. Contests [803]. Context [1540, 2984, 599, 703, 636, 1724, 2061, 1801]. context-aware [599, 703, 636]. contexts [1792]. Contextual [1640]. contextualized [2681]. Continuing [662, 204, 2295]. continuous [2914]. Continuum [343]. contracting [1247, 1553]. Contracts [1125]. contrary [2481]. contributing [1976]. Contribution [1778, 3017, 2389]. contributions [2043]. contributor [407]. contributor-run [407]. Control [258, 739, 1346, 183, 2636, 2841, 1696, 2570, 2694, 2791, 1167, 2903, 2904, 1251, 1990, 12, 356, 1255, 492]. Controlled [712]. Controlling [135, 615]. conundrum [836]. convenience [234]. Conventional [127, 804]. convergence [2582]. convergent [1093]. conversation [2938]. Conversational [184, 189, 78, 188, 1222]. conversations [667]. conversion [2911]. Cooker [807]. Cookies [662, 481, 147, 400, 1741, 513]. Cool [2002]. cooperation [1926]. Cooperative [843, 200]. Coordinating [245]. Coordination [2430]. Coping [493, 1048, 2008]. COPLINK [892]. Copy [739, 1615]. Copying [496]. Copyright [560, 542, 1326, 1962, 969, 620, 2220, 1134, 2265]. Copyrightable [311, 2916]. Copyrighting [1917]. Coralling [1934]. CORBA [2494]. core [544, 2407]. Corner [299, 2567, 2458, 2536, 2504, 2690, 2820]. Corporate [2232, 550, 730, 1498, 249, 1472, 2917, 1701]. Correct [1053]. correcting [2617, 2708]. Correction [377, 2709]. correctness [765]. Corroboration [1028]. Cost [1347, 1609, 213, 612, 797, 2059, 35, 2657, 2323]. Costs [2388, 949, 1004, 672]. COTS [2038]. Could [2348, 2041, 2040, 1169]. Couldn’t [1683]. Count [1163, 1303, 1560, 2002, 1057, 2515, 605]. counterfeiters [200]. Countering [1265]. Counterpoint [2466, 2670, 2834, 2546]. Counting [1780, 1261, 2323]. countries [1728, 2099, 2369, 420]. Country [1982, 1159]. Courriel [1118]. Course [680, 1410, 1962, 562, 995, 1105]. Courses [640, 1276, 1710, 1356, 3007, 624]. courseware [1094]. Coursework [1684]. Court [481, 106, 1612]. courts [2088]. covenant [1601]. Cover [1962]. Covering [1962, 2979]. covers [2334]. Cowboys [1684]. Coyote [515]. CPOs [2751]. Crab [639]. Crafting [468]. cranial [1372]. Crash [299, 785]. Crash-free [299]. craziness [2930]. Create [1028, 1471, 1209]. Created [839]. creates [1543]. Creating [920, 1090, 1194, 1191, 1848, 1485, 1622, 624, 2198, 1642, 2822, 6, 828]. creation [385, 333, 1306, 1492, 2294]. Creative [1456, 2827, 829]. Creativity [823, 1778, 824, 830, 2310, 2340, 160, 826, 825, 142]. Credentials [1347, 2042]. Credibility [1806]. Credit [763, 2308]. crime [1927, 2918]. Criminal [1625, 1272]. Crisis [1738, 2467, 2933, 1963, 2955, 1523]. Criteria [887, 208, 616]. Critical [1028, 1215, 2035, 80, 1221, 2881, 995, 2121, 2412, 2389]. criticality [1290]. critically [2826]. CRM [978]. Cross [2782, 1292, 1110]. Cross-bidding [2782]. cross-cultural [1292]. crosscutting [525, 531]. crossing [379]. Crossroads [2551, 1092, 1171]. crouching [1837].

Crowd [2694, 87]. crowdsourced [2838]. Crowdsourcing [2992]. crucial [2422]. crummy [2244]. Crustacean [838]. Crypto [763, 677]. cryptographic [1992]. cryptography [346, 103]. CS [2422, 1141, 1410, 1237, 1215, 1754, 1709, 1912, 1962, 1935, 2834, 2162, 1263, 1710, 1841, 2551, 2989, 3007, 2990]. CSTA [2988]. CTO [2493, 2610, 2579, 2517, 2870]. Cube [106, 1329, 2404]. Cube-icles [106]. Cultural [834, 1734, 759, 580, 1640, 1292, 780, 74, 1801]. culture [2747, 631, 1109, 287]. Cure [148, 3019]. cure-all [3019]. Currency [319]. Current [887, 435, 1728, 835, 2506, 2037, 2273, 1111, 2018, 2417]. Curricula [1237, 1187, 248]. Curriculum [1215, 2466, 72, 1514, 2958]. Curse [618]. cursor [18]. Curtain [496]. Customer [983, 2003, 1208, 704, 1250, 1013, 940, 1450, 1428, 978, 977, 1028]. Customer-centered [1208]. customer/supplier [1250]. customerization [437]. Customers [454, 1141, 1337]. Customizable [426]. customization [2053]. Customizing [2646]. Cut [257]. cuts [126]. Cutting [2185, 692]. CWOT [982]. Cyber [1269, 983, 982, 2351, 1268, 1426, 2973, 1822, 951, 1814, 2849, 395, 847, 492]. Cyber-commons [2351]. cyber-insurance [2973]. Cyber-Risk [983, 951]. Cyberaccess [1472]. Cybercrime [2736]. cybercrimes [3020]. cyberinfrastructure [2681, 3001]. cybernetics [92]. Cyberpublishing [320]. cybersecurity [2863]. Cyberspace [868, 2347, 314, 2283, 644]. cyberspaces [668]. Cybersphere [1099]. Cyborg [692]. Cyborgs [324]. Cycle [983, 919, 2188, 2059, 2276, 1275, 2298]. Cycles [1982]. Cyronic [807]. czars [40].

D [1619, 1514, 1621, 750, 1371, 1615, 1620]. Dallas [1684]. Damage [2570]. dame [2745]. Dancing [886]. dangerous [2687]. Dark [866, 2736, 2593, 2550]. DARPA [1633, 2296, 2267]. Darwinian [2732]. Dash [397]. Data [1013, 516, 1586, 1983, 1148, 541, 2944, 774, 83, 1104, 2383, 1793, 1150, 689, 769, 1608, 2960, 772, 2565, 1984, 2524, 2656, 926, 81, 1284, 2132, 2611, 883, 2984, 989, 2681, 2141, 2942, 892, 1181, 1799, 2945, 2356, 1273, 1498, 1749, 336, 768, 2299, 1100, 2280, 899, 1156, 2142, 773, 712, 2775, 225, 476, 1998, 898, 1390, 2868, 1879, 902, 487, 1724, 2525, 655, 2927, 2376, 1152, 2648, 162, 228, 1350, 2408, 897, 227, 1310, 2355, 1251, 116, 1606, 1550, 2037, 2970, 1177, 912, 1701, 1179, 1314, 2074, 2916, 952, 1115]. Data-driven [769]. Data-intensive [1150, 2141]. Database [2802, 1734, 496, 2608, 2738, 1604, 246, 2591, 1795]. databases [770, 2839, 391, 800, 742, 2471, 1865, 2584]. Dataless [1633, 1524]. datapath [2314]. David [1684, 2938]. Days [982, 179]. DBMSs [2890]. de-charter [1516, 1517, 1827, 2086]. De-escalating [2951]. DEA [1499]. dead [2830]. deadline [2990]. deadline-driven [2990]. deal [678]. Dealing [2933, 2223]. Dear [983]. dearth [1460]. death [2352, 350]. Debate [867, 564, 939, 2762, 1017, 518, 204, 1016]. Debating [19, 2928]. Debugging [2769, 2974, 1040]. debuts [2665]. decay [908]. Deceived [1190]. decentralization [447]. decentralized [1465]. Deception [740, 2531]. deceptive [1666, 2958, 1272]. Decision [2024, 2113, 570, 2279, 893, 2320, 2145, 1404, 1700, 780, 2231]. decision-support [1404]. decisions [629, 269]. Declaration [515, 520]. Declarative [2976]. decline [463]. decompression [1062]. deconstructing [82]. decrease [2455]. deduction [250]. Deep [2960, 162, 161, 163, 165, 166, 164, 404, 1897, 2169, 2342, 2540]. Deeper [1347]. deeply [347]. defect [1411]. Defects [1734]. Defend [1586]. Defense [1237, 959, 2832, 1269, 1553, 198, 1268, 42]. defenses [1675]. Deficiencies [955]. Define [1586, 1655]. Defined [1862]. Defines [2383, 201]. Defining [542, 1976]. Definition [431, 1053]. definitions [1129]. Definitive [2003]. Degree [618, 1157]. Degrees [2022, 2274]. deja [2221, 2336]. deja-vu [2221]. Deliberating [130]. Delightful [2502]. deliver [2691]. Delivered [2003]. Delivering [1538, 895]. Delivery [680, 242, 1312, 1797, 1593]. Demand [1529, 25, 24, 530]. demise [354, 1304]. Democracies [300]. Democracy [377, 1504, 1778, 1806, 280, 1467, 283, 286, 1096, 284, 224, 1720, 287, 285, 281]. democratic [1057, 288]. demodularization [530]. Demographic [2414]. demonstrating [599]. demonstration [58]. demonstrational [56]. Demystifying [1357]. Denial [257, 85]. denial-of-service [85]. Denning [1077]. density [225]. department [1355, 42]. Departmental [2434]. dependability [914]. dependable [2462, 2740]. Dependency [2011]. Depends [1862, 303]. Deployment [640, 814, 1928, 14]. depression [2931]. Depth [1237]. derived [1701]. Deriving [2813]. describe [1963]. desert [1758]. Deserves [220, 1504]. Design [3, 640, 1435, 1281, 1529, 2188, 573, 3008, 569, 843, 2520, 965, 567, 436, 1154, 671, 1322, 2656, 882, 1543, 1185, 948, 2071, 2622, 1486, 969, 549, 1208, 1392, 1369, 2014, 36, 621, 2440, 625, 1488, 2426, 2768, 2462, 626, 1877, 1643, 391, 157, 844, 2532, 156, 1173, 1555, 949, 1289, 530, 2139, 2167, 527, 2846, 1224, 1809, 879, 1033, 829, 1402, 1882, 2315, 2434, 2548, 1899, 1616, 2595, 952]. Designing [1597, 2141, 2779, 1639, 3009, 2227, 2448, 1372, 267, 572, 1175, 2483, 711, 1899, 2409, 2495, 666, 2446, 103, 2890, 724]. designs [84]. desire [1998]. Desk [318, 2332]. deskill [1736]. desktop [326, 2252]. destination [1750]. detail [1648, 59]. Details [1258]. Detecting [397, 1314, 2711, 1272]. Detection [1743, 81, 1691, 492, 2531]. Determine [1028]. determines [2879]. Determining [797]. deterministic [2808, 2807]. Deterring [2281]. develop [2158]. developed [2099, 2857]. developer [1114, 630, 2778]. Developers [148, 1400, 1494, 2078, 2324, 1731]. Developing [813, 362, 459, 1140, 1096, 1957, 1676, 1181, 2898, 2804, 727, 2369, 420]. Development [763, 1141, 1003, 1119, 1655, 687, 683, 688, 1710, 1282, 241, 1830, 909, 1182, 2488, 1399, 2875, 244, 2316, 652, 1999, 2264, 2364, 2607, 1392, 972, 2012, 2168, 1130, 2874, 2453, 212, 1643, 2838, 2532, 2008, 2010, 2253, 1247, 2814, 409, 2529, 1398, 2914, 2122, 2883, 2276, 2643, 2657, 1976, 2232, 2009, 1395, 2393, 2143, 1427, 556, 1474, 674, 209, 2624, 2948, 1391, 1566, 2324, 1154, 1064, 1394, 1317, 812]. developments [2555]. Device [1639, 1170]. devices [2795, 2743, 992, 2245, 2815, 340, 1693, 1089, 2247, 2744, 432, 1038, 2096]. devil [1183]. Devotion [1934]. Diabolical [618]. Diagnosing [1817]. diagnosis [341]. diagnostics [1451]. diagramming [1100]. Dial [1118]. Dial-A-Dolphin [1118]. dialogue [662, 185, 2608, 12, 171]. diamonds [2839, 2507]. Diary [12]. diaspora [2626]. Did [2727, 1712, 3018]. Didn’t [1435]. differ [1556]. Difference [786, 2603, 790, 2661, 1881]. differences [834, 780, 1522, 778]. Different [1141, 1196]. difficult [2732, 2696]. Diffusion [1569, 2316, 1772, 930, 1570, 1744, 2369]. Digging [2328]. Digital [321, 221, 90, 320, 400, 561, 482, 787, 662, 961, 1054, 1165, 1284, 1389, 1458, 1757, 1562, 1658, 1964, 2042, 1864, 2219, 2308, 2132, 2439, 2400, 1043, 300, 481, 1459, 1684, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1866, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2081, 2126, 2147, 2282, 2303, 2322, 1717, 1786, 403, 1819, 754, 325, 891, 410, 965, 2714, 1280, 679, 2566, 117, 415, 2750, 293, 1541, 313, 414, 411, 2161, 1816, 361, 1243, 907, 2044, 2101, 2173, 2190, 2213, 2233, 2259]. digital [2563, 312, 336, 413, 1828, 899, 1079, 1023, 1645, 712, 1721, 1104, 731, 406, 1354, 407, 2377, 419, 418, 409, 1802, 1715, 412, 405, 1022, 1091, 1821, 424, 904, 2705, 1315, 805, 1448, 416, 423, 1384, 420, 1116, 421, 1718, 87, 620]. digitization [1116]. Dijkstra [810]. Dilemma [1633, 2073, 312]. Dimension [740, 1067]. dimensional [1181]. dimensions [2358, 2666, 2763, 2864, 2573, 2966]. diminishing [1144]. diodes [2884]. direct [192, 2740]. direction [2834]. Directions [1783, 1766, 15, 250]. Directive [969]. dirt [2839]. dirty [812]. disabilities [1404]. Disability [1681]. Disability-related [1681]. disappearing [331, 1539, 1537]. disappointment [2139]. Disaster [2656, 246, 541, 2119, 2114, 1108]. disasters [1558, 2214]. Disband [1709]. disbursed [1930]. Discipline [1387, 1411, 1054, 1906]. Disciplined [1862]. disciplines [1341, 2314]. Disconnection [1982]. disconnects [2144]. Discount [1609]. Discourage [1347]. discover [2960]. Discovery [1056, 1904, 1949, 2817, 2310, 2738]. discursive [644]. Discussing [522, 1994]. diseases [39]. dish [1922]. dishonesty [1714]. Disintermediation [220, 90]. Disk [1964, 2741, 2799, 2404]. Dismiss [1303]. dispatching [997]. Display [64, 948, 2445]. displays [411, 2646]. Disposable [430]. Dispute [1409]. Disputed [2021]. dissemination [2715, 1603]. dissolve [2136]. dissolving [331]. Distance [19, 2269, 1276, 2987]. Distinct [2943]. Distinct-value [2943]. distinguished [2043, 2134, 2403]. Distinguishing [2880].

Distraction [1076]. Distributed [582, 684, 2525, 1990, 2007, 735, 426, 728, 1182, 2875, 1468, 2364, 1297, 2975, 1046, 2012, 2874, 1554, 235, 2008, 2010, 188, 1851, 2009, 2011, 1748]. Distributing [2524]. distribution [2240, 756, 2954, 973, 122, 1091]. distrust [2783]. Disturbing [1479]. disuse [1368]. diverge [2855]. divergence [2814]. diversity [1484, 2763, 2563, 1854, 2235]. dives [2960]. divide [2750, 361, 1354, 632, 1022]. dividend [2770]. DL [408]. DMCA [960, 565]. DMM [2951]. DNA [1322, 39, 496, 866, 2247, 1434, 1982, 912]. DNS [3003]. DNSSEC [2183]. Do [1119, 1435, 1888, 1983, 639, 807, 692, 2947, 1200, 1552, 1613, 1934, 2984, 554, 2191, 2910, 1057, 1697, 2515, 1665, 486, 2145, 605, 2061, 1482, 2457, 1664]. DOA [2545]. Document [2548, 1376, 2207]. documentation [2590]. documents [757, 628, 119]. DOE [2]. Does [2875, 2171, 2299, 1647, 1770, 528, 1659, 1118, 608, 1381, 1881, 151, 2349, 1810, 1697, 1963, 2916, 2888, 2148, 2265]. Doesn’t [1003, 1435, 515, 2833]. dogs [833]. doing [2628]. Dollars [959]. Dolphin [1118]. Domain [1862, 2003, 64, 376, 2253, 2897, 279, 314, 1742, 1305, 531, 2034, 1924]. domain-specific [531]. domains [1871]. Dominance [1346]. dominate [1359]. dominates [39]. domino [1294]. Donald [2468, 2491]. done [2689]. Don’t [454, 660, 1053, 1003, 960, 1237, 1347, 1367, 1457, 1303, 1586, 1633, 1504, 1609, 1888, 2022, 1962, 1935, 1983, 318, 1140, 2973, 159, 2910, 1359, 2971, 1326, 1057]. Doom [107, 1329]. door [217]. doorman [811]. doors [407]. Dot [318, 937, 179, 276, 2661]. Dot-Com [318, 937, 179]. dot-cons [276]. dot-org [2661]. Dot.con [982]. dots [257]. Doubts [640, 1280]. down [2348, 2516]. Downloads [1866, 2081, 2322, 1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2044, 2101, 2126, 2147, 2173, 2190, 2213, 2233, 2259, 2282, 2303]. Downs [1503, 212]. Downturns [1709]. DQ [1724]. Draft [1325]. Drag [453]. Drained [1777]. dreams [370]. Dresser [959]. Dripped [430]. Drive [1028, 2534, 1823]. driven [2279, 651, 1764, 1767, 816, 2293, 769, 2990]. Driver [1076, 1434]. Drivers [1195, 2302]. drives [2886, 2799, 2861, 1880]. Driving [200, 1389, 2847]. DRM [968, 1003, 1028, 966, 970, 1652, 971, 967]. Drop [481]. Drops [430]. drought [893]. Drudgery [404]. Drummer [1366]. Dual [660, 2427]. dual-channel [2427]. duality [2064]. ducation [1098]. Due [1888, 314]. Dumb [107]. dumbing [718, 621]. Dummy [785]. during [2688, 2864, 888]. DVD [91]. Dynamic [2094, 2451, 2231, 1375, 2033, 1988, 2737, 1989, 3016]. dynamics [2251, 2475, 1796, 878, 2476].

e-banking [438, 435, 1197, 1430]. E-Bay [838]. e-books [500, 358, 89]. E-Bullies [1585]. e-business [776, 1187, 976, 1189, 271, 1724, 1765, 387, 709, 1766, 1210, 1761, 1760]. e-businesses [1212]. e-cash [1097]. E-casinos [200]. E-catalog [100]. e-commerce [1929, 591, 1186, 1579, 482, 696, 295, 37, 1208, 2645, 211, 308, 1428, 189, 1198, 1200, 1552, 264, 1203, 941, 1207, 1519, 1744, 143, 386, 710, 1111, 624, 39, 682, 566, 756, 388, 1475]. E-commercializing [933]. e-CRM [978]. e-democracy [1096]. E-ducation [1098]. E-Earache [1386]. E-epistemology [1000]. E-Flush [1076]. e-government [2750, 1342, 1728, 906, 1366, 1525]. e-journals [1295, 481]. e-learning [670, 1318, 2021]. E-loyalty [1106]. e-manners [2627]. e-marketplace [1202, 1675]. e-markets [2302]. e-medical [511]. E-migrating [2260]. E-monitoring [1975]. e-participation [2621]. e-payments [2185]. e-recruiting [2206]. E-Rescue [1708]. e-rulemaking [896]. e-science [1146, 1150, 1703]. e-service [1014, 1206, 1010]. e-services [1011, 2981, 309, 1015, 1007, 1012]. E-speak [1044]. e-strategies [1599]. e-supply [1626]. E-Tending [126]. E-Tooling [739]. e-transformation [1572]. e-village [1239]. e-Voting [1479, 2108, 2304, 1423, 2660]. e-xplained [1044]. E2A [131]. EAI [928]. Ear [299]. Earache [1386]. early [2466, 2161, 2850, 239, 2388]. Ears [717]. Earth [453, 419]. Earthquake [397]. Easier [397]. easing [110]. East [1141]. Easy [937]. eBay [2411, 2658]. ECCC [1802]. eco [1403]. eco-grid [1403]. Economic [2666, 2763, 220, 2864, 2573, 2966, 1925, 2998, 2732, 1102, 149, 1109]. Economical [1018]. Economics [258, 1240, 1296, 1394, 1116]. economies [1070]. economy [1352, 709, 1211, 349, 1716]. ecosystem [313, 2563, 2729]. ecosystems [2559, 139]. edge [1914, 1106]. edges [1769]. Editor [2955, 2332, 2567, 2505, 2568, 2537, 2989, 2821, 2956, 2629, 2723, 2923, 2755, 2787, 2535, 2457, 2754, 2628, 2819, 2689]. Editoral [1139]. Editorial [1, 17, 38, 63, 86, 105, 125, 146, 178, 199, 218, 240, 256, 298, 317, 375, 429, 452, 479, 495, 514, 540, 558, 275, 396, 587, 617, 638, 658, 691, 716, 738, 761, 784, 806, 837, 865, 885, 915, 936, 958, 981, 1001, 1026, 1051, 1075, 1117, 1161, 1213, 1235, 1256, 1279, 1301, 1324, 1345, 1365, 1385, 1408, 1433, 1455, 1477, 1502, 1527, 1559, 1584, 1607, 1631, 1653, 1682, 1707, 1732, 1752, 1776, 1804, 1833, 1860, 1886, 1910, 1933, 1960, 1981, 2001, 2020, 2039, 2066, 2083, 2103, 2129, 2150, 2175, 2193, 2215]. Editorial [2236, 2261, 2284, 2305, 2326, 2360, 2379, 2397, 2418, 2436, 2956]. editorship [2337]. Edmund [2159]. Edsger [810]. educate [2989]. Education [839, 1962, 2667, 370, 2576, 2489, 2765, 1721, 2722, 562, 5, 1027, 734, 1340, 1370, 1614, 2716, 2834, 15, 661, 2162, 1084, 1650, 1276, 1803, 92, 419, 2199, 2872, 764, 2915, 1840, 110, 259, 432, 2989, 2899, 421, 2865]. Educational [3, 431, 2514, 706, 1490, 1289, 1730, 1857]. educators [2989]. effect [1294, 1109, 1293, 1569]. Effective [583, 1845, 1228, 2120, 1013, 913, 1100, 1339, 669, 1877, 2746, 2953, 2890]. effectively [2207]. effectiveness [2186]. effects [2123, 2211]. efficacy [1103, 2205]. efficiency [3004, 1499, 1773]. Efficient [905]. efficiently [2207]. efforts [962]. Egypt [1325]. EHRs [1932]. Eject [639]. Election [1457, 2097, 1936, 2218, 2363, 2669, 1508, 1551, 2109]. Elections [659, 2424, 1164, 466, 1936, 2541, 1419, 1420, 1364, 1421, 292, 1319, 2312]. Electric [1214]. Electronic [1932, 88, 1379, 2092, 21, 2961, 4, 109, 543, 741, 918, 355, 757, 2546, 129, 900, 2962, 2038, 2408, 973, 872, 1010, 874, 608, 423, 2594, 281, 114]. Electronics [453, 866]. Elegant [1410, 1935]. element [159, 160, 157, 156, 158]. elementary [259]. elements [2881]. eleven [2070]. elicitation [1172]. eliminate [1789]. Eliminating [1609]. elite [399]. else [2137]. Elusive [1366, 1106]. Email [88, 640, 1912, 16, 64, 106, 583, 1076, 1791, 831, 546, 1066, 1931, 1770, 494, 767, 1745, 1881]. Email-Only [88]. embarrassing [277]. Embedded [1887, 96, 87, 2275, 725]. Embedding [98, 96, 148, 93, 95, 97, 94]. Embodied [78, 192]. Embrace [763, 2900, 2006]. emergence [2478]. Emergency [2111, 563, 2112, 2430, 2116, 2115, 2117, 663]. Emergent [874, 2711]. Emerging [2463, 2604, 2730, 2998, 1676, 2898, 2445, 773, 771, 1338, 1113, 2428, 118, 1264, 160, 2182, 1444, 2982, 416, 2111]. Emotional [1409]. emperor [537, 619]. Emphasize [1806]. empire [355]. Empirical [2292, 1184, 2875, 2394, 2186, 2911, 1108]. Employee [541, 608, 2918]. employees [1299, 3021]. Employment [1347, 465]. Empowering [1404]. empowers [1571]. Emulex [300, 377]. enable [853, 1761]. enabled [509, 777, 2747, 1159, 1596, 2076, 847]. enablers [1015]. Enabling [2090, 168, 170, 535, 2031, 167, 169, 677, 171, 1333]. encountered [1790]. Encouraging [971, 2095]. Encryption [2021, 87, 2774]. End [1400, 1397, 1862, 179, 1101, 1391, 2060, 367, 348, 1392, 2299, 551, 1192, 927, 1493, 1398, 2800, 2778, 1395, 165, 1394, 2635]. End-user [1400, 1397, 1391, 1392, 2299, 927, 1493, 1398, 1395, 165, 1394]. end-users [2778]. ending [277]. Endorses [1409]. Enduring [799, 101]. Enemy [1068]. energize [1946]. Energy [1708]. Energy-Saving [1708]. Enforce [1167]. enforcement [1822, 892]. Engaging [9, 803]. Engine [2040, 2998, 2573, 2625]. Engineer [938]. Engineering [763, 1119, 1561, 2041, 1862, 1935, 681, 219, 569, 2643, 1397, 2053, 2698, 913, 61, 1950, 694, 1736, 2156, 2221, 1130, 1084, 2518, 1494, 401, 812, 860, 1249, 2052, 2199, 862, 1563, 1974, 2393, 809, 1470, 2046, 697, 1951, 421, 863]. Engineers [1586, 859, 519, 857, 861, 858]. engines [2370, 593]. English [39]. enhanced [2222, 2630]. enhancement [160]. enhances [1694]. Enhancing [2620, 1773, 829, 1450]. enigma [1824]. Enlightening [258]. enlightens [1571]. enlists [1355]. Enough [763, 1215]. Enrollment [1754, 1214, 2455, 2551]. Ensure [1457, 1479]. Ensuring [2717]. Entering [2892]. Enterprise [813, 72, 213, 938, 208, 821, 75, 69, 928, 73, 1948, 76, 1930, 74, 77, 818, 71, 70, 735, 2656, 2521, 1274, 1580, 819, 205, 816, 1747, 1595, 673]. Enterprise-Integration [938]. enterprises [2413, 1592, 1111]. Entertainment [2027, 650, 2202]. Enthusiasm [1028]. entrepreneurial [1622]. entrepreneurs [2273]. Entrepreneurship [2864, 1713]. Environment [572, 313, 1330, 1443, 1010, 1372, 1581, 1403, 1703]. environmental [2895, 1104]. Environments [573, 569, 570, 571, 567, 475, 213, 2779, 1383, 194, 1375, 817, 881, 1446, 1396, 2317, 2592, 1034, 1072]. envision [327]. Envisioning [2004, 2490, 2124]. EPCglobal [2845]. epidemic [1703]. epistemic [2814]. epistemology [1000]. EQ [1724]. equal [380]. equality [925, 1287]. Equate [1935]. equilibrium [2679, 2678]. era [2931, 1786, 995, 92]. Erlang [2704]. ERP [929, 1855, 72, 127, 759, 75, 928, 1359, 73, 1358, 76, 2590, 74, 71, 70]. Error [2709, 2914, 1401, 2708]. error-correcting [2708]. errors [2617, 187, 2616, 2505]. escalating [2951]. Espionage [220, 149]. Essay [1935]. essential [948, 36]. established [2158]. estimation [840, 2942, 2059]. ethic [1919]. Ethical [220, 269]. Ethics [220, 2, 181, 80, 2852, 2544, 2036, 2866]. Ethnocomputing [1801]. Etiquette [1289, 1287, 1291, 1286, 1290]. EU [283]. Euro [18, 397]. Europe [1953, 1079, 2922, 2762, 1838, 1556]. European [282, 969, 2589, 1838, 43, 70]. Evaluating [1520, 1131, 629, 1794, 1428, 1493, 2205, 1363, 2720, 2498, 1497]. Evaluation [1454, 1269, 2499, 448, 2734, 173, 1748]. Even [1778, 2002, 2568, 340, 1427]. event [25, 2651]. events [905, 394]. Eventually [2641]. Ever [258, 1387, 363, 742]. ever-expanding [742]. evermore [2214]. Every [1163, 2041, 1032]. Everybody [147]. Everyday [1303]. Everyone [1435, 1778, 2765, 1573]. Everything [377, 332, 553, 2137, 2349, 1795, 163, 338]. Everywhere [937]. Evidence [1862, 1076, 1819, 2388, 679]. Evolution [2182, 1176, 852, 2559, 1949, 2713, 1721, 2443, 2695, 1463, 998, 755, 769]. Evolutionary [2607, 186]. evolve [627]. evolved [1403]. Evolving [398, 2622, 768, 331, 524, 2240]. exactly [2726]. exaflops [343]. exaggerated [1304]. Examining [2914, 2208, 764, 1522]. example [58, 59, 54, 56, 57, 55, 60]. Examples [1548, 57]. Exascale [2788]. Exceeds [397]. excellence [1590]. Exceptional [1932]. Excess [1560]. exchange [1021, 1379, 2982]. exchanges [1880]. exciting [2535]. Excuses [1326]. execution [2825, 1991]. executions [1087]. Executive [2567, 1836]. executives [1217]. exhaustion [2575]. exhibitions [1287]. Existence [1303, 1979]. existential [92]. Existing [454]. exists [2465]. Exorcised [480]. expanded [1955]. Expanding [107, 742]. expands [1282]. expansion [2266]. Expectations [2022, 1286, 1287, 1065]. Expeditions [1503]. expenditures [1800]. Experience [1053, 2293, 1032, 1646, 2667, 223, 1112, 478, 1694, 325, 154, 1644]. Experience-driven [2293]. Experiences [1639, 1543, 1491, 267, 386]. Experiential [1035]. Experiment [376, 849]. Experimental [2290, 2291, 2293]. experimentation [2330, 827, 2295]. Expert [571, 2253, 1159]. expertise [1393, 2365, 2992]. Experts [1409, 2544]. explain [1925, 2941]. Explaining [780]. exploitation [2621]. exploration [1875, 1874, 487]. Explorations [1924]. Exploratorium [1694]. Exploratory [1870, 1869, 1871]. Exploring [1872, 1873, 2655, 2963, 476, 491, 2593, 1907, 572, 1703, 22, 45]. Exports [659]. Exposing [560]. exposures [1384]. extended [2981]. Extending [938]. extensibility [815]. eXtensible [993]. Extension [2889, 2656, 2852, 2845, 2850, 2847, 2846, 2849, 2848, 2991, 2856, 2957, 1701, 2851, 2924]. extensions [307]. extent [2666]. Exterminator [2617]. External [266, 2325, 2229]. externality [3020]. Externally [1991]. Extracting [1968, 924, 1059]. extraction [2613, 30, 1901]. Extreme [733, 2264]. extremism [450]. Extrovert [1546]. Eye [1684, 739, 1027, 1055, 749]. Eye-Opening [1027]. Eyes [838, 1805, 944].

f [2178]. fab [326]. Face

[39, 453, 2900, 2860, 2874, 1600, 623]. faced [2273]. faceted [1874]. facets [2552]. facilitate [1449]. facilitated [2432]. Facilitating [1021, 471]. Facing [2894, 974]. fact [1453]. Factor [1911, 2002, 3011, 1583, 1212]. factories [84]. Factors [786, 1925, 2075, 2534, 1669, 2948]. Facts [1813]. Fade [1027]. Fail [219, 3018, 884]. failure [1308, 620, 1844]. failures [2784, 1784, 2585, 2102, 1015, 908]. Fair [1028, 1457, 1684, 969, 966, 965, 970, 971, 1216]. fairness [2030]. Faith [2400]. Faith-based [2400]. Fake [1885]. fall [2764, 2494, 463, 941, 2876, 2986]. Falling [2921]. fallout [4]. families [2049, 2318, 1739]. Family [1585, 853]. Fantastic [1346]. Fare [1961, 257]. farewell [2480, 1913]. farm [126]. Fast [1861, 2357]. Fastest [276]. fate [1838]. fatware [498]. fault [2586, 2824]. favorite [1841]. FBI [399]. FCC [1589, 1608]. Fear [1888]. Feasible [983]. Feature [2052, 2050, 2940, 2720]. Feature-oriented [2052]. Features [1156, 783, 2506]. Fed [219]. Federal [1002, 2108, 901, 903, 742]. federated [2098]. FedEx [976, 1012]. FedStats [901]. Feedback [1806]. Feel [1118, 2449]. Fellows [1530, 729, 66, 1260, 2110, 2384, 921, 2697, 1849]. Felt [739]. Felt-Tipped [739]. Felten [520]. female [427]. ferrofluid [2451]. fever [2469]. Few [541]. Fewer [2022]. Fi [991, 2229]. fiction [1453]. field [1348, 897]. Fifty [2331]. fight [111]. Fighting [559, 838, 2836]. Figures [516]. File [692, 2687, 927, 1413]. file-sharing [927]. File-Swapping [692]. Files [515]. Filipino [886]. filtering [1401, 1467]. filters [525]. Finally [1813]. finance [153]. Financial [1313, 1308, 549, 730, 231, 1307, 1627, 1310, 436, 2887, 654, 437]. financing [1314]. Find [1435, 1876, 159, 1254]. Finding [2945, 939, 794, 2507, 2648, 571, 2357, 1870, 1223, 1649]. findings [2287]. fingerprint [1878]. fingers [1482]. Finging [807]. Finland [659]. fire [138]. firm [270]. firms [709, 1144]. First [1529, 430, 1817, 2240, 1535, 2664, 2818, 1802, 2569, 1270, 2989, 458]. first-mover [2818]. Fish [1753]. fisheye [1402]. Fit [580, 1280, 1810, 1187, 2655, 2916]. fit-gap [1187]. fitness [2742]. fits [74, 2614]. fittest [994]. Fitting [1608]. Five [2342, 203, 2835, 1212]. five-factor [1212]. five-minute [2835]. Fix [785, 2320]. fixed [846]. Fixing [2108, 565, 279, 2528]. Fizzles [1386]. Flag [1002]. Flash [2470, 2835, 2743, 2744]. Flatlined [719]. flawed [279]. flaws [989, 1262]. Flesh [866]. flex [1219]. flexibility [2280]. Flexible [2007, 1309, 1346, 984, 817, 598, 2811]. Flight [220]. Floating [717]. Flock [1654]. Floppy [807]. Florida [864]. Flow [2137, 1077, 1100, 794, 932, 2676, 1904]. flows [2556, 444]. fluency [373]. Flush [1076]. fly [845]. Flying [87, 126, 1346]. Focus [1912, 1118, 1428, 1979]. Focusing [948]. Folk [1367, 2003, 964]. Folklore [1813]. Follow [1479]. Following [2926, 2531, 3021]. Follows [1366]. fond [2480]. Food [1861, 2587]. footprints [2531]. force [460, 2763, 473, 472, 461]. forecasting [2011]. Foreign [2, 1379]. forensic [1816, 1756, 1820]. Forensics [2308, 2815, 1817, 1054, 2132, 2439, 1355, 1814, 1821]. foresight [2000]. forest [2986]. forests [894]. Forever [1410]. Forget [454, 1053, 1003, 353]. forgotten [344]. form [2446, 3011]. form-factor [3011]. Formal [2056, 2843, 1368, 2518]. format [411, 1823]. formats [2937]. formed [1863]. forms [2451, 1620]. fortress [179]. Fortune [2751]. Forum [180, 19, 220, 127, 88, 148, 107, 40, 258, 65, 3, 201, 398, 497, 516, 300, 481, 431, 560, 377, 454, 542, 660, 640, 763, 619, 786, 740, 867, 693, 839, 718, 983, 1053, 1141, 1003, 938, 1163, 1077, 887, 1119, 960, 1028, 917, 1410, 1237, 1347, 1435, 1367, 1457, 1215, 1387, 1281, 1258, 1326, 1303, 1479, 1529, 1586, 1633, 1504, 1609, 1754, 1561, 1655, 1709, 1734, 1684, 1778, 1888, 2022, 2041, 1835, 1912, 1806, 1862, 1962, 1935, 1983, 2003, 2068, 2085, 2105, 2151, 2176, 2217, 2238, 2263, 2307, 2130, 2327, 2362]. Forum [2398, 2420, 2438]. forums [2940, 2117]. Forward [1053, 2597, 2823, 2793]. foster [1306]. Foul [1302]. Found [983, 692, 2985]. foundation [1899, 1258]. Four [1796, 480, 2970, 2184, 2948]. fourth [2897, 1944]. Fractal [1595]. fracture [137]. frame [2806]. framed [280]. Framework [481, 929, 2681, 509, 1274, 2817, 951, 210, 1764, 2407, 530, 2621, 1825, 2232, 1042, 1429, 196, 1212, 1497, 1135, 1670, 952]. frameworks [883, 2062, 208, 1244, 211, 207, 205]. fraternal [2815]. fraud [2042, 2016, 2433]. Free [737, 740, 1633, 541, 982, 2002, 484, 402, 1322, 1414, 87, 299, 2872, 1037, 1155, 1602, 1005, 991, 2797]. freedom [1919, 358]. Frequency [453]. frequent [2945, 602]. Friction [1367, 1327]. friendly [1831]. Friends [1632, 2921]. Fries [453, 2714]. Front [639, 2454, 2488, 2545, 2636, 2800]. front-end [2800]. Frontier [543, 4, 109, 741, 918, 1150]. frontiers [13, 1379, 2963, 879]. Fruit [1733]. Fu [2968]. fuels [1998]. Fulfilling [1008]. Full [1038, 1946]. Full-size [1038]. Fuller [220]. fully [1099]. Fun [1503]. Function [1709, 2041, 2915, 1129, 71]. function-based [2915]. functionality [1428, 340]. functions [311]. Fundamentals [1962]. Funding [552, 1935, 1563]. Fungible [2042]. Further [1862]. fusion [81]. Future [2527, 1962, 2619, 39, 397, 318, 3013, 2913, 2810, 1409, 1585, 1230, 1248, 1638, 362, 327, 1145, 670, 2543, 361, 365, 1351, 2004, 877, 231, 880, 2152, 2929, 370, 357, 329, 2662, 2490, 998, 2631, 1086, 250, 835, 543, 2601, 285, 2324, 353, 663, 2338, 2713]. FY01 [600]. FY03 [1218]. FY04 [1469]. FY05 [1781]. FY06 [2072]. FY07 [2367].

Gadget [639]. gadgets [1546]. galactic [850]. Gambling [935]. Game [106, 593, 1032, 1280, 1528, 2021, 1036, 302, 1216, 2289, 1492, 2497, 1033]. Game-like [1036]. GameBots [598]. Gamer [1753]. gamers [2028]. Games [2200, 1077, 1326, 126, 2199, 1503, 2030, 2137, 2029, 134, 9, 133, 2862, 1489, 2921, 2201, 595, 1263, 2204, 748, 2624, 3007, 2492, 2703, 2198, 2495]. gaming [1541, 2368, 596, 2386]. Gap [480, 1080, 1187, 2975, 931, 373]. Garbage [937]. Gastric [886]. gate [1068]. gates [2154, 2638]. gateway [901]. gathering [1267]. Gauging [292]. Geek [1911]. Gender [480, 778, 2763, 1231, 1103]. general [466, 1936, 2218]. generalization [60]. generalizing [59]. generation [2042, 954, 661, 2206, 1073, 1814, 2467, 1821, 724, 507]. generative [2635]. generators [999]. Genesis [2626]. Genetic [2041, 1581]. genome [1443]. geometric [1692]. Geometry [2556, 1602]. George [1456]. Georgia [1418]. Geospatial [893, 897, 903]. Germany [1950, 1951]. gesture [1037]. gestures [1489]. Get [147, 1646, 1052, 2633, 2944, 193, 2964]. Gets [619, 1237, 1366, 1057, 2017, 2682]. Getting [2047, 526, 963, 2273, 2401]. ghost [1672]. ghostbuster [1679]. giant [276, 985]. Gift [1162]. Giga [2040]. Gigahertz [64, 480]. GIM [1792]. Girls [126, 639, 9].

’Give [2432, 481, 763, 938, 1077, 1410, 1367, 1655, 1888, 1912, 2441]. giving [890]. Gizmos [982]. Global [1484, 1925, 2514, 696, 127, 497, 1504, 575, 579, 576, 276, 480, 886, 193, 2883, 581, 1232, 460, 2075, 2270, 1312, 1810, 1146, 2760, 2503, 2532, 1292, 2719, 1352, 1109, 501, 1591, 2412, 2658, 1641, 3014, 289, 1239, 461, 997]. global-scale [1146]. Globalization [1439, 2831, 2463, 2666]. Globally [2012, 2364, 2008, 2010, 1930]. Gloom [107]. Gloom-and-Doom [107]. Gloved [639]. Glow [618]. Go [300, 1141, 430, 937, 1162, 1302, 1961, 1861, 2973, 720]. Goal [1457, 816]. goal-driven [816]. Goals [560, 2064, 949]. Goes [106, 2760, 1721]. Going [481, 219, 1434, 2457]. golden [1926]. gone [1660, 2614]. Gonzo [717]. Good [560, 1806, 1862, 541, 2685, 2015, 2357, 1826, 2799, 1679, 590, 2890, 2271]. good-bye [2890]. Goodness [1529]. Google [1506, 2869, 2830, 2953, 1805]. Googlewhacking [639]. googling [2348]. Got [2611, 1280, 1834, 257]. Gotcha [453]. govern [1811]. governance [2654, 289, 2232]. Governing [2563]. Government [1709, 739, 1366, 1525, 2750, 2146, 907, 895, 899, 1342, 1728, 3015, 891, 904, 906, 2428]. GPS [541, 717]. GPUs [2549]. grab [112]. grade [2716]. grades [1839, 2135, 2402]. Grading [839]. Grads [219]. gradually [1600]. graduate [869, 2127]. Graduates [1410]. grail [694, 2100]. Grammar [1063]. Grammar-based [1063]. Gramr [982]. Graph [1266, 2556, 2555, 2502]. Graph-based [1266]. graph-partitioning [2556]. Graphical [740, 3010]. graphics [1621, 2936, 1371]. Grasping [1317]. grassroots [2117]. Graveyard [318]. Gray [2567, 2580, 2581]. Grayer [982]. graying [1913]. Great [1142, 2897, 2598]. greater [1427]. Greatest [1778]. greatly [1304]. Green [2539]. Greenstone [408].

Greetings [619]. grid [545, 911, 536, 1403, 660, 1582]. Grokster} [1712, 1238]. Groove [1366]. Ground [2636, 1934]. Group [878, 1792, 801]. groups [235, 945]. Groupware [578]. grow [2867]. Growing [276, 1330, 564, 2331, 1913]. Growth [515, 739, 462, 2998, 1774, 1716]. Grunt [559]. GSD [2041]. GSS [195]. guarantees [924]. Guard [717]. Guesses [866]. Guests [1983]. guide [2611, 703]. Guidelines [1224, 2162]. Guitar [1028]. guns [2544]. guns-for-hire [2544].

Habitat [1332]. habits [2120]. Hack [739, 717]. Hacker [1709, 1632, 1534, 1919]. Hackers [541, 886, 1733, 1753, 1585, 1377]. Hacking [1918, 2345, 1921, 1924]. Halt [1077]. Halting [2041]. Hand [1835, 639, 1733, 432]. Handheld [432, 296, 2815, 1089, 1170, 1038, 2096]. Handhelds [559, 547]. Handles [1835, 415, 1768]. Handling [867, 531]. Hands [1215, 959, 1528, 2, 1838, 2241]. Hands-On [1215]. happened [507]. Happy [838, 2329]. Haptic [559, 1372, 48]. harassment [609]. harbor [1536, 1612]. Hard [2799, 1984, 1823, 1615]. Hard-disk [2799]. Hardly [916]. hardness [2873]. hardware [1921, 2808, 2807]. hardware-based [2808]. Hari [1140]. harmful [1524, 1501, 1958, 782, 538, 1137, 863]. harmoniously [1403]. Harvesting [894]. hashing [2358]. Hastily [1863]. Hate [1257, 1665]. having [2524]. HDF5 [2937]. Head [2041, 1503, 1867]. heading [2834]. Headlines [1911]. Health [1932, 1593, 2090, 2187, 2656, 509, 2191, 1723, 115, 1798, 1773, 1233, 1350, 1460, 1787, 875, 2812, 1020]. Healthy [1912, 1118]. Hear [453, 337]. hearing [2644]. hearing- [2644]. heart [1513]. Heaven [1805]. helicopter [2841]. helicopters [2792]. Helix [1249]. Helix-Spindle [1249]. Hello [619, 1434, 538, 2810]. Help [1961, 1861, 2898]. helpdesk [611]. helping [159]. helps [2152, 1590]. Hepatic [1511]. Here [1962, 2047, 605, 1856]. Hermitage [489]. heterogeneous [2431, 2033, 2074]. Hibernating [888]. Hidden [1932, 1199, 506, 2753, 1977]. Hiding [2132]. hierarchical [1377, 210, 1115]. hierarchies [84]. hierarchy [1520, 2254]. High [18, 276, 739, 692, 1355, 2609, 1861, 2358, 1441, 634, 1145, 2952, 664, 1147, 225, 535, 2588, 2617, 1290, 910, 533]. high-criticality [1290]. high-payoff [910]. High-Performance [1861, 2609, 1441, 634, 1145, 535, 533]. high-quality [2952]. High-Tech [276, 739, 692, 1355]. higher [2716, 1721, 1803]. Highest [40]. Highly [2775, 2120, 1988, 1989]. highway [129, 1315]. Hijacking [662, 127]. hint [2482]. Hints [916]. Hip [959]. HIPAA [1350]. HIPAA-potamus [1350]. hire [2544]. Hiring [1503]. Hirings [692]. Historical [2543, 2931, 2764, 422]. History [2021, 2165, 2163, 907, 1535, 1656, 2160]. hit [200, 1723]. hits [2762]. hive [2913]. Hoare [2702]. Hoax [300, 377]. hoc [1545]. Hold [1077, 1347, 1367, 1586, 588, 18]. holder [2073]. Holds [916]. holes [1140]. holistic [1580, 2817, 2206]. holographic [227]. holy [694, 2100]. Homage [1734]. Home [839, 1985, 1027, 1257, 1632, 1478, 1982, 753, 1400, 2981, 299, 849]. Homeland [1434, 1267, 1271, 1264]. Homemade [839]. Homework [1162]. homograph [637]. Hong [2886, 272, 469, 1111]. honored [2697]. honors [2043, 2134, 2403]. Hooked [588]. Hope [431]. Hopes [640]. Horizons [583]. horizontal [1029]. Hospital [2415]. hostile [2142]. Hot [1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1866, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2044, 2081, 2101, 2126, 2147, 2173, 2190, 2213, 2233, 2259, 2282, 2303, 2322, 1412, 1982, 200]. hotspots [1922]. hours [87]. House [785]. Household [717]. Houston [2040]. HP [1982]. HPC [2857]. HR [468, 1883]. HTML [763]. Hubbub [798]. hubs [1578, 1848]. Human [2112, 2952, 740, 1734, 2667, 39, 692, 1286, 1961, 159, 160, 170, 157, 156, 595, 1201, 2080, 158, 2705, 1691]. human-activity [1691]. Human-computer [2112, 1286]. human-level [595]. humanity [2872, 328]. Humanoid [646]. humans [751, 1246]. humble [2567]. humming [1971]. Hunting [39, 942]. Hybrid [1543, 1188]. hype [766]. Hypercomputation [839, 887, 766]. Hyperlinking [2811]. hypermedia [706, 698, 703, 708]. hypertext [702]. hypothesis [2092].

i2i [264]. IBM [40, 2006]. ICANN [481, 376, 130]. iCare [2981]. Ice [1386]. icles [106]. ICT [2187, 2898, 2652, 2621, 1801, 2857, 2595]. ICTs [285]. ID [1674, 659, 3015, 1859, 1885]. IDC [1487]. idea [2696, 1438, 1460, 2614]. ideal [1106]. ideas [2693, 2822, 2273]. identical [2815]. identification [33, 816]. identifiers [415]. identify [910]. identities [1272]. Identity [20, 88, 1935, 639, 586, 2098, 1890]. IEEE [2162]. IEEE-CS [2162]. If [1053, 1457, 2986, 920, 560, 2465, 2752]. ignorance [203]. Ignore [1633]. II [1855, 1458, 2610]. III [599]. ilities [615]. illegal [276]. illusion [2130, 2390]. iLumina [1645]. I’m [1053]. Image [2467, 1257, 1911, 2937, 1194, 1510, 1807, 2955, 1515]. images [1823, 2553, 750]. IMing [1140]. Immaturity [1835]. immersion [1371]. Immersive [1646, 1099, 597]. Immigration [461]. immortality [321]. iMode [1151]. Impact [515, 284, 2750, 1189, 1354, 2414, 2880, 2848, 2300, 2099, 1745, 1116].

Impaired [180, 2644]. imperative [592]. Imperceptible [1609]. Implant [659, 87]. implantable [2795]. implants [1372]. implementation [1252, 814, 476, 511, 1825]. implementations [73, 673]. Implementing [2962, 2317, 1418, 591, 1855, 2075]. Implications [1410, 2447, 1723, 2125, 8, 1110, 1907, 1883]. Implicit [817]. importance [1702]. important [2287, 802, 2078]. Impossible [1586, 1504]. Impractical [542]. impression [2749]. improve [215, 2148, 529, 2296, 1770, 2021]. Improve-ment [2021]. Improved [2654, 3005]. improvement [1750, 392]. Improvements [1733, 1478]. improves [1978]. Improving [734, 1362, 1320, 2672, 2429, 2849, 1820, 801, 427, 896, 1931, 1871]. In-Dash [397]. inaccurate [2381]. inaction [715]. inbox [2089]. incarceration [1562]. incentive [1624]. included [1885]. inclusion [2714]. Inconsistencies [2144]. Inconsistent [1326]. Incorporating [1514]. Increase [1962, 1381]. Increased [2079]. increases [1066]. Increasing [294, 2763, 3020]. increasingly [1351]. Incredible [886]. incremental [2184]. independence [2956]. Independent [1417, 815]. index [510, 2905]. indexing [29, 27]. India [2463, 2004, 462]. Indian [2604, 446, 423]. indispensable [1353]. individual [456, 862, 1590]. indoor [1491]. Industrial [1410, 1912, 2230]. industries [462]. Industry [220, 786, 2604, 1158, 2998, 2816, 974, 1810, 1187, 1153, 32, 1045, 1105, 1773, 490, 721, 777, 467, 1304, 1679, 760, 2686, 360, 2751, 143, 469, 1116]. industry-specific [2751]. inequality [2990]. inevitable [2900, 1662]. Infectious [618]. infidelity [1389, 1458]. infinity [996]. influence [1746, 2560, 1204]. Influences [249, 2747, 2434]. Influencing [786]. Info [40, 276]. Info-czars [40]. Info-tech [276]. Infoglut [1937]. informatics [1056].

Information [2521, 2410, 680, 13, 431, 542, 640, 1077, 1367, 1479, 2003, 144, 2341, 1092, 559, 113, 686, 932, 555, 791, 1138, 1967, 1241, 1979, 1172, 2298, 2452, 563, 1603, 2090, 439, 242, 282, 1096, 656, 2279, 58, 2611, 243, 2075, 1520, 215, 2255, 2112, 1676, 1772, 322, 1872, 1789, 2281, 894, 1968, 1792, 549, 2613, 954, 832, 1244, 2720, 2188, 30, 1875, 486, 1723, 508, 1800, 1901, 388, 2709, 1023, 1874, 34, 2121, 2654, 1788, 1793, 1906, 1794, 2980, 1267, 2275, 705, 2585, 167, 1790, 2747, 345, 2623, 675, 910, 2781, 1265, 1787, 43, 253, 2441, 1042, 1680]. information [116, 310, 556, 1277, 3021, 1242, 1785, 2812, 472, 663, 2111, 1722, 2295, 2676, 2738, 1068, 1791, 1716, 2124, 1871]. information-compatible [2781]. information-retrieval [2738]. Information-rich [1092]. information-sector [472]. information-theoretic [2709]. Informational [1298]. Infrastructure [578, 688, 1230, 1024, 882, 2183, 2456, 1582]. infrastructures [666]. Infringement [1326]. ing [407]. Inhibit [1367]. initiatives [774, 1088, 1563, 2621, 1430]. Injustices [560]. inky [2447]. Innovation [1141, 1367, 1734, 1835, 1895, 2666, 1918, 1283, 2382, 962, 632, 2882, 2719, 1952, 1591, 906, 1943, 2310, 2755, 2855, 2857]. innovations [2622, 677]. Innovative [975, 546, 1543]. Input [2021, 944]. inquirers [2846]. Insect [430]. insecurity [1166]. Inserting [615]. Insider [1364, 2918]. Insight [1047, 1797, 2340, 1875]. insight-based [1875]. insightful [2739]. Insights [548, 1031, 768]. Insist [2041]. insourcing [34]. inspections [801]. Inspiration [2343]. Inspire [2821]. inspired [1847, 2013]. inspires [745]. Inspiring [2274, 2542, 2467]. instant [798, 1638]. instantiation [2231]. institutional [313]. institutions [2189]. instruction [1061, 2409]. Insurance [360, 2973, 951, 1209]. integrate [1902]. Integrated [676, 1580, 27, 121, 2098, 230, 1274, 819, 95, 956]. integrated-circuit [230]. integrates [1491]. Integrating [813, 194, 2743, 800, 675, 2744, 72]. Integration [938, 887, 2521, 1357, 1578, 1148, 113, 928, 1310, 874, 1930, 818, 952]. integrative [196]. integrity [1057, 297, 789, 845]. Intellectual [310, 2906, 352, 2345]. Intelligence [541, 785, 645, 1266, 1875, 170, 56, 47, 204, 2866, 775, 1649, 2124]. Intelligent [786, 1135, 654, 1523, 216, 710, 2124]. intensive [2141, 1369, 1150, 2831]. intent [1516, 1517, 624]. intentional [1286]. Interacting [945, 138]. Interaction [1486, 704, 1288, 2112, 2952, 1896, 1066, 2444, 171]. Interactions [2450, 1877, 2448, 2871]. Interactive [1039, 2041, 1371, 3011, 597, 1700, 648, 1037, 1693, 1835]. interactivity [1651]. interchange [757]. interconnection [1403]. interdisciplinary [2472]. interest [2197]. Interface [823, 1173, 2447, 78, 266, 2443, 1201, 187, 1224, 2449, 1225, 1876]. Interfaces [1347, 943, 1399, 699, 52, 2445, 1221, 51, 331, 192, 1189, 2446, 157, 184, 2788, 189, 161, 56, 186, 49, 1219, 2448, 47, 50, 53, 2890, 48, 46, 2442, 950]. interferes [183]. intermediaries [167]. internal [2281, 2325]. International [2268, 129, 222, 433, 1079, 2668, 868, 1214, 1808, 708, 655, 1744]. internationally [459]. Internet [1768, 2313, 2392, 2407, 2032, 98, 925, 601, 1484, 1151, 1054, 234, 96, 1047, 1759, 502, 1196, 2534, 294, 390, 550, 361, 224, 148, 1410, 1258, 1504, 1778, 1806, 2639, 664, 641, 2016, 930, 1153, 129, 93, 231, 147, 397, 659, 2183, 2433, 1102, 995, 2664, 1570, 10, 1650, 290, 1353, 503, 1667, 1354, 1422, 1195, 1811, 490, 1048, 2672, 2818, 1197, 1571, 2520, 291, 2197, 1720, 1159, 112, 603, 811, 1155, 292, 95, 2959, 1714, 2154, 607, 606, 5, 602, 97, 1581, 1209, 94, 108, 608, 1852, 121, 104, 124, 507, 711, 585, 1569, 2635]. Internet-based [550, 1650, 2818]. Internet-enabled [1159]. Internet/Web [1852]. internetworked [1090]. Interoperability [1253, 1024, 1429]. Interoperation [114]. interorganizational [1896]. interplanetary [503]. interpreted [1063]. interrupted [2491]. intertwined [347]. intervention [2094]. interventions [1799]. Interview [2901, 2968, 2608, 2468, 2491, 2702]. Intestinal [147]. Intra [2076, 2034]. intra-domain [2034]. Intra-organizational [2076]. intractable [2035]. Intranet [2091]. intranets [477]. Intraocular [749]. Intraorganizational [477]. Intricate [126]. Introducing [1159, 2255]. Introduction [2007, 722, 601, 813, 1690, 793, 1145, 823, 1918, 2354, 1331, 575, 1058, 1592, 1082, 1128, 521, 2290, 403, 625, 1639, 1307, 923, 593, 876, 843, 891, 1843, 1568, 851, 1286, 1715, 1988, 965, 2243, 1415, 1219, 1814, 1122, 1967, 665, 567, 152, 132, 1032, 1371, 1462, 1615, 2339, 987, 1241, 1760, 1941, 2027, 1007, 1661, 1440, 2046, 1391, 1785, 1168, 943, 1894, 1869, 2157, 1264, 2198, 381, 460, 304, 435, 93, 512, 280, 225, 290, 69, 441, 54, 24, 172, 310, 46, 2332, 2111, 2442]. introductory [2821]. Intrusion [81, 2142, 1671]. intuition [2878]. intuitive [2426, 1489]. Inventing [357]. invention [2855]. inventory [2620]. Investigating [1815, 2145, 1669, 1729, 2783]. investigation [1184, 2394, 2186, 143]. investing [2980, 8]. investment [1562, 35]. investments [1520, 1363]. Investor [1360, 1498]. Invisa [1456]. Invisa-Bills [1456]. invisible [1542]. invitation [2180, 2229]. inviting [1049]. involvement [2161, 2914]. Involving [2226]. IP [2850, 610]. iPod [2719]. Iraq [2002]. Ireland [1713]. Irk [1753]. irreversible [2537]. IS-user [1926, 632]. isn’t [372, 2513]. isolation [257]. Israel [2998, 739]. Issue [2007, 1690, 1592, 559, 496, 1462, 2027, 2167]. Issues [127, 876, 662, 2182, 435, 1103, 2034, 511, 881, 802, 1292, 1542, 162, 161, 1695, 163, 1089, 165, 643, 166, 1941, 1861, 2273, 164, 1852, 1582]. IT-business [2320, 2321]. IT-enabled [2747, 1596, 2076]. IT-framed [280]. Item [1683]. items [2945]. iterative [2264]. ivory [179].

J [2954]. J-shaped [2954]. J2EE

[1017, 1016]. jack [2971]. Jacked [618].

Jacked-In [618]. jackpot [200]. Jam [1608]. January [2538]. Japan [1370, 1953, 276, 739, 886, 1112, 1708]. Japanese [1778, 1637, 30, 2925, 2564]. Java [536, 684, 1554, 535, 782, 533, 1059, 847]. Jersey [618]. jet [517]. Jim [2567, 2486, 2580, 2581]. job [1316, 2416, 1726]. jobs [1600]. Jockeys [1632]. join [1505, 2752]. joining [1701]. Journal [541, 2240, 2880, 1522]. Journalism [717]. Journals [481, 1500, 696, 1295, 1136, 501, 1746, 1659, 1254, 2982, 2754]. journey [2164, 1750]. Joy [481, 1935, 2827]. Judge [1119]. Judgments [1834]. Juggling [1076]. Just [258, 1003, 1281, 1983, 618, 2383, 2859, 364, 513, 2598, 2605]. just-in-time [364]. Justifying [740].

K-12 [2667]. Katrina [1822]. Keep [481, 1163, 369, 1257]. Keeping [776, 481, 1790, 2896, 2234]. Key [839, 1024, 1013, 1377, 1799, 1540, 994, 1580, 370, 2138, 82]. keyboard [1038]. keys [2774]. keywords [2519]. Kicking [1683]. Kid [2021]. kids [257]. KidsRoom [52]. killed [257]. killer [1745]. killing [1817]. kind [2481, 1384]. kindergarten [102]. Kinds [1141]. kinetic [2448]. King [219]. kissing [2241]. kit [676]. Knock [1560]. Know [1457, 1529, 2003, 159, 554, 500, 316, 2465, 2971, 1183, 102, 1664]. Knowledge [1574, 2413, 917, 1041, 2272, 1627, 251, 1204, 585, 545, 920, 1597, 911, 1465, 892, 2394, 877, 2817, 1021, 913, 1297, 1771, 417, 1449, 824, 1306, 2715, 1088, 2209, 802, 1762, 2372, 2846, 2211, 2831, 1668, 2560, 2496, 2982, 2301, 996, 2738, 1904]. knowledge-intensive [2831]. Knowledge-sharing [1204, 920]. Known [1529]. knows [87, 18]. Knuth [2468, 2491, 2505]. Kode [2465, 2512, 2606, 2934, 2798, 3000, 2731, 2934]. Kong [2886, 272, 469, 1111]. Kyoto [754].

l [148]. lab [395]. label [1647]. Labor [2604, 469]. laboratory [779, 2623]. labyrinth [2196]. Lack [1835]. lacking [200]. lacks [2857]. lags [1425]. laid [1731]. LambdaGrid [1146]. land [112]. landscape [2939]. language [1638, 247, 1895, 27, 2749, 1898, 171, 993]. language-action [1898]. Languages [127, 2165, 1231, 853, 1396, 2737, 2735, 1321, 980]. laptop [2803]. laptops [1987, 2688]. Large [2558, 770, 411, 2356, 2024, 2125, 487, 2227, 2903, 2553, 2376, 1037, 668]. large-format [411]. large-scale [2125, 2227]. larger [356]. Laser [659]. Last [2527, 2713, 3013, 2479, 2557, 2650, 2844, 2745, 1366, 2502, 2526, 2587, 2618, 2680, 2712, 2777, 2979, 3012]. late [1583]. Latency [2029, 1479, 1425, 2030]. later [2835]. Latest [126]. launch [2690, 2538, 2692]. laundering [1314]. Laundry [397]. Law [1003, 106, 2762, 2896, 2635, 2313, 2352, 1822, 2906, 892, 350, 967, 2265, 2345, 1744, 807]. Lawful [887, 1273]. Laws [1655, 2960, 278, 1436, 2366]. Lawyer [220]. lawyers [1351]. Layered [1617, 1618, 530]. layout [1376]. LDAP [955]. LDC [1774]. Lead [376]. leaders [2120, 2926]. leadership [2850]. leading [1728]. leads [1448]. leap [1650]. Learn [1215, 1862, 2906, 975, 1107]. learned [1193, 1487, 822, 2686, 209, 102]. Learner [1052]. Learning [19, 579, 2760, 2711, 2862, 478, 2882, 2021, 2725, 1080, 1914, 604, 25, 670, 1178, 2841, 1131, 2533, 669, 1721, 168, 2987, 1491, 2977, 1704, 1282, 2409, 1318]. left [2202]. legacy [2638, 2542, 2700]. Legal [1028, 1326, 1633, 962, 1993, 2916]. legality [1920]. Legalize [1077]. Legally [149, 21, 456, 620, 809, 1005, 1120, 1238, 1328, 1712, 1536, 1612, 1917, 2006, 1809, 2265, 2088, 2178, 2366, 2464, 2575, 2965, 2830, 2699, 108]. Legends [1934]. Legion [1654]. legislation [67, 1975]. Legislative [1536]. LEGO [3]. lemmings [2755]. Lend [717]. LEO [1586, 1633]. Less [1367, 1684, 959]. Lessons [1326, 1684, 1193, 822, 209, 215, 1921, 10, 1645, 1630, 2686, 2869]. Lest [2774]. Let [1367, 1457, 2041, 2515, 376, 807, 605, 1482, 2535]. Letter [790, 1460, 1590, 1563, 2152, 2288, 2331, 2480, 378, 2503, 2922, 2722, 1531, 1613, 1659, 1483, 1710, 1634, 1687, 1739, 1505, 1840, 1913, 1783, 1813, 1892, 1867, 2988, 2535, 2457, 2754, 2628, 2955, 2819, 2689, 2786, 2566]. Letters [2505, 2568, 2537, 2989, 2821, 2755, 2956, 2787, 2629, 2723, 2923]. level [2053, 595, 289, 743]. Leveraging [2280, 2818]. Liability [2065, 380]. liable [1262]. libel [108]. Liberia [2268]. Liberties [660]. librarian [414]. libraries [117, 411, 403, 413, 1104, 407, 410, 412, 1448, 420]. Library [481, 414, 2044, 2101, 2173, 2190, 2213, 2233, 2259, 1645, 406, 2377, 419, 418, 409, 1802, 405, 424, 416, 1059, 423, 421, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1866, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2081, 2126, 2147, 2282, 2303, 2322, 2566]. library-based [1059]. ’license [2699]. licensed [857]. Licensing [88, 148, 887, 861, 859, 860, 858, 1413]. Lies [1503]. Life [983, 2003, 1753, 1777, 986, 589, 2779, 350, 1283, 2188, 2059, 351, 345, 653, 2276, 1782, 2491, 458, 1275, 2298]. lifelike [6]. lifestyle [1713]. Light [807, 838]. Lights [717, 916]. Like [1468, 1028, 23, 2714, 1452, 1036]. limit [2709, 2650]. Limited [1215, 425]. Limits [1215, 2572, 190]. Line [659, 2047, 2051, 2059, 2048, 2052, 2057, 1589, 2046]. Line-of-Sight [659]. Lineage [1258]. linear [657, 2227]. Lines [1303, 1386, 2454, 2488, 2545, 2636, 2054, 2056, 2055, 2058]. linguistic [1484, 2531]. lining [299]. link [989, 200]. linked [126]. Links [1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1889, 1939, 1965, 1986, 2005, 2025, 2101, 2173, 2190, 2213, 2233, 2259, 2282, 2303, 2322, 1866, 1916, 2044, 2081, 2126, 2147, 2412, 1412]. Linux [1999]. Linux-based [1999]. Liquid [807]. Liskov [2827, 2844, 2761]. List [639]. listings [169]. Literacy [1410, 371, 1343]. literary [828]. literature [1873]. litigious [1351]. Litter [1777]. Little [1236, 1668, 1583, 2970]. Littlest [1302]. livable [1242]. Live [1817, 541, 1816]. Lives [1633, 1891]. Living [839, 960, 2600, 1211]. load [1845]. loads [2500]. loaf [605]. LOC [1261]. local [1822, 742]. Locality [1496, 1636]. Location [2391, 1538, 1176, 384]. Location-based [2391, 1176]. locations [2270]. lock [962]. locomotion [2296]. Log [661, 764, 562, 5, 110, 259, 432]. Logging [87]. Logic [2755, 2838]. Logical [1258]. logistics [442, 1497]. Long [1410, 1992, 1192, 2642, 2899]. Long-Term [1410, 1992]. Look [327, 1347, 1586, 1734, 1862, 141, 1140, 1648, 2549, 2944, 1143, 1240, 2221, 1650, 2465, 2569, 2157]. Looking [926, 2597, 541, 2023, 2793]. Looks [1935]. loose [504]. Lose [1654, 1069]. losers [831]. Loses [496]. losing [2826]. Losses [453]. Lost [148, 1633, 692, 1346, 1733, 344, 2985, 2658]. Love [692, 1257, 1733]. Love/Hate [1257]. Low [1683, 2588]. Low-Tech [1683]. Lowell [1604]. loyalty [1106, 2395, 1201]. Luddite [1961]. lure [1029]. ly [2676].

M [2106, 2421, 1198, 2428]. m-commerce [1198]. M-government [2428]. ma [141, 2082]. Machine [25, 1003, 1028, 2977, 1672, 2952, 204, 2476]. machines [2019, 1057, 2546, 2978, 2600, 2936, 979]. macroeconomic [1749]. Made [397, 496, 2051, 276]. MAESTRO [28]. magazine [2458]. magic [2691]. Magneto [228]. Magneto-optical [228]. Mail [1834, 1378]. Main [2921]. Mainframe [807]. maintainability [1427]. Maintaining [2874, 1851, 1747, 116]. maintenance [751, 1451, 2971]. major [427]. Majoring [1434]. majority [1483]. majorization [2877]. Majors [516]. Make [740, 1028, 1435, 2041, 1912, 854, 1118, 402, 2045, 666, 2535, 1881]. Makeover [1257]. makes [2015, 669, 1245, 2449]. Making [628, 2995, 2603, 890, 790, 269, 2904, 685, 2789, 1242, 1033, 2301, 1831, 2950, 1034, 2663, 2279, 2716, 2024, 2706, 1445, 2921, 76]. Malaria [1982]. Malaysia [1935]. Malware [1165, 1624]. man [2156]. manage [2063, 1065, 1275]. Manageable [1326]. managed [2620, 1843]. Management [983, 1684, 2041, 1637, 965, 3016, 1852, 2210, 1532, 1769, 676, 2098, 2410, 385, 1574, 2681, 2112, 1127, 1465, 1741, 2413, 2942, 2430, 2119, 1121, 939, 1029, 1262, 1351, 1414, 1506, 1713, 1810, 1938, 2004, 2087, 2196, 2264, 2364, 2511, 2732, 2933, 2638, 1789, 2817, 913, 144, 1771, 1046, 1792, 2114, 2780, 1450, 2881, 413, 893, 951, 2054, 401, 1088, 1788, 1793, 1794, 802, 441, 1762, 2052, 1041, 2115, 2741, 791, 2055, 2429, 1765, 1627, 2208, 251, 973, 2811, 2846, 1787, 1203, 2686, 1680, 2831, 2749, 2226, 1113, 805, 196, 2298, 1785, 386, 236, 1791, 2409]. manager [1884]. Managerial [1954, 832]. managers [2973]. Managing [1406, 757, 2030, 1227, 612, 1297, 2382, 1295, 713, 1292, 655, 2927, 2376, 1205, 2143, 673, 443, 799, 1674, 892, 2940, 2188, 1928, 932, 1286, 2364]. Mandate [1215, 1962]. Manhattan [40]. manifestations [266]. Manifesto [320, 1942, 1392, 1531, 1900]. manipulation [192]. manipulations [1419]. manners [2627]. manufacturing [1617, 1618, 2319]. Many [359, 2552, 1461]. map [2715]. Mapping [472, 929, 990]. MapReduce [2356]. March [2690]. marginalize [1356]. Maris [129]. Mark [1163, 618]. Market [1237, 994, 2589, 388, 440, 1359, 2416, 1880, 1641, 1726, 556, 2273, 1951]. market-based [556]. Marketing [1011, 1014, 651, 2911, 1973, 1204, 121]. marketplace [1202, 1675]. Marketplaces [566]. Markets [2463, 2604, 2730, 2998, 2062, 2898, 231, 2092, 2302, 1071, 873, 1556]. Markup [1347, 993, 1321]. Marriages [64, 1693]. marts [2141, 1181]. mas [64]. mass [230]. mass-storage [230]. Master [2828]. Mastering [40, 2133]. Match [1805, 2115]. Matching [457, 510, 1305, 2377, 2074]. Matchmaker [2759]. material [1490]. Math [516, 1141, 1163, 1410, 1215, 785, 1083, 519, 1082, 2669]. Mathematical [938, 1561, 1084, 1620]. mathematics [421]. Mating [318]. matrices [1251]. Matter [2040, 2171, 2833, 1659, 2464, 2888]. matters [2255, 2269, 299, 2509, 2768, 2722, 2548]. mature [1411]. MAUI [1275]. Maurice [2901]. Maxed [740]. Maximal [1410]. Maximizing [550, 3004]. May [40, 664, 2876, 1069]. Mayor [376]. MBA [1325]. McCarthy [1006]. McInternet [2714]. Me [660, 640, 1053, 1410, 1367, 1258, 2022, 717, 1052, 1027, 1982, 763, 2515, 18, 605, 2441]. Meaner [1052]. Meaning [938, 866]. meaningful [1868]. Means [1141, 2022, 1912, 2269, 1352]. Measure [2003, 1316, 1495]. Measurements [31]. Measures [1456, 2279, 2886, 1233]. Measuring [1158, 1367, 2407, 1780, 158]. mechanism [695, 1185]. mechanisms [1627]. media [753, 1896, 1547, 2647, 2646]. mediated [2533, 349]. mediation [2145].

Medical

[541, 2893, 1510, 510, 707, 2795, 1723, 2962, 511, 1514, 341, 1349, 1440, 2594, 1515]. Medicare [1708]. medium [2413, 372, 1111]. medium-sized [2413]. meet [468]. Meeting [592]. meetings [2688]. Meets [1684, 2313, 96, 2942, 2895, 2710]. Mellon [2202]. meltdown [2887]. member [1839, 2135, 2402]. Members [1962, 2043, 2134, 2403]. memoriam [810]. memories [1786]. Memory [219, 11, 2615, 2578, 2671, 2835, 2501, 2474, 2470, 2522, 2617, 2523, 1982, 912]. Memory-chip [219]. MEMS [230]. MEMS-based [230]. Men [496, 2375]. menace [1667]. ment [2021]. Mental [1003, 1047]. Mentally [685]. menu [2423, 1336]. Mercury [515]. Merger [64]. Merging [804, 2351, 1374]. Mersenne [1028]. mess [2133]. Message [886, 372]. messaging [798, 1378, 1638]. messy [2271]. Meta [1392]. Meta-design [1392]. Metadata [1888, 1645, 2299, 2377, 1824]. metaobject [532]. Metaphor [763, 517, 697]. metaphysics [2178]. MetaTM [2523]. meteor [1229]. Meteorites [39]. Meter [1683]. Method [1130, 1181, 972, 1418]. Methodologies [1367, 909, 1598, 2122, 1550]. methodology [2499, 1305, 1499, 2276, 2498, 1582]. Methods [1639, 929, 2225, 1368, 2518, 1494, 2048, 523, 767, 2965, 2738]. metric [2279]. metric-driven [2279]. metrics [2064, 2753, 2091]. Metro [276]. Metropolis [2838]. MGM [1238, 1712]. Micro [2250]. Micro/nanofluidic [2250]. Micromedicine [2790]. Microsoft [1326, 1351, 1938, 2087, 2264]. Microsoft-style [2264]. Middleware [722, 725, 426, 819, 727, 723, 724, 726]. midsize [70]. might [2898]. Mightier [739, 3013]. migrating [1598, 2260]. migrations [75]. milestone [1535]. milestones [1656]. Military [559, 639, 442]. Military-Issue [559]. Millennium [454, 39, 620, 144]. millions [2554]. mind [713, 2913, 142, 1704]. Mindful [1608]. Minds [1163]. mindset [914]. minefields [149]. mining [772, 883, 770, 1996, 336, 768, 774, 83, 773, 476, 2376, 162, 175, 172, 769, 1177, 173, 1314]. Minority [1483, 1831]. minority-friendly [1831]. minority-majority [1483]. Minority-minority [1483]. minute [2835, 1447]. MIS [1723, 1014]. misfits [74]. Misinformation [300, 1000]. misreporting [1152]. Miss [1326]. misses [2797]. Missile [198]. Missing [1028, 2412, 940, 2525]. Misteach [1962]. mistrust [305]. Misuse [1962, 1927, 2281]. mitigate [2529]. Mix [318, 2115]. mix-and-match [2115]. Mixed [619, 1061]. Mixed-width [1061]. Mobile [1741, 786, 1046, 1850, 1639, 1998, 798, 1171, 1311, 1641, 653, 2752, 2386, 1169, 1168, 1179, 1642, 846, 1196, 2184, 1651, 2182, 308, 2602, 992, 1643, 1545, 1173, 2925, 1338, 835, 1521, 897, 1693, 1176, 1644, 1170, 123, 2970, 977, 1172, 1175, 118, 1174, 1523, 1556, 2096]. Mobility [1270, 1640, 410]. Model [2974, 1003, 683, 1767, 1929, 150, 553, 1363, 1312, 2981, 549, 804, 2951, 2515, 498, 819, 1955, 2479, 2059, 2091, 1249, 35, 1188, 1848, 2739, 2838]. model-checking [2479]. Model-driven [1767]. model-integrated [819]. Modeling [127, 1128, 136, 1879, 207, 2224, 2170, 1114, 2591, 1274, 1580, 2225, 2051, 2717, 134, 1131, 531, 817, 2186, 2742, 1510, 1513, 2915, 139, 79, 1361, 1761]. Models [1003, 947, 2050, 1362, 1957, 1047, 2220, 895, 804, 1601, 1619, 1131, 306, 2376, 1338, 854, 2741, 750, 1132]. modern [517, 1953]. Modes [1237]. Modified [2920]. modular [426]. modularity [529]. Molecular [2246, 2475, 2476]. Molecules [572]. momentum [1518]. MonALISA [2903]. MonetDB [2615]. Money [1479, 812, 890, 1314]. mongers [1864]. monitor [1763]. Monitoring [2903, 609, 1975, 1845, 1332, 608, 654]. Monkey [692, 2440]. Monkeying [318]. monoculture [1278]. monopoly [2235]. month [1165]. Moon [541, 1805, 109]. Moore [350, 807, 2770, 1632]. Morphing [64]. Mortality [2107]. mortar [1191]. mosquitoes [330]. ’Most [2415, 258, 1666, 802]. motivate [2203]. motivates [2297]. motivation [157]. Motivational [2595]. Motorola [972]. Motors [1708]. Motto [480]. mountains [1780]. Mouse [480, 141]. mouth [2395]. movement [483]. mover [2818]. Moves [1141, 1585]. Movies [659, 6]. Moving [1049, 2823, 2315, 1135]. MP3 [1153, 397]. mSpace [1871]. Mt. [2788]. Much [88, 590]. mugwump’s [1055]. mugwump’s-eye [1055]. Multi [2053, 2058, 1847, 2170]. Multi-agent [2058, 2170]. Multi-level [2053]. multi-path [1847]. multiagent [598, 1553]. Multibiometric [1220]. Multicast [871]. multicore [3004, 2739, 3005]. multidimensional [2574, 524, 2584]. Multidisciplinary [2314]. multilevel [817, 2915]. multilingual [2425, 1628]. Multimedia [1700, 242, 1642, 15, 705, 1461, 122, 28, 1871]. Multimedia-based [1700]. Multimodal [1223, 1219, 1222, 1221, 2167, 49, 187, 1224, 1871]. Multiparadigm [536]. multiple [804, 1202, 1663]. multiprocessor [2808]. multisensor [81]. multiset [2943]. multisite [73]. Multitasking [2688]. Multivalent [119]. mum [2234]. Municipal [2374]. muscle [2450]. museum [1543, 1694, 489]. Music [1077, 2041, 1969, 541, 618, 1967, 1708, 1043, 2816, 1972, 1153, 1968, 2031, 490, 973, 1091, 1040, 1970, 1116, 1413]. Musings [1908]. Must [1586, 147, 807, 1434, 1351, 853, 111]. mutual [2813]. My [839, 217, 2683, 1892, 2239, 2680]. Myanmar [126]. MyLifeBits [1795]. myocytes [2711]. Mystery [496, 1368]. Myth [1387, 577, 1348, 1144]. Myths [180, 2564, 34, 2373, 1234, 1470].

Nabster [543]. Name [660, 1862, 64, 106, 376, 453, 717, 1632, 1964, 279, 1924, 302]. names [415, 314, 1768]. namespace [1924]. Naming [1835]. NAND [2744]. Nano [639, 1214, 1805, 2040]. Nano-Crab [639]. nanobots [2893]. nanofluidic [2250]. Nanoscience [1453]. nanostructures [2247]. Nanotechnology [1686]. nanotube [2245]. Narrowing [2975]. NASA [683]. Nash [2679, 2678]. National [660, 586, 510, 1751, 3015, 1109, 289, 1485, 421]. national-level [289]. national-security [1751]. Nations [1140, 2898]. nationwide [822]. native [2825, 2885]. Natural [1399, 185, 1396, 2960, 247, 2195, 2552, 171]. naturally [49]. Nature [1734, 1507, 2013]. nature-inspired [2013]. navigating [668]. Navigation [541, 2296, 1036]. Navy [1382]. Neanderthal [1982]. Near [2358, 2357]. Near-optimal [2358]. nearest [2358]. ’nearshore [2269, 2270]. Necessary [2279, 1327, 1291]. Need [258, 1141, 739, 2432, 2249, 468, 984, 519, 1647, 2265, 102]. Needs [300, 481, 542, 1367, 1754, 565, 1751, 1187, 2771, 1679, 2365, 2254]. Neglect [1347]. neglected [1829]. neglecting [2456]. Negotiating [1383]. Negotiation [582]. neighbor [2358, 2357]. neighborhood [1719]. neighbors [1739]. Neo [1798]. Neo-tribes [1798]. nerds [470]. Net [40, 397, 376, 453, 717, 959, 916, 1027, 1162, 1887, 1777, 961, 1013, 2, 18, 126, 2928, 2921, 1388, 268, 916, 170, 172, 44, 1861]. net-based [1013]. Netflix [2925]. Netizens [1585, 1528]. Network [1230, 839, 1436, 226, 2800, 2670, 304, 753, 217, 2028, 2349, 1046, 2845, 2034, 2031, 2762, 2719, 742, 1746, 2102, 95, 274, 1782, 1848, 2560, 338, 2082, 1719, 1625, 2181]. network-centric [2031]. Networked [1107, 2118, 2030, 711]. networker [2946]. Networking [987, 2027, 1334, 1269, 1145, 505, 2164, 2478, 2976]. Networks [1018, 2681, 242, 846, 385, 1597, 779, 1331, 2182, 2024, 1863, 2837, 2142, 2711, 1333, 669, 988, 2571, 2687, 2582, 2031, 1847, 2478, 2396, 835, 2374, 1591, 1296, 1159, 1797, 2849, 1335, 155, 1204, 1332, 118, 871, 1334, 2550, 2181]. neurocomputing [2251]. neutrality [2928, 2762, 2082, 2670]. Never [344]. Newest [453, 717]. News [2993, 2960, 2727, 2759, 2541, 2, 257, 147, 39, 87, 219, 18, 64, 179, 106, 126, 200, 515, 276, 299, 397, 318, 376, 453, 541, 559, 430, 480, 496, 762, 618, 639, 739, 717, 807, 692, 659, 785, 866, 838, 588, 982, 886, 959, 916, 937, 2827, 2929, 2792, 2963, 2509, 2760, 2995, 2826, 2570, 2542, 2510, 2791, 2962, 2996, 2571, 2928, 2729, 2828, 2632, 2507, 2961, 2758, 2825, 2508, 2539, 2572, 2631, 2927, 24, 2790, 2634, 1052, 1027, 1002, 1076, 1162, 1118, 1140, 1257, 1302, 1346, 1409, 1280, 1386, 1236, 1214, 1325, 1434, 1366, 1456, 1503, 1733, 1608, 1753, 1585, 1632, 1683, 1708, 1528, 1478]. News [1560, 1654, 1961, 1861, 1911, 1887, 1982, 1777, 1805, 2002, 2021, 2040, 1834, 1934, 2067, 2084, 2104, 2131, 2153, 2177, 2194, 2216, 2237, 2262, 2285, 2306, 2361, 2380, 2399, 2419, 2437, 2829, 2761, 2725, 2793, 2728, 2540, 2824, 2633, 2726, 2994, 25, 30, 29, 1727, 26]. NewsInEssence [1727]. newspaper [422]. Next [1821, 1214, 634, 2042, 2344, 1377, 1123, 2674, 2961, 2206, 1198, 1378, 287, 724, 507, 1814]. Next-generation [1821, 2042, 2206, 724, 507, 1814]. nice [1291]. night [1380]. nightmare [2, 1349]. NINJA [534]. Nintendo [661]. NIUPEPA [422]. No [640, 1141, 938, 887, 1387, 1326, 1504, 1754, 2041, 1806, 1962, 147, 299, 496, 959, 937, 1302, 1528, 2, 537, 141, 2794, 21, 1856]. node [87]. Noise [838]. Nokialand [659]. nominating [2312]. nominations [2218, 1839, 2106, 2135, 2421, 2402]. Nominees [1164, 2312]. Non [2748, 300, 252, 974, 2653, 1036, 2916]. non-adopters [974, 2653]. non-copyrightable [2916]. non-game [1036]. Non-OOSD [300]. Non-programmed [252]. Non-work [2748]. Nonconsensual [582]. nondeterministic [1087]. noninfringing [183]. nonsecurity [1432]. Nose [1478, 87]. Not-defect [1411]. Not-for-profits [288]. nothing [2219]. Notice [1516, 1517, 1827, 2086]. notification [949]. notion [718, 636, 621]. novice [1957, 1047, 55]. NP [2907]. NSA [541]. NSF [907, 421]. nuances [2670]. nuisance [2313, 1667]. nuke [2]. Number [1028, 1214, 1503, 1478, 999]. Numbers [542, 453, 588, 1002, 1409, 179, 2289, 775, 2777, 20]. Numerical [684, 1554]. Nurture [1935]. Nurturing [1698]. NWRC [2748].

Oath [430]. Object [1114, 248, 2749, 956, 415, 1830, 1362, 36, 2426, 212, 391, 2230, 1137]. Object-oriented [1114, 248, 956, 1830, 2426, 212, 391, 2230, 1137]. Objective [839, 1136]. Objectively [804, 197, 61, 512, 614, 678, 934, 236, 392]. Objects [1561, 684, 1468, 2014, 1524, 2978, 216, 412]. oblivious [1706]. obscurity [1160]. observability [294]. obsolescence [128]. obstacles [1470]. Obvious [2798]. occur [677]. Ocean [127, 79]. oceanarium [140]. Octopus [1097]. ODC [1312]. Ode [2346]. ODP [1253]. Odyssey [39]. Off [88, 839, 2, 1977]. offend [1666]. Office [106]. officers [2866]. officially [276]. officials [290]. Offline [480, 737, 1285]. Offshore [687, 2069, 2813, 2453, 2143, 2099, 2234]. offshore-outsourced [2453]. Offshoring [2983, 1861, 2564, 1813]. often [239]. oil [3019]. OK [1586]. Old [833, 1632, 2007, 2458, 500, 2894, 1908, 709]. old-school [1908]. older [1525]. Olfactory [64]. Olympiad [1634]. OM [65]. Ominous [1528]. on-demand [530]. On-the-fly [845]. One [402, 431, 560, 2041, 1806, 739, 785, 937, 985, 2156, 2803, 1503, 2614, 2916, 804, 2944, 1358, 1447]. one-minute [1447]. ongoing [2089, 2724]. Online [2240, 640, 1410, 1684, 2716, 179, 106, 480, 838, 2753, 3015, 2117, 1325, 283, 286, 1184, 737, 671, 2210, 548, 1185, 1043, 1466, 834, 2028, 2029, 758, 667, 975, 488, 87, 263, 2395, 1190, 1721, 2571, 1700, 1798, 2921, 8, 2682, 2585, 2782, 2123, 1980, 2257, 1973, 1673, 1291, 1727, 2506, 2538, 2569, 2459, 2630, 2992, 2959, 2789, 2823, 2724, 2757, 1356, 270, 267, 700, 1430, 2756, 2926, 1452, 2258, 873, 1360, 1204, 1285, 1212, 2254, 1337, 2919, 2531]. Only [88, 960, 1028, 1561, 1806, 283, 2161, 2578, 2944, 2857]. onto [2744]. OntoClean [629]. Ontological [622, 681, 629, 1249]. Ontologies [2612, 628, 627]. Ontology [625, 2637, 1449, 626, 898, 1764, 1132, 1761]. ontology-driven [1764]. OO [214, 929, 3, 1633, 1130, 1602]. OOSCD [300]. OOSD [300]. Open [1094, 1237, 1281, 1258, 1479, 1633, 2119, 2613, 1244, 407, 2718, 1427, 1409, 2819, 408, 1515, 439, 728, 2514, 1182, 2098, 2118, 2543, 1999, 2940, 1414, 630, 1153, 1143, 1240, 3017, 1678, 2079, 2872, 530, 2006, 2629, 2797, 2406, 1258]. Open-source [408]. Opening [2997, 2624, 1027, 1601]. Opens [455]. Operating [563, 529, 309, 2522, 2523]. operational [1836, 1129]. operations [2943, 2319, 933]. Opinion [839, 1651, 1705, 1605, 566, 835, 82, 2212, 1322, 2565, 1024, 1406, 1094, 1500, 1525, 2064, 2688, 883, 2886, 2148, 1431, 2191, 759, 2852, 913, 144, 1453, 636, 2395, 736, 1299, 36, 2947, 2753, 2172, 1557, 1750, 1524, 1931, 1908, 2377, 2921, 613, 103, 537, 1629, 1884, 2818, 2627, 2416, 2721, 1049, 273, 999, 1958, 450, 1802, 1073, 2987, 2100, 2658, 782, 2266, 1730, 2234, 2785, 1680, 1857, 123, 1254, 1979, 3021, 2258, 2324, 956, 1807, 2127, 1343, 979, 538, 2434, 1137, 2595, 863, 2323]. OPM [1129]. OPM-based [1129]. Opportunities [1633, 1168, 237, 393, 1282]. opportunity [1738, 368]. Opt [64, 303]. opt-in [303]. opt-out [303]. Optical [2249, 1628, 228]. optimal [2358]. Optimistic [2909, 1344]. optimization [2942, 2625, 1344]. optional [1327]. OptIPuter [1149]. Oracle [2683]. oracles [2230]. order [506, 2983]. orders [203]. Oregon [1355]. O’Reilly [1710]. org [2661, 127]. Organic [2446, 2444, 2442, 912, 2447, 2445, 2451, 2448, 2449, 2246]. organism [695]. Organisms [960]. Organization [577, 910, 2319, 1930, 2496, 2011]. organizational [11, 1532, 1178, 1066, 932, 2747, 1623, 2593, 2076, 2229]. organizations [235, 1596, 14]. organizing [2207]. orientations [732]. Oriented [1122, 1725, 1114, 381, 1830, 1127, 248, 1383, 521, 2287, 2426, 212, 391, 2052, 523, 528, 530, 387, 532, 956, 2230, 1137]. Original [1806]. origins [2159, 1079]. ORM [2737]. Orthodonture [257]. Other [148, 1141, 1632, 1964, 1354, 2038, 1263]. Our [2745, 2726, 2826, 1487, 1739, 1758, 177]. ours [329]. ourselves [1739]. out-of-box [223]. outcome [2851]. outcome-based [2851]. outcomes [1293]. outdoor [596, 1491]. Outlawing [542]. Outperform [496, 2947]. outreach [2925]. Outsourced [1255, 2920, 2453, 2035]. Outsourcing [2455, 1076, 1325, 2002, 2813, 2125, 1180, 1292, 998, 2234, 2099, 1313]. Overcoming [2954, 815, 2918]. overflow [1743]. Overhaul [200]. overlay [1847]. Overload [541, 2766, 832]. Overseas [659]. Oversight [1887]. overview [2780]. Own [1457, 1684, 1366, 1838]. owner [2682]. Ownership [1347, 612]. Owning [1805].

Oxley [2125, 1610].

P [2907, 89]. P-books [89]. P2P [925, 1406, 926, 1581]. package [1358]. packaged [1605]. packages [77]. Pacman [1654]. page [1678, 2757]. Pages [559, 2884]. Pain [397]. Painful [1528]. Paint [453, 739, 1409]. painting [1373]. Pair [1978, 2012, 102]. panacea [1349, 2561]. Pandora [2561]. panel [194]. panic [557]. Pants [318]. Paper [938, 219, 376, 2637, 2961]. Paperless [1057, 2546, 233]. papermill [12]. papers [2683, 1460]. parable [1758]. paradigm [1340, 2999, 122, 1974, 1010]. paradigms [2243]. paradox [953, 2533, 334, 45, 2783, 1730, 2224]. Paradoxes [581, 2564]. Parallel [763, 2671, 2252, 2939, 2892, 210, 2807]. Parallelism [660, 2900, 2909, 2908]. Parallels [1709]. parameters [932]. Paranoia [660]. Parliament [1409]. Part [692, 2610, 2517]. participation [1243, 2095, 2652, 2997, 2117, 2621, 2315]. Participatory [377, 1330]. Parties [917]. partitioning [2556, 2677, 2555]. partners [1488, 710]. partnership [2757]. Partnerships [32, 1045, 1105]. party [2065, 1606, 1497]. Pashto [639]. Pashto-Phonics [639]. pass [1060]. Passport [1409]. passports [3002]. Password [1326, 1294]. Passwords [65, 3019]. Past [663, 327, 2164, 2152, 1487, 998, 353, 2338]. Pat [2608]. patch [2881]. Patching [2616]. patchwork [67]. Patent [376, 1478, 2073, 942, 112, 2365, 1328, 2575, 2088]. patentable [311, 2464, 2965]. Patents [1367, 2040, 2178, 2271]. path [2740, 1847]. pathologies [2868]. pathways [1873]. Patience [1028]. patient [2454, 510, 2415]. Patriot [960]. Patron [397]. Patter [453]. pattern [1045]. Patterns [2387, 1362, 2188, 1877, 1729, 1041, 2055, 874].

Pausch [2510]. Paving [2489]. Pay [1141, 1162, 1366, 2956, 18]. Paying [807]. payment [1097]. Payments [992, 2185]. payoff [1102, 910]. Pays [1654]. PBE [127]. PC [974, 502, 1140, 690]. PCs [547]. PDA [866]. PDF [763, 839]. Peak [1934]. pearls [2986]. pearly [2154]. Pedagogical [1163]. Pedagogy [1281]. Pedia [1934]. Peer [2394, 1237, 2365, 922, 923, 1086, 1443, 996]. Peer-based [2394]. Peer-to-peer [922, 923, 1086, 1443, 996]. Peeves [839]. Pen [739, 3013]. Pending [2040]. pendulum [1518]. people [159, 2905, 2146, 1665, 2370, 1078, 1765]. perceived [1207, 2191]. percent [2179]. perception [1696, 50]. Perceptions [480, 580, 1114, 696, 501, 778, 1360]. Perceptual [52, 51, 49, 47, 50, 53, 48, 46]. perfect [2483]. Perfict [1280]. Performance [1578, 686, 1861, 735, 1441, 634, 1145, 1316, 2620, 1147, 535, 448, 2031, 1599, 35, 2672, 1263, 2498, 2300, 2609, 1903, 2739, 533]. performers [2588]. peril [2729, 347]. peril-the [347]. perils [1826, 1152, 273, 539]. Permanence [2606]. Pernicious [1757]. persist [1219]. Persistence [1862]. Person [937, 2871]. Personal [1586, 1787, 89, 1052, 1785, 293, 662, 2045, 907, 1872, 1789, 1792, 1795, 128, 2706, 992, 1788, 1793, 1794, 2392, 92, 165, 166, 1791]. personality [2171, 1276]. Personalization [1725, 704, 220, 1806, 172, 1989, 700, 2094, 174, 701, 2222, 156, 154, 175, 155, 152, 164, 153]. personalize [168, 167]. Personalized [707, 708, 705, 795, 702, 155, 152, 169, 153, 171]. personnel [448]. persons [1404]. Perspective [2475, 220, 1504, 2477, 1725, 2773, 2524, 1642, 2942, 2357, 776, 2910, 2975, 2589, 2944, 2743, 2583, 3010, 2874, 2775, 2807, 2678, 860, 1180, 844, 3008, 927, 2553, 2585, 2648, 467, 2741, 2710, 2522, 2842, 1201, 2805, 186, 2355, 2498, 2616, 2840, 2876, 556, 2500, 2977, 2646, 2708, 2614, 2555, 1898, 2676, 281, 2908]. Perspectives [1992, 868, 1202, 2268, 129, 222, 433, 1079, 1797, 2530, 2076, 416, 1808]. pervasive [2347, 1541]. Pester [2619]. PetaFLOPS [1442]. Petri [1922]. petunias [2007]. PFIRES [1042]. Pharm [1608]. Phase [229]. phenomenal [162]. Philosophical [1410]. philosophy [2014]. Phishing [1864, 1503, 1476, 2278]. Phone [515, 496, 1027, 1257, 87]. Phones [430, 1777, 2191, 2925, 2721, 2752, 2970]. Phonics [639]. photo [1876]. photographs [2554]. Photography [2662, 2553]. Photomapping [515]. Photosensing [1692]. Physical [1479, 2275, 2435, 1541, 1243, 1544, 555, 2477]. Physically [1373]. Physics [133, 336, 135, 2836]. Picks [299]. Picture [693, 1503, 2002, 2503, 1407]. pictures [2427]. piece [2525]. Pillage [147]. pilot [2840]. PIM [1792]. Ping [2968]. pioneer [2634]. Piracy [258, 1927, 106, 453, 1654, 1925, 1043, 1153, 193, 272, 1109, 1232]. pitfalls [1605]. pivotal [2988]. Pixar [6]. Pixel [785]. pixelized [485]. pixies [485]. PKI [955, 16, 37, 451]. Plaît [1118]. place [1609, 2172, 2503, 334]. Plagiarism [1609, 1915, 1577, 1915]. Plagiarize [453]. Plan [1862, 1162, 1409, 2021, 1685, 2863]. Plane [541, 1076, 2979]. planes [517]. planners [2850]. Planning [40, 72, 1947, 75, 69, 1405, 73, 76, 74, 77, 71, 70]. Plans [541, 1685, 1731]. plant [139]. plants [139]. Plasma [64]. Plateau [1528]. platform [846, 2882]. platforms [245, 1333, 2598, 996]. Plato [2014]. play [2107, 736, 1216]. player [2029]. playing [2921]. PlayStation [1052, 1325]. Plea [107, 693]. Please [640, 2459]. plot [1736]. plush [3009]. Poetic [1236]. Point [2466, 220, 2834, 2546, 2670, 990, 2797]. Point/Counterpoint [2466, 2670, 2834, 2546]. Pointers [146, 298, 317, 375, 429, 558, 1, 17, 38, 63, 86, 105, 125, 178, 199, 218, 452, 479, 495, 514, 540, 587, 617, 638, 658, 691, 716, 738, 761, 784, 806, 837, 865, 885, 915, 936, 958, 981, 1001, 1026, 1051, 1075, 1117, 1139, 1161, 1213, 1235, 1256, 1279, 1301, 1324, 1345, 1365, 1385, 1408, 1433, 1455, 1477, 1502, 1527, 1559, 1584, 1607, 1631, 1653, 1682, 1707, 1732, 1752, 1776, 1804, 1833, 1860, 1886, 1910, 1933, 1960, 1981, 2001, 2020, 2039, 2066, 2083, 2103, 2129, 2150, 2175, 2193, 2215, 2236, 2261, 2284]. pointers [2305, 2326, 2360, 2379, 2397, 2418, 2436]. Points [838]. POLA [1167]. Polaris [2583, 1997, 2584]. poles [347]. police [1355]. policies [1306, 2257, 3021]. Policy [1982, 2210, 1915, 1652, 2861, 1977, 1042, 356, 607, 1706, 2659, 2434, 2605, 2513, 2899]. policymakers [2605]. politeness [1287]. Political [106, 1280, 1386]. Politicians [1961]. politics [221, 126, 352]. polity [288]. Poll [397]. Polly [200]. Pondering [485]. Pooh [299]. pool [417]. pools [1846]. Popular [2021, 1245]. popularity [1687]. Populating [489]. Population [453]. pornography [609]. port [539]. portal [2981, 436, 455, 1437]. Portals [1429, 1233]. portfolios [1532]. Portion [480]. Portrait [1560]. ports [1757]. posers [1864]. poses [1663]. position [858]. Positioning [997]. positions [2560]. Positive [1806, 659, 3020]. Posse [1709, 1632]. Possible [866, 1705]. Post [2268, 1060, 1928, 743, 1808, 502]. post-baccalaureate [743]. Post-conflict [2268]. post-deployment [1928]. Post-pass [1060]. post-PC [502]. post-war [1808]. potamus [1350]. Potato [1162]. Potential [1329, 1798, 1605]. potentials [1415]. Potter [1585]. poverty [2485]. Power [376, 736, 1078, 3014, 2256, 2619, 1996, 3004, 746, 1445, 1798, 644, 775, 1718]. Powerful [1077]. Powering [2516]. PowerPoint [2309]. Practical [1053, 1771, 128, 41, 378, 277, 498, 765, 621, 841, 1055, 985, 1143, 1240, 1305, 1368, 1438, 1535, 1736, 1656, 1837, 2023, 1963, 2156, 2221, 2287, 2382, 2440, 591, 2225, 2051, 120, 2873]. Practice [2577, 2578, 2579, 2517, 2549, 3004, 2515, 2516, 2518, 2547, 3002, 3003, 2548, 824, 701, 41, 2156, 2221, 2048, 1555, 672, 1829, 1110, 2941]. Practices [1003, 799, 2920, 1855, 2219, 2716, 1883, 1994, 1728, 2625, 251, 1228, 307]. practitioner [2853, 2287, 1405]. practitioner-oriented [2287]. pragmatic [2006, 1900]. Praise [740, 1077]. precise [2230]. Precision [2022, 1076]. predators [1185, 710]. Predatory [220, 90]. predict [1749]. predictability [2127]. Predicting [2978, 627, 2142]. Predictions [107, 1215, 2977]. Preferences [1777, 1579]. Preferred [635]. prejudice [2512]. Premium [2302]. Prep [2537]. preparation [1203]. preparatory [2576]. Prepare [1410, 2692, 142, 1704, 1591]. preparedness [2117]. preps [2630]. Prerogatives [581]. presence [2822]. present [998, 1825, 2338]. presentation [195]. preservation [2611, 1828, 1023]. Preserving [1390]. President [2922, 2348, 91, 44, 2152, 2288, 2331, 2480, 2664, 2503, 790, 1460, 1531, 1613, 1659, 1483, 1710, 1590, 1634, 1687, 1739, 1563, 1505, 1840, 1913, 1783, 1813, 1892, 1867, 7]. Press [639, 1302]. Pressing [866]. Pressure [807]. prestige [2353]. Prevails [376, 1683]. prevent [2784]. prevention [2142, 1743, 2918]. Preview [106, 2354]. Previewing [2599]. previous [277]. Price [960, 664, 2966, 2302]. pricing [2196]. Pride [2512]. primer [206]. Primitive [782, 2478]. principal [2270]. principle [1636, 2286]. Principles [3, 1435, 1258, 1828, 409, 2746, 1940, 1142, 947, 2660]. print [2757]. print-Web [2757]. Printed [376]. Printed-On [376]. priorities [2879, 606]. Prioritizing [179]. Priority [40]. prism [2124]. Prison [717]. Privacy [1579, 1586, 1504, 219, 559, 496, 2509, 2668, 2222, 1542, 2964, 2574, 1988, 1239, 2866, 2794, 1982, 2210, 1476, 662, 968, 2716, 2969, 975, 179, 842, 308, 1788, 708, 3015, 1200, 1552, 1975, 1695, 1089, 2257, 1989, 2970, 513, 307, 1907, 2756, 1766, 164]. privacy-aware [1989]. Privacy-enhanced [2222]. Private [1018, 1891, 2197]. privilege [2098]. Privileges [692]. prize [87, 2925]. Proactive [1267, 2450, 94]. Probabilistic [2839, 649]. Probabilistically [2707]. Probability [2878, 2617]. Probing [2936]. Problem [2041, 1983, 2003, 2996, 2040, 2447, 2907, 1305, 2894, 2054, 197, 2574, 1940]. problems [636, 21, 988, 1415]. procedures [751, 1692]. Process [481, 917, 1862, 688, 1725, 278, 457, 1520, 2639, 1057, 896, 314, 2881, 1800, 633, 1750, 2168, 2429, 49, 2276, 12, 356, 1979, 1313, 1820, 1763, 801, 392, 492]. process-oriented [1725]. Processes [687, 585, 2007, 652, 714, 1274, 1897, 1167, 932]. processing [247, 2356, 2427, 2944, 2758, 1223, 2560, 1334, 1515]. Processors [180, 2800, 2936]. procurement [100, 2062]. Produce [1862]. Produces [1410]. Product [431, 2054, 440, 686, 1934, 2049, 2047, 2051, 2188, 2168, 2954, 2059, 2056, 2275, 2048, 2052, 2055, 2318, 2058, 2057, 2393, 1275, 2046, 2228, 1616]. Product-market [440]. Production [1962, 2769]. Productivity [1258, 953, 1158, 2148, 1320, 2533, 1066, 2414, 1730, 1381]. Products [938, 2040, 2973, 939, 731, 2317, 2598]. Profession [544, 788, 1564, 1636, 398, 379, 483, 301, 719, 1030, 1142, 940, 1348, 1436, 1283, 1735, 2024, 1863, 1937, 2133, 2195, 2286, 2607, 2401, 2487, 2796, 2897, 2698, 2999, 2375]. Professional [807, 544, 2822, 732, 2036, 2831, 674, 1516, 1710, 1282]. Professionalism [1077, 2149, 862]. professionals [283, 799, 2852, 1883, 2544, 1623, 674, 287, 1277, 623]. professors [1680]. Proff [1862]. profile [2453, 1177, 602].

Profiles [1002, 1733, 1358]. profiling [2684, 1907]. profits [288]. Profs [1753]. Program [2766, 867, 1457, 1983, 2240, 907, 1058, 661, 2455, 1978, 2616, 941, 1087, 2840, 743, 421]. Program-Verification [867]. programmability [3005]. programmable [2247]. programmed [1774, 252]. Programmer [1053, 841, 2015, 128, 41, 378, 277, 498, 765, 621, 1055, 985, 1143, 1240, 1305, 1368, 1438, 1535, 1736, 1656, 1837, 2023, 1963, 2156, 2221, 2287, 2382, 2440, 1908, 2155, 463]. Programmers [1561, 1351, 1348, 484]. Programming [58, 148, 1633, 1778, 1935, 2003, 1231, 59, 54, 450, 56, 57, 803, 55, 60, 1236, 2165, 201, 2264, 2671, 2975, 521, 2012, 1330, 2935, 2203, 2704, 523, 273, 1978, 528, 1396, 1634, 2972, 166, 1993, 532, 956, 102, 1137]. Programs [1806, 591, 1252, 248, 2296, 2807, 1157, 2227, 1493, 1602]. Progress [635, 1111]. Project [1532, 40, 560, 1367, 840, 1769, 3018, 2287, 1339, 401, 1576, 2558, 822, 732, 534, 2778, 1680, 2300, 1210, 1293, 2851, 40, 683]. projection [1099, 1038]. Projects [917, 2022, 1983, 517, 589, 695, 1610, 2920, 1297, 2951, 2023, 2453, 1879, 389, 2008, 1958, 2450, 260, 2848, 2226, 2948, 2035, 1317]. Promise [347, 925, 21, 1008, 420]. Promo [1386]. Promoting [2392, 1994]. promotion [822]. Proof [2041, 1002, 765]. proofs [2707]. Propaganda [1655]. propagation [2204]. property [2906, 310, 352, 2345]. Proposal [220, 2136, 151]. Proposals [497]. prospect [2333]. Prospects [896, 922]. protect [902]. Protecting [243, 1621]. Protection [739, 1134, 380, 2916]. Protocol [1237, 1829, 2183]. protocols [989, 1006, 1147, 532]. prototyping [1620, 1616]. provenance [2408]. Provide [2041]. provider [931]. providing [2605]. provisioning [2592]. provocative [2180]. pseudo [1137]. psychology [2174, 2405, 2359]. Public [2118, 1655, 496, 659, 849, 1024, 1377, 1652, 1742, 1306, 290, 1667, 289, 1596, 356, 2967, 2197]. public-domain [1742]. public-key [1024, 1377]. public-resource [849]. publication [2639, 712]. Publications [1655, 2596, 2189]. Publisher [2458, 2536, 2504, 2690, 2820, 635]. Publishing [585, 402, 696, 2718, 635]. purchasing [1184]. purity [2378]. PURLs [415]. Purpose [2735, 2476, 2495]. Purpose-built [2735]. Pursue [1912, 2274]. Pushing [340]. Put [1237, 1457, 1778]. Putting [701, 3021]. puzzle [1637, 2511]. Puzzled [2502, 2526, 2587, 2618, 2680, 2712, 2777, 2979, 3012, 2878, 2809, 2912]. Puzzles [1247].

Q&A [2479, 2557, 2650, 2844, 2946, 2745]. Quake [599]. Quality [431, 1479, 453, 686, 689, 2875, 1362, 1957, 1926, 2952, 2981, 2028, 758, 2874, 2059, 931, 1136, 1724, 2880, 1978, 808, 1251, 1846, 1067, 196, 1019, 1702]. quality-based [2059]. Quanta}fying [2575]. Quantifying [2980]. quantity [808]. Quantum [2461, 2248]. quarks [1830]. Query [1971, 2942, 2584, 1334]. quest [1854, 2156]. question [1699, 303, 451, 2783, 2992, 2519, 2205]. question/answer [2205]. Questioning [1273]. questions [2794, 2342]. Queue} [1613]. queues [2776]. Quick [299, 286]. quickly [1254]. Quicky [1683]. quiet [859]. quiz [2].

R [2702]. Race [690, 2267]. Radical [2713, 677]. Radio [453, 496, 10, 257]. Rage [318, 1280]. Rages [106]. rails [87]. raise [842]. RAM [2404]. Random [1028, 999]. Randy [2510]. rank [1499]. Ranking [2884, 1136, 2189]. Rankings [219, 1753, 1522]. Rant [220]. Rapid [39, 1620, 1616, 241, 493, 812]. rapidly [929]. Rat [937]. Ratbots [739]. rate [2415]. rates [1109, 1401]. rather [808]. Rating [284, 2528]. Ratings [1805]. Ray [219, 1689]. RBAC [194]. Re [2816, 828, 524]. re-attain [2816]. Re-creating [828]. Re-tuning [2816]. Reach [886]. reaching [369]. read [2891]. Readable [693]. reader [828]. Reading [1586, 828]. Ready [659, 1325, 2994, 1822, 432]. ready-at-hand [432]. Real [1327, 2577, 1410, 1983, 997, 2204, 1751, 2435, 1538, 519, 2351, 1374, 963, 2959, 745, 1859, 725, 123, 138, 1859]. Real-time [997, 2204, 725]. Real-World [1983, 2577, 1538]. realistic [2843, 135]. realities [34]. Reality [740, 1410, 1479, 1912, 752, 120, 2368, 2803, 596, 1372, 747]. Really [300, 1503, 192, 1963, 233, 1200, 1712, 2699, 1381, 102, 1664]. realm [2340]. reason [519, 620]. reasoning [1084, 2651]. reasons [1867, 2595]. rebellion [543]. Rebirth [721]. reboot [2728]. recants [1423]. Receipt [982]. Receive [2432]. Recentering [1735]. Receptionist [1834]. Reciprocity [1926]. recognition [1226, 971, 1628, 2860, 187, 190, 1970]. Recognizing [1590]. Recommendation [2756]. reconfiguring [647]. reconnaissance [491]. Record [1087, 2377]. Record/replay [1087]. recording [229]. Records [1932, 510, 2962, 511, 2594]. recover [1015]. recovering [2505]. Recovery [916, 246, 1844]. recreation [603]. Recruiting [739, 2206]. recruitment [700]. Red [1002, 561]. Redefined [717]. redesigned [2569]. Reducing [1393, 2795, 2433, 1624, 2918]. redundancies [628]. Redux [1028, 788, 2092, 1138, 2547]. reengineering [633, 2275]. Reexamining [11]. reference [1906]. references [908]. refinement [2394]. Reflecting [2705]. reflection [532]. Reflections [2784, 454, 1414, 2350, 1634, 2932, 1025, 2543, 2931, 2764, 2122, 2266]. reflective [723]. Reflects [1835]. reform [1328, 2088, 2899]. reformed [2265].

Reframing [2805]. refresher [1100]. Refreshing [470]. Refuse [1683]. regional [624]. registered [1865]. registration [1807]. Regulating [1809]. regulation [67]. Regulations [1655]. regulatory [896]. Reins [916]. Reinvent [1258]. reissuing [500]. related [2748, 1681, 2350, 1575]. relating [1959]. Relational [2942, 1734, 2471]. relations [1498]. Relationship [1067, 1257, 1182, 1250, 1926, 1316, 1450, 1180, 2657]. relationships [2256, 758, 2320, 1851, 1649, 443, 2777]. Relative [2742, 1746]. Relax [1560]. Relay [690]. releases [902]. relevance [881]. relevancy [2847]. relevancy-based [2847]. Relevant [1912]. reliability [1923]. relic [201]. remarks [2997]. remember [2774]. Remembering [2486]. Remote [1451, 739, 811, 872]. Removal [559]. removing [59]. rendered [850]. Rendering [1115]. Reorg [983, 919]. repair [1451]. Repeating [180]. replace [1318]. replay [2808, 2807, 1087]. Report [2802, 2072, 1164, 2820, 2367, 2640, 1963, 1838, 1781, 600, 1218, 1469, 1861, 2312]. reporting [730, 2914]. Reports [1304]. repositories [2207]. representation [163]. representations [214, 2231]. Representative [377]. Representing [1361]. Reprogramming [2576]. Republish [1586]. Reputation [265]. requests [2940]. require [1082]. Requirement [1028, 1172]. Requirements [1163, 1862, 2393, 1178, 1092, 271]. requires [2909, 862]. requisite [1884]. Rescue [1271, 1708, 2118, 2790]. Rescuing [1739]. Research [660, 1806, 588, 1921, 595, 2734, 1604, 1230, 696, 1640, 1759, 2578, 1942, 2354, 1341, 2287, 2674, 1678, 1994, 1390, 598, 2414, 2490, 593, 2314, 1553, 1150, 1641, 31, 3001, 1460, 1758, 2293, 2189, 828, 1945, 906, 2258, 2389, 1440, 2754, 2990, 2857, 2802]. researchers [2221]. reservation [550]. reserve [1355]. reshapes [1466]. Resizing [1236]. resolution [2949]. Resolving [867, 1633, 628]. resonance [2816]. Resource [1947, 849, 415, 72, 75, 69, 73, 1846, 76, 74, 77, 71, 70]. resources [2514, 418]. respective [2270]. respond [1596]. responder [1270]. Responding [1081, 2854]. Responds [619]. Response [1709, 1935, 1983, 2656, 2430, 1153, 508, 2116, 2117, 1155, 663, 2111, 1523]. Responsibilities [1526, 635]. Responsibility [660, 2192, 2392]. Responsible [787, 1141]. responsiveness [1036]. rest [3008]. Restoration [219]. Restoring [1687]. results [1936, 2230]. resurrection [2764]. retail [446, 1430]. retain [1337]. Retarded [685]. retargeting [2647]. retention [427, 1978]. Rethink [1003, 2187]. Rethinking [2758, 744, 1022, 2667]. Rethought [376]. Retiring [1632]. Retrieval [1967, 1619, 1154, 2295, 2738]. retrieving [1972]. retrospect [2333, 145]. Retrospective [2935]. return [1562, 444]. returns [3020]. reusable [418]. reuse [1796, 1294, 2317, 1071, 255]. reusing [2065]. Revelations [2626]. Reveling [2902]. revenge [2082]. Revenue [692]. Reverse [809]. Review [454, 938, 2161, 2171, 1233]. Reviewer [454]. reviewing [869]. reviews [2954, 2520, 2934]. revision [2904]. revision-control [2904]. revisited [662, 1045, 297, 1832, 1232, 1025]. Revisiting [1892, 2464]. revisits [1612]. Revitalizing [2872, 2458]. Reviving [1841]. revolution [2090, 402, 552, 1721, 2473, 852, 1944]. Revolutionary [1529, 1438]. Revolutionizing [562]. rewarding [277]. rewards [2288]. Rewrite [1052]. RFID [2060, 1690, 1754, 1697, 1696, 1694, 1216, 1909, 2208, 1695, 3002, 2561, 1691, 1689]. RFID-based [1691]. Rhetorical [1835]. RIAA [520]. Rice [860]. rich [1092, 1148]. Ride [1409]. Riding [1911]. right [2172, 2966, 2202]. Rights [1655, 1684, 965, 635, 2320, 177, 805, 484]. Rigor [938, 1778]. rime [2155]. RIP [1076]. Ripping [866]. Rise [1777, 1915, 2764, 2494, 3020, 941]. Rising [1151, 109]. Risk [983, 2041, 1532, 2279, 1562, 234, 2410, 2969, 893, 951, 2453, 1108, 2128, 2234, 2143, 1207, 3021, 513, 1447, 654]. risk-based [2128]. riskier [2235]. Risks [241, 737, 1816, 657, 1323, 1300, 2149, 864, 85, 198, 586, 1681, 2102, 1234, 836, 1606, 513, 539, 1887, 805, 679, 1064, 557, 715, 1183, 1885, 783, 1476, 1751, 2019, 2325, 2435, 2304, 217, 1932, 2191, 1775, 16, 37, 1652, 451, 2192, 2795, 637, 308, 2065, 478, 2456, 1803, 1364, 1630, 1501, 914, 494, 1050, 2283, 2396, 239, 297, 616, 1160, 2038, 2378, 1980, 145, 62, 374, 316, 1000, 935, 1095, 1138, 1407, 1344, 1558, 1526, 2000, 1959, 1832, 1909, 2260, 2350, 2669, 760, 2128, 2235, 2214, 292, 2932, 1859, 274, 395, 1432, 2174, 673, 177, 957]. risks [1706, 1025, 1278, 2035, 1212, 1454, 1293, 104, 124, 428, 1074, 1384, 1255, 1731, 2082, 2417, 884, 980, 2359, 2304, 2192, 2456, 2128, 1606]. RM-ODP [1253]. road [1938, 2087, 2642, 2899]. roadmap [2210, 2047]. Roads [1982]. Roaming [1027, 2229]. rob [1315]. Robo [1632, 1834]. Robo-Jockeys [1632]. Robo-Receptionist [1834]. Robot [18, 639, 886, 1257, 1039, 2267]. Robotic [430]. Robotics [648, 2296, 1271, 649, 650]. Robotoys [64]. Robots [2045, 3, 376, 1133, 1758, 645, 646, 647, 97]. ROI [1562]. Role [1053, 687, 2463, 1096, 2220, 703, 2422, 1444, 1088, 1276, 1552, 527, 2321, 356, 2659, 2723, 71, 2395]. roles [984, 2982]. Roofline [2739]. Room [1962]. Rooms [1257]. Roots [1410, 1347, 1734, 2021, 1321]. Rose [1632]. Rosetta [2785]. RosettaNet [2316]. rough [83, 812, 2507]. roundtable [2493, 2610, 2579, 2517, 2870]. routing [1847]. Royal [1076]. Rubik [2404]. Rule [993, 1744, 1926, 2835, 1536]. rulemaking [896]. Rules [106, 1560, 1288, 1208, 2835]. run [407]. Running [1708]. runtime [1062]. Runway [1777]. rural [2643]. Russia [1810]. Rustlers [1934]. Rx [1325].

Sacked [1585]. Safe

[726, 1536, 1612, 1997, 980]. Safeguarding [2585]. safer [2283]. Safety [1215, 1655, 2454, 80, 1221, 2392, 1388]. Safety-Critical [1215, 80, 1221]. saga [2000]. Saharan [868, 1349]. Said [1076]. Saint [397]. Salaries [588, 1777]. Sale [1347, 147, 2682, 2699]. Salmon [276]. Salt [1118, 1347]. same [1850]. SAND [904]. Sandwich [219]. sanity [2930]. SAP [2947]. Sapphire [980]. Sarbanes [1610, 2125]. satisfaction [2407, 2415]. satisfiability [2873]. Save [1806, 1185, 2686]. saved [2682]. Saving [1708]. Savings [1027]. savvy [1355]. saw [995]. Say [2022, 1162, 2727, 2383]. Saying [2890]. says [2, 1382, 2810]. SCADA [2849]. scalability [197]. Scalable [2776, 899]. Scale [1602, 2639, 1146, 2125, 2558, 2227]. Scale-free [1602]. Scaling [770, 2639, 2492]. Scalpel [106]. Scan [959]. scarlet [378]. Scattering [1654]. scenario [462, 1748]. scenarios [2033, 2371]. Scene [2554]. scenes [2335]. Schelling [1848]. schemes [2528]. Schizophrenia [397]. schmethics [181]. scholarly [3017]. scholars [2189]. Scholarship [542, 183]. school [72, 1908, 1263, 2127]. Schools [1608, 591, 1949, 483, 2533, 1107]. Schoolyard [1585]. Science [1410, 1457, 1734, 1778, 1806, 785, 564, 2021, 2888, 2513, 656, 2292, 1949, 2045, 2733, 1370, 1942, 779, 1549, 1614, 427, 1577, 2475, 398, 1146, 2458, 1564, 1735, 2195, 2897, 1082, 2804, 348, 2290, 2833, 2867, 1369, 2895, 2576, 2472, 781, 1738, 380, 2925, 2718, 2199, 2734, 1150, 1856, 434, 1687, 1563, 1840, 2958, 1268, 2180, 2497, 743, 2967, 1941, 2568, 2821, 2723, 1485, 2988, 2026, 1831, 2127, 1343, 1703, 1904, 421, 2198, 3006, 2895, 2899]. Sciences [938, 1893]. Scientific [1387, 1655, 430, 1649, 402, 1617, 773, 712, 593, 999, 748, 1582]. Scientist [1257]. scientists [1534, 1952, 1838, 2467]. Scope [1962, 2049, 1495, 2851]. Scoping [1890]. score [1969]. Scram [1366]. Scratch [1806, 2972]. Screaming [1683]. Screen [739]. scripts [2703]. scrutiny [2212]. sculpture [2451, 1620]. SDI [180]. SE [107]. Sea [431]. seals [1200, 1552]. Seam [2647]. Search [793, 584, 2573, 1749, 794, 2462, 2370, 2625, 1870, 2100, 795, 2830, 2953, 1869, 1871]. searchable [332]. searchers [1047, 2858]. Searching [1789, 694, 2540, 2146, 2425, 1877]. Seat [1434]. Second [717, 866]. Secrecy [1457, 1432]. Secret [887]. secrets [1005]. Section [2178]. sector [200, 1943, 472]. Secure [16, 1762, 836, 728, 2677, 899, 1050, 2283, 1378, 2660, 305]. securely [2676]. Securing [2806, 304, 65, 690, 1929, 309, 2183]. Security [20, 1920, 989, 481, 431, 660, 1367, 1457, 1586, 1504, 2003, 2845, 64, 430, 639, 659, 988, 511, 306, 2721, 808, 1160, 889, 1004, 1166, 1350, 1461, 1259, 1756, 1587, 1890, 1335, 1089, 872, 987, 1140, 1434, 1683, 1708, 1961, 3016, 2767, 2773, 1406, 2279, 367, 1751, 2019, 662, 2219, 2400, 1520, 2410, 1992, 1784, 1815, 1363, 1262, 144, 2341, 2969, 308, 508, 1800, 2079, 1955, 2034, 2668, 1422, 2654, 1788, 2209, 2980, 1267, 1134, 2964, 2574, 494, 1923, 2373, 1988, 1271, 2805, 2623, 1138, 2100, 2128, 2849, 1922, 1042, 2869, 2174, 1979, 3021, 2866, 2794, 1663, 1766, 1019]. security [2405, 1068, 884, 2390, 1264]. See [739, 807, 1002, 51]. SEED [1443]. Seeds [999, 2629]. Seeing [618, 428]. Seek [1961]. Seeking [1258]. Seeks [2021]. segmentation [25]. Segue [2002]. Segway [2002]. seismic [210]. selection [2525]. Self [1577, 1258, 1609, 647, 2319, 1604, 2993, 2837, 1103, 1847, 2987, 2811, 2247, 1843]. self-adaptive [1847]. self-assembled [2247]. self-assembly [2993]. self-assessment [1604]. self-aware [2837]. self-efficacy [1103]. self-learning [2987]. Self-managed [1843]. self-management [2811]. Self-organization [2319]. Self-Plagiarism [1609, 1577]. Self-reconfiguring [647]. Self-Service [1258]. Selinger [2608]. sell [2073]. seller [873]. sellers [2411]. selling [39, 890]. Semantic [1465, 2785, 1760, 2074, 667, 2975, 627, 2272, 274, 1763, 1766, 1761, 2612, 1767, 1762, 2691, 1582]. Semantics [1946]. Sender [1529]. senior [2403]. Sensate [1547]. Sense [937, 2773, 3010, 2904, 1596, 2663]. sense-and-respond [1596]. Sensing [480, 2385]. Sensitive [739, 2028]. sensor [1331, 1333, 2478, 2396, 1335, 1332, 1334]. sensoring [276]. Sensors [1386, 95, 2663]. sentiments [2726]. Separate [1768]. Separation [2022, 524]. Serchs [1280]. serious [2016, 2201]. serve [1355]. Served [785]. servers [902, 2744]. serves [2804]. Service [385, 1028, 1410, 1258, 686, 1952, 1122, 563, 1642, 1013, 2764, 1926, 2184, 1383, 2780, 1950, 899, 1014, 1998, 931, 1206, 35, 1428, 1627, 85, 1845, 1846, 1010, 1945, 2953, 1943, 1930, 384, 2967, 169, 1970, 1124, 1944, 1761]. Service-Oriented [1122, 1383]. Services [813, 1007, 438, 728, 1468, 1949, 1011, 599, 1127, 509, 1312, 2981, 1942, 1123, 939, 2004, 309, 1308, 895, 1947, 2223, 880, 1953, 1009, 821, 383, 797, 2121, 2391, 2644, 1008, 1307, 2818, 1126, 1338, 2061, 2847, 1311, 1843, 2272, 1017, 2266, 1176, 2831, 1015, 2592, 1946, 700, 1012, 1941, 820, 1593, 1313, 1135, 726, 1016, 437]. Sesame [1987]. Session [1608, 194]. set [83]. SETI [849]. Sets [516, 1708, 487, 1061]. Setting [1560, 1972, 1857]. settlement [2830].

Seven [1799, 232, 1867, 2660]. sexual [609]. Shadow [2685, 588, 1409, 1686]. Shaking [2241]. shall [2683]. Shanghaid [64]. Shannon [2070]. Shape [1619, 573, 746, 2446, 2451, 524]. Shape-based [1619]. shape-changing [2451]. shaped [2954]. Shaping [2929]. sharable [418]. Share [1237, 1479, 2787, 1876]. Shared [568, 572, 2321, 1846]. Sharing [1077, 2822, 1678, 2372, 43, 260, 1043, 920, 1449, 2687, 927, 1204, 1413]. Shark [515]. Sharper [1911]. Shattering [1162]. shave [1867]. Shaw [2938]. Shazam [1970]. shift [1340, 122]. shine [1069]. Shocked [1912]. Shop [1236]. Shopper [1052]. Shopping [219, 234, 1195, 778]. Shortage [497, 1281, 807, 1856]. shortfall [469]. shot [299]. Should [286, 857, 851, 1336, 2906, 1534, 2546, 1339, 1811, 2077, 1427, 449]. Shower [496]. shrink [177]. shrink-wrapping [177]. Shrinking [938, 886]. SI [117, 115]. Sick [1478]. Side [793, 148, 2536, 1490, 1354, 2593, 2477, 827]. Side-Views [827]. Sidebar [2, 257, 147, 219, 276, 588]. Sides [693, 1562]. siege [809]. SIGAPL [2136]. SIGCOMM [2164]. SIGGRAPH [2163, 2665]. Sight [659]. Sign [88, 1734]. Sign-Off [88]. signal [2758]. Signals [147]. Signatures [300]. significant [856, 296, 1535]. signing [261]. SIGs [1508, 2363, 1551, 2097, 2109]. S’il [1118]. Silence [1214, 1887, 918]. silicon [2243]. Silver [299, 1656, 2488, 1629, 1069]. Similar [516]. similarity [2720]. Simple [542, 1705, 1931]. simpler [1705]. Simplicity [1962]. simplified [2356]. simplify [898]. simulated [747]. Simulating [515, 541, 430]. Simulation [886, 139, 1511, 648, 2476]. Simulators [220]. simultaneous [2782]. Singapore [1342]. Singing [886]. Single [1983]. Siren [1040]. Site [151, 1330, 546, 1349, 811, 869, 564, 789, 563, 591, 2064, 303, 920, 1370, 2690, 549, 223, 794, 401, 1245, 130, 767, 489, 501, 6, 941, 42, 2538, 2569, 2599, 743, 173, 458, 663, 624]. Sites [219, 1708, 2094, 2146, 730, 1498, 1208, 1199, 2985, 800, 1627, 176, 1203, 2609]. situation [474]. Situational [2918]. Six [2594, 209]. size [886, 2416, 2657, 2300, 2614, 1038, 2916]. sized [2413]. Skeptical [1835, 970]. skies [126]. Skill [807, 1201]. Skills [1052, 1325, 2667, 459, 1884, 1726, 1228, 2078, 2324]. skin [2444]. Skip [1410, 1457, 1983]. sky [276]. slander [108]. Slanty [2071]. Slight [1835, 692]. Slime [1118]. Slows [541, 1214]. sluggish [864]. slump [2864]. Small [1419, 2482, 2413, 61, 2168, 236, 392, 1111]. smaller [340]. Smalltalk [454]. ’smart [3001, 106, 453, 959, 1544, 755, 1366, 2181, 509, 1741, 2721, 412, 2633]. Smart-Its [1544]. SmartGov [2253]. Smarty [318]. SME [2588]. smell [2499]. SMEs [2063]. smoke [138]. Smoothed [2941]. snake [3019]. snake-oil [3019]. Snap [1834]. Sniffing [1632, 1478]. snipers [1185]. Snooze [1346, 1911]. SOA [2545]. SOAP [1248]. Soar [1528]. Soars [692]. Social [779, 640, 938, 666, 2278, 578, 268, 2733, 758, 1283, 2571, 881, 2582, 923, 999, 2412, 665, 668, 1204, 1801, 20]. society [439, 282, 1717, 280, 1571, 1715, 43, 711, 1718]. sociopolitical [1143]. sockets [2801]. softer [2536]. Software [150, 319, 840, 1217, 1685, 1836, 882, 2488, 2062, 1677, 2220, 2578, 148, 40, 258, 560, 454, 740, 938, 1119, 1367, 1215, 1281, 1258, 1586, 1633, 1655, 1684, 2022, 2041, 1835, 1983, 2003, 1713, 1058, 2051, 972, 866, 308, 687, 1750, 683, 2518, 401, 2807, 1803, 1923, 862, 272, 2057, 2778, 1071, 1474, 1236, 2046, 2271, 1293, 2851, 863, 2007, 1158, 278, 457, 517, 589, 1080, 919, 984, 1261, 1411, 1327, 1532, 1610, 1480, 1780, 1914, 1984, 2069, 2242, 2107, 2179, 2328, 2381, 2832, 2930, 2637, 2666, 426, 1925, 859, 1094, 304, 2920, 1182, 1043, 2875]. software [244, 1466, 2049, 2906, 1250, 1397, 2543, 2764, 1320, 2047, 1227, 2119, 1121, 1262, 1414, 1637, 1810, 2004, 2196, 2364, 994, 630, 2698, 913, 1796, 519, 1771, 61, 512, 614, 678, 2012, 39, 276, 1369, 1647, 2014, 41, 841, 694, 1535, 1736, 1656, 1963, 2156, 2221, 193, 364, 1339, 9, 3017, 1678, 2168, 1103, 2065, 2054, 1084, 1499, 1494, 2740, 382, 1446, 389, 860, 2056, 857, 1292, 2048, 2008, 1247, 363, 1605, 1049, 2166, 2055, 2318, 2529, 1747, 1109, 2872, 1679, 2914, 1975, 2883, 524, 527, 366, 861, 2657, 1976, 260, 1974, 2009, 1401, 1844, 2848, 1067]. software [1427, 447, 1470, 1698, 1232, 2971, 2537, 2629, 2797, 1275, 2078, 2324, 829, 2230, 801, 1154, 1064, 1019, 392, 858, 1995, 408, 1515, 2440, 2378, 2178]. software-based [364]. software-intensive [1369]. solicited [2106, 2421]. Solution [1835, 1608, 1629, 74]. Solutions [651, 2526, 2618, 2712, 3012, 2809, 2912, 768, 1206]. Solutions-driven [651]. solvers [2996]. Solving [1983, 2404, 1940]. Some [636, 465, 2366]. somebody [2727]. sometimes [2985, 884]. Song [1608]. songs [1040]. Sony} [1536, 1612]. sophisticated [1815]. Sorting [841]. souls [2830]. Sound [2905, 542, 1214, 2204]. Source [1237, 1281, 1258, 1479, 1983, 1423, 1094, 2543, 1999, 2119, 630, 1143, 1240, 3017, 1678, 2079, 407, 2872, 1427, 2006, 2629, 2797, 2406, 408, 1515]. Sources [1479, 2160, 116, 2526, 2618, 2712, 3012, 2809, 2912, 2074]. sourcing [1352]. South [467]. sovereignty [354]. SOX [2256]. Space [39, 2591, 555]. spaces [1243, 1374, 2450, 1693]. Spam [1567, 1141, 1529, 1609, 937, 2089, 1280, 1434, 2002, 1074, 2313, 2911, 2781, 1336, 2093, 1629]. Spamalytics [2911]. spammed [2017]. Spamming [1476]. Spare [660, 1258, 1982]. Spatial [952, 904]. spatiotemporal [905]. Speak [1140, 1044, 337]. Speaking [149, 809, 2464, 21, 456, 620, 1005, 1120, 1238, 1328, 1712, 1536, 1612, 1917, 2006, 1809, 2265, 2088, 2178, 2366, 2575, 2965, 2830, 2699, 108].

Special [2007, 1690, 542, 1592, 1462, 1941, 2027, 1325, 2476]. special-purpose [2476]. specialists [1355]. specific [671, 531, 2751]. Specification [814, 816]. specificity [1495]. spectrum [1946]. speculative [2586]. Speech [2788, 186, 737, 311, 1226, 2644, 1155, 1005, 190, 484]. Speech-activated [2788]. speech-impaired [2644]. SpeechWeb [1742]. Speeds [1162]. Spell [1002, 1647]. spell-checking [1647]. spelunking [2547]. Spending [839, 639, 2770, 2879]. spherical [1099]. spied [18]. Spills [1052]. spim [1567]. Spindle [1249]. Spinning [1329]. Spiral [683]. Spirited [693]. Spiritual [345, 589]. Spirituality [717]. spit [1567]. spoken [1288, 185, 27]. Sponge [2002]. sponsor [1836]. Spontaneous [1693]. spoofing [2016]. Sports [866]. Spot [807, 1982]. Spotting [1434]. spreading [2930, 2959]. spreadsheets [2227]. Spritz [1118]. SPUR [1531]. Spy [200]. Spyware [1734, 1662, 1673, 1668, 1661, 1666, 1665, 1667, 1669, 1671, 1663, 1670, 1664]. SQL [1701]. SSME [1952]. Stability [693, 512, 614, 678]. stabilization [1807]. stable [804, 84]. stack [1743]. staffing [474, 1228]. stairs [18]. StAIRway [1805]. stake [1238]. stakeholders [1202, 2848]. Stakes [64]. Stand [560, 515, 1027]. standard [1153]. Standardizing [1818, 43, 260]. Standards [2316, 1166, 2108, 2118, 1416, 774, 43, 260, 1917, 1979, 1566, 884]. Standish [1963]. Stanford [643]. stars [369]. Start [1862]. started [526]. Starting [1456]. Startling [1215]. State [1822, 887, 367, 2515, 2156, 2917, 1555, 1563, 1600]. Statecharts [2706]. stated [1579]. statement [1424, 1892]. Statewide [1865]. Statistical [2877, 2942, 900, 1845]. Statistics [1326, 901, 1749, 903]. Stats [430, 1076, 126]. Status [938, 2018, 2907, 611]. Staying [182, 68, 302, 499, 720, 963, 888, 1078, 1304, 1216, 1388, 1482, 1589, 1686, 1985, 1782, 1891, 2154, 2070, 2241, 2329, 2423]. Steak [1347]. steer [2152]. Steering [2987, 1769]. steganalysis [1426]. steganography [1426]. Step [147, 1123, 985]. Stepping [1366, 1873]. Steps [2837, 2184]. stereotypes [1883]. Stick [1655]. sticky [1597]. Still [839, 887, 1258, 1586, 2041, 2003, 219, 937, 2331, 200]. Stodgy [621]. stone [2444]. stones [1873]. Stop [983, 2130, 2289]. storage [230, 2493, 2517, 1786, 226, 225, 2470, 228, 2741, 1980, 227]. Store [377, 293, 2395, 1194]. stores [2500]. stories [2648]. storing [1823]. Story [693, 1492, 1097, 976, 2842]. storytelling [486, 2203]. Strategic [875, 281]. Strategies [2732, 682, 709, 1450, 440, 1879, 2588, 1599, 2008, 1774, 1641, 1091, 1596, 1071, 1876, 2594, 1070]. Strategy [1637, 468, 1902, 1676, 1121, 939, 1029, 1262, 1351, 1414, 1506, 1713, 1810, 1938, 2004, 2087, 2196, 2264, 2364, 2511, 2732, 2933, 2638, 2850, 406, 1093, 2317, 2781, 2831]. stream [2944]. streaming [2033]. streams [1284, 2945, 1496, 167]. Street [2921, 1673, 1987]. Strength [693, 588]. Stress [1962, 1277]. Strike [1753]. strive [1427]. strong [286, 1096]. Stronger [258]. Structure [1609, 1463, 529, 1897]. structured [2978]. structures [2775, 1309, 2296]. Structuring [529, 2207]. Student [1759, 1141, 739, 1827, 2086, 2533, 1978, 1263, 1517]. students [2274, 1082, 869, 2700, 803, 2537, 2701]. studies [2292]. Study [1935, 683, 1118, 1114, 425, 2875, 2316, 1521, 2724, 643]. Studying [2018]. stupidity [1323]. Style [2040, 2264, 498]. styles [2560]. Sub [1349, 868]. Sub-Saharan [1349, 868]. subject [2464]. subjects [2078]. submit [2683]. subversive [2848]. Success [2075, 40, 2003, 1308, 2064, 1097, 1799, 2873, 1976, 2778, 76, 158, 1210, 2701]. successes [2947]. Successful [1342, 976, 1088, 1622, 2315]. successfully [1855, 1355]. Sucking [496]. suit [499]. summarization [905]. summarizing [1727]. summer [2891]. sun [1151]. Super [886, 1733, 1708, 2021]. Super-size [886]. Superficial [220]. superior [1065, 2949]. Superpower [1052]. Superscaled [1259]. supplier [1250]. supply [2604, 1674, 444, 446, 441, 2227, 2208, 445, 2561, 1113, 1626, 2076, 443]. Support [566, 570, 245, 955, 2940, 2779, 195, 2113, 893, 611, 826, 1404, 1700, 2978, 2807, 2140, 31, 387, 1659, 2231, 830, 2310, 1113, 1898, 801, 1882, 1334, 114, 2409, 1853]. Supported [843, 2394, 844, 1105]. Supporting [1875, 2649, 900, 444, 903, 665, 1846, 255, 827, 1402, 1869, 1299, 441, 1765, 2117, 828]. Supreme [106, 1612]. Sure [560]. surface [3011]. surfaces [1037]. Surfing [639, 1850, 68]. Surgery [1140, 1511, 749]. surveillance [2325, 2969, 842, 2668, 3015, 1199, 1975, 2896, 957]. survey [1430]. survival [994]. Surviving [1576]. Suspension [807]. sustain [894]. Sustainability [2447, 2596]. sustainable [1090, 1774]. swan [1040]. Swapping [692]. Swarm [775]. Sweep [1861]. Swift [2676]. Swift-ly [2676]. swing [1419]. Swings [1683, 1518]. Switching [2562]. symbolism [2014]. Symbols [785]. symmetry [2387]. Synchronized [242]. synchronous [2776]. syndrome [1987]. synergy [1191]. synopses [2943]. synthesis [2204, 1882]. synthetic [2422]. System [180, 1962, 1128, 1129, 1817, 1971, 1097, 529, 309, 2607, 144, 417, 1099, 279, 1156, 2583, 113, 1924, 1928, 75, 1134, 2010, 2206, 419, 1624, 297, 2522, 1553, 2914, 530, 935, 1138, 1959, 2058, 596, 2498, 2523, 1328, 2205, 1277, 2584, 2948, 1033, 1903, 2409]. Systematic [727, 1928]. Systems [680, 1178, 40, 640, 1003, 1028, 1457, 1258, 735, 929, 926, 100, 81, 1182, 2685, 707, 2098, 2075, 2766, 1538, 80, 185, 1013, 215, 230, 2112, 13, 1772, 550, 246, 1417, 2422, 2281, 913, 1416, 3004, 2114, 971, 1369, 2113, 1723, 508, 1424, 1130, 1220, 382, 2186, 212, 1594, 1906, 2838, 1724, 844, 1107, 927, 1669, 2903, 1405, 1952, 1473, 1843, 2747, 864, 1988, 530, 1415, 1832, 1073, 2904, 2140, 1290, 2846, 910, 2293, 1222, 2772, 253, 1680, 265, 1989, 2143, 556, 874, 725, 836, 673, 255, 818, 1593, 1898, 2170, 724, 1402, 663, 2111]. systems [2594, 1882, 1454, 1852, 654, 1418, 1523, 997, 3016].

Table [902, 900]. Tableau [2583]. tabular [902]. tacit [1021]. Tackling [3]. Tag [960, 1216]. Tagging [1690]. Tags [1708, 1692]. tailoring [972]. Taiwan [1609, 1935]. Take [660, 982, 1478, 1982, 2348, 1082, 2675, 1838]. takes [2475, 1620]. Taking [560, 660, 515]. Tale [1753, 62, 1731]. Talk [639, 916, 1409]. Talking [2479, 1482]. Taming [2431, 187]. Tangible [1221, 740, 2443]. TaoBao [2658]. Tapping [217, 1996]. Target [2022, 1377, 1190]. Targeting [390, 560, 619, 1973]. Task [580, 2655]. Task-Technology [580, 2655]. Taste [376]. Tattooed [1733]. taught [449]. taxi [997]. Taxonomy [1019]. Tea [106]. Teach [1028, 1680]. Teachers [785, 200, 110]. Teaching [680, 640, 661, 2765, 869, 546, 1630, 1501, 767, 418]. team [2112, 1320, 2558, 598, 2657]. Teams [576, 457, 2364, 1339, 732, 1267, 3014, 1474, 2301, 2323]. teamwork [2746, 2412]. Tech [1709, 219, 200, 318, 639, 739, 692, 659, 959, 1683, 2040, 18, 276, 1355, 399]. tech-savvy [1355]. Technical [2212, 2773, 1322, 2565, 2524, 1024, 1406, 1094, 1500, 1525, 2064, 2688, 883, 2886, 2148, 1431, 2942, 2357, 2191, 2475, 640, 1258, 2910, 1651, 759, 2852, 913, 144, 1453, 2975, 636, 2395, 2944, 736, 1299, 36, 2947, 2753, 2172, 1557, 1705, 2743, 2583, 3010, 1750, 2874, 2775, 1524, 1931, 1908, 2807, 2678, 2377, 2921, 613, 103, 3008, 537, 1629, 1884, 923, 2553, 2818, 1605, 2627, 2585, 2416, 2721, 566, 1049, 2648, 273, 835, 999, 1958, 450, 1802, 2741, 2710, 2522, 2842, 2805, 1073, 2987, 2100, 2658, 782, 2266, 2355, 1730, 2498, 2234, 2785, 1680, 2616, 2840, 1857, 2876]. Technical [123, 1254, 2500, 2977, 2646, 82, 1979, 3021, 2708, 2258, 2614, 2324, 2555, 956, 1807, 2676, 2127, 1343, 979, 538, 2434, 1137, 2908, 2595, 2477, 863, 2323, 245, 128, 1499, 1695, 1809, 1566]. Techniques [999, 246, 1060, 1058, 842, 1131, 1691, 700]. Techno [1777]. Techno-Chic [1777]. Technological [183, 1163, 962, 2582]. technologies [1725, 284, 2447, 160, 1676, 2534, 2445, 1266, 1229, 168, 170, 27, 167, 1618, 2115, 1398, 2444, 2211, 1113, 169, 28, 2111, 1070, 2562, 1264, 171, 2124, 1853]. technologists [1591, 1526]. Technology [680, 1252, 1003, 887, 1121, 939, 1029, 1262, 1351, 1414, 1506, 1637, 1713, 1810, 1938, 2004, 2087, 2196, 2264, 2364, 2511, 2732, 2933, 2638, 575, 515, 446, 441, 580, 2432, 564, 2831, 2590, 1015, 2559, 2090, 381, 1902, 2120, 1096, 1269, 2998, 2626, 2018, 2255, 554, 2161, 1431, 322, 247, 2458, 1057, 1490, 1092, 2655, 1723, 1488, 2533, 1404, 34, 2479, 440, 3011, 1700, 249, 844, 2275, 347, 2762, 555, 182, 1289, 345, 1774, 764, 387, 1483, 2202, 2530, 1265, 2441, 262, 2896, 2871, 1706, 110, 1807, 2812, 2273, 421, 2635, 2466]. Technology-facilitated [2432]. technology-oblivious [1706]. technonationalism [1566]. technostress [1575]. techonomics [642]. Telecollaborative [574]. telecom [385, 383, 888, 1304]. telecommunication [1045]. telecommunications [381, 32, 2896]. Telecommute [1435]. Telecommuting [496, 2148, 45, 1381]. telematic [383]. telemedicine [1252, 120, 1349, 42]. telephony [185, 653]. Telescope [886, 848]. television [1651, 169]. telework [1112]. Tell [1119]. Telling [1246, 2648]. tells [1945]. templates [2846, 1129]. temporal [1877]. Ten [840]. Tending [126]. tense [2527, 2619, 3013, 2913, 2810, 2713]. tension [1393]. terabytes [1031]. TeraGrid [545]. Term [1410, 1992]. terminate [2680]. Terrible [960]. Terror [866, 2723]. Terrorism [559, 1265, 2205]. terrorist [1314]. Test [785, 1594, 598, 2077]. Testing [599, 1806, 2230, 1411, 1480, 1417, 1227, 2057, 449]. Tetris [126]. Texas [859, 887]. Text [886, 2684, 1996, 2277]. Text-Message [886]. Thank [1529]. Thanks [1754, 541, 937]. theater [597]. theatre [1039]. Theft [88, 1935, 639, 21, 20, 1890]. Their [1457, 2270, 2947, 2370, 2627, 918, 1838, 1065, 2982, 1293]. Them [1435, 2752]. theme [1157]. themes [101]. themselves [2947]. theoretic [2709]. Theoretical [2225, 2122, 2873]. Theory [870, 83, 803, 2497, 2502]. There [1162, 1961, 2047, 1102]. theses [423]. Thesis [2003]. they’re [2985]. Thick [480]. thin [286, 3011]. Thing [560, 590, 1668]. Things [693, 2354, 51, 2482]. Think [1935, 692, 1784, 2826, 2145, 1740]. thinker [2557]. Thinking [1028, 1983, 61, 512, 614, 678, 804, 934, 197, 1882, 236, 392, 657, 2796, 995, 2489, 2080, 1842]. ThinSight [3011]. Third [954, 1073, 552, 544, 2065, 1606, 1497]. Third-generation [954]. third-party [2065, 1606, 1497]. third-wave [544]. Thought [1258, 588, 404]. Thoughts [1478, 465]. Thousand [838]. thousands [62]. Threat [1326, 1504, 296, 1837, 3002]. threaten [1057, 962]. threats [2721, 2787, 1663, 1068]. Three [680, 2270, 1867, 259]. thrill [1660]. Thumbs [739]. tickling [179]. Tied [1028]. ties [2412]. Tiger [1888, 1912, 1837]. Tightening [1977]. Tightest [916]. tightrope [1210]. Time [2187, 714, 180, 2867, 559, 982, 2121, 1027, 2923, 2728, 2591, 641, 257, 2172, 364, 2333, 1740, 2771, 997, 2204, 725, 138, 2988, 2614, 2157].

Time-asymmetry [714]. Time-critical [2121]. Timeliness [1498]. Timely} [730, 1676]. times [2732, 1758, 458]. timing [2877]. Tiny [2040]. Tipped [739]. Tired [147]. titles [500]. Today [983, 1140, 2784]. Toddlers [318]. together [1105, 547, 2535]. Told [1302]. tolerance [2586, 2824]. Tomorrow [1140, 2784, 1591, 2500]. Too [1141, 590, 2524, 664, 741, 1583, 2598]. Tool [1561, 676, 1771, 1620, 1447]. Tooling [739]. Tools [1367, 1586, 668, 2676, 1450, 1445, 2186, 1794, 2652, 3008, 1510, 165, 830, 2310, 166, 829, 1853]. Tooth [785]. Top [2081, 2101, 2126, 2147, 2173, 2190, 2213, 2233, 2259, 2282, 2303, 147, 39, 87, 276, 222, 1254, 2226, 1437, 1459, 1481, 1509, 1533, 1565, 1588, 1611, 1635, 1657, 1688, 1711, 1737, 1755, 1779, 1812, 1858, 1866, 1889, 1916, 1939, 1965, 1986, 2005, 2025, 2044, 2322]. topics [1727]. topology [2550]. Tops [639]. Total [693, 1347, 612, 957]. Touch [1961, 3010, 153]. Touching [2601]. tough [1726, 2836]. touring [149]. tourism [1475]. tourist [703]. tournament [597]. Toxic [659]. Toy [1983, 1958, 2578]. Toy-Program [1983]. toys [3009]. traceability [2529]. Tracing [2318, 1321]. Track [2, 257, 147, 39, 87, 219, 18, 64, 179, 106, 126, 200, 515, 276, 299, 397, 318, 376, 453, 541, 559, 480, 496, 618, 639, 739, 717, 692, 659, 588, 430, 807, 785, 866, 838, 982, 959, 916, 937, 2896, 1052, 1027, 1002, 1076, 1162, 1118, 1140, 1257, 1302, 1346, 1409, 1280, 1386, 1236, 1214, 1325, 1434, 1366, 1456, 1503, 1733, 1608, 1753, 1585, 1632, 1683, 1708, 1528, 1478, 1560, 1654, 1961, 1861, 1911, 1887, 1982, 1777, 1805, 2002, 2021, 2040, 1834, 1934, 2067, 2084, 2104, 2131, 2153, 2177, 2194, 2216, 2237, 2262, 2285, 2306, 2361, 2380, 2399, 2419, 2437]. Tracking [39, 541, 1037]. tracks [762, 886]. Trade [887, 1504, 1005, 1977]. trade-off [1977]. trademarking [44]. trademarks [314]. Tradeoff [559]. trading [1379]. tradition [2295]. Traditional [562, 1360, 3007]. traffic [1273]. Tragedy [1862]. trails [889]. Train [1911]. training [734, 2200, 1316, 1493, 2324]. Transactional [2474, 2522, 2578, 2671, 2523]. Transactions [1126, 2500, 1929, 2501, 1880, 1207]. Transcribing [29]. transcription [30]. transfer [899, 1774, 1483, 2211]. Transformation [585, 1572, 1767, 777, 1948, 1944]. Transforming [1568, 1440]. Transistor [1654]. transition [1205]. transitioning [709]. Translating [866]. translation [2952, 2953]. TransLight [1146]. translucence [666]. Transmission [866]. transparency [1601, 2717, 351, 1587, 2378]. Transponders [717]. Transport [1147]. Trap [2022]. Travel [220, 550]. Treasure [1654]. Treatment [220, 341]. TREC [2295]. TReK [683]. trend [1925]. trends [771, 2266, 289, 2037, 2537, 2428, 2111]. Trespass [2313, 315, 1120]. tribes [1798]. tribute [2580]. tricks [833]. Tricky [219]. trillion [2244]. trio [1999]. troll [2073]. Trouble [2041, 1002, 1478, 2686, 1859]. troubled [389]. Truce [692]. True [741, 2629, 995]. Trust [425, 1956, 1888, 541, 263, 1023, 1955, 271, 1233, 644, 1290, 1519, 1608, 1934, 1285, 305, 1699, 451, 1542, 264, 2783, 3014, 270, 267, 977, 2343, 1025, 1209, 443]. Trust-building [1233]. trust-distrust [2783]. trusted [966]. Trusting [1587, 262, 1095, 1025]. trustworthiness [266, 873]. Trustworthy [1832]. Truth [1302, 1867]. truths [2373, 2798]. Trying [1077]. Ts [259]. Tube [40]. Tug [559]. Tug-of-War [559]. tumble [482]. Tuned [588]. Tunes [1961]. tuning [2816]. turf [68, 1985]. Turing [2106, 2421, 1003, 2041, 2003, 204, 2761, 979]. turn [631, 1861]. Turning [559]. turnip [193]. turnover [631]. Turns [639, 2736, 2757]. TV [807, 886, 959, 2154, 1177]. Twenty [2179]. twins [2815]. Twist [2002]. Twister [1028].

Two [2041, 2808, 2623, 1583, 1685, 1562, 2517, 1358, 62]. Two-factor [1583]. TxLinux [2523]. Type [2383, 980]. type-safe [980]. Types [1162, 457, 1599, 1125, 782]. Typing [838]. Typography [220, 151]. typology [2411].

U. [588]. U.K. [1585, 2021]. U.S [32]. U.S. [1633, 2834, 2546, 276, 397, 376, 2863, 1729, 2847, 1352, 1553, 2063, 2669, 764, 2266, 1328, 2866, 1560, 1777, 1716, 1556]. Ubi [1491]. Ubi-learning [1491]. Ubiquitous [2454, 2508, 882, 1399, 699, 1538, 1243, 2779, 1786, 880, 1548, 878, 934, 881, 876, 366, 2970, 2950]. UCITA [111]. UEML [2231]. ugly [2685, 2799]. Ultimate [346, 2840]. Ultra [225]. Ultra-high-density [225]. UML [1114, 1957, 938, 1962, 1905, 856, 207, 613, 851]. UML2 [853]. Umps [866]. uncertain [1072]. uncertainty [2242, 2286]. Uncommon [616]. unconscious [1480]. unconsciousness [1954]. Unconvincing [1326]. undergraduate [2162, 743]. undergraduate-level [743]. underlying [662]. understand [337, 1552]. Understanding [2559, 1229, 1651, 895, 681, 2653, 1066, 632, 672, 2530, 1170, 1943, 2560, 1174, 2777, 1580, 2472, 1724, 1870, 2321]. undervalued [1067]. underwriters [395]. Undone [1608]. undulating [756]. Unethical [660]. unexpected [1081]. Unfashionable [454]. unforeseen [2035]. unforgettable [2696]. unhappy [399]. unification [1793]. Unified [300, 1765, 978]. Uniform [415, 2036]. Unifying [2937]. uninspired [2026]. Unintended [2229]. Uninvited [1983, 1346]. Unions [1654]. uniqueness [1998]. Universal [371, 2167, 99, 1722, 74, 1343]. universe [2892, 2632]. Universities [1356, 1082]. university [1927, 1680, 1649, 860, 643]. Unknowns [1529, 1805]. Unlawful [542]. Unleashed [1654, 1777]. Unless [2022]. unlikely [856]. unmyths [840]. Unplugged [1608]. Unreal [597]. Unrealistic [2022]. Unsolicited [640, 1120]. Unsolvability [2041]. unsolved [2054]. Unspoken [1288]. unstructured [2277]. untrustworthiness [1095]. unveils [2863]. Updating [1614]. Upgrade [1367, 1855, 2850]. Upped [64]. upper [2650]. Ups [1503, 212]. upward [1343]. Upwardly [2602]. Urban [2385, 1934, 2267]. Usability [1457, 1494, 214, 1525, 2724, 2590, 1470, 99, 1174]. Usable [1788, 2964]. USACM [2659]. usage [900, 175, 1990, 606, 173, 608]. Use [1684, 717, 965, 904, 2750, 2622, 913, 1796, 1771, 969, 975, 966, 970, 18, 1103, 546, 1354, 2370, 1473, 1199, 204, 1170, 607, 2229, 2294, 2595, 1569]. used [955, 1905]. User [65, 1529, 331, 2592, 943, 2094, 1400, 2447, 1640, 52, 1397, 1926, 2445, 51, 1392, 192, 2299, 2446, 2443, 1445, 931, 1643, 157, 632, 156, 2788, 927, 2253, 1493, 1555, 949, 1398, 2914, 49, 366, 2448, 47, 50, 2530, 53, 1224, 1395, 165, 356, 845, 602, 1391, 48, 46, 2442, 2315, 1394]. user-aware [366]. User-Centered [1529, 1445, 1643, 156, 1555]. User-centric [2592]. Users [1457, 1655, 785, 2191, 1101, 2644, 1473, 2778, 2229]. Uses [1028, 971]. Using [1471, 215, 1749, 508, 683, 1499, 898, 1877, 2203, 2529, 547, 524, 978, 2528, 1132, 1177, 1209, 1495, 1761, 929, 525, 1520, 1362, 2141, 529, 1972, 1253, 896, 951, 2554, 2884, 2652, 595, 997, 767, 2651, 172, 2877, 253, 2247, 42, 1848, 532, 912, 1857]. usual [4]. utility [1521].

v [740, 1238, 520]. vacation [2891]. Vacuum [40]. Vagabond [785]. validate [1132]. validating [751]. Validity [454]. Valley [1409]. Value [2003, 1500, 2943, 1471, 550, 451, 2299, 2145, 1309, 812, 2719, 744, 1521, 2575, 873, 2224, 1495]. value-based [2145]. value-for-money [812]. values [1431, 2705]. vanishing [864]. variability [2052]. variations [2318]. Variations2 [1972]. Vasa} [2512].

VCommerce [1193]. VCR [1076]. vector [2978]. vectors [1658, 2720]. vendor [2620]. vendors [1359]. venture [2933]. ventures [1622]. verifiable [1991]. Verification [867, 2974, 2186, 2056, 2843]. versa [2861]. versatile [2189]. Versatility [645]. versus [1453, 2223, 2907, 1874, 2080]. very [2024]. Vested [1983]. via [2620, 87, 811, 1204, 829, 1314, 1603, 129]. viability [1299]. Vibrations [480]. vice [2861]. Vicious [2465, 3000, 2731, 2512, 2606, 2798, 2934]. Vid [2021]. Video [126, 29, 2508, 2032, 26, 1807]. videoconferencing [1956]. Vidyanidhi [423]. View [107, 1257, 2939, 505, 970, 1055, 633, 1696, 448, 2847, 272, 978, 1673]. viewer [1907]. viewers [1177]. Viewpoint [1987, 1056, 23, 1484, 183, 402, 1081, 1868, 1660, 2514, 1638, 2108, 2071, 545, 942, 1468, 2766, 2045, 1759, 2220, 2733, 1784, 22, 870, 1567, 2137, 224, 1614, 2422, 1031, 1534, 1507, 2639, 664, 1006, 519, 1057, 131, 1893, 2833, 279, 2867, 485, 1369, 2330, 2309, 261, 842, 2900, 500, 1966, 962, 1306, 459, 2935, 1192, 812, 2383, 1390, 964, 2155, 1167, 890, 45, 1738, 1811, 2404, 380, 1329, 721, 2490, 92, 1352, 744, 1439, 1263, 2734, 791, 2197, 1591, 112, 2365, 434, 3001, 2289, 2700, 565, 1239, 2441, 1714, 1144, 922, 623, 2180, 520, 2967, 111, 2797, 1485]. Viewpoint [1566, 766, 2026, 2267, 484, 592, 1211, 986, 2701, 1285, 1940, 1842, 1413, 2239]. Viewpoints [2637, 2545, 2636, 2998, 2543, 2638, 2999, 2546, 2573, 2576, 2544, 2997, 2574, 3000, 2575, 2635]. Views [916, 2778, 152, 1402, 827]. vigilance [2396]. vigilante [741]. Village [320, 400, 561, 787, 1165, 1757, 1562, 2132, 221, 90, 662, 961, 1054, 1284, 1389, 1458, 1658, 1964, 2042, 1864, 2219, 2308, 2439, 2400, 1239]. Violation [542]. violence [126]. viral [1204]. Virtual [323, 751, 2019, 1018, 2852, 576, 2845, 496, 785, 573, 2850, 1512, 569, 566, 2847, 570, 571, 2846, 2849, 581, 567, 2848, 1386, 1236, 1887, 2889, 572, 2851, 475, 779, 2727, 140, 2149, 1514, 1194, 235, 2095, 2351, 1373, 1374, 2746, 315, 3014, 1492, 745, 2991, 2856, 2957, 2301, 592, 749, 2492, 1034, 2323, 2924]. virtuality [22]. virtualization [2610, 2579, 2695]. Virus [640, 376, 1708, 1997]. virus-safe [1997]. Viruses [660, 296, 1167, 1501]. Visa [1560]. VisAD [684, 1554]. Visas [739]. visibility [1557]. Visible [740, 666]. Vision [180, 1760, 568, 1489, 2644, 2803, 1257, 1689, 1225]. vision- [2644]. vision-based [1489]. visionary [2567]. Visions [1386]. Visitation [1887]. visitor [1694]. Vista [2875]. Visual [785, 487, 3019, 903, 60, 2739]. Visualization [1450, 684, 1617, 1873, 486, 2649, 1554, 1446, 748, 1620, 2584, 1820, 1625, 1515]. Visualizing [488, 1375, 667]. Vocal [1346]. VocalSearch [1971]. Voice [1347, 191, 1382, 610]. Voices [2487, 2339]. VoiceXML [188]. VoIP [1751, 2589]. volatility [2300]. Volume [850, 227]. Volume-rendered [850]. Volumetric [1513]. Voluntary [1413]. Volunteers [916]. Vote [1163, 1586, 1057, 1419]. Voter [1586]. voters [1865]. Votes [40, 864]. Voting [1479, 1504, 1416, 1057, 376, 2108, 2304, 1417, 2546, 1424, 290, 1422, 1423, 239, 2038, 291, 1415, 872, 2660, 1454, 124, 1418]. Vous [1118]. Voyage [1346]. Voyagers [2649]. voyeurs [2649]. Vs [377, 783, 1579, 1562, 454, 640, 1163, 960, 1237, 1281, 1187, 2588, 2803, 1405, 1049, 1017, 2197, 1531, 1813, 967, 1005, 447, 2754, 1426, 1016]. vu [2336, 2221].

W [810, 2786]. Wading [1284]. Waiting [938]. Wake [2499]. Walking [87, 276]. Wall [126, 2615, 870]. Want [839, 1236, 1473]. Wanted [785, 1306, 732]. wants [200, 991]. WAP [1237]. War [559, 866, 1528, 1389, 2723, 1808]. Wardrobe [619]. warehouse [2565, 2656, 1799, 2299, 2280]. warehouses [116]. warehousing [1156, 1104, 1879, 1550]. warfare [1426]. warming [2514]. Warner [180]. Warning [127, 496, 2118, 1647]. Warnings [397]. wars [765, 1074, 1834]. Was [1387, 1662]. Washing [1028]. Washington [181, 67, 518, 399, 742, 590]. Watch [808, 1166, 1756, 1587, 889, 1004, 1350, 1461, 1259, 1890]. Watching [496, 1861]. water [136]. Watermark [1134, 1027]. Watermark-based [1134]. waters [202]. wave [552, 544]. Waves [496, 618]. Way [431, 887, 1028, 584, 1434, 2489, 2446, 110]. ways [2515, 605]. wealth [1109]. wearable [1545, 1089]. Wearables [339]. Weather [1934]. Weaving [1388, 126]. Web [2425, 126, 376, 1105, 2096, 2094, 2141, 2146, 2613, 1496, 2, 2169, 2272, 2599, 2037, 2609, 335, 1903, 1701, 2124, 506, 702, 100, 2064, 20, 787, 662, 243, 699, 1468, 96, 706, 698, 1127, 509, 295, 2677, 300, 867, 983, 1141, 1119, 1028, 1123, 1308, 2690, 975, 549, 730, 1498, 1749, 1208, 1009, 39, 200, 515, 453, 559, 639, 717, 692, 1055, 584, 1557, 1901, 2822, 1850, 794, 2472, 2612, 306, 249, 2884, 1245, 627, 1276, 713, 1008, 2825, 35, 1767, 1762, 1050, 491, 1126, 1472, 777, 188, 2061, 1729, 2847, 1199, 2985, 1017, 489, 175, 800, 796, 172, 2805, 176]. Web [1845, 1203, 2673, 2211, 2630, 2757, 2519, 1900, 845, 2180, 1671, 513, 778, 307, 173, 908, 1162, 2691, 458, 820, 1649, 3016, 1135, 1852, 1016, 2540, 2994, 1124, 1761]. Web-based [100, 1308, 1749, 306, 1276, 713, 188, 2211, 778]. Web-connected [126]. Web-enabled [509, 777]. Web-enhanced [2630]. Web-supported [1105]. Webbing [289]. Webmind [2810]. Webs [1649]. website [2917]. WEBVIEW [1701]. Weighing [718]. Welcome [1734, 1983]. We’re [1935]. western [1572, 300, 1359]. Westfall [619]. Whatever [789, 507]. wheelchair [18]. Wheels [1236]. ’Where [2457, 909, 414, 1347, 1435, 739, 2710, 2876, 792, 2683, 2092, 2160, 543, 2465, 2883, 2457, 1810]. Which [2565, 1684, 2235]. whistle [389]. White [1924]. White-hat [1924]. whitewater [2168]. Whither [2801]. whiz [2488]. Who [1431, 220, 300, 660, 1754, 1888, 1862, 1262, 630, 376, 1339, 2017, 2719, 2077, 1315, 1162, 1366, 1934, 301, 1057, 18, 2221, 1101, 459, 2682, 1811]. whom [1339, 459, 2077]. whose [2614]. Wi [991, 2229]. Wi-Fi [991, 2229]. Wide [96, 2031, 848]. wide-area [2031]. widely [728]. wider [633]. Widespread [2102]. width [1061]. WiFi [1658]. Wiki [1683, 1806, 1994]. Wikipedia [1775, 2239]. Wikipedians [2297]. Wild [1456]. wilderness [433]. Wildlife [515]. Wile [515]. Wilkes [2901]. Will [2527, 2543, 376, 613, 363, 2680, 920, 627, 1315]. win [1712, 353]. windows [2591, 2875, 1938, 1997]. winners [2829, 831]. Winning [639]. wins [2761, 2181]. winter [888]. wiping [1964]. Wired [2415]. Wireless [1987, 1389, 1458, 1237, 1734, 1862, 1331, 1826, 106, 1922, 987, 1151, 2182, 1333, 2478, 2373, 1729, 2847, 835, 2374, 1388, 1073, 1296, 2266, 1335, 1089, 95, 1692, 2592, 118, 871, 335, 1603, 2396]. Wiring [433]. Wisdom [1028, 2510, 2986]. wise}Pad [2644]. Within [1326]. without [1817, 2094, 2230]. withstand [358]. Wobble [515]. woes [2990]. woman [2634]. Women [497, 516, 2, 496, 1102, 2675, 2375, 1856, 1280, 1236, 464, 434, 2786, 2857]. wonderful [1668]. won’t [2691]. Wood [1560]. woods [2385]. Word [180, 959, 1366, 2395]. Words [2427, 2481, 1322]. Work [740, 917, 1734, 376, 559, 639, 692, 982, 843, 1140, 1585, 1327, 460, 2748, 2763, 2779, 877, 628, 87, 299, 1055, 1339, 2995, 363, 1200, 528, 2811, 473, 2491, 2301, 592, 472, 608, 461]. Worker [497, 1281, 1439]. workers [468, 1352, 2099, 1716]. workflow [194]. workflows [2811]. Working [588, 1488].

Workloads [2294]. Workplace [318, 601, 1975, 603]. Works [916]. workspace [847]. World [619, 1983, 480, 496, 1456, 2730, 2248, 2435, 1538, 1690, 2577, 2425, 2804, 1148, 1544, 222, 364, 746, 2983, 538, 2857, 96, 848]. World-Wide [848]. worlds [2351, 1492, 2492]. Worldwide [682]. Worm [561, 1237]. Worms [660, 1501]. worse [131]. Worst [2, 2625]. Worth [2041, 2002, 1995]. Worthy [220]. Would [2041, 2714, 2568]. Wound [1118]. wrapping [177]. Wrist [453]. writeprint [1878]. writing [2822]. Wrong [1053, 1528, 2834, 2257, 2794]. Wrongs [1655]. WYXIWYG [1646].

x [1689, 219, 64]. X-mas [64]. x-ray [1689, 219]. XBRL [2388]. XML [2469]. XP [1997]. xplained [1044].

Y2K [482, 215, 88, 839, 131, 18, 541, 41, 4]. Yahoo [154]. year [792, 2458]. Years [2504, 2344, 2341, 2331, 2835, 2674, 1802, 764, 209, 2371]. ’Yes [2689]. Yield [40]. young [2504, 1488]. You’re [147]. yourself [499]. You’ve [257]. Yukon [433].

Zealand [465]. Zeppelins [517, 619]. zeros [359]. zone [984]. Zyzzyva [2586].

References

Crawford:2000:EPa

[1] Diane Crawford. Editorial pointers. Communications of the ACM, 43(1): 5, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                              org/pubs/articles/journals/cacm/
                                              2000-43-1/p5-crawford/p5-crawford.
                                              pdf;   http://www.acm.org/pubs/

citations/journals/cacm/2000-43-
1/p5-crawford/.

Fox:2000:NTCa

[2] Robert Fox. News track: Celestial backbone; foreign hands off China’s net; women and the web; DOE says no nuke access; worst net nightmare; sidebar: Ethics quiz. Communications of the ACM, 43(1):9, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p9-fox/p9-fox.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2000-43-1/p9-fox/.

Crawford:2000:FTO

[3] Diane Crawford, Brian K. E. Balke, Steve Wilstrom, Jef Raskin, and Pat McGee. Forum: Tackling OO design principles; educational concerns; building LEGO robots. Communications of the ACM, 43(1):11–13, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-1/p11-crawford/p11-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2000-43-1/p11-crawford/.

Meeks:2000:EFBa

[4] Brock N. Meeks. Electronic frontier: bugging out: Y2K fallout or business as usual? Communications of the ACM, 43(1):15, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/

                                        2000-43-1/p15-meeks/p15-meeks.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2000-43-1/p15-meeks/

.

Soloway:2000:LEK

[5] Elliot Soloway, Cathleen Norris, Phyllis Blumenfeld, Barry Fishman, Joseph Krajcik, and Ronald Marx. Log on education: K–12 and the Internet. Communications of the ACM, 43(1): 19, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-1/p19-krajcik/.

Porter:2000:SCL

[6] Tom Porter and Galyn Susman. On site: creating lifelike characters in Pixar movies. Communications of the ACM, 43(1):25, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2000-43-1/p25-porter/p25-porter.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2000-43-
                                        1/p25-porter/.

Simons:2000:PBB

[7] Barbara Simons. From the President: building Big Brother. Communications of the ACM, 43(1):31, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2000-43-1/p31-simons/p31-simons.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2000-43-
                                        1/p31-simons/.

Konana:2000:IOI

[8] Prabhudev Konana, Nirup M. Menon, and Sridhar Balasubramanian. The implications of online investing. Communications of the ACM, 43(1):34, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p34-konana/p34-konana.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
1/p34-konana/.

Gorriz:2000:EGC

[9] Cecilia M. Gorriz and Claudia Medina. Engaging girls with computers through software games. Communications of the ACM, 43(1):42, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p42-gorriz/p42-gorriz.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
1/p42-gorriz/.

Hargittai:2000:RLI

[10] Eszter Hargittai. Radio’s lessons for the Internet. Communications of the ACM, 43(1):50, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p50-hargittai/p50-hargittai.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
1/p50-hargittai/.

Ackerman:2000:ROM

[11] Mark S. Ackerman and Christine A. Hadverson. Reexamining organizational memory. Communications of the ACM, 43(1):58, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-1/p58-ackerman/p58-
ackerman.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2000-43-1/p58-ackerman/.

Robinson:2000:DDP

[12] Mike Robinson, Mikko Kovalainen, and Esa Auramäki. Diary as dialogue in papermill process control. Communications of the ACM, 43(1):65, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-1/p65-robinson/p65-
robinson.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2000-43-1/p65-robinson/.

Chandra:2000:ISF

[13] Jagdish Chandra, Salvatore T. March, Satyen Mukherjee, Will Pape, R. Ramesh, H. Raghav Rao, and Ray O. Waddoups. Information systems frontiers. Communications of the ACM, 43(1): 71, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-1/p71-chandra/p71-
chandra.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-1/p71-chandra/.

Sharma:2000:CDO

[14] Srinarayan Sharma and Arun Rai. CASE deployment in IS organizations. Communications of the ACM, 43(1): 80, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p80-sharma/p80-sharma.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
1/p80-sharma/.

Gonzalez:2000:ADM

[15] Ruben Gonzalez, Greg Cranitch, and Jun Jo. Academic directions of multimedia education. Communications of the ACM, 43(1):89, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-1/p89-gonzalez/p89-
gonzalez.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2000-43-1/p89-gonzalez/.

Ellison:2000:IRRa

[16] Carl Ellison and Bruce Schneier. Inside risks: risks of PKI: Secure email. Communications of the ACM, 43(1): 160, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-1/p160-ellison/p160-ellison.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
1/p160-ellison/.

Crawford:2000:EPb

[17] Diane Crawford. Editorial pointers. Communications of the ACM, 43(2):

5, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p5-crawford/.

Fox:2000:NTMa

[18] Robert Fox. News track: More pay now; Euro net use; the cursor who spied me; robot wheelchair climbs stairs.; the boss knows; Y2K aftermath; high tech, hold the computer. Communications of the ACM, 43(2): 9, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p9-fox/.

Crawford:2000:FDD

[19] Diane Crawford, Terry D. Decker, Francisco M. De La Vega, Jeffrey Shallit, John A. Wills, Dennis J. Frailey, Brad Montgomery, Carole M. Milazzo, W. J. Liederbach, David Mould, and Ewen McNeill. Forum: Debating distance learning. Communications of the ACM, 43(2):11–15, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p11-crawford/.

Berghel:2000:ITS

[20] Hal Berghel. Identity theft, Social Security Numbers, and the Web. Communications of the ACM, 43(2):17, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p17-berghel/.

Grosso:2000:LSP

[21] Andrew Grosso. Legally speaking: the promise and problems of the No Electronic Theft Act. Communications of the ACM, 43(2):23, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p23-grosso/.

Castel:2000:VEV

[22] F. Castel. Viewpoint: exploring virtuality. Communications of the ACM, 43(2):27, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-2/p27-castel/.

Allen:2000:VNL

[23] Fran Allen, Paula Hawthorn, and Barbara Simons. Viewpoint: not now, not like this. Communications of the ACM, 43(2):29, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p29-allen/.

Maybury:2000:NDI

[24] Mark Maybury. News on demand: introduction. Communications of the ACM, 43(2):32, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p32-maybury/.

Boykin:2000:MLE

[25] Stanley Boykin and Andrew Merlino. Machine learning of event segmentation for news on demand. Communications of the ACM, 43(2):35, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p35-boykin/.

Wactlar:2000:CVA

[26] Howard D. Wactlar, Alexander G. Hauptmann, Michael G. Christel, Ricky A. Houghton, and Andreas M. Olligschlaeger. Complementary video and audio analysis for broadcast news archives. Communications of the ACM, 43(2):42, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-2/p42-wactlar/.

Kubala:2000:ITI

[27] Francis Kubala, Sean Colbath, Daben Liu, Amit Srivastava, and John Makhoul. Integrated technologies for indexing spoken language. Communications of the ACM, 43(2):48, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p48-kubala/.

Team:2000:MCM

[28] The SRI MAESTRO Team. MAE-STRO: conductor of multimedia analysis technologies. Communications of the ACM, 43(2):57, February 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p57-the_sri_maestro_
team/.

Gauvain:2000:TBN

[29] Jean-Luc Gauvain, Lori Lamel, and Gilles Adda. Transcribing broadcast news for audio and video indexing. Communications of the ACM, 43(2): 64, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p64-gauvain/.

Furui:2000:JBN

[30] Sadaoki Furui, Katsutoshi Ohtsuki, and Zhi-Peng Zhang. Japanese broadcast news transcription and information extraction. Communications of the ACM, 43(2):71, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p71-furui/.

Pallett:2000:MSR

[31] David S. Pallett, John S. Garofolo, and Jonathan G. Fiscus. Measurements in support of research accomplishments. Communications of the ACM, 43(2): 75, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p75-pallett/.

Grover:2000:PUT

[32] Varun Grover and Pradeep Vaswani. Partnerships in the U.S. telecommunications industry. Communications of the ACM, 43(2):80, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-2/p80-grover/.

Jain:2000:BI

[33] Anil Jain, Lin Hong, and Sharath Pankanti. Biometric identification. Communications of the ACM, 43(2): 90, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-2/p90-jain/.

Hirschheim:2000:MRI

[34] Rudy Hirschheim and Mary Lacity. The myths and realities of information technology insourcing. Communications of the ACM, 43(2):99, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-2/p99-hirschheim/.

Larsen:2000:CPM

[35] Kai R. T. Larsen and Peter A. Bloniarz. A cost and performance model for Web service investment. Communications of the ACM, 43(2):109, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-2/p109-larsen/.

Gillibrand:2000:TOE

[36] David Gillibrand. Technical opinion: essential business object design. Communications of the ACM, 43(2):117, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p117-gillibrand/.

Ellison:2000:IRRb

[37] Carl Ellison and Bruce Schneier. Inside risks: risks of PKI: e-commerce. Communications of the ACM, 43(2):152, February 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-2/p152-ellison/.

Crawford:2000:EPc

[38] Diane Crawford. Editorial pointers. Communications of the ACM, 43(3): 5, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p5-crawford/.

Fox:2000:NTFa

[39] Robert Fox. News track: A future space odyssey; calculating DNA; rapid tracking of diseases; hunting meteorites; “millennium” banished; English dominates Web; top selling software; the human face of E-commerce. Communications of the ACM, 43(3):9– 10, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p9-fox/.

Crawford:2000:FPS

[40] Diane Crawford, Richard H. Veith, Lauri Pirttiaho, Ralph Miller, Hal Berghel, Paul E. McKenney, Tom Moran, Antanas V. Dundzila, and Frank Cuccias. Forum: Planning for software project success; the Manhattan Project and info-czars; mastering the basics the highest priority; more net may not yield more votes; are IBM 3083s vacuum tube systems. Communications of the ACM, 43(3):11– 15, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p11-crawford/.

Glass:2000:PPY

[41] Robert L. Glass. Practical programmer: Y2K and believing in software practice. Communications of the ACM, 43(3):17–18, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p17-glass/.

Rodger:2000:SUT

[42] James A. Rodger and Parag C. Pendharkar. On site: using telemedicine in the Department of Defense. Communications of the ACM, 43(3):19– 20, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p19-rodger/.

Rada:2000:SSSa

[43] Roy Rada and John Ketchell. Sharing standards: standardizing the European information society. Communications of the ACM, 43(3):21– 25, March 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p21-rada/.

Simons:2000:PTN

[44] Barbara Simons. From the president: trademarking the Net. Communications of the ACM, 43(3):27– 28, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p27-simons/.

Khalifa:2000:VET

[45] Mohamed Khalifa and Robert Davison. Viewpoint: exploring the telecommuting paradox. Communications of the ACM, 43(3):29–31, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p29-khalifa/.

Turk:2000:PUI

[46] Matthew Turk and George Robertson. Perceptual user interfaces (introduction). Communications of the ACM, 43(3):32–34, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-3/p32-turk/.

Pentland:2000:PUI

[47] Alex Pentland. Perceptual user interfaces: perceptual intelligence. Communications of the ACM, 43(3):35– 44, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-3/p35-pentland/.

Tan:2000:PUI

[48] Hong Z. Tan. Perceptual user interfaces: haptic interfaces. Communications of the ACM, 43(3):40– 41, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-3/p40-tan/.

Oviatt:2000:PUI

[49] Sharon Oviatt and Philip Cohen. Perceptual user interfaces: multimodal interfaces that process what comes naturally. Communications of the ACM, 43(3):45–53, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-3/p45-oviatt/.

Picard:2000:PUI

[50] Rosalind W. Picard. Perceptual user interfaces: affective perception. Communications of the ACM, 43(3):50– 51, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-3/p50-picard/.

Crowley:2000:PUI

[51] James L. Crowley, Joëlle Coutaz, and François Bérard. Perceptual user interfaces: things that see. Communications of the ACM, 43(3):54, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p54-crowley/.

Bobick:2000:PUI

[52] Aaron F. Bobick, Stephen S. Intille, James W. Davis, Freedom Baird, Claudio S. Pinhanez, Lee W. Campbell, Yuri A. Ivanov, Arjan Schütte, and Andrew Wilson. Perceptual user interfaces: the KidsRoom. Communications of the ACM, 43(3):60–61, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p60-bobick/.

Reeves:2000:PUI

[53] Byron Reeves and Clifford Nass. Perceptual user interfaces: perceptual bandwidth. Communications of the ACM, 43(3):65–70, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p65-reeves/.

Lieberman:2000:PEI

[54] Henry Lieberman. Programming by example (introduction). Communications of the ACM, 43(3):72–74, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p72-lieberman/.

Smith:2000:PEN

[55] David Canfield Smith, Allen Cypher, and Larry Tesler. Programming by example: novice programming comes of age. Communications of the ACM, 43(3):75–81, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-3/p75-smith/.

Myers:2000:PEI

[56] Brad A. Myers, Richard McDaniel, and David Wolber. Programming by example: intelligence in demonstrational interfaces. Communications of the ACM, 43(3):82–89, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p82-myers/.

Repenning:2000:PEP

[57] Alexander Repenning and Corrina Perrone. Programming by example: programming by analogous examples. Communications of the ACM, 43(3):90–97, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-3/p90-repenning/.

Bauer:2000:PEP

[58] Mathias Bauer, Dietmar Dengler, Gabriele Paul, and Markus Meyer. Programming by example: programming by demonstration for information agents. Communications of the ACM, 43(3):98–103, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p98-bauer/.

Kahn:2000:PEG

[59] Ken Kahn. Programming by example: generalizing by removing detail. Communications of the ACM, 43(3):104– 106, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p104-kahn/.

StAmant:2000:PEV

[60] Robert St. Amant, Henry Lieberman, Richard Potter, and Luke Zettlemoyer. Programming by example: visual generalization in programming by example. Communications of the ACM, 43(3):107–114, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p107-st._amant/.

Fayad:2000:TOS

[61] Mohamed E. Fayad, Mauri Laitinen, and Robert P. Ward. Thinking objectively: software engineering in the small. Communications of the ACM, 43(3):115–118, March 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-3/p115-fayad/.

Neumann:2000:IRT

[62] Peter G. Neumann. Inside risks: a tale of two thousands. Communications of the ACM, 43(3):144, March 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-3/p144-neumann/.

Crawford:2000:EPd

[63] Diane Crawford. Editorial pointers. Communications of the ACM, 43(4): 5, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p5-crawford/.

Fox:2000:NTMb

[64] Robert Fox. News track: Merger marriages; call for security; gigahertz and beyond; morphing robotoys for X-mas; plasma display; domain name stakes upped; olfactory opt in; China’s email shanghaid. Communications of the ACM, 43(4):9– 10, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p9-fox/.

Crawford:2000:FSU

[65] Diane Crawford, Paul Abrahams, Bill Patterson, and Anthony Fedzano. Forum: Securing user passwords; combining OM and OG. Communications of the ACM, 43(4):11–12, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p11-crawford/.

McGuinness:2000:AF

[66] Rosemary McGuinness. ACM fellows. Communications of the ACM, 43(4):

13–14, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p13-mcguinness/.

Munro:2000:WPL

[67] Neil Munro. From Washington: a patchwork of legislation and regulation. Communications of the ACM, 43(4): 15–17, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p15-munro/.

McGinity:2000:SCS

[68] Meg McGinity. Staying connected: surfing the turf. Communications of the ACM, 43(4):19–21, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p19-mcginity/.

Kumar:2000:ERP

[69] Kuldeep Kumar and Jos van Hillegersberg. Enterprise resource planning: introduction. Communications of the ACM, 43(4):22–26, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p22-kumar/.

vanEverdingen:2000:ERP

[70] Yvonne van Everdingen, Jos van Hillegersberg, and Eric Waarts. Enterprise resource planning: ERP adoption by European midsize companies. Communications of the ACM, 43(4):

27–31, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p27-van_everdingen/

.

Willcocks:2000:ERP

[71] Leslie P. Willcocks and Richard Sykes. Enterprise resource planning: the role of the CIO and IT function in ERP. Communications of the ACM, 43(4): 32–38, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p32-willcocks/.

Becerra-Fernandez:2000:ERP

[72] Irma Becerra-Fernandez, Kenneth E. Murphy, and Steven J. Simon. Enterprise resource planning: integrating ERP in the business school curriculum. Communications of the ACM, 43(4): 39–41, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p39-becerra-fernandez/

.

Markus:2000:ERP

[73] M. Lynne Markus, Cornelis Tanis, and Paul C. van Fenema. Enterprise resource planning: multisite ERP implementations. Communications of the ACM, 43(4):42– 46, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p42-markus/.

Soh:2000:ERP

[74] Christina Soh, Sia Siew Kien, and Joanne Tay-Yap. Enterprise resource planning: cultural fits and misfits: is ERP a universal solution? Communications of the ACM, 43(4):47– 51, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p47-soh/.

Kremers:2000:ERP

[75] Mark Kremers and Han van Dissel. Enterprise resource planning: ERP system migrations. Communications of the ACM, 43(4):53–56, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p53-kremers/.

Scheer:2000:ERP

[76] August-Wilhelm Scheer and Frank Habermann. Enterprise resource planning: making ERP a success. Communications of the ACM, 43(4):57– 61, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p57-scheer/.

Sprott:2000:ERP

[77] David Sprott. Enterprise resource planning: componentizing the enterprise application packages. Communications of the ACM, 43(4):63– 69, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p63-sprott/.

Cassell:2000:ECI

[78] Justine Cassell. Embodied conversational interface agents+. Communications of the ACM, 43(4):70– 78, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p70-cassell/.

Semtner:2000:OCM

[79] Albert Semtner. Ocean and climate modeling. Communications of the ACM, 43(4):80–89, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-4/p80-semtner/.

Bowen:2000:ESC

[80] Jonathan Bowen. The ethics of safety-critical systems. Communications of the ACM, 43(4):91–97, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p91-bowen/.

Bass:2000:IDS

[81] Tim Bass. Intrusion detection systems and multisensor data fusion. Communications of the ACM, 43(4):99– 105, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-4/p99-bass/.

Singh:2000:TOD

[82] Munindar P. Singh and Mona Singh. Technical opinion: deconstructing the “any” key. Communications of the ACM, 43(4):107–108, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-4/p107-singh/.

Grzymala-Busse:2000:DMR

[83] Jerzy W. Grzymala-Busse and Wojciech Ziarko. Data mining and rough set theory. Communications of the ACM, 43(4):108–109, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p108-grzymala-busse/

.

Steimann:2000:ACH

[84] Friedrich Steimann. Abstract class hierarchies, factories, and stable designs. Communications of the ACM, 43(4):109–111, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-4/p109-steimann/.

Neumann:2000:IRD

[85] Peter G. Neumann. Inside risks: denial-of-service attacks. Communications of the ACM, 43(4):136, April 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-4/p136-neumann/.

Crawford:2000:EPe

[86] Diane Crawford. Editorial pointers. Communications of the ACM, 43(5): 5, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p5-crawford/.

Fox:2000:NTFb

[87] Robert Fox. News track: Flying the rails; logging online hours at work; top prize: embedded encryption; Digital Nose knows; walking again via chip implant; cell-phone-free class; another node in the crowd. Communications of the ACM, 43(5): 9, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p9-fox/.

Crawford:2000:FMA

[88] Diane Crawford, Nancy Mead, Thomas W. Moran, Jason Steffler, Anonymous, William Dan Terry, Merrill Warkentin, John Michael Williams, Vir V. Phoha, Vic Basili, Fred Brooks, Tom DeMarco, Ernst Denert, Koichi Kishida, Manny Lehman, and Elliot Soloway. Forum: Much ado about licensing and electronic identity theft and email-only classes and sign-off on Y2K. Communications of the ACM, 43(5):11– 15, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p11-crawford/.

Press:2000:PCP

[89] Larry Press. Personal computing: from P-books to E-books. Communications of the ACM, 43(5):17, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p17-press/.

Berghel:2000:DVP

[90] Hal Berghel. Digital village: predatory disintermediation. Communications of the ACM, 43(5):23, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p23-berghel/.

Simons:2000:PDD

[91] Barbara Simons. From the president: to DVD or not to DVD. Communications of the ACM, 43(5): 31, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p31-simons/.

Mann:2000:VEE

[92] Steve Mann. Viewpoint: existential education in the era of personal cybernetics. Communications of the ACM, 43 (5):33, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p33-mann/.

Estrin:2000:EII

[93] Deborah Estrin, Ramesh Govindan, and John Heidemann. Embedding the Internet: introduction. Communications of the ACM, 43(5):38, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p38-estrin/.

Tennenhouse:2000:EIP

[94] David Tennenhouse. Embedding the Internet: proactive computing. Communications of the ACM, 43(5): 43, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p43-tennenhouse/.

Pottie:2000:EIW

[95] G. J. Pottie and W. J. Kaiser. Embedding the Internet: wireless integrated network sensors. Communications of the ACM, 43(5):51, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p51-pottie/.

Borriello:2000:EIE

[96] Gaetano Borriello and Roy Want. Embedding the Internet: embedded computation meets the World Wide Web. Communications of the ACM, 43(5): 59, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p59-borriello/.

Sukhatme:2000:EIE

[97] Gaurav S. Sukhatme and Maja J. Mataric. Embedding the Internet: embedding robots into the Internet. Communications of the ACM, 43(5): 67, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p67-sukhatme/.

Abelson:2000:EIA

[98] Harold Abelson, Don Allen, Daniel Coore, Chris Hanson, George Homsy, Thomas F. Knight, Radhika Nagpal, Erik Rauch, Gerald Jay Sussman, and Ron Weiss. Embedding the Internet: amorphous computing. Communications of the ACM, 43(5): 74, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p74-abelson/.

Shneiderman:2000:UU

[99] Ben Shneiderman. Universal usability. Communications of the ACM, 43 (5):84, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p84-shneiderman/.

Baron:2000:WBC

[100] John P. Baron, Michael J. Shaw, and Andrew D. Bailey. Web-based Ecatalog systems in B2B procurement. Communications of the ACM, 43(5): 93, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-5/p93-baron/.

Cline:2000:EBT

[101] Marshall Cline and Mike Girou. Enduring business themes. Communications of the ACM, 43(5):101, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-5/p101-cline/.

Williams:2000:ARN

[102] Laurie A. Williams and Robert R. Kessler. All I really need to know about pair programming I learned in kindergarten. Communications of the ACM, 43(5):108, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-5/p108-williams/.

Landau:2000:TOD

[103] Susan Landau. Technical opinion: designing cryptography for the new century. Communications of the ACM, 43 (5):115, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-5/p115-landau/.

Weinstein:2000:IRI

[104] Lauren Weinstein and Peter G. Neumann. Inside risks: Internet risks. Communications of the ACM, 43(5): 144, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-5/p144-weinstein/.

Crawford:2000:EPf

[105] Diane Crawford. Editorial pointers. Communications of the ACM, 43(6): 5, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p5-crawford/.

Fox:2000:NTS

[106] Robert Fox. News track: The Supreme Court goes online; coffee? tea? email?; political coffers; smart scalpel; office cube-icles; law preview; piracy rages; wireless rules; the name game. Communications of the ACM, 43(6): 9–10, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p9-fox/.

Crawford:2000:FPD

[107] Diane Crawford, Morton Grosser, Don Walter, Steve Grief, Bernard Abramson, Robert LeMay, and Rob Du-Wors. Forum: A plea for dumb computers; gloom-and-doom predictions; expanding the SE view. Communications of the ACM, 43(6):11– 14, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p11-crawford/.

Townsend:2000:LSL

[108] Anthony M. Townsend, Robert J. Aalberts, and Steven A. Gibson. Legally speaking: libel and slander on the Internet. Communications of the ACM, 43(6):15–17, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p15-townsend/.

Meeks:2000:EFBb

[109] Brock N. Meeks. Electronic frontier: Bad moon rising. Communications of the ACM, 43(6):19–21, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p19-meeks/.

Soloway:2000:LETa

[110] Elliot Soloway, Henry J. Becker, Cathleen Norris, and Neal Topp. Log on education: teachers and technology: easing the way. Communications of the ACM, 43(6):23– 26, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p23-soloway/.

Stallman:2000:VWW

[111] Richard Stallman. Viewpoint: Why we must fight UCITA. Communications of the ACM, 43(6):27–28, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p27-stallman/.

OReilly:2000:VIP

[112] Tim O’Reilly. Viewpoint: The Internet patent land grab. Communications of the ACM, 43(6):29– 31, June 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p29-o_reilly/.

Hasselbring:2000:ISI

[113] Wilhelm Hasselbring. Information system integration. Communications of the ACM, 43(6):32–38, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p32-hasselbring/.

Yang:2000:ISE

[114] Jian Yang and Mike P. Papazoglou. Interoperation support for electronic business. Communications of the ACM, 43(6):39–47, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p39-yang/.

Grimson:2000:SCH

[115] Jane Grimson, William Grimson, and Wilhelm Hasselbring. The SI challenge in health care. Communications of the ACM, 43(6):48–55, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p48-grimson/.

Rundensteiner:2000:MDW

[116] Elke A. Rundensteiner, Andreas Koeller, and Xin Zhang. Maintaining data warehouses over changing information sources. Communications of the ACM, 43(6):57–62, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p57-rundensteiner/

.

Adam:2000:SDL

[117] Nabil R. Adam, Vijayalakshmi Atluri, and Igg Adiwijaya. SI in digital libraries. Communications of the ACM, 43(6):64–72, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-6/p64-adam/.

Varshney:2000:EMW

[118] Upkar Varshney and Ron Vetter. Emerging mobile and wireless networks. Communications of the ACM, 43(6):73–81, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-6/p73-varshney/.

Phelps:2000:MD

[119] Thomas A. Phelps and Robert Wilensky. Multivalent documents. Communications of the ACM, 43(6):82– 90, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-6/p82-phelps/.

Huston:2000:TPR

[120] Terry L. Huston and Janis L. Huston. Is telemedicine a practical reality? Communications of the ACM, 43(6):

91–95, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p91-huston/.

Watson:2000:IIM

[121] Richard T. Watson, George M. Zinkhan, and Leyland F. Pitt. Integrated Internet marketing. Communications of the ACM, 43(6):97– 102, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p97-watson/.

Parr:2000:PSD

[122] Gerard Parr and Kevin Curran. A paradigm shift in the distribution of multimedia. Communications of the ACM, 43(6):103–109, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p103-parr/.

Schoder:2000:TOR

[123] Detlef Schoder and Torsten Eymann. Technical opinion: The real challenges of mobile agents. Communications of the ACM, 43(6):111– 112, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p111-schoder/.

Weinstein:2000:IRR

[124] Lauren Weinstein. Inside risks: Risks of Internet voting. Communications of the ACM, 43(6):128, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-6/p128-weinstein/.

Crawford:2000:EPg

[125] Diane Crawford. Editorial pointers. Communications of the ACM, 43(7): 5, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p5-crawford/.

Fox:2000:NTTa

[126] Robert Fox. News track: E-tending the farm; latest agricultural stats; Myanmar cuts net; video games, violence linked; flying the web-connected skies; intricate politics of weaving; girls and computers; wall of Tetris. Communications of the ACM, 43(7): 9–10, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p9-fox/.

Crawford:2000:FGW

[127] Diane Crawford, Mark Wallace, Albert Semter, Alireza Ebrahimi, Pat McGee, Rommert J. Casimir, and Frank Land. Forum: Global warning in ocean and climate modeling; add PBE to conventional languages; abbreviation hijacking; more to ERP than Org issues. Communications of the ACM, 43(7): 11–14, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p11-crawford/.

Glass:2000:PPP

[128] Robert L. Glass. Practical programmer: On personal technical obsolescence. Communications of the ACM, 43(7):15–17, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p15-glass/.

Ein-Dor:2000:IPM

[129] Phillip Ein-Dor, Seymour E. Goodman, and Peter Wolcott. International perspectives: From Via Maris to electronic highway: the Internet in Canaan. Communications of the ACM, 43(7):19– 23, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p19-ein-dor/.

Kornfeld:2000:SDI

[130] Dori Kornfeld. On site: Deliberating on ICANN. Communications of the ACM, 43(7):25–26, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p25-kornfeld/.

Earnest:2000:VEW

[131] Les Earnest. Viewpoint: E2A is worse than Y2K. Communications of the ACM, 43(7):27–28, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p27-earnest/.

Rosenbloom:2000:I

[132] Andrew Rosenbloom. Introduction. Communications of the ACM, 43(7): 30–32, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p30-rosenbloom/.

Hecker:2000:PCG

[133] Chris Hecker. Physics in computer games. Communications of the ACM, 43(7):34–39, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-7/p34-hecker/.

Funge:2000:CMG

[134] John Funge. Cognitive modeling for games and animation. Communications of the ACM, 43(7):40– 48, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p40-funge/.

Popovic:2000:CPR

[135] Zoran Popović. Controlling physics in realistic character animation. Communications of the ACM, 43(7):50– 58, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p50-popovic/.

Foster:2000:MWC

[136] Nick Foster and Dimitris Metaxas. Modeling water for computer animation. Communications of the ACM, 43(7):60–67, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-7/p60-foster/.

OBrien:2000:AF

[137] James F. O’Brien and Jessica K. Hodgins. Animating fracture. Communications of the ACM, 43(7):68– 75, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p68-o_brien/.

Stam:2000:ISF

[138] Jos Stam. Interacting with smoke and fire in real time. Communications of the ACM, 43(7):76–83, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p76-stam/.

Prusinkiewicz:2000:SMP

[139] Przemyslaw Prusinkiewicz. Simulation modeling of plants and plant ecosystems. Communications of the ACM, 43(7):84–93, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p84-prusinkiewicz/

.

Frohlich:2000:VO

[140] Torsten Fröhlich. The virtual oceanarium. Communications of the ACM, 43(7):94–101, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p94-frohlich/.

Segen:2000:LMN

[141] Jakub Segen and Senthil Kumar. Look ma, no mouse! Communications of the ACM, 43(7):102– 109, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p102-segen/.

Smith:2000:PYM

[142] Deborah K. Smith, David B. Paradice, and Steven M. Smith. Prepare your mind for creativity. Communications of the ACM, 43(7):110– 116, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p110-smith/.

Storey:2000:CIC

[143] Veda C. Storey, Detmar W. Straub, Kathy A. Stewart, and Richard J. Welke. A conceptual investigation of the e-commerce industry. Communications of the ACM, 43(7):117– 123, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-7/p117-storey/.

Dhillon:2000:TOI

[144] Gurpreet Dhillon and James Backhouse. Technical opinion: Information system security management in the new millennium. Communications of the ACM, 43(7):125– 128, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p125-dhillon/.

Neumann:2000:IRR

[145] Peter G. Neumann. Inside risks: risks in retrospect. Communications of the ACM, 43(7):144, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-7/p144-neumann/.

Crawford:2000:EPh

[146] Diane Crawford. Editorial pointers. Communications of the ACM, 43(8): 5, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p5-crawford/.

Fox:2000:NTE

[147] Robert Fox. News track: Everybody must get cloned; A step beyond cookies; the chicken and the chip; air communication breakdown; no sale; intestinal pillage; sidebar: Top 10 signals you’re tired of the Internet business. Communications of the ACM, 43(8):9– 10, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p9-fox/.

Crawford:2000:FOS

[148] Diane Crawford, Robert Levine, John A. Wills, and James L. Rash.

Forum: The other side of embedding the Internet; A cure for lost programming companionship;l licensing software developers. Communications of the ACM, 43(8):11–14, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p11-crawford/.

Grosso:2000:LSE

[149] Andrew Grosso. Legally speaking: the economic espionage act: touring the minefields. Communications of the ACM, 43(8):15–18, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p15-grosso/.

Armour:2000:BSC

[150] Phillip G. Armour. The business of software: the case for a new business model. Communications of the ACM, 43(8):19–22, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p19-armour/.

Berleant:2000:SDT

[151] Daniel Berleant. On site: does typography affect proposal assessment? Communications of the ACM, 43(8):24– 25, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p24-berleant/.

Riecken:2000:IPV

[152] Doug Riecken. Introduction: personalized views of personalization. Communications of the ACM, 43(8):26– 28, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p26-riecken/.

Wells:2000:BPF

[153] Nigel Wells and Jeff Wolfers. The business of personalization: finance with a personalized touch. Communications of the ACM, 43(8):30–34, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p30-wells/.

Manber:2000:BPE

[154] Udi Manber, Ash Patel, and John Robison. The business of personalization: experience with personalization of Yahoo! Communications of the ACM, 43(8):35–39, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p35-manber/.

Riecken:2000:BPP

[155] Doug Riecken. The business of personalization: personalized communication networks. Communications of the ACM, 43(8):41–42, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p41-riecken/.

Kramer:2000:HEU

[156] Joseph Kramer, Sunil Noronha, and John Vergo. The human element: a user-centered design approach to personalization. Communications of the ACM, 43(8):44–48, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p44-kramer/.

Karat:2000:HEA

[157] John Karat, Clare-Marie Karat, and Jacob Ukelson. The human element: affordances, motivation, and the design of user interfaces. Communications of the ACM, 43(8):49– 51, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p49-karat/.

Schonberg:2000:HEM

[158] Edith Schonberg, Thomas Cofino, Robert Hoch, Mark Podlaseck, and Susan L. Spraragen. The human element: measuring success. Communications of the ACM, 43(8):53–57, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p53-schonberg/.

Belkin:2000:HEH

[159] Nicolas J. Belkin. The human element: helping people find what they don’t know. Communications of the ACM, 43(8):58–61, August 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p58-belkin/.

Candy:2000:HEC

[160] Linda Candy and Ernest Edmonds. The human element: creativity enhancement with emerging technologies. Communications of the ACM, 43(8):62–65, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-8/p62-candy/.

Minsky:2000:DIC

[161] Marvin Minsky. Deep issues: commonsense-based interfaces. Communications of the ACM, 43(8):66– 73, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p66-minsky/.

McCarthy:2000:DIP

[162] John McCarthy. Deep issues: phenomenal data mining. Communications of the ACM, 43(8):75–79, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p75-mccarthy/.

Pednault:2000:DIR

[163] Edwin P. D. Pednault. Deep issues: representation is everything. Communications of the ACM, 43(8):80– 83, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p80-pednault/.

Volokh:2000:DIP

[164] Eugene Volokh. Deep issues: personalization and privacy. Communications of the ACM, 43(8):84–88, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p84-volokh/.

Riecken:2000:DIP

[165] Doug Riecken. Deep issues: personal end-user tools. Communications of the ACM, 43(8):89–91, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-8/p89-riecken/.

Smith:2000:DIB

[166] David Canfield Smith. Deep issues: building personal tools by programming. Communications of the ACM, 43(8):92–95, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-8/p92-canfield_smith/.

Maglio:2000:ETI

[167] Paul Maglio and Rob Barrett. Enabling technologies: intermediaries personalize information streams. Communications of the ACM, 43(8):96– 101, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p96-maglio/.

Hirsh:2000:ETL

[168] Haym Hirsh, Chumki Basu, and Brian D. Davison. Enabling technologies: learning to personalize. Communications of the ACM, 43(8):102– 106, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p102-hirsh/.

Smyth:2000:ETP

[169] Barry Smyth and Paul Cotter. Enabling technologies: a personalized television listings service. Communications of the ACM, 43(8):107–111, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p107-smyth/.

Kantor:2000:ETC

[170] Paul B. Kantor, Endre Boros, Benjamin Melamed, Vladimir Meñkov, Bracha Shapira, and David J. Neu. Enabling technologies: capturing human intelligence in the Net. Communications of the ACM, 43(8):112– 115, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p112-kantor/.

Zadrozny:2000:ETN

[171] Wlodek Zadrozny, M. Budzikowska, J. Chai, N. Kambhatla, S. Levesque, and N. Nicolov. Enabling technologies: natural language dialogue for personalized interaction. Communications of the ACM, 43(8):116–120, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p116-zadrozny/.

Mulvenna:2000:PNU

[172] Maurice D. Mulvenna, Sarabjot S. Anand, and Alex G. Büchner. Personalization on the Net using Web mining: introduction. Communications of the ACM, 43(8):122–125, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p122-mulvenna/.

Spiliopoulou:2000:WUM

[173] Myra Spiliopoulou. Web usage mining for Web site evaluation. Communications of the ACM, 43(8):127– 134, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p127-spiliopoulou/

.

Cingil:2000:BAP

[174] Ibrahim Cingil, Asuman Dogac, and Ayca Azgin. A broader approach to personalization. Communications of the ACM, 43(8):136–141, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p136-cingil/.

Mobasher:2000:APB

[175] Bamshad Mobasher, Robert Cooley, and Jaideep Srivastava. Automatic personalization based on Web usage mining. Communications of the ACM, 43(8):142–151, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p142-mobasher/.

Perkowitz:2000:AWS

[176] Mike Perkowitz and Oren Etzioni. Adaptive Web sites. Communications of the ACM, 43(8):152–158, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p152-perkowitz/.

Simons:2000:IRS

[177] Barbara Simons. Inside risks: shrinkwrapping our rights. Communications of the ACM, 43(8):168, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-8/p168-simons/.

Crawford:2000:EPi

[178] Diane Crawford. Editorial pointers. Communications of the ACM, 43(9):5, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-9/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
9/p5-crawford/.

Fox:2000:NTP

[179] Robert Fox. News track: Prioritizing privacy; end of dot-com days; online ivory tickling; privacy fortress; asteroids by the numbers. Communications of the ACM, 43(9):9–10, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p9-fox/.

Crawford:2000:FAT

[180] Diane Crawford, John Letaw, Mark Wallace, Les Earnest, and Peter Neely. Forum: Is AOL–Time Warner a closed system? repeating SDI myths; better word processors for the vision impaired. Communications of the ACM, 43(9):11–12, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p11-crawford/.

Munro:2000:WES

[181] Neil Munro. From Washington: Ethics, schmethics. Communications of the ACM, 43(9):13–15, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p13-munro/.

McGinity:2000:SCB

[182] Meg McGinity. Staying connected: Body of technology. Communications of the ACM, 43(9):17–19, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p17-mcginity/.

Appel:2000:VTA

[183] Andrew W. Appel and Edward W. Felten. Viewpoint: Technological access control interferes with noninfringing scholarship. Communications of the ACM, 43(9):21–23, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p21-appel/.

Lai:2000:CI

[184] Jennifer Lai. Conversational interfaces. Communications of the ACM, 43(9):24–27, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-9/p24-lai/.

Boyce:2000:NSD

[185] Susan J. Boyce. Natural spoken dialogue systems for telephony applications. Communications of the ACM, 43(9):29–34, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p29-boyce/.

Nass:2000:SIE

[186] Clifford Nass and Li Gong. Speech interfaces from an evolutionary perspective. Communications of the ACM, 43 (9):36–43, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p36-nass/.

Oviatt:2000:TRE

[187] Sharon Oviatt. Taming recognition errors with a multimodal interface. Communications of the ACM, 43(9):45–51, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p45-oviatt/.

Lucas:2000:VWB

[188] Bruce Lucas. VoiceXML for Webbased distributed conversational applications. Communications of the ACM, 43(9):53–57, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p53-lucas/.

Lucente:2000:CIC

[189] Mark Lucente. Conversational interfaces for e-commerce applications. Communications of the ACM, 43(9): 59–61, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p59-lucente/.

Shneiderman:2000:LSR

[190] Ben Shneiderman. The limits of speech recognition. Communications of the ACM, 43(9):63–65, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p63-shneiderman/.

Markowitz:2000:VB

[191] Judith A. Markowitz. Voice biometrics. Communications of the ACM, 43 (9):66–73, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-9/p66-markowitz/.

Fishkin:2000:EUI

[192] Kenneth P. Fishkin, Anuj Gujar, Beverly L. Harrison, Thomas P. Moran, and Roy Want. Embodied user interfaces for really direct manipulation. Communications of the ACM, 43(9):74–80, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-9/p74-fishkin/p74-
fishkin.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-9/p74-fishkin/.

Gopal:2000:GSP

[193] Ram D. Gopal and G. Lawrence Sanders. Global software piracy: you can’t get blood out of a turnip. Communications of the ACM, 43(9):82–89, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p82-gopal/.

Ferraiolo:2000:IRW

[194] David Ferraiolo. Integrating RBAC into workflow and collaborative environments (panel session). Communications of the ACM, 43(9):89–90, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p89-ferraiolo/.

Davison:2000:GPS

[195] Robert M. Davison and Robert O. Briggs. GSS for presentation support. Communications of the ACM, 43 (9):91–97, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p91-davison/.

Stylianou:2000:IFQ

[196] Antonis C. Stylianou and Ram L. Kumar. An integrative framework for IS quality management. Communications of the ACM, 43(9):99–104, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p99-stylianou/.

Laitinen:2000:TOP

[197] Mauri Laitinen, Mohamed E. Fayad, and Robert P. Ward. Thinking objectively: The problem with scalability. Communications of the ACM, 43(9):105–107, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-9/p105-laitinen/.

Neumann:2000:IRM

[198] Peter G. Neumann. Inside risks: Missile defense. Communications of the ACM, 43(9):128, September 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-9/p128-neumann/.

Crawford:2000:EPj

[199] Diane Crawford. Editorial pointers. Communications of the ACM, 43(10): 5, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p5-crawford/.

Fox:2000:NTTb

[200] Robert Fox. News track: Tech sector hot for teachers; overhaul still lacking; spy cooperative; E-casinos hit jackpot; driving out counterfeiters; Polly wants a Web link. Communications of the ACM, 43(10):9–10, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p9-fox/.

Crawford:2000:FWD

[201] Diane Crawford, Michael Gardner, Andrew Bender, Frank Linton, Pete Patton, David Parnas, Fran Allen, Paula Hawthorn, and Barbara Simons. Forum: what defines a programming relic? Communications of the ACM, 43(10):11–12, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p11-crawford/.

Meeks:2000:BBB

[202] Brock N. Meeks. Being blank in bright waters. Communications of the ACM, 43(10):13–16, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-10/p13-meeks/.

Armour:2000:FOI

[203] Phillip G. Armour. The five orders of ignorance. Communications of the ACM, 43(10):17–20, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-10/p17-armour/.

Ronald:2000:WUT

[204] Edmund M. A. Ronald and Moshe Sipper. What use is a Turing chatterbox? continuing the debate on machine intelligence. Communications of the ACM, 43(10):21–23, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-10/p21-ronald/.

Larsen:2000:CEF

[205] Grant Larsen. Component-based enterprise frameworks. Communications of the ACM, 43(10):24–26, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-10/p24-larsen/.

Hopkins:2000:CP

[206] Jon Hopkins. Component primer. Communications of the ACM, 43 (10):27–30, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-10/p27-hopkins/.

Kobryn:2000:MCF

[207] Cris Kobryn. Modeling components and frameworks with UML. Communications of the ACM, 43(10):31–38, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p31-kobryn/.

Fayad:2000:EFC

[208] Mohamed E. Fayad, David S. Hamu, and Davide Brugali. Enterprise frameworks characteristics, criteria, and challanges. Communications of the ACM, 43(10):39–46, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p39-fayad/.

Sparling:2000:LLT

[209] Michael Sparling. Lessons learned: through six years of componentbased development. Communications of the ACM, 43(10):47–53, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p47-sparling/.

Jian:2000:HFP

[210] Lu Jian, Li Yingjun, Ma Xiaoxing, Cai Min, Tao Xianping, Zhang Guanqun, and Liu Jianzhong. A hierarchical framework: for parallel seismic applications. Communications of the ACM, 43(10):55–59, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p55-jian/.

Fingar:2000:CFE

[211] Peter Fingar. Component-based frameworks for e-commerce. Communications of the ACM, 43(10):61–67, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p61-fingar/.

Johnson:2000:UDO

[212] Richard A. Johnson. The ups and downs of object-oriented systems development. Communications of the ACM, 43(10):68–73, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p68-johnson/.

Bhattacharjee:2000:ECE

[213] Sudip Bhattacharjee and R. Ramesh. Enterprise computing environments and cost assessment. Communications of the ACM, 43(10):74–82, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p74-bhattacharjee/

.

Agarwal:2000:UR

[214] Ritu Agarwal, Prabuddha De, Atish P. Sinha, and Mohan Tanniru. On the usability of OO representations. Communications of the ACM, 43(10):83– 89, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p83-agarwal/.

Brown:2000:ULY

[215] Garland Brown, Marshall Fisher, Ned Stoll, Dave Beeksma, Mark Black, Ron Taylor, Choe Seok Yon, Aaron J. Williams, William Bryant, and Bernard J. Jansen. Using the lessons of Y2K to improve information systems architecture. Communications of the ACM, 43(10):90–97, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p90-brown/.

Maamar:2000:TIB

[216] Zakaria Maamar and Jeff Sutherland. Toward intelligent business objects. Communications of the ACM, 43(10):99–102, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-10/p99-maamar/.

Blaze:2000:IRT

[217] Matt Blaze and Steven M. Bellovin. Inside risks: Tapping on my network door. Communications of the ACM, 43(10):136, October 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-10/p136-blaze/.

Crawford:2000:EPk

[218] Diane Crawford. Editorial pointers. Communications of the ACM, 43(11): 5, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p5-crawford/.

Fox:2000:NTFc

[219] Robert Fox. News track: Fed sites fail on privacy; memory-chip sandwich; better X-ray tech; paper still king; tricky restoration; going coffin shopping; sidebar: Engineering grads rankings. Communications of the ACM, 43(11):9–10, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2000-43-11/p9-fox/p9-fox.pdf;
http://www.acm.org/pubs/citations/
journals/cacm/2000-43-11/p9-fox/

.

Crawford:2000:FEP

[220] Diane Crawford, John Bashinski, Eugene Volokh, Brian Balke, John Michael Williams, Neville Holmes, Daniel Berleant, John J. Bosley, Jeffrey Rosenwald, Christopher G. Prince, Chris Hecker, and Terry Denton. Forum: The ethics of personalization; the economic espionage act: A lawyer’s perspective; proposal typography deserves fuller treatment; superficial ethical rant; who is worthy; A point about flight simulators; predatory disintermediation in the travel industry. Communications of the ACM, 43(11):11–15, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p11-crawford/.

Berghel:2000:DVD

[221] Hal Berghel. Digital village: digital politics 2000. Communications of the ACM, 43(11):17–22, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p17-berghel/.

Goodman:2000:IPC

[222] Seymour Goodman, Timothy Kelly, Michael Minges, and Larry Press. International perspectives: computing at the top of the world. Communications of the ACM, 43(11):23–27, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p23-goodman/.

Fouts:2000:SQB

[223] Jason W. Fouts. On site: an “out-of-box” experience. Communications of the ACM, 43(11):28–29, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p28-fouts/.

Clift:2000:VID

[224] Steven Clift. Viewpoint: an Internet of democracy. Communications of the ACM, 43(11):31–32, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-11/p31-clift/.

Hesselink:2000:UDS

[225] Lambertus Hesselink. Ultra-high-density data storage: introduction. Communications of the ACM, 43(11): 33–34, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2000-43-11/p33-hesselink/.

Gibson:2000:NAS

[226] Garth A. Gibson and Rodney Van Meter. Network attached storage architecture. Communications of the ACM, 43(11):37–45, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-11/p37-gibson/.

Orlov:2000:VHD

[227] Sergei S. Orlov. Volume holographic data storage. Communications of the ACM, 43(11):46–54, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-11/p46-orlov/.

McDaniel:2000:MDS

[228] Terry McDaniel. Magneto-optical data storage. Communications of the ACM, 43(11):56–63, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p56-mcdaniel/.

vanHouten:2000:PCR

[229] Henk van Houten and Wouter Leibbrandt. Phase change recording. Communications of the ACM, 43(11):64–71, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p64-van_houten/.

Carley:2000:MIM

[230] L. Richard Carley, Gregory R. Ganger, and David F. Nagle. MEMS-based integrated-circuit mass-storage systems. Communications of the ACM, 43 (11):72–80, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-11/p72-carley/.

Fan:2000:IFF

[231] Ming Fan, Jan Stallaert, and Andrew B. Whinston. The Internet and the future of financial markets. Communications of the ACM, 43(11):82–88, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p82-fan/.

Ginzboorg:2000:SCC

[232] Philip Ginzboorg. Seven comments on charging and billing. Communications of the ACM, 43(11):89–92, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p89-ginzboorg/.

Liu:2000:PRM

[233] Ziming Liu and David G. Stork. Is paperless really more? Communications of the ACM, 43(11):94–97, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p94-liu/.

Bhatnagar:2000:RCI

[234] Amit Bhatnagar, Sanjog Misra, and H. Raghav Rao. On risk, convenience, and Internet shopping behavior. Communications of the ACM, 43(11):98–105, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p98-bhatnagar/.

Kock:2000:BVO

[235] Ned Kock. Benefits for virtual organizations from distributed groups. Communications of the ACM, 43(11):107–112, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-11/p107-kock/.

Ward:2000:TOM

[236] Robert P. Ward, Mauri Laitinen, and Mohamed E. Fayad. Thinking objectively: management in the small. Communications of the ACM, 43(11):113–116, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p113-ward/.

Anonymous:2000:COa

[237] Anonymous. Career opportunities. Communications of the ACM, 43(11): 117–168, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:2000:COb

[238] Anonymous. Calendar. Communications of the ACM, 43(11):169–174, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2000:IRV

[239] Rebecca Mercuri. Inside risks: voting automation (early and often?). Communications of the ACM, 43(11):176, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p176-mercuri/.

Crawford:2000:Ea

[240] Diane Crawford. Editorial. Communications of the ACM, 43(11es), November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11/p5-crawford/.

Agarwal:2000:RRA

[241] Ritu Agarwal, Jayesh Prasad, Mohan Tanniru, and John Lynch. Risks of rapid application development. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-11es/a1-agarwal/a1-
agarwal.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-11es/a1-agarwal/. Article No. 1.

Ali:2000:SDM

[242] Zafar Ali and Arif Ghafoor. Synchronized delivery of multimedia information over ATM networks. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a2-ali/. Article

No. 2.

Bertino:2000:PIW

[243] Elisa Bertino, Elena Pagani, Gian Paolo Rossi, and Pierangela Samarati. Protecting information on the Web. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a3-bertino/. Ar-

ticle No. 3.

Blackburn:2000:CSD

[244] Joseph Blackburn, Gary Scudder, and Luk N. Van Wassenhove. Concurrent software development. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-11es/a4-blackburn/. Article No.

4.

Bodker:2000:CTS

[245] Susanne Bødker. Coordinating technical support platforms. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-11es/a5-bodker/. Article No. 5.

Choy:2000:DRT

[246] Manhoi Choy, Hong Va Leong, and Man Hon Wong. Disaster recovery techniques for database systems. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a6-choy/. Arti-

cle No. 6.

Dahl:2000:CNL

[247] Deborah A. Dahl, Lewis M. Norton, and K. W. Scholz. Commercialization of natural language processing technology. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/

                                        cacm/2000-43-11es/a7-dahl/. Arti-

cle No. 7.

Douglas:2000:OOC

[248] David E. Douglas and Bill C. Hardgrave. Object-oriented curricula in academic programs. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2000-
                                        43-11es/a8-douglas/. Article No. 8.

Kambil:2000:ICA

[249] Ajit Kambil, Arnold Kamis, Marios Koufaris, and Henry C. Lucas. Influences on the corporate adoption of Web technology. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-11es/a9-kambil/. Ar-

ticle No. 9.

Loveland:2000:ADA

[250] D. W. Loveland. Automated deduction: achievements and future directions. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-11es/a10-loveland/.

Article No. 10.

OLeary:2000:KMB

[251] Daniel E. O’Leary and Peter Selfridge. Knowledge management for best practices. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a11-o_leary/.

Article No. 11.

Partridge:2000:NPC

[252] Derek Partridge. Non-programmed computation. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2000-43-11es/a12-partridge/

. Article No. 12.

Ragowsky:2000:BUI

[253] Arik Ragowsky, Niv Ahituv, and Seev Neumann. The benefits of using information systems. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2000-
43-11es/a13-ragowsky/. Article No.

13.

Scott:2000:CCC

[254] Louis Scott, Levente Horvath, and Donald Day. Characterizing case constraints. Communications of the ACM, 43(11es), 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a14-scott/. Ar-

ticle No. 14.

Sugumaran:2000:SRS

[255] Vijayan Sugumaran, Mohan Tanniru, and Veda C. Storey. Supporting reuse in systems analysis. Communications of the ACM, 43 (11es), 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-11es/a15-sugumaran/

. Article No. 15.

Crawford:2000:Eb

[256] Diane Crawford. Editorial. Communications of the ACM, 43(12):5, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2000-43-12/p5-crawford/p5-
crawford.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2000-43-12/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2000-43-
12/p5-crawford/.

Fox:2000:NTCb

[257] Robert Fox. News track: Cut kids computer time; computers killed the radio; 3D orthodonture; denial of isolation; you’ve got Braille; all’s fare; sidebar: Connecting the dots. Communications of the ACM, 43(12):9–10, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p9-fox/.

Crawford:2000:FSP

[258] Diane Crawford, David Davenport, Michael P. Conlon, Wallace Koehler, Henry Gladney, Andrew W. Appel, Edward W. Felten, Mark Wallace, Fred Ballard, Rich Morin, and Steve Green. Forum: Software piracy is not just economics anymore; need stronger argument for access control; most enlightening ever. Communications of the ACM, 43(12):11–13, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p11-crawford/.

Soloway:2000:LETb

[259] Elliot Soloway, Cathleen Norris, Phyllis Blumenfeld, Ron Marx, Joe Krajcik, and Barry Fishman. Log on education: the three Ts of elementary education. Communications of the ACM, 43(12):15–19, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p15-soloway/.

Rada:2000:SSSb

[260] Roy Rada and John Craparo. Sharing standards: standardizing software projects. Communications of the ACM, 43(12):21–25, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p21-rada/.

Gelbord:2000:VSY

[261] Boaz Gelbord. Viewpoint: signing your 011001010. Communications of the ACM, 43(12):27–28, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p27-gelbord/.

Rosenbloom:2000:TT

[262] Andrew Rosenbloom. Trusting technology. Communications of the ACM, 43(12):31–32, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-12/p31-rosenbloom/.

Friedman:2000:TO

[263] Batya Friedman, Peter H. Khan, and Daniel C. Howe. Trust online. Communications of the ACM, 43(12):34–40, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-12/p34-friedman/.

Olson:2000:ITE

[264] Judith S. Olson and Gary M. Olson. i2i trust in e-commerce. Communications of the ACM, 43(12):41–44, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-12/p41-olson/.

Resnick:2000:RS

[265] Paul Resnick, Ko Kuwabara, Richard Zeckhauser, and Eric Friedman. Reputation systems. Communications of the ACM, 43(12):45–48, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2000-43-12/p45-resnick/.

Cassell:2000:EMT

[266] Justine Cassell and Timothy Bickmore. External manifestations of trustworthiness in the interface. Communications of the ACM, 43(12):50–56, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p50-cassell/.

Shneiderman:2000:DTO

[267] Ben Shneiderman. Designing trust into online experiences. Communications of the ACM, 43(12):57–59, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p57-shneiderman/

.

Uslaner:2000:SCN

[268] Eric M. Uslaner. Social capital and the net. Communications of the ACM, 43(12):60–64, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p60-uslaner/.

Kreie:2000:MED

[269] Jennifer Kreie and Timothy Paul Cronan. Making ethical decisions. Communications of the ACM, 43(12):66–71, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p66-kreie/.

Schoder:2000:BFT

[270] Detlef Schoder and Pai-Ling Yin. Building firm trust online. Communications of the ACM, 43(12):73–79, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p73-schoder/.

Jones:2000:TRE

[271] Sara Jones, Marc Wilikens, Philip Morris, and Marcelo Masera. Trust requirements in e-business. Communications of the ACM, 43(12):81–87, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p81-jones/.

Moores:2000:SPV

[272] Trevor Moores and Gurpreet Dhillon. Software piracy: a view from Hong Kong. Communications of the ACM, 43(12):88–93, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p88-moores/.

Malik:2000:TOP

[273] Masud Ahmad Malik. Technical opinion: On the perils of programming. Communications of the ACM, 43(12):95–97, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p95-malik/.

Schneier:2000:IRS

[274] Bruce Schneier. Inside risks: semantic network attacks. Communications of the ACM, 43(12):168, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2000-43-12/p168-schneier/.

See letters [300, 377].

Crawford:2001:Ea

[275] Diane Crawford. Editorial. Communications of the ACM, 44(1):5, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p5-crawford/.

Fox:2001:NTCa

[276] Robert Fox. News track: Codebreaking software officially illegal; clogging sky channels; info-tech giant Japan; global dot-cons; salmon sensoring; boots made for walking; sidebar: Top 10 fastest growing, high-tech metro areas in the U.S. Communications of the ACM, 44(1):9–10, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

delivery.acm.org/10.1145/360000/
357492/p9-fox.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-1/p9-fox/.

Glass:2001:PPE

[277] Robert L. Glass. Practical programmer: an embarrassing, yet rewarding, ending to a previous column. Communications of the ACM, 44(1):11, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p11-glass/.

Armour:2001:BSL

[278] Phillip G. Armour. The business of software: the laws of software process. Communications of the ACM, 44(1): 15, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p15-armour/.

Foner:2001:VFF

[279] Lenny Foner. Viewpoint: fixing a flawed domain name system. Communications of the ACM, 44(1):19, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p19-foner/.

Gronlund:2001:DIF

[280] Åke Grönlund. Democracy in an ITframed society: introduction. Communications of the ACM, 44(1):22, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p22-gronlund/.

Watson:2001:SPE

[281] Richard T. Watson and Bryan Mundy. A strategic perspective of electronic democracy. Communications of the ACM, 44(1):27, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p27-watson/.

Anttiroiko:2001:TEI

[282] Ari-Veikko Anttiroiko. Toward the European information society. Communications of the ACM, 44(1):31, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p31-anttiroiko/.

AAgren:2001:ODE

[283] Per-Olof Ågren. Is online democracy in the EU for professionals only? Communications of the ACM, 44(1): 36, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p36-agren/.

Becker:2001:RIN

[284] Ted Becker. Rating the impact of new technologies on democracy. Communications of the ACM, 44(1):39, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p39-becker/.

Snellen:2001:IBF

[285] Ignace Snellen. ICTs, bureaucracies, and the future of democracy. Communications of the ACM, 44(1):45, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p45-snellen/.

AAstrom:2001:SDO

[286] Joachim Åström. Should democracy online be quick, strong, or thin? Communications of the ACM, 44(1):49, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p49-astrom/.

Schuler:2001:CPN

[287] Doug Schuler. Computer professionals and the next culture of democracy. Communications of the ACM, 44(1): 52, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p52-schuler/.

Taylor:2001:PDP

[288] John A. Taylor and Eleanor Burt. Notfor-profits in the democratic polity. Communications of the ACM, 44(1): 58, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p58-taylor/.

Porte:2001:WGG

[289] Todd M. La Porte, Chris C. Demchak, and Christian Friis. Webbing governance: global trends across national-level public agencies. Communications of the ACM, 44(1):63, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-1/p63-la_porte/.

Hoffman:2001:IVP

[290] Lance J. Hoffman and Lorrie Cranor. Internet voting for public officials: introduction. Communications of the ACM, 44(1):69, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p69-hoffman/.

Mohen:2001:CIV

[291] Joe Mohen and Julia Glidden. The case for Internet voting. Communications of the ACM, 44(1):72, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p72-mohen/.

Phillips:2001:GRI

[292] Deborah M. Phillips and Hans A. von Spakovsky. Gauging the risks of Internet elections. Communications of the ACM, 44(1):73, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p73-phillips/. See

correction [377].

Bell:2001:PDS

[293] Gordon Bell. A personal digital store. Communications of the ACM, 44(1): 86, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p86-bell/.

Chen:2001:IOI

[294] Thomas M. Chen. Increasing the observability of Internet behavior. Communications of the ACM, 44(1):93, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p93-chen/.

Chaudhury:2001:WCC

[295] Abhijit Chaudhury, Debasish Mallick, and H. Raghav Rao. Web channels in e-commerce. Communications of the ACM, 44(1):99, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p99-chaudhury/.

Foley:2001:HVS

[296] Simon N. Foley and Robert Dumigan. Are handheld viruses a significant threat? Communications of the ACM, 44(1):105, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p105-foley/.

Mercuri:2001:IRS

[297] Rebecca T. Mercuri and Peter G. Neumann. Inside risks: system integrity revisited. Communications of the ACM, 44(1):160, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-1/p160-mercuri/.

Crawford:2001:EPa

[298] Diane Crawford. Editorial pointers. Communications of the ACM, 44(2): 5, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p5-crawford/.

Fox:2001:NTCb

[299] Robert Fox. News track: Crashfree consortium; Pooh corner; bedside matters; silver lining; no home work; ear shot; quick picks. Communications of the ACM, 44(2):9–10, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p9-fox/.

Crawford:2001:FHC

[300] Diane Crawford, Mick Zraly, Hal Berghel, Ken Pugh, Mat H. West, Conrad Weisert, Terry Steyaert, and Richard Johnson. Forum: How can the Web advance Western democracies? who needs digital signatures; misinformation and the Emulex hoax; OOSCD not really so unified; go back to non-OOSD. Communications of the ACM, 44(2):11–13, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p11-crawford/. See

[274].

Denning:2001:PIW

[301] Peter J. Denning. The profession of IT: who are we? Communications of the ACM, 44(2):15–19, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p15-denning/.

McGinity:2001:SCN

[302] Meg McGinity. Staying connected: name is the game. Communications of the ACM, 44(2):21–23, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p21-mcginity/.

Bellman:2001:SOO

[303] Steven Bellman, Eric J. Johnson, and Gerald L. Lohse. On site: opt-in or optout?: it depends on the question. Communications of the ACM, 44(2):25–27, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p25-bellman/.

Bashir:2001:SNS

[304] Imran Bashir, Enrico Serafini, and Kevin Wall. Securing network software applications: introduction. Communications of the ACM, 44(2):28–30, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p28-bashir/.

Viega:2001:TMS

[305] John Viega, Tadayoshi Kohno, and Bruce Potter. Trust (and mistrust) in secure applications. Communications of the ACM, 44(2):31–36, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p31-viega/.

Joshi:2001:SMW

[306] James B. D. Joshi, Walid G. Aref, Arif Ghafoor, and Eugene H. Spafford. Security models for Webbased applications. Communications of the ACM, 44(2):38–44, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p38-joshi/.

Smith:2001:PPW

[307] Richard M. Smith, Michael Brittain, Ivan Fetch, Hailin Wu, and David M. Martin. The privacy practices of Web browser extensions. Communications of the ACM, 44(2):45–50, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p45-smith/.

Ghosh:2001:SSP

[308] Anup K. Ghosh and Tara M. Swaminatha. Software security and privacy risks in mobile e-commerce. Communications of the ACM, 44(2):51–57, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p51-ghosh/.

Dalton:2001:OSA

[309] Chris Dalton and Tse Huong Choo. An operating system approach to securing e-services. Communications of the ACM, 44(2):58–64, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p58-dalton/.

Samuelson:2001:IPI

[310] Pamela Samuelson. Intellectual property for an information age: introduction. Communications of the ACM, 44(2):66–68, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-2/p66-samuelson/.

Burk:2001:CFP

[311] Dan L. Burk. Copyrightable functions and patentable speech. Communications of the ACM, 44(2):69–75, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p69-burk/.

Davis:2001:DD

[312] Randall Davis. The digital dilemma. Communications of the ACM, 44(2): 77–83, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-2/p77-davis/.

Benkler:2001:BIE

[313] Yochai Benkler. The battle over the institutional ecosystem in the digital environment. Communications of the ACM, 44(2):84–90, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p84-benkler/.

Froomkin:2001:CTD

[314] A. Michael Froomkin. The collision of trademarks, domain names, and due process in cyberspace. Communications of the ACM, 44(2):91–97, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p91-froomkin/.

ORourke:2001:VTA

[315] Maureen A. O’Rourke. Is virtual trespass an apt analogy? Communications of the ACM, 44(2):98–103, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p98-o_rourke/.

Neumann:2001:IRW

[316] Peter G. Neumann. Inside risks: what to know about. Communications of the ACM, 44(2):136, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-2/p136-neumann/.

Crawford:2001:EPb

[317] Diane Crawford. Editorial pointers. Communications of the ACM, 44(3): 5, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-2/p5-crawford/p5-crawford.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        2/p5-crawford/.

Fox:2001:NTF

[318] Robert Fox. News track: The future of the tech workplace and from the desk of rage and smarty pants and toddlers and computers don’t mix and dot-com graveyard and monkeying around and mating call. Communications of the ACM, 44(3):9– 10, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p9-fox/.

Armour:2001:BSS

[319] Phillip G. Armour. The business of software: Software as currency. Communications of the ACM, 44(3):13–14, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2001:DVCa

[320] Hal Berghel. Digital village: A cyberpublishing manifesto. Communications of the ACM, 44(3):17–20, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2001:DI

[321] Gordon Bell and Jim Gray. Digital immortality. Communications of the ACM, 44(3):28–30, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-3/p28-bell/.

Colwell:2001:CCI

[322] Rita R. Colwell. Closing the circle of information technology. Communications of the ACM, 44(3):31– 34, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p31-colwell/.

Badler:2001:VB

[323] Norman I. Badler. Virtual beings. Communications of the ACM, 44(3):33–35, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p33-badler/.

Norman:2001:C

[324] Donald A. Norman. Cyborgs. Communications of the ACM, 44(3):36– 37, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p36-norman/.

Jain:2001:DE

[325] Ramesh Jain. Digital experience. Communications of the ACM, 44(3):38– 40, March 2001. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p38-jain/.

Jacobson:2001:DF

[326] Joseph Jacobson. The desktop fab. Communications of the ACM, 44 (3):41–43, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p41-jacobson/.

Bricklin:2001:LPE

[327] Dan Bricklin. Look to the past to envision the future. Communications of the ACM, 44(3):44, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p44-bricklin/.

Selker:2001:AH

[328] Ted Selker. Affecting humanity. Communications of the ACM, 44(3):45, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p45-selker/.

Kappelman:2001:FO

[329] Leon Kappelman. The future is ours. Communications of the ACM, 44(3):46–47, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p46-kappelman/.

Bove:2001:AM

[330] V. Michael Bove. Astronauts and mosquitoes. Communications of the ACM, 44(3):48–49, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-3/p48-bove/.

Dam:2001:UID

[331] Andries van Dam. User interfaces: disappearing, dissolving, and evolving. Communications of the ACM, 44(3):50–52, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p50-van\_dam/.

Brewer:2001:WES

[332] Eric A. Brewer. When everything is searchable. Communications of the ACM, 44(3):53–54, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p53-brewer/.

Cooper:2001:BCA

[333] Martin Cooper. Bandwidth and the creation of awareness. Communications of the ACM, 44(3):55– 57, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p55-cooper/.

Horan:2001:PP

[334] Thomas A. Horan. The paradox of place. Communications of the ACM, 44(3):58–59, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-3/p58-horan/.

Vetter:2001:WW

[335] Ron Vetter. The wireless web. Communications of the ACM, 44(3):60– 61, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p60-vetter/.

Fayyad:2001:DPD

[336] Usama Fayyad. The digital physics of data mining. Communications of the ACM, 44(3):62–65, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p62-fayyad/.

Lai:2001:WCS

[337] Jennifer Lai. When computers speak, hear, and understand. Communications of the ACM, 44(3):66–67, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p66-lai/.

Waldo:2001:WNE

[338] Jim Waldo. When the network is everything. Communications of the ACM, 44(3):68–69, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/

pubs/citations/journals/cacm/2001-
44-3/p68-waldo/.

Schwartz:2001:W

[339] Steven J. Schwartz. Wearables in 2048. Communications of the ACM, 44(3):70–71, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p70-schwartz/.

Miner:2001:PFE

[340] Cameron Miner. Pushing functionality into even smaller devices. Communications of the ACM, 44(3):72– 73, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p72-miner/.

Johnson:2001:CBM

[341] Christopher R. Johnson. Computational bioimaging for medical diagnosis and treatment. Communications of the ACM, 44(3):74–75, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p74-johnson/.

Cohen:2001:CB

[342] Jacques Cohen. Computers and biology. Communications of the ACM, 44(3):76–79, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p76-cohen/.

Sterling:2001:CCA

[343] Thomas Sterling. Continuum computer architecture for exaflops computation. Communications of the ACM, 44(3):78–80, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p78-sterling/.

Crowcroft:2001:NLN

[344] Jon Crowcroft. Never lost, never forgotten. Communications of the ACM, 44(3):81, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p81-crowcroft/.

Muller:2001:SLI

[345] Michael J. Muller, Ellen Christiansen, Bonnie Nardi, and Susan Dray. Spiritual life and information technology. Communications of the ACM, 44(3):82–83, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p82-muller/.

Diffie:2001:UC

[346] Whitfield Diffie. Ultimate cryptography. Communications of the ACM, 44(3):84–86, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p84-diffie/.

Kurzweil:2001:PPD

[347] Ray Kurzweil. Promise and perilthe deeply intertwined poles of 21st century technology. Communications of the ACM, 44(3):88–91, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p88-kurzweil/.

Dijkstra:2001:ECS

[348] Edsger W. Dijkstra. The end of computing science? Communications of the ACM, 44(3):92, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p92-dijkstra/. Copy-

right held by author.

Varian:2001:CE

[349] Hal R. Varian. The computermediated economy. Communications of the ACM, 44(3):93, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-3/p93-varian/.

Cringely:2001:BAD

[350] Robert X. Cringely. Be absolute for death: life after Moore’s law. Communications of the ACM, 44(3):94, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p94-cringely/.

Meeks:2001:ATT

[351] Brock N. Meeks. Accountability through transparency: life in 2050. Communications of the ACM, 44 (3):96–97, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p96-meeks/.

Samuelson:2001:TNP

[352] Pamela Samuelson. Toward a new politics of intellectual property. Communications of the ACM, 44(3):98– 99, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p98-samuelson/.

Tsichritzis:2001:FPW

[353] Dennis Tsichritzis. Forget the past to win the future. Communications of the ACM, 44(3):100–101, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p100-tsichritzis/

.

Grosso:2001:DS

[354] Andrew Grosso. The demise of sovereignty. Communications of the ACM, 44(3):102–103, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-3/p102-grosso/.

Townsend:2001:EE

[355] Anthony M. Townsend and James T. Bennett. Electronic empire. Communications of the ACM, 44(3):104– 105, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p104-townsend/.

Schwartz:2001:LRP

[356] Ari Schwartz. A larger role in the public policy process for user control. Communications of the ACM, 44(3):106– 107, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p106-schwartz/.

Holtzblatt:2001:IF

[357] Karen Holtzblatt. Inventing the future. Communications of the ACM, 44(3):108–110, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p108-holtzblatt/.

Stallman:2001:CFW

[358] Richard Stallman. Can freedom withstand e-books? Communications of the ACM, 44(3):111, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p111-stallman/.

Denning:2001:MZA

[359] Peter J. Denning. Many zeros ahead. Communications of the ACM, 44(3): 112, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p112-denning/.

Schneier:2001:ICI

[360] Bruce Schneier. Insurance and the computer industry. Communications of the ACM, 44(3):114–115, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p114-schneier/.

Chon:2001:FID

[361] Kilnam Chon. The future of the Internet digital divide. Communications of the ACM, 44(3):116–117, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p116-chon/.

Booch:2001:DF

[362] Grady Booch. Developing the future. Communications of the ACM, 44(3):118–121, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p118-booch/.

Lieberman:2001:WSE

[363] Henry Lieberman and Christopher Fry. Will software ever work? Communications of the ACM, 44(3):122– 124, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p122-lieberman/.

Gorenberg:2001:JSW

[364] Mark Gorenberg and Ann Winblad. A just-in-time software-based world.

Communications of the ACM, 44(3): 125, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p125-gorenberg/.

Constantine:2001:BF

[365] Larry L. Constantine. Back to the future. Communications of the ACM, 44(3):126, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-3/p126-constantine/

.

Pancake:2001:UBU

[366] Cherri Pancake. The ubiquitous beauty of user-aware software. Communications of the ACM, 44(3):130, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p130-pancake/.

Bellovin:2001:CSE

[367] Steven M. Bellovin. Computer security — an end state? Communications of the ACM, 44(3):131– 132, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p131-bellovin/.

Riecken:2001:COC

[368] Doug Riecken. A commonsense opportunity for computing. Communications of the ACM, 44(3):133, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p133-riecken/.

Ganesan:2001:KRS

[369] Ravi Ganesan. Keep (over) reaching for the stars. Communications of the ACM, 44(3):134–135, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p134-ganesan/.

Glenn:2001:EKF

[370] John Glenn. Education is the key to future dreams. Communications of the ACM, 44(3):136–138, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p136-glenn/.

Borg:2001:ULC

[371] Anita Borg. Universal literacy — a challenge for computing in the 21st century. Communications of the ACM, 44(3):139–141, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-3/p139-borg/.

Schank:2001:CIM

[372] Roger C. Schank. The computer isn’t the medium, it’s the message. Communications of the ACM, 44(3):142– 143, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-3/p142-schank/.

Resnick:2001:CFG

[373] Mitchel Resnick. Closing the fluency gap. Communications of the ACM, 44(3):144–145, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-3/p144-resnick/.

Neumann:2001:IRC

[374] Peter G. Neumann and David L. Parnas. Inside risks: Computers: Boon or bane? Communications of the ACM, 44(3):168, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:EPc

[375] Diane Crawford. Editorial pointers. Communications of the ACM, 44(4): 5, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-3/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
3/p5-crawford/.

Fox:2001:NTI

[376] Robert Fox. News track: ICANN’s domain name experiment; paper voting prevails; printed-on battery power; campus net access rethought; who let the robots out?; U.S. patent lead; taste the web; have virus, will work for mayor. Communications of the ACM, 44(4):9–10, April 2001. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-4/p9-fox/.

Crawford:2001:FPV

[377] Diane Crawford, Thomas Tiahrt, Moti Ben-Ari, Matt West, Hans A. von Spakovsky, and Deborah Phillips. Forum: Participatory vs. representative democracy; why store everything?; Emulex hoax; correction. Communications of the ACM, 44(4):13– 15, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-4/p13-crawford/. See

[274, 292].

Glass:2001:PPA

[378] Robert L. Glass. Practical programmer: academics, and the scarlet letter “A”. Communications of the ACM, 44(4):17–18, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-4/p17-glass/.

Denning:2001:PICa

[379] Peter J. Denning. The profession of IT: crossing the chasm. Communications of the ACM, 44(4):21– 25, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-4/p21-denning/.

Lanning:2001:VBS

[380] David Lanning. Viewpoint: on bias, science, equal protection, and liability. Communications of the ACM, 44(4): 27–29, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p27-lanning/.

Albayrak:2001:ATT

[381] Sahin Albayrak. Agent-oriented technology for telecommunications: introduction. Communications of the ACM, 44(4):30–33, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-4/p30-albayrak/.

Jennings:2001:AAB

[382] Nicholas R. Jennings. An agentbased approach for building complex software systems. Communications of the ACM, 44(4):35–41, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p35-jennings/.

Fricke:2001:ATS

[383] Stefan Fricke, Karsten Bsufka, Jan Keiser, Torge Schmidt, Ralf Sesseler, and Sahin Albayrak. Agentbased telematic services and telecom applications. Communications of the ACM, 44(4):43–48, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p43-fricke/.

Singh:2001:CSL

[384] Munindar P. Singh, Bin Yu, and Mahadevan Venkatraman. Communitybased service location. Communications of the ACM, 44(4):49–54, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-4/p49-singh/.

Brunner:2001:SCM

[385] Marcus Brunner, Bernhard Plattner, and Rolf Stadler. Service creation and management in active telecom networks. Communications of the ACM, 44(4):55–61, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-4/p55-brunner/.

Valera:2001:CME

[386] Francisco Valera, Jorge E. López de Vergara, José I. Moreno, Víctor A. Villagrá, and Julio Berrocal. Communication management experiences in e-commerce. Communications of the ACM, 44(4):63–69, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-4/p63-valera/.

Papazoglou:2001:ATS

[387] Mike P. Papazoglou. Agent-oriented technology in support of e-business. Communications of the ACM, 44(4): 71–77, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p71-papazoglou/.

Grover:2001:EIM

[388] Varun Grover and James T. C. Teng. E-commerce and the information market. Communications of the ACM, 44(4):79–86, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p79-grover/.

Keil:2001:BWT

[389] Mark Keil and Daniel Robey. Blowing the whistle on troubled software projects. Communications of the ACM, 44(4):87–93, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-4/p87-keil/.

Cheyne:2001:TAI

[390] Tanya L. Cheyne and Frank E. Ritter. Targeting audiences on the Internet. Communications of the ACM, 44(4): 94–98, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p94-cheyne/.

Kappel:2001:BDA

[391] G. Kappel, S. Rausch-Schott, W. Retschitzegger, and M. Sakkinen. Bottomup design of active object-oriented databases. Communications of the ACM, 44(4):99–104, April 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/citations/
journals/cacm/2001-44-4/p99-kappel/

.

Ward:2001:TOS

[392] Robert P. Ward, Mohamed E. Fayad, and Mauri Laitinen. Thinking objectively: software process improvement in the small. Communications of the ACM, 44(4):105–107, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p105-ward/.

Crawford:2001:CO

[393] Diane Crawford. Career opportunities. Communications of the ACM, 44(4):109–124, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/journals/
cacm/2001-44-4/p109-crowford/.

Crowford:2001:CE

[394] Diane Crowford. Calendar of events. Communications of the ACM, 44(4): 125, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p125-crowford/.

Schneier:2001:IRC

[395] Bruce Schneier. Inside risks: cyber underwriters lab. Communications of the ACM, 44(4):128, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-4/p128-schneier/.

Crawford:2001:Eb

[396] Diane Crawford. Editorial. Communications of the ACM, 44(5):5, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-4/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
4/p5-crawford/.

Fox:2001:NTE

[397] Robert Fox. News track: Euro net access exceeds U.S.; in-dash MP3; poll on the future; beam of pain; laundry made easier; detecting schizophrenia; Internet earthquake warnings; patron net saint. Communications of the ACM, 44(5):9– 10, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-5/p9-fox/p9-fox.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-5/p9-fox/.

Crawford:2001:FEC

[398] Diane Crawford, John F. Dalphin, Andy Oram, Louis A. Russ, and Peter Denning. Forum: Evolving commitment to the profession of computer science. Communications of the ACM, 44(5):11–12, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p11-crawford/p11-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-5/p11-crawford/.

Munro:2001:WUB

[399] Neil Munro. From Washington: The unhappy but beneficial coexistence of the FBI and the tech elite. Communications of the ACM, 44(5):15– 18, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p15-munro/p15-munro.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-5/p15-munro/

.

Berghel:2001:DVCb

[400] Hal Berghel. Digital village: Caustic cookies. Communications of the ACM, 44(5):19–22, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p19-berghel/p19-
                                        berghel.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-5/p19-berghel/.

Howard:2001:SSE

[401] Alan Howard. On site: Software engineering project management. Communications of the ACM, 44(5):23– 24, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p23-howard/p23-howard.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p23-howard/.

Apt:2001:VOM

[402] Krzysztof R. Apt. Viewpoint: One more revolution to make: free scientific publishing. Communications of the ACM, 44(5):25–28, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p25-apt/p25-apt.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p25-apt/.

Fox:2001:DLI

[403] Edward A. Fox and Gary Marchionini. Digital libraries: Introduction. Communications of the ACM, 44(5): 30–32, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p30-fox/p30-fox.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p30-fox/.

Crane:2001:DDT

[404] Gregory Crane, Robert F. Chavez, Anne Mahoney, Thomas L. Milbank, Jeffrey A. Rydberg-Cox, David A. Smith, and Clifford E. Wulfman. Drudgery and deep thought. Communications of the ACM, 44(5):34– 40, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-5/p34-crane/p34-crane.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-5/p34-crane/

.

Nevill-Manning:2001:BDL

[405] Craig Nevill-Manning. The biological digital library. Communications of the ACM, 44(5):41–42, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p41-nevill-manning/p41-
                                        nevill-manning.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        cacm/2001-44-5/p41-nevill-manning/

.

Inouye:2001:DSL

[406] Alan S. Inouye. A digital strategy for the library of congress. Communications of the ACM, 44(5):43, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p43-inouye/p43-inouye.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p43-inouye/.

Jones:2001:OSD

[407] Paul Jones. Open (source)ing the doors for contributor-run digital libraries. Communications of the ACM, 44(5): 45–46, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p45-jones/p45-jones.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-5/p45-jones/

.

Witten:2001:GOD

[408] Ian H. Witten, David Bainbridge, and Stefan Boddie. Greenstone: Opensource DL software. Communications of the ACM, 44(5):47, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-5/p47-witten/p47-witten.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p47-witten/.

McCray:2001:PDL

[409] Alexa T. McCray and Marie E. Gallagher. Principles for digital library development. Communications of the ACM, 44(5):48–54, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-5/p48-mccray/p48-mccray.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p48-mccray/.

Marshall:2001:DLM

[410] Catherine C. Marshall, Gene Golovchinsky, and Morgan N. Price. Digital libraries and mobility. Communications of the ACM, 44(5):55– 56, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p55-marshall/p55-
marshall.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-5/p55-marshall/.

Brown:2001:BLD

[411] Michael S. Brown, W. Brent Seales, Stephen B. Webb, and Christopher O. Jaynes. Building large-format displays for digital libraries. Communications of the ACM, 44(5):57– 59, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p57-brown/p57-brown.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-5/p57-brown/

.

Nelson:2001:BSO

[412] Michael L. Nelson and Kurt Maly. Buckets: smart objects for digital libraries. Communications of the ACM, 44(5):60–62, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p60-nelson/p60-nelson.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p60-nelson/.

Gladney:2001:AMD

[413] H. M. Gladney and Arthur Cantu. Authorization management for digital libraries. Communications of the ACM, 44(5):63–65, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p63-gladney/p63-gladney.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p63-gladney/.

Borgman:2001:WLD

[414] Christine L. Borgman. Where is the librarian in the digital library? Communications of the ACM, 44(5):66– 68, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p66-borgman/p66-

borgman.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-5/p66-borgman/.

Arms:2001:URN

[415] William Y. Arms. Uniform resource names: handles, PURLs, and digital object identifiers. Communications of the ACM, 44(5):68, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-5/p68-arms/p68-arms.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-5/p68-arms/

.

Tibbo:2001:APE

[416] Helen R. Tibbo. Archival perspectives on the emerging digital library. Communications of the ACM, 44(5): 69–70, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-5/p69-tibbo/p69-tibbo.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-5/p69-tibbo/

.

Duval:2001:AKP

[417] Erik Duval, Eddy Forte, Kris Cardinaels, Bart Verhoeven, Rafael Van Durm, Koen Hendrikx, Maria Wentland Forte, Norbert Ebel, Maciej Macowicz, Ken Warkentyne, and Florence Haenni. The Ariadne knowledge pool system. Communications of the ACM, 44(5):72–78, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/

                                        2001-44-5/p72-duval/p72-duval.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-5/p72-duval/

.

McArthur:2001:TSD

[418] David McArthur, Sarah Giersch, Bill Graves, Charles R. Ward, Richard Dillaman, Russell Herman, Gabriel Lugo, James Reeves, Ron Vetter, Deborah Knox, and Scott Owen. Toward a sharable digital library of reusable teaching resources. Communications of the ACM, 44(5):79, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p79-mcarthur/p79-
                                        mcarthur.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-5/p79-mcarthur/.

Marlino:2001:DLE

[419] Mary Marlino, Tamara Sumner, David Fulker, Cathryn Manduca, and David Mogk. The digital library for Earth system education: building community, building the library. Communications of the ACM, 44(5):80– 81, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p80-marlino/p80-
                                        marlino.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-5/p80-marlino/.

Witten:2001:PDL

[420] Ian H. Witten, Michel Loots, Maria F. Trujillo, and David Bainbridge. The promise of digital libraries in developing countries. Communications of the ACM, 44(5):82–85, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-5/p82-witten/p82-witten.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p82-witten/.

Zia:2001:NNS

[421] Lee L. Zia. The NSF national science, mathematics, engineering, and technology education digital library program. Communications of the ACM, 44(5): 83, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-5/p83-zia/.

Apperley:2001:NHN

[422] Mark Apperley, Sally Jo Cunningham, Te Taka Keegan, and Ian H. Witten. NIUPEPA: a historical newspaper collection. Communications of the ACM, 44(5):86–87, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p86-apperley/p86-
apperley.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-5/p86-apperley/.

Urs:2001:VID

[423] Shalini R. Urs and K. S. Raghavan. Vidyanidhi: Indian digital library of electronic theses. Communications of the ACM, 44(5):88– 89, May 2001. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p88-urs/p88-urs.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p88-urs/.

Rous:2001:ADL

[424] Bernard Rous. The ACM digital library. Communications of the ACM, 44(5):90–91, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        cacm/2001-44-5/p90-rous/p90-rous.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p90-rous/.

Baster:2001:BCC

[425] Greg Baster, Prabhudev Konana, and Judy E. Scott. Business components: a case study of Bankers Trust Australia Limited. Communications of the ACM, 44(5):92– 98, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-5/p92-baster/p92-baster.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        5/p92-baster/.

Astley:2001:CMM

[426] Mark Astley, Daniel Sturman, and Gul Agha. Customizable middleware for modular distributed software. Communications of the ACM, 44(5):99– 107, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-5/p99-astley/p99-astley.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p99-astley/.

Cohoon:2001:TIF

[427] J. McGrath Cohoon. Toward improving female retention in the computer science major. Communications of the ACM, 44(5):108–114, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-5/p108-cohoon/p108-
cohoon.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-5/p108-cohoon/.

Weinstein:2001:IRB

[428] Lauren Weinstein. Inside risks: be seeing you! Communications of the ACM, 44(5):128, May 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-5/p128-weinstein/p128-weinstein.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
5/p128-weinstein/.

Crawford:2001:EPd

[429] Diane Crawford. Editorial pointers. Communications of the ACM, 44(6): 5, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-6/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/

citations/journals/cacm/2001-44-
6/p5-crawford/.

Fox:2001:NTSa

[430] Robert Fox. News track: Security stats; robotic insect; disposable phones; scientific oath; simulating dripped drops; an aviation first; coffee-cam to go black. Communications of the ACM, 44(6):9– 10, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-6/p9-fox/p9-fox.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-6/p9-crawford/.

Crawford:2001:FMT

[431] Diane Crawford, Ilya Yakovlev, Yochai Benkler, Paul McKenney, and Donn B. Parker. Forum: More than one way to product quality information; hope for educational sea change; A consistent definition of information security. Communications of the ACM, 44(6): 11–14, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-6/p11-crawford/p11-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-6/p11-crawford/.

Soloway:2001:LEH

[432] Elliot Soloway, Cathleen Norris, Phyllis Blumenfeld, Barry Fishman, Joseph Krajcik, and Ron Marx. Log on education: Handheld devices are ready-at-hand. Communications of the ACM, 44(6):15–20, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-6/p15-soloway/p15-
soloway.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-6/p15-soloway/.

Goodman:2001:IPW

[433] Seymour E. Goodman, James B. Gottstein, and Diane S. Goodman. International perspectives: Wiring the wilderness in Alaska and the Yukon. Communications of the ACM, 44(6): 21–26, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-6/p21-goodman/p21-
goodman.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-6/p21-goodman/.

Palma:2001:VWW

[434] Paul de Palma. Viewpoint: Why women avoid computer science. Communications of the ACM, 44(6):27– 30, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-6/p27-de\_palma/p27-de\_
palma.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-6/p27-de\_palma/;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-6/p27-de_palma/.

Dewan:2001:CIE

[435] Rajiv Dewan and Abraham Seidmann. Current issues in ebanking: introduction. Communications of the ACM, 44(6):31–32, June 2001. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p31-dewan/p31-dewan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-6/p31-dewan/

.

Saatcioglu:2001:DFP

[436] Kemal Saatcioglu, Jan Stallaert, and Andrew B. Whinston. Design of a financial portal. Communications of the ACM, 44(6):33–38, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p33-saatcioglu/p33-saatcioglu.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        6/p33-saatcioglu/.

Wind:2001:CQF

[437] Yoram (Jerry) Wind. The challenge of “customerization” in financial services. Communications of the ACM, 44(6): 39–44, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p39-wind/p39-wind.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-6/p39-wind/

.

Altinkemer:2001:BES

[438] Kermal Altinkemer. Bundling ebanking services. Communications of the ACM, 44(6):45–47, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p45-altinkemer/p45-altinkemer.

pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
6/p45-altinkemer/.

Ahituv:2001:OIS

[439] Niv Ahituv. The open information society. Communications of the ACM, 44(6):48–52, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p48-ahituv/p48-ahituv.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
6/p48-ahituv/.

Holland:2001:PTS

[440] Christopher P. Holland and John B. Westwood. Product-market and technology strategies in banking. Communications of the ACM, 44(6):53– 57, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-6/p53-holland/p53-
holland.pdf;   http://www.acm.org/
pubs/citations/journals/cacm/2001-
44-6/p53-holland/.

Kumar:2001:TSS

[441] Kuldeep Kumar. Technology for supporting supply chain management: introduction. Communications of the ACM, 44(6):58–61, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p58-kumar/p58-kumar.pdf;
http://www.acm.org/pubs/citations/

journals/cacm/2001-44-6/p58-kumar/

.

Simon:2001:AML

[442] Steve John Simon. The art of military logistics. Communications of the ACM, 44(6):62–66, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p62-simon/p62-simon.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-6/p62-simon/

.

Welty:2001:MTC

[443] Bill Welty and Irma Becerra-Fernandez. Managing trust and commitment in collaborative supply chain relationships. Communications of the ACM, 44(6):67–73, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/2001-44-6/p67-welty/p67-welty.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
6/p67-welty/.

Hillegersberg:2001:SRF

[444] Jos van Hillegersberg, Rob Zuidwijk, Jo van Nunen, and Diane van Eijk. Supporting return flows in the supply chain. Communications of the ACM, 44(6):74–79, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p74-van\_hillegersberg/
p74-van\_hillegersberg.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-6/p74-zuidwijk/.

Ramachandran:2001:CAC

[445] Girish Ramachandran and Sanjay Tiwari. Challenges in the air cargo supply chain. Communications of the ACM, 44(6):80–82, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p80-ramachandran/p80-
ramachandran.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-6/p80-ramachandran/

.

Krishnan:2001:TIR

[446] R. Krishnan. Technology in the Indian retail supply chain. Communications of the ACM, 44(6):83– 87, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-6/p83-krishnan/p83-
krishnan.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-6/p83-krishnan/.

Schuff:2001:CVD

[447] David Schuff and Robert St. Louis. Centralization vs. decentralization of application software. Communications of the ACM, 44(6):88–94, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p88-schuff/p88-schuff.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
6/p88-schuff/.

Klein:2001:NVP

[448] Gary Klein, James J. Jiang, and Marion G. Sobol. A new view of IS personnel performance evaluation. Communications of the ACM, 44(6):95– 102, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p95-klein/p95-klein.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-6/p95-jiang/

.

Shepard:2001:MTS

[449] Terry Shepard, Margaret Lamb, and Diane Kelly. More testing should be taught. Communications of the ACM, 44(6):103–108, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p103-shepard/p103-shepard.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        6/p103-shepard/.

McCormick:2001:TOP

[450] Michael McCormick. Technical opinion: Programming extremism. Communications of the ACM, 44(6):109– 119, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-6/p109-mccormick/p109-mccormick.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        6/p109-mccormick/.

Forno:2001:IRP

[451] Richard Forno and William Feinbloom. Inside risks: PKI: a question of trust and value. Communications of the ACM, 44(6):120, June 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-6/p120-forno/p120-forno.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
6/p120-forno/.

Crawford:2001:EPf

[452] Diane Crawford. Editorial pointers. Communications of the ACM, 44(7): 5, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p5-crawford/p5-crawford.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p5-crawford/p5-crawford.pdf;
http://www.acm.org/pubs/citations/
journals/cacm/2001-44-7/p5-crawford/

.

Fox:2001:NTNa

[453] Robert Fox. News track: Net population’s newest numbers; brush-on smart paint; losses in the face of piracy; losses in the name of quality; hear the Earth’s patter; radio frequency fries electronics; click, drag, plagiarize, gotcha; Web as close as your wrist. Communications of the ACM, 44 (7):9, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p9-fox/p9-fox.html;
http://www.acm.org/pubs/articles/
journals/cacm/2001-44-7/p9-fox/
p9-fox.pdf;   http://www.acm.org/

                                        pubs/citations/journals/cacm/2001-
                                        44-7/p9-fox/.

Crawford:2001:FRB

[454] Diane Crawford, Richard G. Canning, Todd Hodes, Bill Leal, George Nagy, Alan Knight, G. Kappel, S. Rausch-Schott, W. Retschitzegger, M. Sakkinen, Luke Hohmann, and Nils Jensen. Forum: Reviewer bias vs. review validity; millennium reflections; unfashionable Smalltalk; don’t forget existing customers; the “character” of software. Communications of the ACM, 44(7): 11–13, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-7/p11-crawford/p11-
                                        crawford.html;   http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p11-crawford/p11-crawford.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        7/p11-crawford/.

White:2001:AOP

[455] John White. ACM opens portal. Communications of the ACM, 44(7):14–16, 28, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p14-white/p14-white.html;
                                         http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-7/p14-white/
                                        p14-white.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-7/p14-white/.

Grosso:2001:LSI

[456] Andrew Grosso. Legally speaking: the individual in the new age.

Communications of the ACM, 44(7): 17, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p17-grosso/p17-grosso.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p17-grosso/p17-grosso.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p17-grosso/.

Armour:2001:BSM

[457] Phillip G. Armour. The business of software: Matching process to types of teams. Communications of the ACM, 44(7):21, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p21-armour/p21-armour.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p21-armour/p21-armour.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p21-armour/.

Stafford-Fraser:2001:SLT

[458] Quentin Stafford-Fraser. On site: The life and times of the first Web Cam. Communications of the ACM, 44(7): 25, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p25-stafford-fraser/p25-
stafford-fraser.html;   http://
www.acm.org/pubs/articles/journals/
cacm/2001-44-7/p25-stafford-fraser/
p25-stafford-fraser.pdf;   http:
//www.acm.org/pubs/citations/journals/

                                        cacm/2001-44-7/p25-stafford-fraser/

.

Hazelhurst:2001:VDI

[459] Scott Hazelhurst. Viewpoint: Developing IT skills internationally: who’s developing whom? Communications of the ACM, 44(7):27, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p27-hazelhurst/p27-hazelhurst.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        7/p27-hazelhurst/.

Arnold:2001:GIW

[460] David Arnold and Fred Niederman. The global IT work force: introduction. Communications of the ACM, 44 (7):30, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p30-arnold/p30-arnold.
                                        html;   http://www.acm.org/pubs/
                                        articles/journals/cacm/2001-44-
                                        7/p30-arnold/p30-arnold.pdf;   http:
                                        //www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-7/p30-arnold/.

West:2001:IGI

[461] Lawrence A. West and Walter A. Bogumil. Immigration and the global IT work force. Communications of the ACM, 44(7):34, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p34-west/p34-west.html;
                                         http://www.acm.org/pubs/articles/

journals/cacm/2001-44-7/p34-west/
p34-west.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-7/p34-west/.

Gupta:2001:GSI

[462] Phalguni Gupta. Growth scenario of IT industries in India. Communications of the ACM, 44(7):40, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p40-gupta/p40-gupta.html;
http://www.acm.org/pubs/articles/
journals/cacm/2001-44-7/p40-gupta/
p40-gupta.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-7/p40-gupta/.

Mander:2001:DFA

[463] Keith Mander. The decline and fall of the American programmer. Communications of the ACM, 44(7): 43, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p43-mander/p43-mander.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p43-mander/p43-mander.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p43-mander/.

Hellens:2001:AWI

[464] Liisa von Hellens and Sue Nielsen. Australian women in IT. Communications of the ACM, 44(7):46, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/

2001-44-7/p46-von\_hellens/p46-
von\_hellens.html;   http://www.
acm.org/pubs/articles/journals/
cacm/2001-44-7/p46-von\_hellens/
p46-von\_hellens.pdf;   http://
www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p46-von\_hellens/
;   http://www.acm.org/pubs/citations/
journals/cacm/2001-44-7/p46-von_
hellens/.

Sallis:2001:STI

[465] Philip J. Sallis. Some thoughts on IT employment in New Zealand. Communications of the ACM, 44(7): 53, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-7/p53-sallis/p53-sallis.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p53-sallis/p53-sallis.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p53-sallis/.

Crawford:2001:AGE

[466] Diane Crawford. ACM general elections. Communications of the ACM, 44 (7):54, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-7/p54-crawford/p54-
crawford.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p54-crawford/p54-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
7/p54-crawford/.

Marshall:2001:PII

[467] Linda Marshall. A perspective on the IT industry in South Africa. Communications of the ACM, 44(7): 55, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-7/p55-marshall/p55-
marshall.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p55-marshall/p55-marshall.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
7/p55-marshall/.

Agarwal:2001:CHS

[468] Ritu Agarwal and Thomas W. Ferratt. Crafting an HR strategy to meet the need for IT workers. Communications of the ACM, 44(7): 58, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-7/p58-agarwal/p58-
agarwal.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p58-agarwal/p58-agarwal.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
7/p58-agarwal/.

Wong:2001:LSH

[469] Kam-Fai Wong. Labor shortfall in Hong Kong’s IT industry. Communications of the ACM, 44(7):65, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p65-wong/p65-wong.html;
 http://www.acm.org/pubs/articles/
journals/cacm/2001-44-7/p65-wong/
p65-wong.pdf;   http://www.acm.

                                        org/pubs/citations/journals/cacm/
                                        2001-44-7/p65-wong/.

Klawe:2001:RN

[470] Maria Klawe. Refreshing the nerds. Communications of the ACM, 44(7): 67, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p67-klawe/p67-klawe.html;
                                         http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-7/p67-klawe/
                                        p67-klawe.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-7/p67-klawe/.

Moore:2001:FCC

[471] Jo Ellen Moore, Susan E. Yager, Mary Sumner, and Galen B. Crow. Facilitating career changes into IT. Communications of the ACM, 44(7): 70, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p70-moore/p70-moore.html;
                                         http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-7/p70-moore/
                                        p70-moore.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-7/p70-moore/.

Trauth:2001:MIW

[472] Eileen M. Trauth. Mapping informationsector work to the work force. Communications of the ACM, 44(7):74, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p74-trauth/p74-trauth.
                                        html;   http://www.acm.org/pubs/

articles/journals/cacm/2001-44-
7/p74-trauth/p74-trauth.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p74-trauth/.

Shan:2001:IWF

[473] Wang Shan. The IT work force in China. Communications of the ACM, 44(7):76, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/2001-44-7/p76-shan/p76-shan.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p76-shan/p76-shan.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-7/p76-shan/.

Anishchanko:2001:ISS

[474] Uladzimir Anishchanko. The IT staffing situation in Belarus. Communications of the ACM, 44(7):77, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-7/p77-anishchanko/
p77-anishchanko.html;   http://
www.acm.org/pubs/articles/journals/
cacm/2001-44-7/p77-anishchanko/
p77-anishchanko.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-7/p77-anishchanko/.

Benford:2001:CVE

[475] Steve Benford, Chris Greenhalgh, Tom Rodden, and James Pycock. Collaborative virtual environments. Communications of the ACM, 44(7): 79, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-7/p79-benford/p79-
                                        benford.html;   http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p79-benford/p79-benford.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        7/p79-benford/.

Hirji:2001:EDM

[476] Karim K. Hirji. Exploring data mining implementation. Communications of the ACM, 44(7):87, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-7/p87-hirji/p87-hirji.html;
                                         http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-7/p87-hirji/
                                        p87-hirji.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-7/p87-hirji/.

Lai:2001:ICI

[477] Vincent S. Lai. Intraorganizational communication with intranets. Communications of the ACM, 44(7):95, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-7/p95-lai/p95-lai.
                                        html;   http://www.acm.org/pubs/
                                        articles/journals/cacm/2001-44-
                                        7/p95-lai/p95-lai.pdf;   http://
                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-7/p95-lai/.

Horning:2001:IRL

[478] Jim Horning. Inside risks: Learning from experience. Communications of the ACM, 44(7):112, July 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-7/p112-horning/p112-horning.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
7/p112-horning/p112-horning.pdf;
http://www.acm.org/pubs/citations/
journals/cacm/2001-44-7/p112-horning/

.

Crawford:2001:EPg

[479] Diane Crawford. Editorial pointers. Communications of the ACM, 44(8): 5, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-8/p5-crawford/p5-crawford.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
8/p5-crawford/p5-crawford.pdf;
http://www.acm.org/pubs/citations/
journals/cacm/2001-44-8/p5-crawford/

.

Fox:2001:NTSb

[480] Robert Fox. News track: Sensing global vibrations; the big four online; gender gap perceptions; portion of the world offline; cat … and mouse; 20 Gigahertz chip 3 atoms thick; Apple motto exorcised. Communications of the ACM, 44(8):9–10, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-8/p9-fox/p9-fox.html;
http://www.acm.org/pubs/articles/
journals/cacm/2001-44-8/p9-fox/
p9-fox.pdf;   http://www.acm.org/

pubs/citations/journals/cacm/2001-
44-8/p9-fox/.

Crawford:2001:FIF

[481] Diane Crawford, Richar Hill, George Ellingham, Lawrence C. Paulson, Adriano Comai, Bill Brykczynski, Saul I. Gass, Ferdinand Wagner, and Peter Wolstenholme. Forum: The ICANN framework better than going to court; drop the adjective [digital library]; keep E-journals affordable; needs and adoption of process; give security accreditation a chance; keeping out cookies; the joy of coding. Communications of the ACM, 44(8):11–13, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p11-crawford/p11-
crawford.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-8/p11-crawford/p11-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
8/p11-crawford/.

Berghel:2001:DYE

[482] Hal Berghel. Digital: The Y2K e-commerce tumble. Communications of the ACM, 44(8):15, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p15-berghel/p15-
berghel.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-8/p15-berghel/p15-berghel.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
8/p15-berghel/.

Denning:2001:PII

[483] Peter J. Denning. The profession of IT: The IT schools movement. Communications of the ACM, 44(8):19, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p19-denning/p19-
denning.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-8/p19-denning/p19-denning.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
8/p19-denning/.

Touretzky:2001:VFS

[484] David S. Touretzky. Viewpoint: Free speech rights for programmers. Communications of the ACM, 44(8):23, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-8/p23-touretzky/p23-touretzky.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
8/p23-touretzky/p23-touretzky.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-8/p23-touretzky/

.

Freedman:2001:VPP

[485] Eric M. Freedman. Viewpoint: Pondering pixelized pixies. Communications of the ACM, 44(8):27, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p27-freedman/p27-
freedman.html;   http://www.acm.
org/pubs/articles/journals/cacm/

                                        2001-44-8/p27-freedman/p27-freedman.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        8/p27-freedman/.

Gershon:2001:WSC

[486] Nahum Gershon and Ward Page. What storytelling can do for information visualization. Communications of the ACM, 44(8):31, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-8/p31-gershon/p31-
                                        gershon.html;   http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-8/p31-gershon/p31-gershon.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        8/p31-gershon/.

Keim:2001:VEL

[487] Daniel A. Keim. Visual exploration of large data sets. Communications of the ACM, 44(8):38, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-8/p38-keim/p38-keim.html;
                                         http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-8/p38-keim/
                                        p38-keim.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-8/p38-keim/.

Eick:2001:VOA

[488] Stephen G. Eick. Visualizing online activity. Communications of the ACM, 44(8):45, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/2001-44-8/p45-eick/p45-eick.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
8/p45-eick/p45-eick.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-8/p45-eick/.

Mintzer:2001:PHM

[489] Fred Mintzer, Gordon W. Braudaway, Francis P. Giordano, Jack C. Lee, Karen A. Magerlein, Silvana D’Auria, Amnon Ribak, Gil Shapir, Fabio Schiattarella, John Tolva, and Andrey Zelenkov. Populating the Hermitage Museum’s new Web site. Communications of the ACM, 44(8):52, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p52-mintzer/p52-
mintzer.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-8/p52-mintzer/p52-mintzer.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
8/p52-mintzer/.

Lam:2001:ICM

[490] Calvin K. M. Lam and Bernard C. Y. Tan. The Internet is changing the music industry. Communications of the ACM, 44(8):62, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p62-lam/p62-lam.
html;   http://www.acm.org/pubs/
articles/journals/cacm/2001-44-
8/p62-lam/p62-lam.pdf;   http://

                                        www.acm.org/pubs/citations/journals/
                                        cacm/2001-44-8/p62-lam/.

Lieberman:2001:EWR

[491] Henry Lieberman, Christopher Fry, and Louis Weitzman. Exploring the Web with reconnaissance agents. Communications of the ACM, 44(8):69, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-8/p69-lieberman/p69-lieberman.
                                        html;   http://www.acm.org/pubs/
                                        articles/journals/cacm/2001-44-
                                        8/p69-lieberman/p69-lieberman.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-8/p69-lieberman/

.

Ye:2001:PCA

[492] Nong Ye, Joseph Giordano, and John Feldman. A process control approach to cyber attack detection. Communications of the ACM, 44(8):76, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-8/p76-ye/p76-ye.html;
                                        http://www.acm.org/pubs/articles/
                                        journals/cacm/2001-44-8/p76-ye/
                                        p76-ye.pdf;   http://www.acm.org/
                                        pubs/citations/journals/cacm/2001-
                                        44-8/p76-ye/.

Benamati:2001:CRC

[493] John Benamati and Albert L. Lederer. Coping with rapid changes in IT. Communications of the ACM, 44(8): 83, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-8/p83-benamati/p83-
benamati.html;   http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-8/p83-benamati/p83-benamati.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
8/p83-benamati/.

Levi:2001:IRR

[494] Albert Levi and Çetin Kaya Koç. Inside risks: Risks in email security. Communications of the ACM, 44(8): 112, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-8/p112-levi/p112-levi.html;
http://www.acm.org/pubs/articles/
journals/cacm/2001-44-8/p112-levi/
p112-levi.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-8/p112-levi/.

Crawford:2001:EPh

[495] Diane Crawford. Editorial pointers. Communications of the ACM, 44(9):5, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-9/p5-crawford/p5-crawford.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
9/p5-crawford/.

Fox:2001:NTT

[496] Robert Fox. News track: Telecommuting loses appeal; privacy issue for DNA database; no copying, no warning; big ‘boss’ is watching; men outperform women in virtual world; cell phone radio waves made public; the mystery of the sucking shower curtain. Communications of the ACM, 44(9):9–10, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-9/p9-fox/p9-fox.pdf;   http:
//www.acm.org/pubs/citations/journals/
cacm/2001-44-9/p9-fox/.

Crawford:2001:FHA

[497] Diane Crawford, Oleg Dulin, David Arnold, Fred Niederman, Ralpha Castain, and Bill Rivera. Forum: How to address the global IT worker shortage; proposals for attracting women back into IT. Communications of the ACM, 44(9):11, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-9/p11-crawford/p11-
crawford.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-9/p11-crawford/.

Glass:2001:PPM

[498] Robert L. Glass. Practical programmer: of model changeovers, style, and fatware. Communications of the ACM, 44(9):17–20, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-9/p17-glass/p17-glass.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-9/p17-glass/

.

McGinity:2001:SCS

[499] Meg McGinity. Staying connected: suit yourself. Communications of the ACM, 44(9):21–24, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/articles/journals/
cacm/2001-44-9/p21-mcginity/p21-
mcginity.pdf;   http://www.acm.
org/pubs/citations/journals/cacm/
2001-44-9/p21-mcginity/.

Ginsberg:2001:VWK

[500] Jane C. Ginsberg. Viewpoint: what to know before reissuing old titles as e-books. Communications of the ACM, 44(9):25, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/citations/journals/
cacm/2001-44-9/p25-ginsberg/.

Mylonopoulos:2001:SGP

[501] Nikolaos A. Mylonopoulos and Vasilis Theoharakis. On site: global perceptions of IS journals. Communications of the ACM, 44(9):29–33, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-9/p29-mylonopoulos/p29-
mylonopoulos.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-9/p29-mylonopoulos/

.

Cerf:2001:BPI

[502] Vinton Cerf. Beyond the post-PC Internet. Communications of the ACM, 44(9):34–37, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-9/p34-cerf/p34-cerf.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-9/p34-cerf/

.

Hooke:2001:II

[503] Adrian Hooke. The interplanetary Internet. Communications of the ACM, 44(9):38–40, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-9/p38-hooke/p38-hooke.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-9/p38-hooke/

.

Kleinrock:2001:BL

[504] Leonard Kleinrock. Breaking loose. Communications of the ACM, 44(9): 41–46, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

                                        www.acm.org/pubs/articles/journals/
                                        cacm/2001-44-9/p41-kleinrock/p41-
                                        kleinrock.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-9/p41-kleinrock/.

Corrales:2001:AVN

[505] José A. Corrales. An Asturian view of networking 2015. Communications of the ACM, 44(9):47–54, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/articles/journals/
                                        cacm/2001-44-9/p47-corrales/p47-
                                        corrales.pdf;   http://www.acm.
                                        org/pubs/citations/journals/cacm/
                                        2001-44-9/p47-corrales/.

Adamic:2001:WHO

[506] Lada A. Adamic and Bernardo A. Huberman. The Web’s hidden order. Communications of the ACM, 44 (9):55–60, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/articles/journals/
cacm/2001-44-9/p55-adamic/p55-adamic.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
9/p55-adamic/.

Weiser:2001:WHN

[507] Mark Weiser. Whatever happened to the next-generation Internet? Communications of the ACM, 44(9):61–69, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-9/p61-weiser/p61-weiser.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
9/p61-weiser/.

Gordon:2001:UIS

[508] Lawrence A. Gordon and Martin P. Loeb. Using information security as a response to competitor analysis systems. Communications of the ACM, 44(9):70–75, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-9/p70-gordon/p70-gordon.
pdf;   http://www.acm.org/pubs/
citations/journals/cacm/2001-44-
9/p70-gordon/.

Chan:2001:WFS

[509] Alvin T. S. Chan, Jiannong Cao, Henry Chan, and Gilbert Young. A Webenabled framework for smart card applications in health services. Communications of the ACM, 44(9):76–82, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-9/p76-chan/p76-chan.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-9/p76-chan/

.

Bell:2001:MRN

[510] Glenn B. Bell and Anil Sethi. Matching records in a national medical patient index. Communications of the ACM, 44(9):83–88, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/articles/journals/cacm/
                                        2001-44-9/p83-bell/p83-bell.pdf;
                                         http://www.acm.org/pubs/citations/
                                        journals/cacm/2001-44-9/p83-bell/

.

Huston:2001:SII

[511] Terry Huston. Security issues for implementation of e-medical records. Communications of the ACM, 44(9):89–94, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

                                        org/pubs/articles/journals/cacm/
                                        2001-44-9/p89-huston/p89-huston.
                                        pdf;   http://www.acm.org/pubs/
                                        citations/journals/cacm/2001-44-
                                        9/p89-huston/.

Fayad:2001:TOI

[512] Mohammed E. Fayad and Adam Altman. Thinking objectively: an introduction to software stability. Communications of the ACM, 44(9):95, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/articles/journals/cacm/
2001-44-9/p95-fayad/p95-fayad.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-9/p95-fayad/

.

Sit:2001:IRW

[513] Emil Sit and Kevin Fu. Inside risks: Web cookies: not just a privacy risk. Communications of the ACM, 44(9):120, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic).  URL http://www.acm.
org/pubs/articles/journals/cacm/
2001-44-9/p120-sit/p120-sit.pdf;
 http://www.acm.org/pubs/citations/
journals/cacm/2001-44-9/p120-sit/

.

Crawford:2001:EPi

[514] Diane Crawford. Editorial pointers. Communications of the ACM, 44(10):5, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2001:NTA

[515] Robert Fox. News track: ACM files declaration; taking a bite out of shark attacks; China Web growth; photomapping technology; simulating Mercury’s wobble; cell phone impact on wildlife; Wile E. Coyote doesn’t stand a chance. Communications of the ACM, 44(10):9–10, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:FHB

[516] Diane Crawford, Ted Mihalsin, William Bogstad, Daniel Kiem, J. William Cupp, and Sandy Ressler. Forum: How big is “big” when it comes to data sets? similar figures for women math majors; commendation. Communications of the ACM, 44(10):11–12, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2001:BSZ

[517] Phillip Armour. The business of software: Zeppelins and jet planes: a metaphor for modern software projects. Communications of the ACM, 44 (10):13–15, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Munro:2001:WCB

[518] Neil Munro. From Washington: considering the broadband debate. Communications of the ACM, 44(10):17– 19, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Devlin:2001:VRR

[519] Keith Devlin. Viewpoint: the real reason why software engineers need math. Communications of the ACM, 44(10):21–22, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simons:2001:VAD

[520] Barbara Simons. Viewpoint: the ACM declaration in Felten v. RIAA. Communications of the ACM, 44(10):23– 26, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elrad:2001:AOP

[521] Tzilla Elrad, Robert E. Filman, and Atef Bader. Aspect-oriented programming: Introduction. Communications of the ACM, 44(10):29–32, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elrad:2001:DAA

[522] Tzilla Elrad, Mehmet Aksits, Gregor Kiczales, Karl Lieberherr, and Harold Ossher. Discussing aspects of AOP. Communications of the ACM, 44(10):33–38, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lieberherr:2001:AOP

[523] Karl Lieberherr, Doug Orleans, and Johan Ovlinger. Aspect-oriented programming with adaptive methods. Communications of the ACM, 44 (10):39–41, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ossher:2001:UMS

[524] Harold Ossher and Peri Tarr. Using multidimensional separation of concerns to (re)shape evolving software. Communications of the ACM, 44(10):43–50, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bergmans:2001:CCC

[525] Lodewijk Bergmans and Mehmet Aksits. Composing crosscutting concerns using composition filters. Communications of the ACM, 44(10):51–57, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kiczales:2001:GSA

[526] Gregor Kiczales, Erik Hilsdale, Jim Hugunin, Mik Kersten, Jeffrey Palm, and William Griswold. Getting started with ASPECTJ. Communications of the ACM, 44(10):59–65, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pace:2001:ARA

[527] J. Andrés Díaz Pace and Marcelo R. Campo. Analyzing the role of aspects in software design. Communications of the ACM, 44(10):66–73, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Murphy:2001:DAO

[528] Gail C. Murphy, Robert J. Walker, Elisa L. A. Baniassad, Martin P. Robillard, Albert Lai, and Mik A. Kersten. Does aspect-oriented programming work? Communications of the ACM, 44(10):75–77, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coady:2001:SOS

[529] Yvonne Coady, Gregor Kiczales, Mike Feeley, Norm Hutchinson, and Joon Suan Ong. Structuring operating system aspects: using AOP to improve OS structure modularity. Communications of the ACM, 44(10):79–82, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Netinant:2001:LAB

[530] Paniti Netinant, Tzilla Elrad, and Mohamed E. Fayad. A layered approach to building open aspect-oriented systems: a framework for the design of on-demand system demodularization. Communications of the ACM, 44(10):83–85, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:2001:HCC

[531] Jeff Gray, Ted Bapty, Sandeep Neema, and James Tuck. Handling crosscutting constraints in domain-specific modeling. Communications of the ACM, 44 (10):87–93, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sullivan:2001:AOP

[532] Gregory T. Sullivan. Aspect-oriented programming using reflection and metaobject protocols. Communications of the ACM, 44(10):95–97, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pancake:2001:HPJ

[533] Cherri Pancake and Christian Lengauer. High-performance Java. Communications of the ACM, 44(10):98–101, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moreira:2001:NP

[534] José E. Moreira, Samuel P. Midkiff, Manish Gupta, Pedro V. Artigas, Peng Wu, and George Almasi. The NINJA project. Communications of the ACM, 44(10):102–109, October 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kielmann:2001:EJH

[535] Thilo Kielmann, Philip Hatcher, Luc Bougé, and Henri E. Bal. Enabling Java for high-performance computing. Communications of the ACM, 44(10): 110–117, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Getov:2001:MCJ

[536] Vladimir Getov, Gregor von Laszewski, Michael Philippsen, and Ian Foster. Multiparadigm communications in Java for grid computing. Communications of the ACM, 44(10):118–125, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledgard:2001:TOE

[537] Henry F. Ledgard. Technical opinion: The emperor with no clothes. Communications of the ACM, 44(10): 126–128, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Westfall:2001:TOH

[538] Ralph Westfall. Technical opinion: Hello, world considered harmful. Communications of the ACM, 44(10):129– 130, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Somogyi:2001:IRP

[539] Stephan Somogyi and Bruce Schneier. Inside risks: The perils of port 80. Communications of the ACM, 44(10):

168, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:EPj

[540] Diane Crawford. Editorial pointers. Communications of the ACM, 44(11):5, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2001:NTNb

[541] Robert Fox. News track: NSA’s data overload slows intelligence; disaster plans, thanks to Y2K; simulating the birth of the moon; looking for a few good hackers; GPS for plane navigation; free medical journal access; music to live by; tracking employee trust. Communications of the ACM, 44(11):9–10, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:FSC

[542] Diane Crawford, Scott E. Hamilton, David Touretzky, Raymond Michiels, David Parnas, Peter Denning, and Alan Creak. Forum: Simple copyright violation or unlawful circumvention? outlawing numbers impractical; defining information; special needs and sound scholarship. Communications of the ACM, 44(11):11–15, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meeks:2001:EFB

[543] Brock N. Meeks. Electronic frontier: Blanking on rebellion: where the future is “Nabster”. Communications of the ACM, 44(11):17–20, November 2001.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2001:PICb

[544] Peter J. Denning and Robert Dunham. The profession of IT: The core of the third-wave professional. Communications of the ACM, 44(11): 21–25, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berman:2001:VTK

[545] Fran Berman. Viewpoint: From TeraGrid to knowledge grid. Communications of the ACM, 44(11):27–28, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:2001:SIU

[546] Albert Huang. On site: Innovative use of email for teaching. Communications of the ACM, 44(11):29–32, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Myers:2001:UHP

[547] Brad A. Myers. Using handhelds and PCs together. Communications of the ACM, 44(11):34–41, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bapna:2001:IAO

[548] Ravi Bapna, Paulo Goes, and Alok Gupta. Insights and analyses of online auctions. Communications of the ACM, 44(11):42–50, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ettredge:2001:WSD

[549] Michael Ettredge, Vernon J. Richardson, and Susan Scholz. A Web site design model for financial information. Communications of the ACM, 44 (11):51–56, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chircu:2001:MVI

[550] Alina M. Chircu, Robert J. Kauffman, and Doug Keskey. Maximizing the value of Internet-based corporate travel reservation systems. Communications of the ACM, 44(11):57–63, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:2001:EC

[551] John Gray. The end of career. Communications of the ACM, 44(11):65–69, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Backus:2001:FCR

[552] John Backus. Funding the computing revolution’s third wave. Communications of the ACM, 44(11):70–76, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carriero:2001:CME

[553] Nicholas Carriero and David Gelernter. A computational model of everything. Communications of the ACM, 44 (11):77–81, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burgin:2001:HWK

[554] Mark Burgin. How we know what technology can do. Communications of the ACM, 44(11):82–88, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lucas:2001:ITP

[555] Henry C. Lucas. Information technology and physical space. Communications of the ACM, 44(11):89–96, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sawyer:2001:MBP

[556] Steve Sawyer. A market-based perspective on information systems development. Communications of the ACM, 44 (11):97–102, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2001:IRR

[557] Lauren Weinstein and Peter G. Neumann. Inside risks: Risks of panic. Communications of the ACM, 44 (11):152, November 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:EPk

[558] Diane Crawford. Editorial pointers. Communications of the ACM, 44(12):5, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2001:NTP

[559] Robert Fox. News track: Privacy tradeoff fighting terrorism; information removal; military-issue handhelds; haptic tug-of-war; turning back 10 billion (Web) pages of time; brainy baboons; grunt work. Communications of the ACM, 44(12):9–10, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2001:FOS

[560] Diane Crawford, Scott W. Ambler, John Jaros, Jim Highsmith, Kurt Guntheroth, Eric A. Weiss, and Steven Pothier. Forum: One sure thing if good clean code; exposing agency injustices; targeting software project goals; taking a stand on copyright. Communications of the ACM, 44(12):11–13, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2001:DVCc

[561] Hal Berghel. Digital village: The Code Red worm. Communications of the ACM, 44(12):15–19, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schank:2001:LER

[562] Roger C. Schank. Log on education: Revolutionizing the traditional classroom course. Communications of the ACM, 44(12):21–24, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yee:2001:SOE

[563] Ka-Ping Yee. On site: Operating an emergency information service. Communications of the ACM, 44 (12):25–28, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peha:2001:SGD

[564] Jon M. Peha. On site: The growing debate over science and technology advice for Congress. Communications of the ACM, 44(12):29–31, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phoha:2001:VDN

[565] Vir V. Phoha. Viewpoint: The DMCA needs fixing. Communications of the ACM, 44(12):33–34, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maamar:2001:TOT

[566] Zakaria Maamar, Eric Dorion, and Catherine Daigle. Technical opinion: Toward virtual marketplaces for Ecommerce support. Communications of the ACM, 44(12):35–38, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ragusa:2001:CVD

[567] James M. Ragusa and Grace M. Bochenek. Collaborative virtual design environments: Introduction. Communications of the ACM, 44(12):40–43, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2001:SV

[568] Randall C. Smith. Shared vision. Communications of the ACM, 44(12): 45–48, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loftin:2001:DEV

[569] R. Bowen Loftin. Design engineering in virtual environments. Communications of the ACM, 44(12):49–50, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maybury:2001:CVE

[570] Mark Maybury. Collaborative virtual environments for analysis and decision support. Communications of the ACM, 44(12):51–54, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maybury:2001:EFC

[571] Mark Maybury, Ray D’Amore, and David House. Expert finding for collaborative virtual environments. Communications of the ACM, 44(12):55–56, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Su:2001:SVE

[572] Simon Su and R. Bowen Loftin. A shared virtual environment for exploring and designing molecules. Communications of the ACM, 44(12):57–58, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horvath:2001:CSC

[573] Imre Horváth and Zoltán Rusák. Collaborative shape conceptualization in virtual design environments. Communications of the ACM, 44(12):59–63, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Luciano:2001:ATA

[574] Cristian Luciano, Pat Banerjee, and Sanjay Mehrotra. 3D animation of telecollaborative anthropomorphic avatars. Communications of the ACM, 44(12):64–67, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davison:2001:GAC

[575] Robert Davison and Gert-Jan De Vreede. Global applications of collaborative technology: Introduction. Communications of the ACM, 44(12): 68–70, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dube:2001:GVT

[576] Line Dubé and Guy Paré. Global virtual teams. Communications of the ACM, 44(12):71–73, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Newell:2001:MBO

[577] Sue Newell, Shan L. Pan, Robert D. Galliers, and Jimmy C. Huang. The myth of the boundaryless organization. Communications of the ACM, 44 (12):74–76, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kelly:2001:GSI

[578] Séamas Kelly and Matthew Jones. Groupware and the social infrastructure of communication. Communications of the ACM, 44(12):77–79, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeSanctis:2001:BGL

[579] Gerardine DeSanctis, Matthew Wright, and Lu Jiang. Building a global learning community. Communications of the ACM, 44(12):80–82, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Massey:2001:CPT

[580] Anne P. Massey, Mitzi Montoya-Weiss, Caisy Hung, and V. Ramesh. Cultural perceptions of task-technology fit. Communications of the ACM, 44 (12):83–84, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Qureshi:2001:PPG

[581] Sajda Qureshi and Ilze Zigurs. Paradoxes and prerogatives in global virtual collaboration. Communications of the ACM, 44(12):85–88, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evaristo:2001:NND

[582] Roberto Evaristo. Nonconsensual negotiation in distributed collaboration. Communications of the ACM, 44(12):89, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gopal:2001:ANH

[583] Ram D. Gopal, Zhiping Walter, and Arvind K. Tripathi. Admediation: New horizons in effective email advertising. Communications of the ACM, 44 (12):91–96, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glover:2001:WSY

[584] Eric J. Glover, Steve Lawrence, Michael D. Gordon, William P. Birmingham, and C. Lee Giles. Web search—your way. Communications of the ACM, 44(12):97–102, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhao:2001:IPT

[585] J. Leon Zhao and Vincent H. Resh. Internet publishing and transformation of knowledge processes. Communications of the ACM, 44(12):103–109, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2001:IRR

[586] Peter G. Neumann and Lauren Weinstein. Inside risks: Risks of national identity cards. Communications of the ACM, 44(12):176, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPa

[587] Diane Crawford. Editorial pointers. Communications of the ACM, 45(1): 5, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTW

[588] Robert Fox. News track: Working shadow; hold that thought; access advantage; tuned out or in?; strength in numbers; research salaries; bye big blues; sidebar: Hooked U. Communications of the ACM, 45(1):9–10, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2002:BSS

[589] Phillip G. Armour. The business of software: The spiritual life of projects. Communications of the ACM, 45(1):11–14, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Munro:2002:WTM

[590] Neil Munro. From Washington: Too much of a good thing. Communications of the ACM, 45(1):15–17, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bartholome:2002:SPA

[591] Lloyd W. Bartholome and David H. Olsen. On site: A practical approach for implementing e-commerce programs in business schools. Communications of the ACM, 45(1):19–21, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Townsend:2002:VMV

[592] Anthony M. Townsend, Anthony R. Hendrickson, and Samuel M. DeMarie. Viewpoint: Meeting the virtual work imperative. Communications of the ACM, 45(1):23–26, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewis:2002:GES

[593] Michael Lewis and Jeffrey Jacobson. Game engines in scientific research: Introduction. Communications of the ACM, 45(1):27–31, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewis:2002:NC

[594] Michael Lewis. The new cards. Communications of the ACM, 45(1):30–31, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Laird:2002:RHL

[595] John E. Laird. Research in human-level AI using computer games. Communications of the ACM, 45(1):32–35, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Piekarski:2002:AOA

[596] Wayne Piekarski and Bruce Thomas. ARQuake: the outdoor augmented reality gaming system. Communications of the ACM, 45(1):36–38, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jacobson:2002:UTI

[597] Jeffrey Jacobson and Zimmy Hwang. Unreal tournament for immersive interactive theater. Communications of the ACM, 45(1):39–42, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaminka:2002:GFT

[598] Gal A. Kaminka, Manuela M. Veloso, Steve Schaffer, Chris Sollitto, Rogelio Adobbati, Andrew N. Marshall, Andrew Scholer, and Sheila Tejada. GameBots: a flexible test bed for multiagent team research. Communications of the ACM, 45(1):43–45, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bylund:2002:TDC

[599] Markus Bylund and Fredrik Espinoza. Testing and demonstrating contextaware services with Quake III Arena. Communications of the ACM, 45 (1):46–48, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2002:AAR

[600] CACM Staff. The ACM annual report FY01. Communications of the ACM, 45(1):49–52, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anandarajan:2002:IAW

[601] Murugan Anandarajan. Internet abuse in the workplace: Introduction. Communications of the ACM, 45(1):53–54, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stanton:2002:CPF

[602] Jeffrey M. Stanton. Company profile of the frequent Internet user. Communications of the ACM, 45(1):55–59, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oravec:2002:CAI

[603] Jo Ann Oravec. Constructive approaches to Internet recreation in the workplace. Communications of the ACM, 45(1):60–63, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Belanger:2002:AL

[604] France Belanger and Craig Van Slyke. Abuse or learning? Communications of the ACM, 45(1):64–65, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lim:2002:HDL

[605] Vivien K. G. Lim, Thompson S. H. Teo, and Geok Leng Loo. How do I loaf here? let me count the ways. Communications of the ACM, 45(1):66–70, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simmers:2002:AIU

[606] Claire A. Simmers. Aligning Internet usage with business priorities. Communications of the ACM, 45(1):71–74, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siau:2002:AIU

[607] Keng Siau, Fiona Fui-Hoon Nah, and Limei Teng. Acceptable Internet use policy. Communications of the ACM, 45(1):75–79, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Urbaczewski:2002:DEM

[608] Andrew Urbaczewski and Leonard M. Jessup. Does electronic monitoring of employee Internet usage work? Communications of the ACM, 45(1):80–83, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Panko:2002:MPS

[609] Raymond R. Panko and Hazel Glenn Beh. Monitoring for pornography and sexual harassment. Communications of the ACM, 45(1):84–87, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varshney:2002:VI

[610] Upkar Varshney, Andy Snow, Matt Mc-Givern, and Christi Howard. Voice over IP. Communications of the ACM, 45(1):89–96, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Govindarajulu:2002:SHS

[611] Chittibabu Govindarajulu. The status of helpdesk support. Communications of the ACM, 45(1):97–100, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

David:2002:MYT

[612] Julie Smith David, David Schuff, and Robert St. Louis. Managing your total IT cost of ownership. Communications of the ACM, 45(1):101–106, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kobryn:2002:TOW

[613] Cris Kobryn. Technical opinion: Will UML 2.0 be agile or awkward? Communications of the ACM, 45(1):107– 110, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayad:2002:TOA

[614] Mohamed Fayad. Thinking objectively: Accomplishing software stability. Communications of the ACM, 45 (1):111–115, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Filman:2002:IIC

[615] Robert E. Filman, Stuart Barrett, Diana D. Lee, and Ted Linden. Inserting ilities by controlling communications. Communications of the ACM, 45(1):116–122, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2002:IRU

[616] Rebecca Mercuri. Inside risks: Uncommon criteria. Communications of the ACM, 45(1):172, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPb

[617] Diane Crawford. Editorial pointers. Communications of the ACM, 45(2):5–6, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTBa

[618] Robert Fox. News track: A blessing and a curse; jacked-in jersey; not just music; infectious glow; all-seeing; on your mark; brain waves; diabolical degree. Communications of the ACM, 45(2):9–10, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FHW

[619] Diane Crawford, Byron Weber Becker, Jaime Nino, Hugh McGuire, Gil Colgate, Mark Lutton, Vladimir Safonov, Ralph Westfall, Bradley K. Sherman, Ed Lowry, and Kurt Guntheroth. Forum: ‘hello, world’ gets mixed greetings; Westfall responds; the Emperor’s wardrobe; targeting zeppelins. Communications of the ACM, 45(2):11–17, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grosso:2002:LSW

[620] Andrew Grosso. Legally speaking: Why the Digital Millennium Copyright Act is a failure of reason. Communications of the ACM, 45(2):19–23, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2002:PPSa

[621] Robert L. Glass. Practical programmer: Stodgy by design, and the notion of ‘dumbing up’. Communications of the ACM, 45(2):25–27, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Castel:2002:OC

[622] Felipe Castel. Ontological computing. Communications of the ACM, 45 (2):29–30, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:2002:VAC

[623] Ben Shneiderman. Viewpoint: ACM’s computing professionals face new challenges. Communications of the ACM, 45(2):31–34, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2002:SCC

[624] Qing Bian Zhang and Patrick Y. K. Chau. On site: Creating e-commerce courses with regional intent. Communications of the ACM, 45(2):35–37, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gruninger:2002:OAD

[625] Michael Gruninger and Jintae Lee. Ontology applications and design: Introduction. Communications of the ACM, 45(2):39–41, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holsapple:2002:CAO

[626] Clyde W. Holsapple and K. D. Joshi. A collaborative approach to ontology design. Communications of the ACM, 45(2):42–47, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2002:PHO

[627] Henry Kim. Predicting how ontologies for the semantic Web will evolve. Communications of the ACM, 45(2):48–54, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Everett:2002:MOW

[628] John O. Everett, Daniel G. Bobrow, Reinhard Stolle, Richard Crouch, Valeria de Paiva, Cleo Condoravdi, Martin van den Berg, and Livia Polanyi. Making ontologies work for resolving redundancies across documents. Communications of the ACM, 45(2):55–60, February 2002. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Guarino:2002:EOD

[629] Nicola Guarino and Christopher Welty. Evaluating ontological decisions with OntoClean. Communications of the ACM, 45(2):61–65, February 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dempsey:2002:WOS

[630] Bert J. Dempsey, Debra Weiss, Paul Jones, and Jane Greenberg. Who is an open source software developer? Communications of the ACM, 45(2):67–72, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moore:2002:HTA

[631] Jo Ellen Moore and Lisa A. Burke. How to turn around ‘turnover culture’ in IT. Communications of the ACM, 45(2):73–78, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kettinger:2002:UUD

[632] William J. Kettinger and Choong C. Lee. Understanding the IS-user divide in IT innovation. Communications of the ACM, 45(2):79–84, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grant:2002:WVB

[633] Delvin Grant. A wider view of business process reengineering. Communications of the ACM, 45(2):85–90, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2002:WNH

[634] Gordon Bell and Jim Gray. What’s next in high-performance computing? Communications of the ACM, 45(2): 91–95, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Snodgrass:2002:RRA

[635] Richard T. Snodgrass. Rights and responsibilities in ACM publishing: Progress on ACM’s becoming the preferred publisher. Communications of the ACM, 45(2):97–101, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2002:TOS

[636] Thomas Erickson. Technical opinion: Some problems with the notion of context-aware computing. Communications of the ACM, 45(2):102–104, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gabrilovich:2002:IRH

[637] Evgeniy Gabrilovich and Alex Gontmakher. Inside risks: The homograph attack. Communications of the ACM, 45(2):128, February 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPc

[638] Diane Crawford. Editorial pointers. Communications of the ACM, 45(3):5, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTBb

[639] Robert Fox. News track: Boosting military tech spending; press ‘eject’ button and call security; identity tops theft list; Pashto-phonics; talk to the (gloved) hand; winning over girls on the gadget front; A robot to do a nanocrab’s work; when Web surfing turns to Googlewhacking. Communications of the ACM, 45(3):9–10, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FDH

[640] Diane Crawford, Richard Staehli, Donald J. Reifer, Anonymous, William L. Anderson, Atanas Radenski, Rick Simkin, Ram D. Gopal, Zhiping Walter, Arvind K. Tripathi, and Phil Armour. Forum: Doubts and hopes for AOP; social and technical aspects of information systems design and deployment; classroom teaching vs. online courses; no unsolicited email ads for me, please; A chilling virus. Communications of the ACM, 45(3):11–14, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2002:IT

[641] Peter Denning. Internet time out. Communications of the ACM, 45 (3):15–18, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2002:CIT

[642] Meg McGinity. Call it techonomics. Communications of the ACM, 45 (3):19–21, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

SCCNPP:2002:SSP

[643] The Stanford Student Computer and Network Privacy Project. A study of student privacy issues at Stanford University. Communications of the ACM, 45(3):23–25, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mitra:2002:TAD

[644] Ananda Mitra. Trust, authenticity, and discursive power in cyberspace. Communications of the ACM, 45(3):27–29, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sukhatme:2002:RIV

[645] Gaurav S. Sukhatme and Maja J. Matarić. Robots: Intelligence, versatility, adaptivity. Communications of the ACM, 45(3):30–32, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brooks:2002:HR

[646] Rodney Brooks. Humanoid robots. Communications of the ACM, 45 (3):33–38, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rus:2002:SRR

[647] Daniela Rus, Zack Butler, Keith Kotay, and Marsette Vona. Self-reconfiguring robots. Communications of the ACM, 45(3):39–45, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khatib:2002:RIS

[648] Oussama Khatib, Oliver Brock, Kyong-Sok Chang, François Conti, Diego Ruspini, and Luis Sentis. Robotics and interactive simulation. Communications of the ACM, 45(3):46–51, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thrun:2002:PR

[649] Sebastian Thrun. Probabilistic robotics. Communications of the ACM, 45(3):52–57, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Veloso:2002:ER

[650] Manuela M. Veloso. Entertainment robotics. Communications of the ACM, 45(3):59–63, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grenci:2002:SDM

[651] Richard T. Grenci and Peter A. Todd. Solutions-driven marketing. Communications of the ACM, 45(3):64–71, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cameron:2002:CDP

[652] John Cameron. Configurable development processes. Communications of the ACM, 45(3):72–77, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Palen:2002:MTC

[653] Leysia Palen. Mobile telephony in a connected life. Communications of the ACM, 45(3):78–82, March 2002.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2002:IAF

[654] Huaiqing Wang, John Mylopoulos, and Stephen Liao. Intelligent agents and financial risk monitoring systems. Communications of the ACM, 45(3):83–88, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lai:2002:MID

[655] Vincent S. Lai and Wingyan Chung. Managing international data communications. Communications of the ACM, 45(3):89–93, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bajcsy:2002:CSS

[656] Ruzena Bajcsy and Craig Reynolds. Computer science: the science of and about information and computation. Communications of the ACM, 45(3):94–98, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2002:IRR

[657] Peter J. Denning and James J. Horning. Inside risks: Risks of linear thinking. Communications of the ACM, 45(3):120, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPd

[658] Diane Crawford. Editorial pointers. Communications of the ACM, 45(4):5, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTSa

[659] Robert Fox. News track: Security boom; line-of-sight laser communication; toxic tech exports; Internet not ready for public elections; chip implant for positive ID; clicking overseas to the movies; Finland? how about Nokialand? Communications of the ACM, 45(4):9–10, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:F

[660] Diane Crawford, Edward J. Beroset, John Jaros, C. J. Date, Robert Plotkin, Rick Simkin, Thomas W. Moran, Hugh Kawabata, Neil Munro, Jason Funk, Pal Valckenaers, Bob Rinker, Brock Meeks, Dennis de Champeaux, John A. Wills, and Gabriel Mateescu. Forum: Don’t take civil liberties in the name of national security; spare me the paranoia; taking responsibility for worms and viruses; who’s research is unethical?; dual parallelism in the Grid. Communications of the ACM, 45(4):11–16, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guzdial:2002:TNG

[661] Mark Guzdial and Elliot Soloway. Log on education: Teaching the Nintendo generation to program. Communications of the ACM, 45(4):17–21, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2002:HW

[662] Hal Berghel. Digital village: Hijacking the Web: Cookies revisited: Continuing the dialogue on personal security and underlying privacy issues. Communications of the ACM, 45(4):23–27, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Turoff:2002:SPF

[663] Murray Turoff. On site: Past and future emergency response information systems. Communications of the ACM, 45(4):29–32, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davenport:2002:VAI

[664] David Davenport. Viewpoint: Anonymity on the Internet: why the price may be too high. Communications of the ACM, 45(4):33–35, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Preece:2002:ISC

[665] Jenny Preece. Introduction: Supporting community and building social capital. Communications of the ACM, 45(4):37–39, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2002:STD

[666] Thomas Erickson, Christine Halverson, Wendy A. Kellogg, Mark Laff, and Tracee Wolf. Social translucence: designing social infrastructures that make collective activity visible. Communications of the ACM, 45(4):40–44, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Donath:2002:SAV

[667] Judith Donath. A semantic approach to visualizing online conversations. Communications of the ACM, 45(4):45–49, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2002:TNL

[668] Marc Smith. Tools for navigating large social cyberspaces. Communications of the ACM, 45(4):51–55, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hiltz:2002:WML

[669] Starr Roxanne Hiltz and Murray Turoff. What makes learning networks effective? Communications of the ACM, 45(4):56–59, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruckman:2002:FLC

[670] Amy Bruckman. The future of elearning communities. Communications of the ACM, 45(4):60–63, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andrews:2002:ASO

[671] Dorine C. Andrews. Audience-specific online community design. Communications of the ACM, 45(4):64–68, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Millen:2002:UBC

[672] David R. Millen, Michael A. Fontaine, and Michael J. Muller. Understanding the benefit and costs of communities of practice. Communications of the ACM, 45(4):69–73, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scott:2002:MRE

[673] Judy E. Scott and Iris Vessey. Managing risks in enterprise systems implementations. Communications of the ACM, 45(4):74–81, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schambach:2002:PDC

[674] Thomas Schambach and J. Ellis Blanton. The professional development challenge for IT professionals. Communications of the ACM, 45(4):83–87, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nardi:2002:ICI

[675] Bonnie A. Nardi, Steve Whittaker, Ellen Isaacs, Mike Creech, Jeff Johnson, and John Hainsworth. Integrating communication and information through ContactMap. Communications of the ACM, 45(4):89–95, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ben-Menachem:2002:IIM

[676] Mordechai Ben-Menachem and Roy Gelbard. Integrated IT management tool kit. Communications of the ACM, 45(4):96–102, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weber:2002:ECH

[677] Arnd Weber. Enabling crypto: how radical innovations occur. Communications of the ACM, 45(4):103–107, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayad:2002:TOH

[678] Mohamed E. Fayad. Thinking objectively: How to deal with software stability. Communications of the ACM, 45(4):109–112, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stringer-Calvert:2002:DE

[679] David W. J. Stringer-Calvert. Inside risks: Digital evidence. Communications of the ACM, 45(4):128, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boulet:2002:CTD

[680] Marie-Michèle Boulet, Faouzi Ben Jebara, Fathi Bemmira, and Serge Boudreault. A comparison of three delivery systems for teaching an information technology course. Communications of the ACM, 45(4ve): 129–135, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p129-boulet.

pdf. Web only: not in printed issue.

Devedzic:2002:UOE

[681] Vladan Devedžić. Understanding ontological engineering. Communications of the ACM, 45(4ve): 136–144, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p136-devedzic.

pdf. Web only: not in printed issue.

Georgiou:2002:SAW

[682] Christos J. Georgiou and Petros S. Stefaneas. Strategies for accelerating the worldwide adoption of E-commerce. Communications of the ACM, 45 (4ve):145–151, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p145-georgiou.

pdf. Web only: not in printed issue.

Hendrix:2002:NTP

[683] T. Dean Hendrix and Michelle P. Schneider. NASA’s TReK Project: A case study in using the spiral model of software development. Communications of the ACM, 45 (4ve):152–159, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p152-hendrix.

pdf. Web only: not in printed issue.

Hibbard:2002:JDO

[684] William Hibbard, Curtis Rueden, Steve Emmerson, Tom Rink, David Glowacki, Tom Whittaker, Don Murray, David Fulker, and John Anderson. Java distributed objects for numerical visualization in VisAD. Communications of the ACM, 45 (4ve):160–170, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p160-hibbard.

pdf. Web only: not in printed issue.

Robertson:2002:MCA

[685] Gretchen L. Robertson and Deborah Hix. Making the computer accessible to mentally retarded adults. Communications of the ACM, 45 (4ve):171–183, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p171-hix.pdf.

Web only: not in printed issue.

Kahn:2002:IQB

[686] Beverly K. Kahn, Diane M. Strong, and Richard Y. Wang. Information quality benchmarks: Product and service performance. Communications of the ACM, 45(4ve): 184–192, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p184-kahn.pdf.

Web only: not in printed issue.

Gopal:2002:RSP

[687] Anandasivam Gopal, Tridas Mukhopadhyay, and Mayuram S. Krishnan. The role of software processes and communication in offshore software development. Communications of the ACM, 45(4ve):193–200, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p193-mukhopadhyay.

pdf. Web only: not in printed issue.

Patnayakuni:2002:DIC

[688] Ravi Patnayakuni and Arun Rai. Development infrastructure characteristics and process capability. Communications of the ACM, 45(4ve): 201–210, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/portal.
acm.org/cacm/ve2002/p201-patnayakuni.

pdf. Web only: not in printed issue.

Pipino:2002:DQA

[689] Leo L. Pipino, Yang W. Lee, and Richard Y. Wang. Data quality assessment. Communications of the ACM, 45(4ve):211–218, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org/cacm/0402/portal.
                                        acm.org/cacm/ve2002/p211-pipino.

pdf. Web only: not in printed issue.

Zviran:2002:SPA

[690] Moshe Zviran. Securing PC applications: The relay race approach. Communications of the ACM, 45 (4ve):219–227, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

                                        //www.acm.org/cacm/0402/portal.
                                        acm.org/cacm/ve2002/p219-zviran.

pdf. Web only: not in printed issue.

Crawford:2002:EPe

[691] Diane Crawford. Editorial pointers. Communications of the ACM, 45(5):5, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTP

[692] Robert Fox. News track: Part cyborg, part human; airline (Web) revenue soars; cutting work privileges; slight change in high-tech hirings; monkey think, monkey do; campus fileswapping truce; love found, love lost. Communications of the ACM, 45(5): 9–10, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FSS

[693] Diane Crawford, Steven Warshawsky, Kamesh Balasubramanian, H. Paul Zeiger, Richard Ho, David Cymbala, John C. Nash, Bertrand Meyer, and Mordechai Ben-Menachem. Forum: The spirited sides of a story; strength in stability; total picture; A readable plea; the more things change. … Communications of the ACM, 45(5):11–13, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2002:SHG

[694] Robert L. Glass. Searching for the holy grail of software engineering. Communications of the ACM, 45(5):15–16, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2002:OMP

[695] Phillip G. Armour. The organism and the mechanism of projects. Communications of the ACM, 45(5):17–20, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bharati:2002:GPJ

[696] Pratyush Bharati and Peter Tarasewich. Global perceptions of journals publishing e-commerce research. Communications of the ACM, 45(5):21–26, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2002:BEM

[697] Wei-Lung Wang. Beware the engineering metaphor. Communications of the ACM, 45(5):27–29, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brusilovsky:2002:AHA

[698] Peter Brusilovsky and Mark T. Maybury. From adaptive hypermedia to the adaptive Web. Communications of the ACM, 45(5):30–33, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Billsus:2002:AIU

[699] Daniel Billsus, Clifford A. Brunk, Craig Evans, Brian Gladish, and Michael Pazzani. Adaptive interfaces for ubiquitous Web access. Communications of the ACM, 45(5):34–38, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smyth:2002:PTO

[700] Barry Smyth, Keith Bradley, and Rachael Rafter. Personalization techniques for online recruitment services. Communications of the ACM, 45(5): 39–40, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fink:2002:PPP

[701] Josef Fink, Jürgen Koenemann, Stephan Noller, and Ingo Schwab. Putting personalization into practice. Communications of the ACM, 45(5): 41–42, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andre:2002:AHP

[702] Elisabeth André and Thomas Rist. From adaptive hypertext to personalized Web companions. Communications of the ACM, 45(5):43–46, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheverst:2002:RAH

[703] Keith Cheverst, Keith Mitchell, and Nigel Davies. The role of adaptive hypermedia in a context-aware tourist GUIDE. Communications of the ACM, 45(5):47–51, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ardissono:2002:PBC

[704] Liliana Ardissono, Anna Goy, Giovanna Petrone, and Marino Segnan. Personalization in business-to-customer interaction. Communications of the ACM, 45(5):52–53, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Light:2002:PMI

[705] Marc Light and Mark T. Maybury. Personalized multimedia information access. Communications of the ACM, 45(5):54–59, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bra:2002:AEH

[706] Paul De Bra. Adaptive educational hypermedia on the Web. Communications of the ACM, 45(5):60–61, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bental:2002:PAS

[707] Diana Bental and Alison Cawsey. Personalized and adaptive systems for medical consumer applications. Communications of the ACM, 45(5):62–63, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kobsa:2002:PHI

[708] Alfred Kobsa. Personalized hypermedia and international privacy. Communications of the ACM, 45(5):64–67, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pinker:2002:STO

[709] Edieal J. Pinker, Abraham Seidmann, and Reginald C. Foster. Strategies for transitioning ‘old economy’ firms to e-business. Communications of the ACM, 45(5):76–83, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:2002:ICA

[710] Christian Wagner and Efraim Turban. Are intelligent e-commerce agents partners or predators? Communications of the ACM, 45(5):84–90, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wellman:2002:DIN

[711] Barry Wellman. Designing the Internet for a networked society. Communications of the ACM, 45(5):91–96, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Helly:2002:CPD

[712] John J. Helly, T. Todd Elvins, Don Sutton, David Martinez, Scott E. Miller, Steward Pickett, and Aaron M. Ellison. Controlled publication of digital scientific data. Communications of the ACM, 45(5):97–101, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kock:2002:MWB

[713] Ned Kock. Managing with Webbased IT in mind. Communications of the ACM, 45(5):102–106, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheung:2002:TAB

[714] Michael Tow Cheung and Ziqi Liao. Time-asymmetry in business processes. Communications of the ACM, 45 (5):107–108, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2002:IRR

[715] Lauren Weinstein. Inside risks: Risks of inaction. Communications of the ACM, 45(5):120, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPf

[716] Diane Crawford. Editorial pointers. Communications of the ACM, 45(6):5, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTF

[717] Robert Fox. News track: Floating transponders; gonzo journalism redefined; China second in Web use; household appliances, lend me your ears; GPS as prison guard; newest hack: Blinking lights; its name is ASIMO; beware net spirituality. Communications of the ACM, 45(6):9–10, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FWN

[718] Diane Crawford, Peter Magsig, Marc Shapiro, Paul Zeiger, Pete McDougall, Earl Furman, James Bach, Sandy Ressler, and Daniel Savarese. Forum: Weighing in on the notion of ‘dumbing up’. Communications of the ACM, 45(6):11–13, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2002:PIF

[719] Peter J. Denning. The profession of IT: Flatlined. Communications of the ACM, 45(6):15–19, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2002:SCB

[720] Meg McGinity. Staying connected: Broadband to go. Communications of the ACM, 45(6):21–24, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lawson:2002:VRC

[721] Harold W. Lawson. Viewpoint: Rebirth of the computer industry. Communications of the ACM, 45(6):25–29, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agha:2002:IAM

[722] Gul A. Agha. Introduction: Adaptive middleware. Communications of the ACM, 45(6):30–32, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kon:2002:CRM

[723] Fabio Kon, Fabio Costa, Gordon Blair, and Roy H. Campbell. The case for reflective middleware. Communications of the ACM, 45(6):33–38, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tripathi:2002:CDN

[724] Anand Tripathi. Challenges designing next-generation middleware systems. Communications of the ACM, 45(6):39–42, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmidt:2002:MRT

[725] Douglas C. Schmidt. Middleware for real-time and embedded systems. Communications of the ACM, 45(6):43–48, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Venkatasubramanian:2002:SCM

[726] Nalini Venkatasubramanian. Safe ‘composability’ of middleware services. Communications of the ACM, 45(6): 49–52, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Issarny:2002:SAD

[727] Valérie Issarny, Christos Kloukinas, and Apostolos Zarras. Systematic aid for developing middleware architectures. Communications of the ACM, 45(6):53–58, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bacon:2002:TOS

[728] Jean Bacon and Ken Moody. Toward open, secure, widely distributed services. Communications of the ACM, 45(6):59–64, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:AF

[729] Diane Crawford. ACM fellows. Communications of the ACM, 45(6):65, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ettredge:2002:TFR

[730] Michael Ettredge, Vernon J. Richardson, and Susan Scholz. Timely financial reporting at corporate Web sites? Communications of the ACM, 45(6):67–71, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hui:2002:CDP

[731] Kai Lung Hui and Patrick Y. K. Chau. Classifying digital products. Communications of the ACM, 45(6):73–79, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klein:2002:WPT

[732] Gary Klein, James J. Jiang, and Debbie B. Tesch. Wanted: project teams with a blend of is professional orientations. Communications of the ACM, 45(6):81–87, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mark:2002:EC

[733] Gloria Mark. Extreme collaboration. Communications of the ACM, 45(6):89–93, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Benbunan-Fich:2002:IET

[734] Raquel Benbunan-Fich. Improving education and training with IT. Communications of the ACM, 45(6):94–99, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aries:2002:CPA

[735] James A. Aries, Subhankar Banerjee, Marc S. Brittan, Eric Dillon, Janusz S. Kowalik, and John P. Lixvar. Capacity and performance analysis of distributed enterprise systems. Communications of the ACM, 45(6):100–105, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gelsinger:2002:TOP

[736] Patrick P. Gelsinger. Technical opinion: Power play. Communications of the ACM, 45(6):106, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2002:FSO

[737] Ross Anderson. Inside risks: Free speech online and offline. Communications of the ACM, 45(6):120, June 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPg

[738] Diane Crawford. Editorial pointers. Communications of the ACM, 45(7):5, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTE

[739] Robert Fox. News track: Eye on sensitive student visas; E-tooling government IT recruiting; remote control ratbots; more hack attacks in Israel; felttipped pen mightier than CD copy protection; need a screen? paint one; where high-tech CEOs see growth; all thumbs in Japan. Communications of the ACM, 45(7):9–10, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FPW

[740] Diane Crawford, Aram Gharib, Ken Kahn, B. Jagannathan, Richard Stallman, and Jim Densmore. Forum: In praise of work’s human dimension; make it visible, tangible, graphical; justifying anonymity; free software reality v. deception; already maxed out. Communications of the ACM, 45(7):11–12, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meeks:2002:EFT

[741] Brock N. Meeks. Electronic frontier: True blue and vigilante, too. Communications of the ACM, 45(7):13–15, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Munro:2002:WEE

[742] Neil Munro. From Washington: The ever-expanding network of local and federal databases. Communications of the ACM, 45(7):17–19, July 2002.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simonson:2002:SPB

[743] Shai Simonson. On site: A postbaccalaureate undergraduate-level program in computer science. Communications of the ACM, 45(7):21–24, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martinsons:2002:VRV

[744] Maris G. Martinsons and Valdis Martinsons. Viewpoint: Rethinking the value of IT, again. Communications of the ACM, 45(7):25–26, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2002:HVI

[745] Andrew Rosenbloom. How the virtual inspires the real. Communications of the ACM, 45(7):28–31, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hillis:2002:PSW

[746] W. Daniel Hillis. The power to shape the world. Communications of the ACM, 45(7):32–35, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2002:RSA

[747] Alvy Ray Smith. The reality of simulated actors. Communications of the ACM, 45(7):36–39, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rhyne:2002:CGS

[748] Theresa-Marie Rhyne. Computer games and scientific visualization.

Communications of the ACM, 45(7): 40–44, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:2002:ISV

[749] Clemens Wagner, Markus Schill, and Reinhard Männer. Intraocular surgery on a virtual eye. Communications of the ACM, 45(7):45–49, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pollefeys:2002:IM

[750] Marc Pollefeys and Luc Van Gool. From images to 3D models. Communications of the ACM, 45(7):50–55, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Badler:2002:VHV

[751] Norman I. Badler, Charles A. Erignac, and Ying Liu. Virtual humans for validating maintenance procedures. Communications of the ACM, 45(7):56–63, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Billinghurst:2002:CAR

[752] Mark Billinghurst and Hirokazu Kato. Collaborative augmented reality. Communications of the ACM, 45(7):64–70, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2002:CHM

[753] Gordon Bell and Jim Gemmell. A call for the home media network. Communications of the ACM, 45(7):71–75, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ishida:2002:DCK

[754] Toru Ishida. Digital city Kyoto. Communications of the ACM, 45(7):76–81, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shelfer:2002:SCE

[755] Katherine M. Shelfer and J. Drew Procaccino. Smart card evolution. Communications of the ACM, 45(7):83–88, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gallaugher:2002:CUD

[756] John M. Gallaugher. E-commerce and the undulating distribution channel. Communications of the ACM, 45(7):89–95, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Banerjee:2002:MEI

[757] Snehamay Banerjee and Ram L. Kumar. Managing electronic interchange of business documents. Communications of the ACM, 45(7):96–102, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cummings:2002:QOS

[758] Jonathon N. Cummings, Brian Butler, and Robert Kraut. The quality of online social relationships. Communications of the ACM, 45(7):103–108, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davison:2002:TOC

[759] Robert Davison. Technical opinion: Cultural complications of ERP. Communications of the ACM, 45(7):109– 111, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Norman:2002:BCI

[760] Donald A. Norman. Inside risks: Beyond the computer industry. Communications of the ACM, 45(7):120, July 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPh

[761] Diane Crawford. Editorial pointers. Communications of the ACM, 45(8):5, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NT

[762] Robert Fox. News tracks. Communications of the ACM, 45(8):9–10, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FEE

[763] Diane Crawford, Srinivas Nedunuri, Adrian Bowyer, Arnd Weber, Greg A. Woods, and Mark Adler. Forum: Embrace the engineering metaphor; credit for crypto’s parallel development; enough PDF: Give me HTML. Communications of the ACM, 45 (8):11–14, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Norrisand:2002:EYT

[764] Cathleen Norrisand, Elliot Soloway, and Terry Sullivan. Log on education: Examining 25 years of technology in U.S. education. Communications of the ACM, 45(8):15–18, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2002:PCW

[765] Robert L. Glass. Practical programmer: The proof of correctness wars. Communications of the ACM, 45(8):19–21, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teuscher:2002:VHH

[766] Christof Teuscher and Moshe Sipper. Viewpoint: Hypercomputation: hype or computation? Communications of the ACM, 45(8):23–24, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malaga:2002:AMW

[767] Ross A. Malaga. On site: Additional methods when using email for teaching. Communications of the ACM, 45(8):25–27, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayyad:2002:EDM

[768] Usama Fayyad and Ramasamy Uthurusamy. Evolving data into mining solutions for insights. Communications of the ACM, 45(8):28–31, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smyth:2002:DDE

[769] Padhraic Smyth, Daryl Pregibon, and Christos Faloutsos. Data-driven evolution of data mining algorithms. Communications of the ACM, 45(8):33–37, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bradley:2002:SMA

[770] Paul Bradley, Johannes Gehrke, Raghu Ramakrishnan, and Ramakrishnan Srikant. Scaling mining algorithms to large databases. Communications of the ACM, 45(8):38–43, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kohavi:2002:ETB

[771] Ron Kohavi, Neal J. Rothleder, and Evangelos Simoudis. Emerging trends in business analytics. Communications of the ACM, 45(8):45–48, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Apte:2002:BAD

[772] Chidanand Apte, Bing Liu, Edwin P. D. Pednault, and Padhraic Smyth. Business applications of data mining. Communications of the ACM, 45(8):49–53, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Han:2002:ESA

[773] Jiawei Han, Russ B. Altman, Vipin Kumar, Heikki Mannila, and Daryl Pregibon. Emerging scientific applications in data mining. Communications of the ACM, 45(8):54–58, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grossman:2002:DMS

[774] Robert L. Grossman, Mark F. Hornick, and Gregor Meyer. Data mining standards initiatives. Communications of the ACM, 45(8):59–61, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tarasewich:2002:SIP

[775] Peter Tarasewich and Patrick R. Mc-Mullen. Swarm intelligence: power in numbers. Communications of the ACM, 45(8):62–67, August 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coltman:2002:KBP

[776] Tim Coltman, Timothy M. Devinney, Alopi S. Latukefu, and David F. Midgley. Keeping e-business in perspective. Communications of the ACM, 45(8):69–73, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Looney:2002:WET

[777] Clayton A. Looney and Debabroto Chatterjee. Web-enabled transformation of the brokerage industry. Communications of the ACM, 45(8):75–81, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slyke:2002:GDP

[778] Craig Van Slyke, Christie L. Comunale, and France Belanger. Gender differences in perceptions of Webbased shopping. Communications of the ACM, 45(8):82–86, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chin:2002:SNV

[779] George Chin, James Myers, and David Hoyt. Social networks in the virtual science laboratory. Communications of the ACM, 45(8):87–92, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Quaddus:2002:ECD

[780] M. A. Quaddus and Lai Lai Tung. Explaining cultural differences in decision conferencing. Communications of the ACM, 45(8):93–98, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:2002:CCS

[781] Lawrence G. Jones and Arthur L. Price. Changes in computer science accreditation. Communications of the ACM, 45(8):99–103, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ourosoff:2002:PTJ

[782] Nick Ourosoff. Technical opinion: Primitive types in Java considered harmful. Communications of the ACM, 45(8):105–106, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Acar:2002:IRR

[783] Tolga Acar and John R. Michener. Inside risks: Risks in features vs. assurance. Communications of the ACM, 45(8):112, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPi

[784] Diane Crawford. Editorial pointers. Communications of the ACM, 45(9):5, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTSb

[785] Robert Fox. News track: Symbols for vagabond users; chipped tooth; pixel fix on visual intelligence; blinking bumper; virtual house calls; more than one billion served; math and science teachers wanted; crash test dummy. Communications of the ACM, 45(9):9–10, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FIF

[786] Diane Crawford, Al Endres, Jason Funk, John Florentin, Gregg Irwin, Bud Lawson, Alvin P. Mullery, Christian Wagner, and Efraim Turban. Forum: Influencing factors in the computer industry; the difference between intelligent agents and mobile agents. Communications of the ACM, 45(9): 11–14, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2002:DVR

[787] Hal Berghel. Digital village: Responsible Web caching. Communications of the ACM, 45(9):15–20, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2002:PIC

[788] Peter J. Denning. The profession of IT: Career redux. Communications of the ACM, 45(9):21–26, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:2002:SWB

[789] Paul F. Ross. On site: Whatever became of integrity? Communications of the ACM, 45(9):27–28, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klawe:2002:LPM

[790] Maria M. Klawe. Letter from the President: Making a difference through computing communities. Communications of the ACM, 45(9):29–30, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:2002:VIM

[791] Ralph R. Miller. Viewpoint: Information management in the aftermath of 9/11. Communications of the ACM, 45 (9):31–33, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davies:2002:YAW

[792] Simon Davies. A year after 9/11: where are we now? Communications of the ACM, 45(9):35–39, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brewer:2002:ICS

[793] Eric A. Brewer. Introduction: The consumer side of search. Communications of the ACM, 45(9):40–41, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hearst:2002:FFW

[794] Marti Hearst, Ame Elliott, Jennifer English, Rashmi Sinha, Kirsten Swearingen, and Ka-Ping Yee. Finding the flow in Web site search. Communications of the ACM, 45(9):42–49, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pitkow:2002:PS

[795] James Pitkow, Hinrich Schütze, Todd Cass, Rob Cooley, Don Turnbull, Andy Edmonds, Eytan Adar, and Thomas Breuel. Personalized search. Communications of the ACM, 45(9):50–55, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mowshowitz:2002:BW

[796] Abbe Mowshowitz and Akira Kawaguchi. Bias on the Web. Communications of the ACM, 45(9):56–60, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gerlach:2002:DCI

[797] James Gerlach, Bruce Neumann, Edwin Moldauer, Martha Argo, and Daniel Frisby. Determining the cost of IT services. Communications of the ACM, 45(9):61–67, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Isaacs:2002:MIM

[798] Ellen Isaacs, Alan Walendowski, and Dipti Ranganathan. Mobile instant messaging through Hubbub. Communications of the ACM, 45(9):68–72, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agarwal:2002:EPM

[799] Ritu Agarwal and Thomas W. Ferratt. Enduring practices for managing IT professionals. Communications of the ACM, 45(9):73–79, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morrison:2002:IWS

[800] Mike Morrison, Joline Morrison, and Anthony Keys. Integrating Web sites and databases. Communications of the ACM, 45(9):81–86, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tyran:2002:ISI

[801] Craig K. Tyran and Joey F. George. Improving software inspections with group process support. Communications of the ACM, 45(9):87–92, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

King:2002:MII

[802] William R. King, Peter V. Marks, Jr., and Scott McCoy. The most important issues in knowledge management. Communications of the ACM, 45(9):93–97, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shilov:2002:EST

[803] Nikolay V. Shilov and Kwangkeun Yi. Engaging students with theory through ACM Collegiate Programming Contests. Communications of the ACM, 45 (9):98–101, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fayad:2002:TOM

[804] Mohamed Fayad and Shasha Wu. Thinking objectively: Merging multiple conventional models in one stable model. Communications of the ACM, 45(9):102–106, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stamp:2002:IRR

[805] Mark Stamp. Inside risks: Risks of digital rights management. Communications of the ACM, 45(9):120, September 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPj

[806] Diane Crawford. Editorial pointers. Communications of the ACM, 45(10):5, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTL

[807] Robert Fox. News track: Liquid light; let your finging do the paying; IT professional pressure cooker; A spot in the clouds; must see floppy TV; mainframe skill shortage; Moore’s Law and cyronic suspension. Communications of the ACM, 45(10):9–10, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2002:SWC

[808] Rebecca T. Mercuri. Security watch: Computer security: quality rather than quantity. Communications of the ACM, 45(10):11–14, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2002:LSR

[809] Pamela Samuelson. Legally speaking: Reverse engineering under siege. Communications of the ACM, 45(10):15– 20, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyer:2002:MEW

[810] R. S. Boyer, W. Feijen, D. Gries, C. A. R. Hoare, J. Misra, J. Moore, and H. Richards. In memoriam: Edsger W. Dijkstra 1930–2002. Communications of the ACM, 45(10):21–22, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Palacios:2002:SRA

[811] Rafael Palacios. On site: Remote automatic doorman via the Internet. Communications of the ACM, 45 (10):23–25, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howard:2002:VRA

[812] Alan Howard. Viewpoint: Rapid Application Development: rough and dirty or value-for-money engineering? Communications of the ACM, 45(10):27–29, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arsanjani:2002:IDI

[813] Ali Arsanjani. Introduction: Developing and integrating enterprise components and services. Communications of the ACM, 45(10):30–34, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crnkovic:2002:SID

[814] Ivica Crnkovic, Brahim Hnich, Torsten Jonsson, and Zeynep Kiziltan. Specification, implementation, and deployment of components. Communications of the ACM, 45(10):35–40, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meijer:2002:OIE

[815] Erik Meijer and Clemens Szyperski. Overcoming independent extensibility challenges. Communications of the ACM, 45(10):41–44, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levi:2002:GDA

[816] Keith Levi and Ali Arsanjani. A goaldriven approach to enterprise component identification and specification. Communications of the ACM, 45(10):45–52, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jardim-Goncalves:2002:IMM

[817] Ricardo Jardim-Gonçalves and Adolfo Steiger-Garção. Implicit multilevel modeling in flexible business environments. Communications of the ACM, 45(10):53–57, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sutherland:2002:EAI

[818] Jeff Sutherland and Willem-Jan van den Heuvel. Enterprise application integration and complex adaptive systems. Communications of the ACM, 45(10):59–64, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gokhale:2002:AMI

[819] Aniruddha Gokhale, Douglas C. Schmidt, Balachandran Natarajan, and Nanbor Wang. Applying modelintegrated computing to component middleware and enterprise applications. Communications of the ACM, 45(10):65–70, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stal:2002:WSB

[820] Michael Stal. Web services: beyond component-based computing. Communications of the ACM, 45(10):71– 76, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fremantle:2002:ES

[821] Paul Fremantle, Sanjiva Weerawarana, and Rania Khalaf. Enterprise services. Communications of the ACM, 45(10):77–82, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2002:LLN

[822] Soo Dong Kim. Lessons learned from a nationwide CBD promotion project. Communications of the ACM, 45(10):83–87, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burleson:2002:ICI

[823] Winslow Burleson and Ted Selker. Introduction: Creativity and interface. Communications of the ACM, 45(10):88–90, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Edmonds:2002:CAP

[824] Ernest Edmonds and Linda Candy. Creativity, art practice, and knowledge. Communications of the ACM, 45(10):91–95, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mamykina:2002:CC

[825] Lena Mamykina, Linda Candy, and Ernest Edmonds. Collaborative creativity. Communications of the ACM, 45(10):96–99, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greene:2002:CAS

[826] Sharon L. Greene. Characteristics of applications that support creativity. Communications of the ACM, 45 (10):100–104, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Terry:2002:SES

[827] Michael Terry and Elizabeth D. Mynatt. Supporting experimentation with Side-Views. Communications of the ACM, 45(10):106–108, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roast:2002:RCR

[828] Chris Roast, Innes Ritchie, and Stephanie Thomas. Re-creating the reader: supporting active reading in literary research. Communications of the ACM, 45(10):109–111, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thomas:2002:ECD

[829] John C. Thomas, Alison Lee, and Catalina Danis. Enhancing creative design via software tools. Communications of the ACM, 45(10):112–115, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:2002:CST

[830] Ben Shneiderman. Creativity support tools. Communications of the ACM, 45 (10):116–120, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:2002:EWL

[831] E. Vance Wilson. Email winners and losers. Communications of the ACM, 45(10):121–126, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farhoomand:2002:MIO

[832] Ali F. Farhoomand and Don H. Drury. Managerial information overload. Communications of the ACM, 45 (10):127–131, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nelson:2002:ODN

[833] H. James Nelson, Deborah J. Armstrong, and Mehdi Ghods. Old dogs and new tricks. Communications of the ACM, 45(10):132–137, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chau:2002:CDO

[834] Patrick Y. K. Chau, Melissa Cole, Anne P. Massey, Mitzi Montoya-Weiss, and Robert M. O’Keefe. Cultural differences in the online behavior of consumers. Communications of the ACM, 45(10):138–143, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malladi:2002:TOC

[835] Rajeswari Malladi and Dharma P. Agrawal. Technical opinion: Current and future applications of mobile and wireless networks. Communications of the ACM, 45(10):144–146, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneider:2002:IRS

[836] Fred B. Schneider. Inside risks: Secure systems conundrum. Communications of the ACM, 45(10):160, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPk

[837] Diane Crawford. Editorial pointers. Communications of the ACM, 45(11):5, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTT

[838] Robert Fox. News track: A thousand points of light; typing with your eyes; cleaning crustacean; the E-Bay of Argentina; online abduction alerts; fighting airport noise; happy birthday :- ). Communications of the ACM, 45 (11):9–10, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FWM

[839] Diane Crawford, Sherif Yacoub, John Stracke, Gordon Bell, Jim Gemmell, Bob Ellis, Myles Losch, Ronnie Ward, Anthony Ralston, Rich Henders, Steve Arnold, and Mark Lutton. Forum: I want my home network; consumercreated content key to broadband; still living off Y2K spending; grading education; objective opinion; PDF peeves; homemade hypercomputation. Communications of the ACM, 45(11): 11–14, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2002:BST

[840] Phillip Armour. Business of software: Ten unmyths of project estimation. Communications of the ACM, 45 (11):15–18, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2002:PPSb

[841] Robert L. Glass. Practical programmer: Sorting out software complexity. Communications of the ACM, 45 (11):19–21, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gelbord:2002:VNS

[842] Boaz Gelbord and Gert Roelofsen. Viewpoint: New surveillance techniques raise privacy concerns. Communications of the ACM, 45(11):23–24, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maamar:2002:ICS

[843] Zakaria Maamar and Weiming Shen. Introduction: Computer-supported cooperative work in design. Communications of the ACM, 45(11):25–26, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klein:2002:CSP

[844] Mark Klein, Hiroki Sayama, Peyman Faratin, and Yaneer Bar-Yam. A complex systems perspective on computersupported collaborative design technology. Communications of the ACM, 45 (11):27–31, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sedaghat:2002:FWC

[845] Soroush Sedaghat, Josef Pieprzyk, and Ehsan Vossough. On-the-fly Web content integrity check boosts users’ confidence. Communications of the ACM, 45(11):33–37, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bergenti:2002:CPF

[846] Federico Bergenti, Agostino Poggi, and Matteo Somacher. A collaborative platform for fixed and mobile networks. Communications of the ACM, 45(11):39–44, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2002:JEC

[847] Lihui Wang, Brian Wong, Weiming Shen, and Sherman Lang. A Java 3D-enabled cyber workspace. Communications of the ACM, 45(11):45–49, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:2002:WWT

[848] Jim Gray and Alex Szalay. The World-Wide Telescope. Communications of the ACM, 45(11):50–55, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2002:SEP

[849] David P. Anderson, Jeff Cobb, Eric Korpela, Matt Lebofsky, and Dan Werthimer. SETI@home: an experiment in public-resource computing. Communications of the ACM, 45 (11):56–61, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Genetti:2002:VRG

[850] Jon Genetti. Volume-rendered galactic animations. Communications of the ACM, 45(11):62–66, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:2002:IWU

[851] Joaquin Miller. Introduction: What UML should be. Communications of the ACM, 45(11):67–69, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Selic:2002:ER

[852] Bran Selic, Guus Ramackers, and Cris Kobryn. Evolution, not revolution. Communications of the ACM, 45 (11):70–72, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duddy:2002:UME

[853] Keith Duddy. UML2 must enable a family of languages. Communications of the ACM, 45(11):73–75, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mellor:2002:MMA

[854] Stephen J. Mellor. Make models be assets. Communications of the ACM, 45 (11):76–78, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frank:2002:CCC

[855] William Frank and Kevin P. Tyson. Be clear, clean, concise. Communications of the ACM, 45(11):79–81, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dori:2002:WSU

[856] Dov Dori. Why significant UML change is unlikely. Communications of the ACM, 45(11):82–85, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knight:2002:SSE

[857] John C. Knight and Nancy G. Leveson. Should software engineers be licensed? Communications of the ACM, 45(11):87–90, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:2002:APL

[858] John White and Barbara Simons. ACM’s position on the licensing of software engineers. Communications of the ACM, 45(11):91, November 2002.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bagert:2002:TLS

[859] Donald J. Bagert. Texas licensing of software engineers: all’s quiet, for now. Communications of the ACM, 45 (11):92–94, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kennedy:2002:RUP

[860] Ken Kennedy and Moshe Y. Vardi. A Rice University perspective on software engineering licensing. Communications of the ACM, 45(11):94–95, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:2002:LSE

[861] David Lorge Parnas. Licensing software engineers in Canada. Communications of the ACM, 45(11):96–98, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCalla:2002:SER

[862] Gord McCalla. Software engineering requires individual professionalism. Communications of the ACM, 45(11): 98–101, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deChampeaux:2002:TOS

[863] Dennis de Champeaux. Technical opinion: Software engineering considered harmful. Communications of the ACM, 45(11):102–104, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2002:IRF

[864] Rebecca Mercuri. Inside risks: Florida 2002: sluggish systems, vanishing votes. Communications of the ACM, 45(11):136, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:EPl

[865] Diane Crawford. Editorial pointers. Communications of the ACM, 45(12):5, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2002:NTSc

[866] Robert Fox. News track: Software to combat terror; translating the art of war; pressing flesh for transmission; ripping and burning DNA; new meaning of “PDA”; advanced electronics second guesses umps; sports in the dark possible. Communications of the ACM, 45(12):9–10, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2002:FRP

[867] Diane Crawford, C. W. Neville, James H. Fetzer, Rodney M. Bates, Paul E. McKenney, Ann Sobel, Michael Clarkson, Todd Palino, and Hal Berghel. Forum: Resolving the program-verification debate; handling Web content. Communications of the ACM, 45(12):11–15, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mbarika:2002:IPC

[868] Victor Mbarika, Mike Jensen, and Peter Meso. International perspectives: Cyberspace across sub-Saharan Africa. Communications of the ACM, 45(12):17–21, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Palsberg:2002:STR

[869] Jens Palsberg and Scott J. Baxter. On site: Teaching reviewing to graduate students. Communications of the ACM, 45(12):22–24, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Castel:2002:VTT

[870] Felipe Castel. Viewpoint: Theory, theory on the wall. Communications of the ACM, 45(12):25–26, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varshney:2002:MWN

[871] Upkar Varshney. Multicast over wireless networks. Communications of the ACM, 45(12):31–37, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubin:2002:SCR

[872] Aviel D. Rubin. Security considerations for remote electronic voting. Communications of the ACM, 45 (12):39–44, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Strader:2002:VST

[873] Troy J. Strader and Sridhar N. Ramaswami. The value of seller trustworthiness in C2C online markets. Communications of the ACM, 45(12): 45–49, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schlueter-Langdon:2002:EPI

[874] Christoph Schlueter-Langdon and Michael J. Shaw. Emergent patterns of integration in electronic channel systems. Communications of the ACM, 45 (12):50–55, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raghupathi:2002:SIA

[875] Wullianallur Raghupathi and Joseph Tan. Strategic IT applications in health care. Communications of the ACM, 45 (12):56–61, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lyytinen:2002:IIC

[876] Kalle Lyytinen and Youngjin Yoo. Introduction: Issues and challenges in ubiquitous computing. Communications of the ACM, 45(12):62–65, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davis:2002:AAC

[877] Gordon B. Davis. Anytime/anyplace computing and the future of knowledge work. Communications of the ACM, 45 (12):67–73, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grudin:2002:GDU

[878] Jonathan Grudin. Group dynamics and ubiquitous computing. Communications of the ACM, 45(12):74–78, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siewiorek:2002:NFA

[879] Daniel P. Siewiorek. New frontiers of application design. Communications of the ACM, 45(12):79–82, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fano:2002:FBS

[880] Andrew Fano and Anatole Gershman. The future of business services in the age of ubiquitous computing. Communications of the ACM, 45 (12):83–87, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jessup:2002:RSI

[881] Leonard M. Jessup and Daniel Robey. The relevance of social issues in ubiquitous computing environments. Communications of the ACM, 45(12):88–91, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Banavar:2002:SID

[882] Guruduth Banavar and Abraham Bernstein. Software infrastructure and design challenges for ubiquitous computing applications. Communications of the ACM, 45(12):92–96, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berzal:2002:TOC

[883] Fernando Berzal, Ignacio Blanco, Juan-Carlos Cubero, and Nicolas Marin. Technical opinion: Component-based data mining frameworks. Communications of the ACM, 45(12):97–100, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wool:2002:IRW

[884] Avishai Wool. Inside risks: Why security standards sometimes fail. Communications of the ACM, 45(12):144, December 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPa

[885] Diane Crawford. Editorial pointers. Communications of the ACM, 46(1): 5, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:NTI

[886] Robert Fox. News tracks: Incredible shrinking computing; super-size telescope; global reach around; gastric simulation; Filipino text-message TV; caging hackers; Japan’s singing, dancing robot. Communications of the ACM, 46(1):9–10, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FLW

[887] Diane Crawford, Barbara Simons, Eugene H. Spafford, Christof Schmalenbach, László Karafiát, Dave Isaacs, Ken Kennedy, and Moshe Vardi. Forum: A lawful way to acquire a trade secret; criteria for integration technology; the current state of hypercomputation; still no licensing in Texas. Communications of the ACM, 46(1):11–12, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2003:SCH

[888] Meg McGinity. Staying connected: Hibernating during telecom winter. Communications of the ACM, 46(1):13–15, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2003:SWAa

[889] Rebecca T. Mercuri. Security watch: On auditing audit trails. Communications of the ACM, 46(1):17–20, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karp:2003:VMM

[890] Alan H. Karp. Viewpoint: Making money selling content that others are giving away. Communications of the ACM, 46(1):21–22, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marchionini:2003:IDG

[891] Gary Marchionini, Hanan Samet, and Larry Brandt. Introduction: Digital government. Communications of the ACM, 46(1):24–27, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2003:CML

[892] Hsinchun Chen, Daniel Zeng, Homa Atabakhsh, Wojciech Wyzga, and Jenny Schroeder. COPLINK: managing law enforcement data and knowledge. Communications of the ACM, 46(1):28–34, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goddard:2003:GDS

[893] Steve Goddard, Sherri K. Harms, Stephen E. Reichenbach, Tsegaye Tadesse, and William J. Waltman. Geospatial decision support for drought risk management. Communications of the ACM, 46(1):35–37, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delcambre:2003:HIS

[894] Lois Delcambre, Timothy Tolle, and The Forest Project Team. Harvesting information to sustain forests. Communications of the ACM, 46(1):38–39, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dawes:2003:UNM

[895] Sharon S. Dawes and Lise Préfontaine. Understanding new models of collaboration for delivering government services. Communications of the ACM, 46(1):40–42, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fountain:2003:PIR

[896] Jane E. Fountain. Prospects for improving the regulatory process using e-rulemaking. Communications of the ACM, 46(1):43–44, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nusser:2003:GIM

[897] Sarah Nusser, Leslie Miller, Keith Clarke, and Michael Goodchild. Geospatial IT for mobile field data collection. Communications of the ACM, 46(1):45–46, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hovy:2003:UOS

[898] Eduard Hovy. Using an ontology to simplify data access. Communications of the ACM, 46(1):47–49, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golubchik:2003:BSS

[899] Leana Golubchik, William C. Cheng, Cheng-Fu Chou, Samir Khuller, Hanan Samet, and C. Justin Wan. Bistro: a scalable and secure data transfer service for digital government applications. Communications of the ACM, 46(1):50–51, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hert:2003:SSE

[900] Carol A. Hert, Elizabeth D. Liddy, Ben Shneiderman, and Gary Marchionini. Supporting statistical electronic table usage by citizens. Communications of the ACM, 46(1):52–54, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dippo:2003:FGF

[901] Cathryn Dippo. FedStats: the gateway to federal statistics. Communications of the ACM, 46(1):55, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karr:2003:TSP

[902] Alan F. Karr, Adrian Dobra, and Ashish P. Sanil. Table servers protect confidentiality in tabular data releases. Communications of the ACM, 46(1):57–58, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

MacEachren:2003:SVA

[903] Alan M. MacEachren, Frank Hardisty, Xiping Dai, and Linda Pickle. Supporting visual analysis of federal geospatial statistics. Communications of the ACM, 46(1):59–60, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samet:2003:USS

[904] Hanan Samet, Houman Alborzi, František Brabec, Claudio Esperança, Gísli R. Hjaltason, Frank Morgan, and Egemen Tanin. Use of the SAND spatial browser for digital government applications. Communications of the ACM, 46(1):61–64, January 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agouris:2003:ESS

[905] Peggy Agouris and Anthony Stefanidis. Efficient summarization of spatiotemporal events. Communications of the ACM, 46(1):65–66, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scherlis:2003:IRI

[906] William L. Scherlis and Jon Eisenberg. IT research, innovation, and e-government. Communications of the ACM, 46(1):67–68, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ciment:2003:PHN

[907] Melvyn Ciment. A personal history of the NSF digital government program. Communications of the ACM, 46(1):69–70, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2003:DFW

[908] Diomidis Spinellis. The decay and failures of Web references. Communications of the ACM, 46(1):71–77, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Avison:2003:WND

[909] David E. Avison and Guy Fitzgerald. Where now for development methodologies? Communications of the ACM, 46(1):78–82, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peffers:2003:HIN

[910] Ken Peffers and Charles E. Gengler. How to identify new high-payoff information systems for the organization. Communications of the ACM, 46(1):83–88, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cannataro:2003:KG

[911] Mario Cannataro and Domenico Talia. The knowledge grid. Communications of the ACM, 46(1):89–93, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wong:2003:ODM

[912] Pak Chung Wong, Kwong kwok Wong, and Harlan Foote. Organic data memory using the DNA approach. Communications of the ACM, 46(1):95–98, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desouza:2003:TOB

[913] Kevin C. Desouza. Technical opinion: Barriers to effective use of knowledge management systems in software engineering. Communications of the ACM, 46(1):99–101, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lesk:2003:IRM

[914] Michael Lesk. Inside risks: The mindset of dependability. Communications of the ACM, 46(1):136, January 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPb

[915] Diane Crawford. Editorial pointers. Communications of the ACM, 46(2):5, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:NTN

[916] Robert Fox. News track: NET works for volunteers; lights out back in; China holds tightest net reins; hints of recovery; talk is hardly cheap; ’02 views. Communications of the ACM, 46(2):9–10, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FWP

[917] Diane Crawford, Simon Bennett, Lesley Clark, Dan Klarmann, Gernot Salzer, Dan Hestand, Richard Molpus, and Rebecca Mercuri. Forum: Why projects as knowledge acquisition work; complex, but not that complex; blame the process, not the parties. Communications of the ACM, 46(2):11–13, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meeks:2003:EFC

[918] Brock N. Meeks. Electronic frontier: Conspicuous in their silence. Communications of the ACM, 46(2):15–16, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2003:BSR

[919] Phillip Armour. The business of software: The reorg cycle. Communications of the ACM, 46(2):19–22, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brazelton:2003:SCK

[920] Jessica Brazelton and G. Anthony Gorry. On site: Creating a knowledgesharing community: if you build it, will they come? Communications of the ACM, 46(2):23–25, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:AF

[921] Diane Crawford. ACM Fellows. Communications of the ACM, 46(2):26, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schoder:2003:VPP

[922] Detlef Schoder and Kai Fischbach. Viewpoint: Peer-to-peer prospects. Communications of the ACM, 46(2): 27–29, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lethin:2003:ITS

[923] Richard Lethin. Introduction: Technical and social components of peer-to-peer computing. Communications of the ACM, 46(2):30–32, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kubiatowicz:2003:EGC

[924] John Kubiatowicz. Extracting guarantees from chaos. Communications of the ACM, 46(2):33–38, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agre:2003:PPI

[925] Philip E. Agre. P2P and the promise of Internet equality. Communications of the ACM, 46(2):39–42, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Balakrishnan:2003:LDP

[926] Hari Balakrishnan, M. Frans Kaashoek, David Karger, Robert Morris, and Ion Stoica. Looking up data in P2P systems. Communications of the ACM, 46(2):43–48, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:EUP

[927] Jintae Lee. An end-user perspective on file-sharing systems. Communications of the ACM, 46(2):49–53, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:EIE

[928] Jinyoul Lee, Keng Siau, and Soongoo Hong. Enterprise integration with ERP and EAI. Communications of the ACM, 46(2):54–60, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arinze:2003:FUO

[929] Bay Arinze and Murugan Anandarajan. A framework for using OO mapping methods to rapidly configure ERP systems. Communications of the ACM, 46(2):61–65, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dutta:2003:AID

[930] Amitava Dutta and Rahul Roy. Anticipating Internet diffusion. Communications of the ACM, 46(2):66–71, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jiang:2003:CUP

[931] James J. Jiang, Gary Klein, Debbie Tesch, and Hong-Gee Chen. Closing the user and provider service quality gap. Communications of the ACM, 46(2):72–76, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krovi:2003:IFP

[932] Ravindra Krovi, Akhilesh Chandra, and Balaji Rajagopalan. Information flow parameters for managing organizational processes. Communications of the ACM, 46(2):77–82, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shi:2003:CBO

[933] Xinping Shi and Philip C. Wright. E-commercializing business operations. Communications of the ACM, 46(2):

83–87, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Islam:2003:TOT

[934] Nayeem Islam and Mohamed Fayad. Thinking objectively: Toward ubiquitous acceptance of ubiquitous computing. Communications of the ACM, 46 (2):89–92, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2003:IRG

[935] Peter G. Neumann. Inside risks: Gambling on system accountability. Communications of the ACM, 46(2):120, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPc

[936] Diane Crawford. Editorial pointers. Communications of the ACM, 46(3):5, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:NTT

[937] Robert Fox. News track: Thanks, but no thanks; still choking on spam; cameras, cameras, everywhere; rat’s sense; one person’s garbage; easy dot-com and go. Communications of the ACM, 46(3):9–10, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FES

[938] Diane Crawford, Dennis de Champeaux, David Lorge Parnas, Craig Partridge, Julio Lerm, and Oladokun Olajoyegbe. Forum: Extending and shrinking UML; give meaning to the status of software engineer; the art of the paper review; no waiting for enterprise-integration products; mathematical rigor for the social sciences. Communications of the ACM, 46(3):11–13, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2003:TSMa

[939] Michael Cusumano. Technology strategy and management: Finding your balance in the products and services debate. Communications of the ACM, 46(3):15–17, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2003:PIM

[940] Peter J. Denning and Robert Dunham. The profession of IT: The missing customer. Communications of the ACM, 46(3):19–23, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rob:2003:SRF

[941] Mohammad Rob. On site: The rise and fall of an e-commerce program. Communications of the ACM, 46(3):25–26, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Besaha:2003:VBH

[942] Bob Besaha. Viewpoint: Bounty hunting in the patent base. Communications of the ACM, 46(3):27–29, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vertegaal:2003:IAU

[943] Roel Vertegaal. Introduction: Attentive user interfaces. Communications of the ACM, 46(3):30–33, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhai:2003:WEA

[944] Shumin Zhai. What’s in the eyes for attentive input. Communications of the ACM, 46(3):34–39, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shell:2003:IGC

[945] Jeffrey S. Shell, Ted Selker, and Roel Vertegaal. Interacting with groups of computers. Communications of the ACM, 46(3):40–46, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maglio:2003:AA

[946] Paul P. Maglio and Christopher S. Campbell. Attentive agents. Communications of the ACM, 46(3):47–51, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horvitz:2003:MAC

[947] Eric Horvitz, Carl Kadie, Tim Paek, and David Hovel. Models of attention in computing and communication: from principles to applications. Communications of the ACM, 46(3):52–59, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baudisch:2003:FEC

[948] Patrick Baudisch, Doug DeCarlo, Andrew T. Duchowski, and Wilson S. Geisler. Focusing on the essential: considering attention in display design. Communications of the ACM, 46(3):60–66, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCrickard:2003:AND

[949] D. Scott McCrickard and C. M. Chewar. Attuning notification design to user goals and attention costs. Communications of the ACM, 46(3):67–72, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

delRMillan:2003:ABI

[950] José del R. Millán. Adaptive brain interfaces. Communications of the ACM, 46(3):74–80, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gordon:2003:FUI

[951] Lawrence A. Gordon, Martin P. Loeb, and Tashfeen Sohail. A framework for using insurance for cyber-risk management. Communications of the ACM, 46(3):81–85, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

daRochaBarretoPinto:2003:SDI

[952] Gustavo da Rocha Barreto Pinto, Sergio Palma J. Medeiros, Jano Moreira de Souza, Julia Celia Mercedes Strauch, and Carlete Rosana Ferreira Marques. Spatial data integration in a collaborative design framework. Communications of the ACM, 46(3):86–90, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2003:NPP

[953] Mark C. Anderson, Rajiv D. Banker, and Sury Ravindran. The new productivity paradox. Communications of the ACM, 46(3):91–94, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evernden:2003:TGI

[954] Roger Evernden and Elaine Evernden. Third-generation information architecture. Communications of the ACM, 46(3):95–98, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chadwick:2003:DLW

[955] David Chadwick. Deficiencies in LDAP when used to support PKI. Communications of the ACM, 46(3):99–104, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Templeton:2003:TOO

[956] Gary F. Templeton. Technical opinion: Object-oriented programming of integrated circuits. Communications of the ACM, 46(3):105–108, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simons:2003:IRR

[957] Barbara Simons and Eugene H. Spafford. Inside risks: Risks of total surveillance. Communications of the ACM, 46(3):120, March 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPd

[958] Diane Crawford. Editorial pointers. Communications of the ACM, 46(4):5, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:NTM

[959] Robert Fox. News track: More tech dollars for defense; no hands, no break; more net = less TV; brain scan; A word to the hip; smart dresser. Communications of the ACM, 46(4):9–10, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FTP

[960] Diane Crawford, Sandy Ressler, Michael DeBellis, Mark Wallace, Arthur M. Rosenberg, Robert Levine, and Barry I. Soroka. Forum: The ‘terrible price’ of the DMCA; the Patriot Act is only the beginning; communicating vs. computing; don’t tag living organisms. Communications of the ACM, 46(4):11–14, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2003:DVA

[961] Hal Berghel and Kim Womack. Digital village: Anonymizing the net. Communications of the ACM, 46(4):15–20, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grove:2003:VLT

[962] Jeff Grove. Viewpoint: Legal and technological efforts to lock up content threaten innovation. Communications of the ACM, 46(4):21–22, April 2003.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2003:SCG

[963] Meg McGinity. Staying connected: Getting real. Communications of the ACM, 46(4):23–26, April 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jain:2003:VFC

[964] Ramesh Jain. Viewpoint: Folk computing. Communications of the ACM, 46(4):27–29, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mulligan:2003:IDR

[965] Deirdre K. Mulligan. Introduction: Digital rights management and fair use by design. Communications of the ACM, 46(4):30–33, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2003:FUD

[966] John S. Erickson. Fair use, DRM, and trusted computing. Communications of the ACM, 46(4):34–39, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2003:DVL

[967] Pamela Samuelson. DRM and, or, vs. the law. Communications of the ACM, 46(4):41–45, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2003:DP

[968] Julie E. Cohen. DRM and privacy. Communications of the ACM, 46(4):46–49, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dusollier:2003:FUD

[969] Séverine Dusollier. Fair use by design in the European Copyright Directive of 2001. Communications of the ACM, 46(4):51–55, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Felten:2003:SVD

[970] Edward W. Felten. A skeptical view of DRM and fair use. Communications of the ACM, 46(4):56–59, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:ERF

[971] Barbara L. Fox and Brian A. LaMacchia. Encouraging recognition of fair uses in DRM systems. Communications of the ACM, 46(4):61–63, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fitzgerald:2003:SDM

[972] Brian Fitzgerald, Nancy L. Russo, and Tom O’Kane. Software development method tailoring at Motorola. Communications of the ACM, 46(4):64–70, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pachet:2003:CME

[973] François Pachet. Content management for electronic music distribution. Communications of the ACM, 46(4):71–75, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:2003:BNA

[974] Susan A. Brown and Viswanath Venkatesh. Bringing non-adopters along: the challenge facing the PC industry. Communications of the ACM, 46(4):76–80, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Earp:2003:IWU

[975] Julia B. Earp and David Baumer. Innovative Web use to learn about consumer behavior and online privacy. Communications of the ACM, 46(4):81–83, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farhoomand:2003:BSB

[976] Ali F. Farhoomand, Pauline S. P. Ng, and William L. Conley. Building a successful e-business: the FedEx story. Communications of the ACM, 46(4):84–89, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siau:2003:BCT

[977] Keng Siau and Zixing Shen. Building customer trust in mobile commerce. Communications of the ACM, 46(4):91–94, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pan:2003:UCU

[978] Shan L. Pan and Jae-Nam Lee. Using e-CRM for a unified view of the customer. Communications of the ACM, 46(4):95–99, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegner:2003:TOC

[979] Peter Wegner and Dina Goldin. Technical opinion: Computation beyond Turing machines. Communications of the ACM, 46(4):100–102, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2003:IRS

[980] Andrew Wright. Inside risks: On Sapphire and type-safe languages. Communications of the ACM, 46(4):120, April 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPe

[981] Diane Crawford. Editorial pointers. Communications of the ACM, 46(5):5, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2003:NTA

[982] Robert Fox. News track: Antiterror gizmos; grayer days ahead; take your receipt; gramr CWOT; dot.con; work free on free time; cyber arches. Communications of the ACM, 46(5):9–10, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FAW

[983] Diane Crawford, Urs Hoelzle, Michael Lewchuk, Lawrence A. Stabile, Donn B. Parker, Lawrence A. Gordon, Martin P. Loeb, Tashfeen Sohail, and Harry Rudin. Forum: Archiving the Web is feasible today; stop the reorg cycle; customer found; cyber-risk management not feasible; dear life. Communications of the ACM, 46(5):11–13, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2003:BSZ

[984] Phillip Armour. The business of software: In the zone: the need for flexible roles. Communications of the ACM, 46(5):15–19, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2003:PPO

[985] Robert L. Glass. Practical programmer: One giant step backward. Communications of the ACM, 46(5):21–23, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsichritzis:2003:VLB

[986] Dennis Tsichritzis. Viewpoint: Life beyond the bubbles. Communications of the ACM, 46(5):25–27, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubin:2003:IWN

[987] Aviel D. Rubin. Introduction: Wireless networking security. Communications of the ACM, 46(5):28–30, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Housley:2003:SPB

[988] Russ Housley and William Arbaugh. Security problems in 802.11-based networks. Communications of the ACM, 46(5):31–34, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cam-Winget:2003:SFD

[989] Nancy Cam-Winget, Russ Housley, David Wagner, and Jesse Walker. Security flaws in 802.11 data link protocols. Communications of the ACM, 46(5):35–39, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Byers:2003:APM

[990] Simon Byers and Dave Kormann. 802.11b access point mapping. Communications of the ACM, 46(5):41–46, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmidt:2003:WWF

[991] Terry Schmidt and Anthony Townsend. Why wi-fi wants to be free. Communications of the ACM, 46(5):47–52, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herzberg:2003:PBM

[992] Amir Herzberg. Payments and banking with mobile personal devices. Communications of the ACM, 46(5):53–58, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:ERM

[993] Jae Kyu Lee and Mye M. Sohn. The eXtensible Rule Markup Language. Communications of the ACM, 46(5):59–64, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

David:2003:AKS

[994] Julie Smith David, William E. Mc-Carthy, and Brian S. Sommer. Agility: the key to survival of the fittest in the software market. Communications of the ACM, 46(5):65–69, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graham:2003:CTS

[995] Leah Graham and Panagiotis Takis Metaxas. “of course it’s true; I saw it on the Internet!”: critical thinking in the Internet era. Communications of the ACM, 46(5):70–75, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tiwana:2003:AIP

[996] Amrit Tiwana. Affinity to infinity in peer-to-peer knowledge platforms. Communications of the ACM, 46(5):76–80, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liao:2003:RTT

[997] Ziqi Liao. Real-time taxi dispatching using Global Positioning Systems. Communications of the ACM, 46(5): 81–83, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:IOE

[998] Jae-Nam Lee, Minh Q. Huynh, Ron Chi-Wai Kwok, and Shih-Ming Pi. IT outsourcing evolution: past, present, and future. Communications of the ACM, 46(5):84–89, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marsaglia:2003:TOS

[999] George Marsaglia. Technical opinion: Seeds for random number generators: Techniques for choosing seeds for social and scientific applications of random number generators. Communications of the ACM, 46(5):90–93, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2003:IRE

[1000] Peter G. Neumann. Inside risks: E-epistemology and misinformation. Communications of the ACM, 46(5): 104, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPf

[1001] Diane Crawford. Editorial pointers. Communications of the ACM, 46(6):5, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTF

[1002] CACM staff. News track: Federal red flag; proof in numbers; bots for what?; spell checked; now you see it; profiles in trouble. Communications of the ACM, 46(6):9–10, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FDR

[1003] Diane Crawford, Mark Strefik, Rory Short, Simone Santinie, Peter Wegner, and Dina Goldin. Forum: In DRM, rethink business practices, not just technology and the law; don’t forget mental models in systems development; what the Turing machine doesn’t model. Communications of the ACM, 46(6):11–13, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2003:SWAb

[1004] Rebecca T. Mercuri. Security watch:

Analyzing security costs. Communications of the ACM, 46(6):15–18, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2003:LST

[1005] Pamela Samuelson. Legally speaking:

Trade secrets vs. free speech. Communications of the ACM, 46(6):19–23, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeMarco:2003:VMP

[1006] Tom DeMarco. Viewpoint: The Mc-Carthy protocols. Communications of the ACM, 46(6):24–25, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stafford:2003:IS

[1007] Thomas F. Stafford. Introduction:

E-services. Communications of the ACM, 46(6):26–28, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kreger:2003:FWS

[1008] Heather Kreger. Fulfilling the Web services promise. Communications of the ACM, 46(6):29–30, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ferris:2003:WWS

[1009] Christopher Ferris and Joel Farrell. What are Web services? Communications of the ACM, 46(6):31, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rust:2003:SNP

[1010] Roland T. Rust and P. K. Kannan. Eservice: a new paradigm for business in the electronic environment. Communications of the ACM, 46(6):36–42, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bolton:2003:MCS

[1011] Ruth N. Bolton. Marketing challenges of e-services. Communications of the ACM, 46(6):43–44, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Song:2003:SF

[1012] Hongjun Song. E-services at FedEx. Communications of the ACM, 46(6): 45–46, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brohman:2003:DCK

[1013] M. Kathryn Brohman, Richard T. Watson, Gabriele Piccoli, and A. Parasurama. Data completeness: a key to effective net-based customer service systems. Communications of the ACM, 46(6):47–51, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:2003:MMS

[1014] K. Douglas Hoffman. Marketing + MIS = e-service. Communications of the ACM, 46(6):53–55, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shaw:2003:TER

[1015] Neal G. Shaw and Christopher W.

Craighead. Technology enablers to recover from failures in e-services. Communications of the ACM, 46(6):56–57, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williams:2003:WSD

[1016] Joseph Williams. The Web services debate: J2EE vs. .NET. Communications of the ACM, 46(6):58–63, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:2003:WSD

[1017] Gerry Miller. The Web services debate: .NET vs. J2EE. Communications of the ACM, 46(6):64–67, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ben-Ameur:2003:NEV

[1018] Walid Ben-Ameur and Hervé Kerivin.

New economical virtual private networks. Communications of the ACM, 46(6):69–73, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2003:TSC

[1019] Huaiqing Wang and Chen Wang. Taxonomy of security considerations and software quality. Communications of the ACM, 46(6):75–78, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:2003:AHC

[1020] E. Vance Wilson. Asynchronous health care communication. Communications of the ACM, 46(6):79–84, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desouza:2003:FTK

[1021] Kevin C. Desouza. Facilitating tacit knowledge exchange. Communications of the ACM, 46(6):85–88, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payton:2003:RDD

[1022] Fay Cobb Payton. Rethinking the digital divide. Communications of the ACM, 46(6):89–91, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hart:2003:TPD

[1023] Peter E. Hart and Ziming Liu. Trust in the preservation of digital information. Communications of the ACM, 46(6):93–97, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Backhouse:2003:TOT

[1024] James Backhouse, Carol Hsu, and Aidan McDonnell. Technical opinion: Toward public-key infrastructure interoperability. Communications of the ACM, 46(6):98–100, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2003:IRR

[1025] Diomidis Spinellis. Inside risks: Reflections on trusting trust revisited. Communications of the ACM, 46(6):112, June 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPg

[1026] Diane Crawford. Editorial pointers.

Communications of the ACM, 46(7):5, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTC

[1027] CACM Staff. News track: Cinematic watermark; eye-opening education; roaming time; stand by me; savings bonds fade to net; phone home. Communications of the ACM, 46(7): 9–10, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FWC

[1028] Diane Crawford, Simone Santini, Ralph Castain, William F. Dowling, John Cook, Simon Dobson, Peter J. Denning, Robert Dunham, Jef Raskin, and Dennis Tsichritzis. Forum: When is a computer more like a guitar than a washing machine?; corroboration the only way to determine Web accuracy; how to teach critical thinking about Web content; create a random number service based on the Mersenne Twister; make fair uses a legal requirement in DRM systems; “The Missing Customer” redux; enthusiasm, drive, wisdom, patience not tied to age. Communications of the ACM, 46(7):11–13, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2003:TSMb

[1029] Michael Cusumano. Technology strategy and management: Beware the lure of the horizontal. Communications of the ACM, 46(7):15–17, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2003:PIA

[1030] Peter J. Denning. The profession of IT: Accomplishment. Communications of the ACM, 46(7):19–23, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Colwell:2003:VTI

[1031] Rita R. Colwell. Viewpoint: From terabytes to insights. Communications of the ACM, 46(7):25–27, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2003:IGE

[1032] Andrew Rosenbloom. Introduction: A game experience in every application. Communications of the ACM, 46(7):28–31, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Swartout:2003:MGS

[1033] William Swartout and Michael van Lent. Making a game of system design. Communications of the ACM, 46(7):32–39, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitton:2003:MVE

[1034] Mary C. Whitton. Making virtual environments compelling. Communications of the ACM, 46(7):40–47, July 2003.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jain:2003:EC

[1035] Ramesh Jain. Experiential computing. Communications of the ACM, 46(7):48–55, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tsang:2003:GLN

[1036] Michael Tsang, George Fitzmaurice, Gord Kurtenbach, and Azam Khan. Game-like navigation and responsiveness in non-game applications. Communications of the ACM, 46(7):56–61, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paradiso:2003:TCF

[1037] Joseph A. Paradiso. Tracking contact and free gesture across large interactive surfaces. Communications of the ACM, 46(7):62–69, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tomasi:2003:FSP

[1038] Carlo Tomasi, Abbas Rafii, and Ilhami Torunoglu. Full-size projection keyboard for handheld devices. Communications of the ACM, 46(7):70–75, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Breazeal:2003:IRT

[1039] Cynthia Breazeal, Andrew Brooks, Jesse Gray, Matt Hancher, John McBean, Dan Stiehl, and Joshua Strickon. Interactive robot theatre. Communications of the ACM, 46(7): 76–85, July 2003. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Vickers:2003:SSS

[1040] Paul Vickers and James L. Alty. Siren songs and swan songs debugging with music. Communications of the ACM, 46(7):86–93, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

May:2003:KMP

[1041] Daniel May and Paul Taylor. Knowledge management with patterns. Communications of the ACM, 46(7):94–99, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rees:2003:PPF

[1042] Jackie Rees, Subhajyoti Bandyopadhyay, and Eugene H. Spafford. PFIRES: a policy framework for information security. Communications of the ACM, 46(7):101–106, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bhattacharjee:2003:DMO

[1043] Sudip Bhattacharjee, Ram D. Gopal, and G. Lawrence Sanders. Digital music and online sharing: software piracy 2.0? Communications of the ACM, 46(7):107–111, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karp:2003:SX

[1044] Alan H. Karp. E-speak e-xplained. Communications of the ACM, 46 (7):112–118, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grover:2003:TIR

[1045] Varun Grover and Khawaja Saeed.

The telecommunication industry revisited: the changing pattern of partnerships. Communications of the ACM, 46(7):119–125, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Du:2003:MAD

[1046] Timon C. Du, Eldon Y. Li, and An-Pin Chang. Mobile agents in distributed network management. Communications of the ACM, 46(7):127–132, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brandt:2003:IMM

[1047] D. Scott Brandt and Lorna Uden. Insight into mental models of novice Internet searchers. Communications of the ACM, 46(7):133–136, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:CIC

[1048] Younghwa Lee, Zoonky Lee, and Kai R. T. Larsen. Coping with Internet channel conflict. Communications of the ACM, 46(7):137–142, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maamar:2003:TOM

[1049] Zakaria Maamar and Paul Labbé.

Technical opinion: Moving vs. inviting software agents. Communications of the ACM, 46(7):143–144, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levi:2003:IRH

[1050] Albert Levi. Inside risks: How secure is secure Web browsing? Communications of the ACM, 46(7):152, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPh

[1051] Diane Crawford. Editorial pointers. Communications of the ACM, 46(8):5, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTB

[1052] ACM Staff. News track; blood spills builds skills; canary in a chip; PlayStation superpower; personal shopper; learner and meaner; get me rewrite! Communications of the ACM, 46(8):9– 10, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FBF

[1053] Diane Crawford, Ralph Castain, Viktor Decyk, Ben Schwarz, Michael Metcalf, Rob Dublin, Robert L. Glass, Bob Morrison, Frans Swaalf, and Martin Henz. Forum: Backward and forward with the practical programmer; don’t forget experience in role definition; correct me if I’m wrong; article or ad? Communications of the ACM, 46(8):11–13, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2003:DVD

[1054] Hal Berghel. Digital village: The discipline of Internet forensics. Communications of the ACM, 46(8):15–20, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2003:PPM

[1055] Robert L. Glass. Practical programmer: A mugwump’s-eye view of Web work. Communications of the ACM, 46(8):21–23, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agresti:2003:VDI

[1056] William W. Agresti. Viewpoint: Discovery informatics. Communications of the ACM, 46(8):25–28, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dill:2003:VVT

[1057] David L. Dill, Bruce Schneier, and Barbara Simons. Viewpoint: Voting and technology: who gets to count your vote? paperless voting machines threaten the integrity of democratic process by what they don’t do. Communications of the ACM, 46(8):29–31, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeSutter:2003:IST

[1058] Bjorn De Sutter and Koen De Bosschere. Introduction: Software techniques for program compaction. Communications of the ACM, 46(8):32–34, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tip:2003:ELB

[1059] Frank Tip, Peter F. Sweeney, and Chris Laffra. Extracting library-based Java applications. Communications of the ACM, 46(8):35–40, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeBus:2003:PPC

[1060] Bruno De Bus, Daniel Kästner, Dominique Chanet, Ludo Van Put, and Bjorn De Sutter. Post-pass compaction techniques. Communications of the ACM, 46(8):41–46, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krishnaswamy:2003:MWI

[1061] Arvind Krishnaswamy and Rajiv Gupta. Mixed-width instruction sets. Communications of the ACM, 46 (8):47–52, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Debray:2003:CCD

[1062] Saumya Debray and William S. Evans. Cold code decompression at runtime. Communications of the ACM, 46(8):54–60, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evans:2003:GBC

[1063] William S. Evans and Christopher W. Fraser. Grammar-based compression of interpreted code. Communications of the ACM, 46(8):61–66, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vitharana:2003:RCC

[1064] Padmal Vitharana. Risks and challenges of component-based software development. Communications of the ACM, 46(8):67–72, August 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Potter:2003:HCM

[1065] Richard E. Potter. How CIOs manage their superiors’ expectations. Communications of the ACM, 46(8):74–79, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jackson:2003:UEI

[1066] Thomas W. Jackson, Ray Dawson, and Darren Wilson. Understanding email interaction increases organizational productivity. Communications of the ACM, 46(8):80–84, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Russell:2003:RQU

[1067] Bruce Russell and Sangit Chatterjee. Relationship quality: the undervalued dimension of software quality. Communications of the ACM, 46(8):85–89, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitman:2003:EGT

[1068] Michael E. Whitman. Enemy at the gate: threats to information security. Communications of the ACM, 46(8):91–95, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deVreede:2003:HSB

[1069] Gert-Jan de Vreede, Robert M. Davison, and Robert O. Briggs. How a silver bullet may lose its shine. Communications of the ACM, 46(8):96–101, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walsh:2003:AAA

[1070] Kenneth R. Walsh. Analyzing the application ASP concept: technologies, economies, and strategies. Communications of the ACM, 46(8):103–107, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ravichandran:2003:SRS

[1071] T. Ravichandran and Marcus A. Rothenberger. Software reuse strategies and component markets. Communications of the ACM, 46(8):109–114, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhiying:2003:CUE

[1072] Zhou Zhiying. CMM in uncertain environments. Communications of the ACM, 46(8):115–119, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nicopolitidis:2003:TOT

[1073] P. Nicopolitidis, G. I. Papadimitriou, M. S. Obaidat, and A. S. Pomportsis. Technical opinion: Third generation and beyond wireless systems. Communications of the ACM, 46(8):120–124, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2003:IRS

[1074] Lauren Weinstein. Inside risks: Spam wars. Communications of the ACM, 46(8):136, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPi

[1075] Diane Crawford. Editorial pointers. Communications of the ACM, 46(9):7, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTJ

[1076] CACM staff. News track: Juggling outsourcing stats; he said, she said; royal E-flush; plane precision; email evidence; VCR RIP? driver to distraction. Communications of the ACM, 46 (9):11–12, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FLM

[1077] Diane Crawford, Richard M. Stallman, Richard Proudfoot, Sudip Bhattacharjee, Ram D. Gopal, G. Lawrence Sanders, Tom Pittman, Richard R. Brooks, and Dennis E. Hamilton. Forum: Legalize music sharing now; hold the games; give up trying to halt the flow of information; praise for Denning’s powerful chord of professionalism. Communications of the ACM, 46 (9):13–15, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2003:SCP

[1078] Meg McGinity. Staying connected:

Power to the people. Communications of the ACM, 46(9):17–20, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goodman:2003:IPO

[1079] Seymour Goodman. International perspectives: The origins of digital computing in Europe. Communications of the ACM, 46(9):21–25, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2003:BSC

[1080] Phillip Armour. The business of software: Closing the learning application gap. Communications of the ACM, 46 (9):27–31, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arens:2003:VRU

[1081] Yigal Arens and Paul S. Rosenbloom. Viewpoint: Responding to the unexpected. Communications of the ACM, 46(9):33–35, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Devlin:2003:IWU

[1082] Keith Devlin. Introduction: Why universities require computer science students to take math. Communications of the ACM, 46(9):36–39, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruce:2003:WM

[1083] Kim B. Bruce, Robert L. Scot Drysdale, Charles Kelemen, and Allen Tucker. Why math? Communications of the ACM, 46(9):40–44, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henderson:2003:MRS

[1084] Peter B. Henderson. Mathematical reasoning in software engineering education. Communications of the ACM, 46 (9):45–50, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Almstrum:2003:WAC

[1085] Vicki L. Almstrum. What is the attraction to computing? Communications of the ACM, 46(9):51–55, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loo:2003:FPP

[1086] Alfred W. Loo. The future of peerto-peer computing. Communications of the ACM, 46(9):56–61, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ronsse:2003:RRN

[1087] Michiel Ronsse, Koen De Bosschere, Mark Christiaens, Jacques Chassin de Kergommeaux, and Dieter Kranzlmüller. Record/replay for nondeterministic program executions. Communications of the ACM, 46(9):62–67, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kankanhalli:2003:RIS

[1088] Atreyi Kankanhalli, Fransiska Tanudidjaja, Juliana Sutanto, and Bernard C. Y. Tan. The role of IT in successful knowledge management initiatives. Communications of the ACM, 46 (9):69–73, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pietro:2003:SPI

[1089] Roberto Di Pietro and Luigi V. Mancini. Security and privacy issues of handheld and wearable wireless devices. Communications of the ACM, 46 (9):74–79, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farhoomand:2003:CSC

[1090] Ali F. Farhoomand and Pauline S. P. Ng. Creating sustainable competitive advantage through internetworked communities. Communications of the ACM, 46(9):83–88, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Premkumar:2003:ADS

[1091] G. Prem Premkumar. Alternate distribution strategies for digital music. Communications of the ACM, 46(9): 89–95, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fichman:2003:IRC

[1092] Robert G. Fichman and Mary J. Cronin. Information-rich commerce at a crossroads: business and technology adoption requirements. Communications of the ACM, 46(9):96–102, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liao:2003:CCS

[1093] Ziqi Liao and Michael Tow Cheung. Concurrent-convergent strategy in IT consulting. Communications of the ACM, 46(9):103–104, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baldi:2003:TOO

[1094] Stefan Baldi, Hauke Heier, and Anett Mehler-Bicher. Technical opinion: Open courseware and open source software. Communications of the ACM, 46 (9):105–107, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2003:IRR

[1095] Peter G. Neumann. Inside risks: Risks in trusting untrustworthiness. Communications of the ACM, 46(9):120, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anttiroiko:2003:BSD

[1096] Ari-Veikko Anttiroiko. Building strong e-democracy: the role of technology in developing democracy for the information age. Communications of the ACM, 46(9):121–128, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chau:2003:OCP

[1097] Patrick Y. K. Chau and Simpson Poon.

Octopus: an e-cash payment system success story. Communications of the ACM, 46(9):129–133, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2003:D

[1098] John Erickson and Keng Siau. Education. Communications of the ACM, 46(9):134–140, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fernandes:2003:CFI

[1099] Kiran J. Fernandes, Vinesh Raja, and Julian Eyre. Cybersphere: the fully immersive spherical projection system. Communications of the ACM, 46(9): 141–146, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:2003:RDF

[1100] Lee A. Freeman. A refresher in data flow diagramming: an effective aid for analysts. Communications of the ACM, 46(9):147–151, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Govindarajulu:2003:EUW

[1101] Chittibabu Govindarajulu. End users: who are they? Communications of the ACM, 46(9):152–159, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goss:2003:WIT

[1102] Ernest Goss and Uma Gupta. Women and the Internet: is there an economic payoff? Communications of the ACM, 46(9):160–166, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hartzel:2003:HSE

[1103] Kathleen Hartzel. How self-efficacy and gender issues affect software adoption and use. Communications of the ACM, 46(9):167–171, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holowczak:2003:DWE

[1104] Richard D. Holowczak, Nabil R. Adam, Francisco J. Artigas, and Irfan Bora. Data warehousing in environmental digital libraries. Communications of the ACM, 46(9):172–178, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kock:2003:WSC

[1105] Ned Kock, Camille Auspitz, and Brad King. Web-supported course partnerships: bringing industry and academia together. Communications of the ACM, 46(9):179–183, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Devaraj:2003:LEI

[1106] Sarv Devaraj, Ming Fan, and Rajiv Kohli. E-loyalty: elusive ideal or competitive edge? Communications of the ACM, 46(9):184–191, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Laffey:2003:NSS

[1107] James M. Laffey, Dale Musser, Herbert Remidez, and Josh Gottdenker. Networked systems for schools that learn. Communications of the ACM, 46(9): 192–200, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lewis:2003:EAI

[1108] William Lewis, Jr., Richard T. Watson, and Ann Pickren. An empirical assessment of IT disaster risk. Communications of the ACM, 46(9): 201–206, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moores:2003:ENC

[1109] Trevor T. Moores. The effect of national culture and economic wealth on global software piracy rates. Communications of the ACM, 46(9):207–215, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneberger:2003:CCI

[1110] Scott L. Schneberger and Ephraim R. McLean. The complexity cross: implications for practice. Communications of the ACM, 46(9):216–225, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yeung:2003:CPC

[1111] Jeff Hoi Yan Yeung, J. P. Shim, and Andy Yin Kit Lai. Current progress of e-commerce adoption: small and medium enterprises in Hong Kong. Communications of the ACM, 46(9): 226–232, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Higa:2003:TEJ

[1112] Kunihiko Higa and Bongsik Shin. The telework experience in Japan. Communications of the ACM, 46(9):233–242, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2003:ETS

[1113] Nitin Singh. Emerging technologies to support supply chain management. Communications of the ACM, 46(9): 243–247, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agarwal:2003:OOM

[1114] Ritu Agarwal and Atish P. Sinha.

Object-oriented modeling with UML: a study of developers’ perceptions. Communications of the ACM, 46(9): 248–256, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanWijk:2003:RHD

[1115] Jarke J. van Wijk, Frank van Ham, and Huub van de Wetering. Rendering hierarchical data. Communications of the ACM, 46(9):257–263, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhu:2003:EDB

[1116] Kevin Zhu and Bryan MacQuarrie. The economics of digital bundling: the impact of digitization and bundling on the music industry. Communications of the ACM, 46(9):264–270, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPj

[1117] Diane Crawford. Editorial pointers. Communications of the ACM, 46(10):5, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTSa

[1118] ACM Staff. News track: Salt in the wound; coming into focus; study in slime; that’s courriel, S’il vous plaît; A healthy spritz; how does that make you feel?; dial-A-dolphin. Communications of the ACM, 46(10):9–10, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FTI

[1119] Diane Crawford, Peter Sommers, Hal Berghel, W. Terry Hardgrave, Chris Morris, and Jim Haynes. Forum: Tell it to the judge; what software engineering can do for Web development. Communications of the ACM, 46(10):11–13, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2003:LSU

[1120] Pamela Samuelson. Legally speaking: Unsolicited communications as trespass? Communications of the ACM, 46(10):15–20, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2003:CCS

[1121] Michael Cusumano. Technology strategy and management: Company character and the software business. Communications of the ACM, 46(10):21– 23, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Papazoglou:2003:ISO

[1122] M. P. Papazoglou and D. Georgakopoulos. Introduction: Service-oriented computing. Communications of the ACM, 46(10):24–28, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Curbera:2003:NSW

[1123] Francisco Curbera, Rania Khalaf, Nirmal Mukhi, Stefan Tai, and Sanjiva Weerawarana. The next step in Web services. Communications of the ACM, 46(10):29–34, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yang:2003:WSC

[1124] Jian Yang. Web service componentization. Communications of the ACM, 46(10):35–40, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meredith:2003:CT

[1125] L. G. Meredith and Steve Bjorg. Contracts and types. Communications of the ACM, 46(10):41–47, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Little:2003:TWS

[1126] Mark Little. Transactions and Web services. Communications of the ACM, 46(10):49–54, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Casati:2003:BOM

[1127] Fabio Casati, Eric Shan, Umeshwar Dayal, and Ming-Chien Shan. Business-oriented management of Web services. Communications of the ACM, 46(10):55–60, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dori:2003:ICM

[1128] Dov Dori. Introduction: Conceptual modeling and system architecting. Communications of the ACM, 46 (10):62–65, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soderborg:2003:SFA

[1129] Nathan R. Soderborg, Edward F. Crawley, and Dov Dori. System function and architecture: OPMbased definitions and operational templates. Communications of the ACM, 46(10):67–72, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henderson-Sellers:2003:MEO

[1130] Brian Henderson-Sellers. Method engineering for OO systems development. Communications of the ACM, 46(10):73–78, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gemino:2003:EMT

[1131] Andrew Gemino and Yair Wand. Evaluating modeling techniques based on models of learning. Communications of the ACM, 46(10):79–84, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shanks:2003:UOV

[1132] Graeme Shanks, Elizabeth Tansley, and Ron Weber. Using ontology to validate conceptual models. Communications of the ACM, 46(10):85–89, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moravec:2003:RAA

[1133] Hans Moravec. Robots, after all. Communications of the ACM, 46(10):90– 97, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kwok:2003:WBC

[1134] Sai Ho Kwok. Watermark-based copyright protection system security. Communications of the ACM, 46(10):98– 101, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

VanDenHeuvel:2003:IWS

[1135] Willem-Jan Van Den Heuvel and Zakaria Maamar. Intelligent Web services moving toward a framework to compose. Communications of the ACM, 46 (10):103–109, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katerattanakul:2003:OQR

[1136] Pairin Katerattanakul, Bernard Han, and Soongoo Hong. Objective quality ranking of computing journals. Communications of the ACM, 46(10): 111–114, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Xing:2003:TOP

[1137] Cong cong Xing and Boumediene Belkhouche. Technical opinion: On pseudo object-oriented programming considered harmful. Communications of the ACM, 46(10):115–117, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2003:ISS

[1138] Peter G. Neumann. Inside risks: Information system security redux. Communications of the ACM, 46(10):136, October 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPk

[1139] Diane Crawford. Editoral pointers. Communications of the ACM, 46(11):5, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTSb

[1140] ACM Staff. News track: Security holes; don’t look back; developing PC nations; hari today, brain surgery tomorrow?; IMing at work; now speak. Communications of the ACM, 46(11):9–10, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FCS

[1141] Diane Crawford, Joseph F. Traub, Reidar Conradi, Steve Johnson, Clinton Mah, Frédéric Guéron, David lowe, Bill Willcox, Rommert J. Casimir, Kevin Mackie, Mark Wallace, and Andries van Dam. Forum: CS student’s need other kinds of math, too; Web development not so different after all; for responsible spam, pay as you go; no innovation means no customers; computer allure moves east. Communications of the ACM, 46(11):11–13, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2003:PIG

[1142] Peter J. Denning. The profession of IT: Great principles of computing. Communications of the ACM, 46 (11):15–20, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2003:PPS

[1143] Robert L. Glass. Practical programmer: A sociopolitical look at open source. Communications of the ACM, 46(11):21–23, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sampson:2003:VMD

[1144] Geoffrey Sampson. Viewpoint: The myth of diminishing firms. Communications of the ACM, 46(11):25–28, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:2003:IBF

[1145] Maxine D. Brown. Introduction:

Blueprint for the future of highperformance networking. Communications of the ACM, 46(11):30–33, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DeFanti:2003:TGS

[1146] Tom DeFanti, Cees de Laat, Joe Mambretti, Kees Neggers, and Bill St. Arnaud. TransLight: a global-scale LambdaGrid for e-science. Communications of the ACM, 46(11):34–41, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Falk:2003:TPH

[1147] Aaron Falk, Ted Faber, Joseph Bannister, Andrew Chien, Robert Grossman, and Jason Leigh. Transport protocols for high performance. Communications of the ACM, 46(11):42–49, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Foster:2003:DIB

[1148] Ian Foster and Robert L. Grossman. Data integration in a bandwidth-rich world. Communications of the ACM, 46(11):50–57, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smarr:2003:O

[1149] Larry L. Smarr, Andrew A. Chien, Tom DeFanti, Jason Leigh, and Philip M. Papadopoulos. The OptIPuter. Communications of the ACM, 46(11):58–67, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Newman:2003:DIS

[1150] Harvey B. Newman, Mark H. Ellisman, and John A. Orcutt. Dataintensive e-science frontier research. Communications of the ACM, 46(11): 68–77, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barnes:2003:RSI

[1151] Stuart J. Barnes and Sid L. Huff. Rising sun: iMode and the wireless Internet. Communications of the ACM, 46 (11):78–84, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lorence:2003:PDM

[1152] Daniel P. Lorence. The perils of data misreporting. Communications of the ACM, 46(11):85–88, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Easley:2003:MOS

[1153] Robert F. Easley, John G. Michel, and Sarv Devaraj. The MP3 open standard and the music industry’s response to Internet piracy. Communications of the ACM, 46(11):90–96, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vitharana:2003:DRA

[1154] Padmal Vitharana, Fatemah Mariam Zahedi, and Hemant Jain. Design, retrieval, and assembly in componentbased software development. Communications of the ACM, 46(11):97–102, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peace:2003:BFS

[1155] A. Graham Peace. Balancing free speech and censorship: academia’s response to the Internet. Communications of the ACM, 46(11):104–109, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gorla:2003:FCD

[1156] Narasimhaiah Gorla. Features to consider in a data warehousing system. Communications of the ACM, 46(11): 111–115, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landry:2003:CTI

[1157] Jeffrey P. Landry, J. Harold Pardue, Herbert E. Longenecker, Jr., and David F. Feinstein. A common theme for IT degree programs. Communications of the ACM, 46(11):

117–120, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anselmo:2003:MPS

[1158] Donald Anselmo and Henry Ledgard. Measuring productivity in the software industry. Communications of the ACM, 46(11):121–125, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nordfors:2003:ICI

[1159] David Nordfors, Michel Bajuk, Lena Norberg, Jochen Brinkmann, and Dan Forbush. Introducing a country to Internet-enabled expert networks. Communications of the ACM, 46(11): 127–132, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2003:SO

[1160] Rebecca T. Mercuri and Peter G. Neumann. Inside risks: Security by obscurity. Communications of the ACM, 46(11):160, November 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:EPl

[1161] Diane Crawford. Editorial pointers. Communications of the ACM, 46(12):5, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2003:NTP

[1162] ACM Staff. News track: Pay as you go plan; you say “potato”; who types there?; homework and the Web; à la cart; shattering net speeds; gift bots. Communications of the ACM, 46(12):9–10, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2003:FKP

[1163] Diane Crawford, Simone Santini, Bill Buxton, Joanna Wolfe, Carl J. Mueller, and Halina Kaminski. Forum: Keep the pedagogical ahead of the technological; minds over math; requirements vs. components; mark every vote count. Communications of the ACM, 46(12):11–13, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Committee:2003:NER

[1164] ACM Nominating Committee. Nominees for elections and report of the ACM Nominating Committee. Communications of the ACM, 46(12):14, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2003:DVM

[1165] Hal Berghel. Digital village: Malware month. Communications of the ACM, 46(12):15–19, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2003:SWS

[1166] Rebecca T. Mercuri. Security watch:

Standards insecurity. Communications of the ACM, 46(12):21–25, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karp:2003:VEP

[1167] Alan H. Karp. Viewpoint: Enforce POLA on processes to control viruses. Communications of the ACM, 46(12):27–29, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Urbaczewski:2003:IMC

[1168] Andrew Urbaczewski, Joseph S. Valacich, and Leonard M. Jessup. Introduction: Mobile commerce: Opportunities and challenges. Communications of the ACM, 46(12):30–32, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stafford:2003:MCW

[1169] Thomas F. Stafford and Mark L. Gillenson. Mobile commerce: what it is and what it could be. Communications of the ACM, 46(12):33–34, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sarker:2003:UMH

[1170] Suprateek Sarker and John D. Wells. Understanding mobile handheld device use and adoption. Communications of the ACM, 46(12):35–40, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jarvenpaa:2003:MCC

[1171] Sirkka L. Jarvenpaa, Karl Reiner Lang, Yoko Takeda, and Virpi Kristiina Tuunainen. Mobile commerce at crossroads. Communications of the ACM, 46(12):41–44, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sun:2003:IRE

[1172] Jun Sun. Information requirement elicitation in mobile commerce. Communications of the ACM, 46(12):45–47, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2003:IDM

[1173] Young Eun Lee and Izak Benbasat.

Interface design for mobile commerce. Communications of the ACM, 46(12): 48–52, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Venkatesh:2003:UUM

[1174] Viswanath Venkatesh, V. Ramesh, and Anne P. Massey. Understanding usability in mobile commerce. Communications of the ACM, 46(12):53–56, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tarasewich:2003:DMC

[1175] Peter Tarasewich. Designing mobile commerce applications. Communications of the ACM, 46(12):57–60, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rao:2003:EML

[1176] Bharat Rao and Louis Minakakis. Evolution of mobile location-based services. Communications of the ACM, 46 (12):61–65, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spangler:2003:UDM

[1177] William E. Spangler, Mordechai Gal-Or, and Jerrold H. May. Using data mining to profile TV viewers. Communications of the ACM, 46(12):66–72, December 2003. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2003:SRO

[1178] Jim Q. Chen, Ted E. Lee, Ruidong Zhang, and Yue Jeff Zhang. Systems requirements for organizational learning. Communications of the ACM, 46 (12):73–78, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yan:2003:MDC

[1179] Xu Yan. Mobile data communications in China. Communications of the ACM, 46(12):80–85, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kishore:2003:RPI

[1180] Rajiv Kishore, H. R. Rao, K. Nam, S. Rajagopalan, and A. Chaudhury. A relationship perspective on IT outsourcing. Communications of the ACM, 46(12):86–92, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chenoweth:2003:MDD

[1181] Tim Chenoweth, David Schuff, and Robert St. Louis. A method for developing dimensional data marts. Communications of the ACM, 46(12):93–98, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bayrak:2003:RBD

[1182] Coskun Bayrak and Chad Davis. The relationship between distributed systems and open software development. Communications of the ACM, 46(12): 99–102, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2003:IRD

[1183] Lauren Weinstein. Inside risks: The devil you know. Communications of the ACM, 46(12):144, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ahuja:2003:EIO

[1184] Manju Ahuja, Babita Gupta, and Pushkala Raman. An empirical investigation of online consumer purchasing behavior. Communications of the ACM, 46(12):145–151, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bapna:2003:WSB

[1185] Ravi Bapna. When snipers become predators: can mechanism design save online auctions? Communications of the ACM, 46(12):152–158, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Basu:2003:AC

[1186] Amit Basu and Steve Muylle. Authentication in e-commerce. Communications of the ACM, 46(12):159–166, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davis:2003:FGA

[1187] Sid Davis, Keng Siau, and Kumar Dhenuvakonda. A fit-gap analysis of e-business curricula vs. industry needs. Communications of the ACM, 46(12): 167–177, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Prasarnphanich:2003:HCB

[1188] Pattarawan Prasarnphanich and Mark L. Gillenson. The hybrid clicks and bricks business model. Communications of the ACM, 46(12):178–185, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gosain:2003:ICB

[1189] Sanjay Gosain, Arvind Malhotra, Omar A. El Sawy, and Fadi Chehade. The impact of common e-business interfaces. Communications of the ACM, 46(12):186–195, December 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grazioli:2003:DUT

[1190] Stefano Grazioli and Sirkka L. Jarvenpaa. Deceived: under target online. Communications of the ACM, 46 (12):196–205, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saeed:2003:CSC

[1191] Khawaja A. Saeed, Varun Grover, and Yujong Hwang. Creating synergy with a clicks and mortar approach. Communications of the ACM, 46(12): 206–212, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holmes:2003:VEL

[1192] Andrew Holmes and John Ryan. Viewpoint: the end of the long boom. Communications of the ACM, 46(12): 213–216, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2003:LLV

[1193] Philip M. Johnson, Mette L. Moffett, and Brian T. Pentland. Lessons learned from VCommerce. Communications of the ACM, 46(12):217–225, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katerattanakul:2003:CVS

[1194] Pairin Katerattanakul and Keng Siau.

Creating a virtual store image. Communications of the ACM, 46(12): 226–232, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khalifa:2003:DIS

[1195] Mohamed Khalifa and Moez Limayem.

Drivers of Internet shopping. Communications of the ACM, 46(12):233–239, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chae:2003:WDA

[1196] Minhee Chae and Jinwoo Kim. What’s so different about the mobile Internet? Communications of the ACM, 46 (12):240–247, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liao:2003:CIB

[1197] Ziqi Liao and Michael Tow Cheung. Challenges to Internet e-banking. Communications of the ACM, 46(12): 248–250, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maamar:2003:CCM

[1198] Zakaria Maamar. Commerce, ecommerce, and m-commerce: what comes next? Communications of the ACM, 46(12):251–257, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martin:2003:HSW

[1199] David Martin, Hailin Wu, and Adil Alsaid. Hidden surveillance by Web sites: Web bugs in contemporary use. Communications of the ACM, 46(12): 258–264, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moores:2003:DPS

[1200] Trevor T. Moores and Gurpreet Dhillon. Do privacy seals in ecommerce really work? Communications of the ACM, 46(12):265–271, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Murray:2003:HCP

[1201] Kyle B. Murray and Gerald Häubl. A human capital perspective of skill acquisition and interface loyalty. Communications of the ACM, 46(12):272–278, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Premkumar:2003:PMM

[1202] G. Prem Premkumar. Perspectives of the e-marketplace by multiple stakeholders. Communications of the ACM, 46(12):279–288, December 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Proctor:2003:CPM

[1203] Robert W. Proctor, Kim-Phuong L.

Vu, Lawrence J. Najjar, Misha W. Vaughan, and Gavriel Salvendy. Content preparation and management for e-commerce Web sites. Communications of the ACM, 46(12):289–299, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Subramani:2003:KSI

[1204] Mani R. Subramani and Balaji Rajagopalan. Knowledge-sharing and influence in online social networks via viral marketing. Communications of the ACM, 46(12):300–307, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ranganathan:2003:MTB

[1205] C. Ranganathan, Vern Goode, and Arkalgud Ramaprasad. Managing the transition to bricks and clicks. Communications of the ACM, 46(12): 308–316, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2003:CBS

[1206] Dan Jong Kim, Manish Agrawal, Bharat Jayaraman, and H. Raghav Rao. A comparison of B2B e-service solutions. Communications of the ACM, 46(12):317–324, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salam:2003:CPR

[1207] A. F. Salam, H. R. Rao, and C. C.

Pegels. Consumer-perceived risk in ecommerce transactions. Communications of the ACM, 46(12):325–331, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fang:2003:CCR

[1208] Xiaowen Fang and Gavriel Salvendy. Customer-centered rules for design of ecommerce Web sites. Communications of the ACM, 46(12):332–336, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tang:2003:UIC

[1209] Fang-Fang Tang, Marcel G. Thom, Liang Toon Wang, Joai Ching Tan, Wai Yee Chow, and Xiaoyun Tang. Using insurance to create trust on the Internet. Communications of the ACM, 46(12):337–344, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tiwana:2003:TBP

[1210] Amrit Tiwana and Ephraim R. McLean. The tightrope to e-business project success. Communications of the ACM, 46(12):345–350, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Townsend:2003:VLB

[1211] Anthony M. Townsend and James T. Bennett. Viewpoint: living and bidding in an auction economy. Communications of the ACM, 46(12):351–353, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vaidyanathan:2003:FFF

[1212] Ganesh Vaidyanathan and Sarv Devaraj. A five-factor framework for analyzing online risks in e-businesses. Communications of the ACM, 46(12): 354–361, December 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPa

[1213] Diane Crawford. Editorial pointers.

Communications of the ACM, 47(1): 5, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTSa

[1214] ACM Staff. News track: Sound of silence; ‘and for out next nano number … ’; international enrollment slows; the body electric; calling all buoys; checked out. Communications of the ACM, 47(1):9–10, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FMS

[1215] Diane Crawford, H. Joel Jeffrey, Hans Moravec, Joseph Frisina, Douglas Lyon, Jef Raskin, John Craig, Joy Getha, and Art Gittleman. Forum: More startling AI predictions; mandate safety-critical software; limited math limits the CS curriculum; hands-on how we learn; not enough joy of computing. Communications of the ACM, 47(1):11–13, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2004:SCR

[1216] Meg McGinity. Staying connected:

RFID: is this game of tag fair play? Communications of the ACM, 47 (1):15–18, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2004:BSW

[1217] Phillip Armour. The business of software: When executives code. Communications of the ACM, 47(1):19–22, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:AARa

[1218] CACM Staff. ACM annual report for FY03. Communications of the ACM, 47(1):23–29, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oviatt:2004:IMI

[1219] Sharon Oviatt, Trevor Darrell, and Myron Flickner. Introduction: Multimodal interfaces that flex, adapt, and persist. Communications of the ACM, 47(1):30–33, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jain:2004:MS

[1220] Anil K. Jain and Arun Ross. Multibiometric systems. Communications of the ACM, 47(1):34–40, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2004:TMI

[1221] Philip R. Cohen and David R. McGee. Tangible multimodal interfaces for safety-critical applications. Communications of the ACM, 47(1):41–46, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pieraccini:2004:MCS

[1222] Roberto Pieraccini, Krishna Dayanidhi, Jonathan Bloom, Jean-Gui Dahan, Michael Phillips, Bryan R. Goodman, and K. Venkatesh Prasad. Multimodal conversational systems for automobiles. Communications of the ACM, 47(1):47–49, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nock:2004:MPF

[1223] Harriet J. Nock, Giridharan Iyengar, and Chalapathy Neti. Multimodal processing by finding common cause. Communications of the ACM, 47(1):51–56, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reeves:2004:GMU

[1224] Leah M. Reeves, Jennifer Lai, James A.

Larson, Sharon Oviatt, T. S. Balaji, Stéphanie Buisine, Penny Collings, Phil Cohen, Ben Kraal, Jean-Claude Martin, Michael McTear, T. V. Raman, Kay M. Stanney, Hui Su, and Qian Ying Wang. Guidelines for multimodal user interface design. Communications of the ACM, 47(1):57–59, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Turk:2004:CVI

[1225] Matthew Turk. Computer vision in the interface. Communications of the ACM, 47(1):60–67, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Deng:2004:CAS

[1226] Li Deng and Xuedong Huang. Challenges in adopting speech recognition. Communications of the ACM, 47(1):69–75, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2004:MCS

[1227] Cynthia F. Cohen, Stanley J. Birkin, Monica J. Garfield, and Harold W. Webb. Managing conflict in software testing. Communications of the ACM, 47(1):76–81, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwarzkopf:2004:EPI

[1228] Albert B. Schwarzkopf, Roberto J. Mejias, ‘Jon (Sean) Jasperson, Carol S. Saunders, and Hermann Gruenwald. Effective practices for IT skills staffing. Communications of the ACM, 47 (1):83–88, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cumberland:2004:UMB

[1229] B. Craig Cumberland, Joseph S. Valacich, and Leonard M. Jessup. Understanding meteor burst communications technologies. Communications of the ACM, 47(1):89–92, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aiken:2004:NCR

[1230] Robert J. Aiken, Javad Boroumand, and Stephen Wolff. Network and computing research infrastructure: back to the future. Communications of the ACM, 47(1):93–98, January 2004.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dattero:2004:PLG

[1231] Ronald Dattero and Stuart D. Galup.

Programming languages and gender. Communications of the ACM, 47(1): 99–102, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shin:2004:GSP

[1232] Seung Kyoon Shin, Ram D. Gopal, G. Lawrence Sanders, and Andrew B. Whinston. Global software piracy revisited. Communications of the ACM, 47(1):103–107, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Luo:2004:TBM

[1233] Wenhong Luo and Mohammad Najdawi. Trust-building measures: a review of consumer health portals. Communications of the ACM, 47(1): 108–113, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ranum:2004:IRB

[1234] Marcus J. Ranum. Inside risks: Believing in myths. Communications of the ACM, 47(1):144, January 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPb

[1235] Diane Crawford. Editorial pointers.

Communications of the ACM, 47(2):5, February 2004.

Staff:2004:NTR

[1236] ACM Staff. News track: Resizing America; little shop of software; what women want in wheels; virtual autopsies; coming of age; poetic programming. Communications of the ACM, 47(2):9–10, February 2004.

Crawford:2004:FDD

[1237] Diane Crawford, Steve Wildstrom, Hal Berghel, Charles Burnette, Jiming Liu, Peter J. Denning, Carl Binding, and Robert Swarr. Forum: Defense in depth gets the worm; put cognitive modes in CS and its curricula; don’t blame WAP [Wireless Application Protocol]; market share vs. peer approval in open source. Communications of the ACM, 47(2):11–13, February 2004.

Samuelson:2004:LSWa

[1238] Pamela Samuelson. Legally speaking: What’s at stake in MGM v. Grokster? Communications of the ACM, 47(2): 15–20, February 2004.

Pottie:2004:VPG

[1239] Gregory J. Pottie. Viewpoint: Privacy in the global e-village. Communications of the ACM, 47(2):21–23, February 2004.

Glass:2004:PPL

[1240] Robert L. Glass. Practical programmer: A look at the economics of open source. Communications of the ACM, 47(2):25–27, February 2004.

Sairamesh:2004:IIC

[1241] Jakka Sairamesh, Alison Lee, and Loretta Anania. Introduction: Information cities. Communications of the ACM, 47(2):28–31, February 2004.

Sproull:2004:MIC

[1242] Lee Sproull and John F. Patterson. Making information cities livable. Communications of the ACM, 47(2): 33–37, February 2004.

Churchill:2004:BDP

[1243] Elizabeth Churchill, Andreas Girgensohn, Les Nelson, and Alison Lee. Blending digital and physical spaces for ubiquitous community participation. Communications of the ACM, 47 (2):38–44, February 2004.

Ferguson:2004:OFI

[1244] Donald Ferguson, Jakka Sairamesh, and Stuart Feldman. Open frameworks for information cities. Communications of the ACM, 47(2):45–49, February 2004.

Kavassalis:2004:WMW

[1245] Petros Kavassalis, Stelios Lelis, Mahmoud Rafea, and Seif Haridi. What makes a Web site popular? Communications of the ACM, 47(2):50–55, February 2004.

vonAhn:2004:THC

[1246] Luis von Ahn, Manuel Blum, and John Langford. Telling humans and computers apart automatically. Communications of the ACM, 47(2):56–60, February 2004.

Lichtenstein:2004:PSD

[1247] Yossi Lichtenstein. Puzzles in software development contracting. Communications of the ACM, 47(2):61–65, February 2004.

Albrecht:2004:HCF

[1248] Conan C. Albrecht. How clean is the future of SOAP? Communications of the ACM, 47(2):66–68, February 2004.

Kishore:2004:HSM

[1249] Rajiv Kishore, Hong Zhang, and R. Ramesh. A helix-spindle model for ontological engineering. Communications of the ACM, 47(2):69–75, February 2004.

Brereton:2004:SCS

[1250] Pearl Brereton. The software customer/supplier relationship. Communications of the ACM, 47(2):77–81, February 2004.

Pierce:2004:ADQ

[1251] Elizabeth M. Pierce. Assessing data quality with control matrices. Communications of the ACM, 47(2):82–86, February 2004.

Chau:2004:TIT

[1252] Patrick Y. K. Chau and Paul Jen-Hwa Hu. Technology implementation for telemedicine programs. Communications of the ACM, 47(2):87–92, February 2004.

Egyhazy:2004:IAU

[1253] Csaba Egyhazy and Raj Mukherji. Interoperability architecture using RM-ODP. Communications of the ACM, 47(2):93–97, February 2004.

Schwartz:2004:TOH

[1254] Ruth Bolotin Schwartz and Michele C. Russo. Technical opinion: How to quickly find articles in the top IS journals. Communications of the ACM, 47 (2):98–101, February 2004.

Weinstein:2004:IRO

[1255] Lauren Weinstein. Inside risks: Outsourced and out of control. Communications of the ACM, 47(2):120, February 2004.

Crawford:2004:EPc

[1256] Diane Crawford. Editorial pointers.

Communications of the ACM, 47(3):5, March 2004.

Staff:2004:NTB

[1257] CACM Staff. News track: Blurred vision; (keep) phone home; love/hate relationship; image makeover; robot scientist; rooms with a view. Communications of the ACM, 47(3):9–10, March 2004.

Crawford:2004:FPL

[1258] Diane Crawford, Peter G. Raeth, Mark Wallace, David Hawley, Donald Anselmo, Henry Ledgard, Dwayne Nelson, Yitzchak Gottlieb, Alex Simonelis, Coskun Bayrak, and Chad Davis. Forum: Principles of logical thought before technical details; why reinvent when we have open source?; still seeking software productivity; spare me the self-service; lineage of the Internet’s Open Systems Foundation. Communications of the ACM, 47(3):11–13, March 2004.

Mercuri:2004:SWS

[1259] Rebecca T. Mercuri. Security watch:

Superscaled security. Communications of the ACM, 47(3):15–18, March 2004.

Staff:2004:AF

[1260] CACM Staff. ACM fellows. Communications of the ACM, 47(3):19–20, March 2004.

Armour:2004:BSB

[1261] Phillip G. Armour. The business of software: Beware of counting LOC. Communications of the ACM, 47(3): 21–24, March 2004.

Cusumano:2004:TSMa

[1262] Michael A. Cusumano. Technology strategy and management: Who is liable for bugs and security flaws in software? Communications of the ACM, 47(3):25–27, March 2004.

Messerly:2004:VHC

[1263] John G. Messerly. Viewpoint: How computer games affect CS (and other) students’ school performance. Communications of the ACM, 47(3):29–31, March 2004.

Yen:2004:ETH

[1264] John Yen. Emerging technologies for homeland security: Introduction. Communications of the ACM, 47(3):32–35, March 2004.

Popp:2004:CTT

[1265] Robert Popp, Thomas Armour, Ted Senator, and Kristen Numrych. Countering terrorism through information technology. Communications of the ACM, 47(3):36–43, March 2004.

Coffman:2004:GBT

[1266] Thayne Coffman, Seth Greenblatt, and Sherry Marcus. Graph-based technologies for intelligence analysis. Communications of the ACM, 47(3):45–47, March 2004.

Kogut:2004:PIG

[1267] Paul Kogut, John Yen, Yui Leung, Shuang Sun, Rui Wang, Ted Mielczarek, and Ben Hellar. Proactive information gathering for homeland security teams. Communications of the ACM, 47(3):48–50, March 2004.

Saydjari:2004:CDA

[1268] O. Sami Saydjari. Cyber defense: art to science. Communications of the ACM, 47(3):52–57, March 2004.

Bajcsy:2004:CDT

[1269] R. Bajcsy, T. Benzel, M. Bishop, B. Braden, C. Brodley, S. Fahmy, S. Floyd, W. Hardaker, A. Joseph, G. Kesidis, K. Levitt, B. Lindell, P. Liu, D. Miller, R. Mundy, C. Neuman, R. Ostrenga, V. Paxson, P. Porras, C. Rosenberg, J. D. Tygar, S. Sastry, D. Sterne, and S. F. Wu. Cyber defense technology networking and evaluation. Communications of the ACM, 47 (3):58–61, March 2004.

Sawyer:2004:MFR

[1270] Steve Sawyer, Andrea Tapia, Leonard Pesheck, and John Davenport. Mobility and the first responder. Communications of the ACM, 47(3):62–65, March 2004.

Murphy:2004:RRH

[1271] Robin R. Murphy. Rescue robotics for homeland security. Communications of the ACM, 47(3):66–68, March 2004.

Wang:2004:ADD

[1272] Gang Wang, Hsinchun Chen, and Homa Atabakhsh. Automatically detecting deceptive criminal identities. Communications of the ACM, 47(3): 70–76, March 2004.

Escudero-Pascual:2004:QLA

[1273] Alberto Escudero-Pascual and Ian Hosein. Questioning lawful access to traffic data. Communications of the ACM, 47(3):77–82, March 2004.

Dalal:2004:TIF

[1274] Nikunj P. Dalal, Manjunath Kamath, William J. Kolarik, and Eswar Sivaraman. Toward an integrated framework for modeling enterprise processes. Communications of the ACM, 47(3): 83–87, March 2004.

Suardi:2004:HMY

[1275] Luigi Suardi. How to manage your software product life cycle with MAUI. Communications of the ACM, 47(3): 89–94, March 2004.

Kim:2004:RPW

[1276] Eyong B. Kim and Marc J. Schniederjans. The role of personality in Web-based distance education courses. Communications of the ACM, 47(3): 95–98, March 2004.

Sethi:2004:WCS

[1277] Vikram Sethi, Ruth C. King, and James Campbell Quick. What causes stress in information system professionals? Communications of the ACM, 47 (3):99–102, March 2004.

Stamp:2004:IRR

[1278] Mark Stamp. Inside risks: Risks of monoculture. Communications of the ACM, 47(3):120, March 2004.

Crawford:2004:EPd

[1279] Diane Crawford. Editorial pointers. Communications of the ACM, 47(4):5, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTPa

[1280] CACM Staff. News track: Political agendas; spam rage; digital doubts; women got game; perfict fit; awkshun serchs. Communications of the ACM, 47(4):9–10, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FOS

[1281] Diane Crawford, David Monniaux, Ralph Johnson, Clark E. Hartsock, III, Ken Nickerson, Robert L. Glass, Marc Bellusci, and Alex Simonelis. Forum: Open source vs. capitalism and communism; what worker shortage?; design for pedagogy not just for software. Communications of the ACM, 47(4):11–13, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:2004:APD

[1282] John White. ACM’s Professional Development Centre expands learning opportunities. Communications of the ACM, 47(4):14, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2004:PIS

[1283] Peter J. Denning. The profession of IT:

The social life of innovation. Communications of the ACM, 47(4):15–19, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2004:DVWa

[1284] Hal Berghel and Natasa Brajkovska. Digital village: Wading into alternate data streams. Communications of the ACM, 47(4):21–27, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Uslaner:2004:VTO

[1285] Eric M. Uslaner. Viewpoint: Trust online, trust offline. Communications of the ACM, 47(4):28–29, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Miller:2004:IHC

[1286] Christopher A. Miller. Introduction: Human-computer etiquette: managing expectations with intentional agents. Communications of the ACM, 47(4): 30–34, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nass:2004:EEE

[1287] Clifford Nass. Etiquette equality: exhibitions and expectations of computer politeness. Communications of the ACM, 47(4):35–37, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bickmore:2004:URS

[1288] Timothy W. Bickmore. Unspoken rules of spoken interaction. Communications of the ACM, 47(4):38–44, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mishra:2004:EDE

[1289] Punya Mishra and Kathryn A. Hershey. Etiquette and the design of educational technology. Communications of the ACM, 47(4):45–49, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parasuraman:2004:TEH

[1290] Raja Parasuraman and Christopher A.

Miller. Trust and etiquette in highcriticality automated systems. Communications of the ACM, 47(4):51–55, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Preece:2004:EON

[1291] Jenny Preece. Etiquette online: from nice to necessary. Communications of the ACM, 47(4):56–61, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krishna:2004:MCC

[1292] S. Krishna, Sundeep Sahay, and Geoff Walsham. Managing cross-cultural issues in global software outsourcing. Communications of the ACM, 47(4):62–66, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wallace:2004:SPR

[1293] Linda Wallace and Mark Keil. Software project risks and their effect on outcomes. Communications of the ACM, 47(4):68–73, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ives:2004:DEP

[1294] Blake Ives, Kenneth R. Walsh, and Helmut Schneider. The domino effect of password reuse. Communications of the ACM, 47(4):75–78, April 2004.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hovav:2004:MAJ

[1295] Anat Hovav and Paul Gray. Managing academic e-journals. Communications of the ACM, 47(4):79–82, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nicopolitidis:2004:EWN

[1296] Petros Nicopolitidis, Georgios Papadimitriou, Mohammed S. Obaidat, and Andreas S. Pomportsis. The economics of wireless networks. Communications of the ACM, 47(4):83–86, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desouza:2004:MKD

[1297] Kevin C. Desouza and J. Roberto Evaristo. Managing knowledge in distributed projects. Communications of the ACM, 47(4):87–91, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Li:2004:ICI

[1298] Xiaotong Li. Informational cascades in IT adoption. Communications of the ACM, 47(4):93–97, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gerdes:2004:TOV

[1299] John Gerdes, Jr. Technical opinion: The viability of supporting anonymous employees. Communications of the ACM, 47(4):98–100, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horning:2004:IRC

[1300] Jim Horning. Inside risks: Coincidental risks. Communications of the ACM, 47(4):112, April 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPe

[1301] Diane Crawford. Editorial pointers.

Communications of the ACM, 47(5):5, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTC

[1302] CACM Staff. News track: Concerts to go: No foul; press this!; littlest champ; truth be told. Communications of the ACM, 47(5):9–10, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FWC

[1303] Diane Crawford, Michael J. Wax, Tom Pittman, Madhu Banavati, Jef Raskin, Rebecca T. Mercuri, Bob Ellis, and Paul Nicholson. Forum: Why count lines of code?; computers and consciousness; computers, ACM, and everyday existence; don’t dismiss computers in the classroom. Communications of the ACM, 47(5):11–13, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2004:SCA

[1304] Meg McGinity. Staying connected:

It’s alive: Reports of the demise of the telecom industry have been greatly exaggerated. Communications of the ACM, 47(5):15–18, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2004:PPMa

[1305] Robert L. Glass. Practical programmer: Matching methodology to problem domain. Communications of the ACM, 47(5):19–21, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grove:2004:VWP

[1306] Jeff Grove. Viewpoint: Wanted: public policies that foster creation of knowledge. Communications of the ACM, 47(5):23–25, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kumar:2004:INA

[1307] Kuldeep Kumar and Jos van Hillegersberg. Introduction: New architectures for financial services. Communications of the ACM, 47(5):26–30, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dandapani:2004:SFW

[1308] Krishnan Dandapani. Success and failure in Web-based financial services. Communications of the ACM, 47(5): 31–33, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Homann:2004:FVS

[1309] Ulrich Homann, Michael Rill, and Andreas Wimmer. Flexible value structures in banking. Communications of the ACM, 47(5):34–36, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pan:2004:AAF

[1310] Alberto Pan and Ángel Viña. An alternative architecture for financial data integration. Communications of the ACM, 47(5):37–40, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mallat:2004:MBS

[1311] Niina Mallat, Matti Rossi, and Virpi Kristiina Tuunainen. Mobile banking services. Communications of the ACM, 47(5):42–46, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chandrasekaran:2004:OGI

[1312] N. Chandrasekaran and Geert Ensing. ODC: a global IT services delivery model. Communications of the ACM, 47(5):47–49, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tas:2004:FSB

[1313] Jeroen Tas and Shyam Sunder. Financial services business process outsourcing. Communications of the ACM, 47(5):50–52, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zdanowicz:2004:DML

[1314] John S. Zdanowicz. Detecting money laundering and terrorist financing via data mining. Communications of the ACM, 47(5):53–55, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slewe:2004:WWR

[1315] Ton Slewe and Mark Hoogenboom. Who will rob you on the digital highway? Communications of the ACM, 47(5):56–60, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Devaraj:2004:HMR

[1316] Sarv Devaraj and S. Ramesh Babu. How to measure the relationship between training and job performance. Communications of the ACM, 47(5): 62–67, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Xia:2004:GCD

[1317] Weidong Xia and Gwanhoo Lee. Grasping the complexity of IS development projects. Communications of the ACM, 47(5):68–74, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2004:CLR

[1318] Dongsong Zhang, J. Leon Zhao, Lina Zhou, and Jay F. Nunamaker, Jr. Can e-learning replace classroom learning? Communications of the ACM, 47(5):75–79, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Segal:2004:AE

[1319] Gerald Segal. 2004 ACM elections. Communications of the ACM, 47(5): 80–89, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chiang:2004:IST

[1320] I. Robert Chiang and Vijay S. Mookerjee. Improving software team productivity. Communications of the ACM, 47(5):89–93, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Toshniwal:2004:TRM

[1321] Rishi Toshniwal and Dharma P.

Agrawal. Tracing the roots of markup languages. Communications of the ACM, 47(5):95–98, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arita:2004:TOC

[1322] Masanori Arita. Technical opinion:

Comma-free design for DNA words. Communications of the ACM, 47(5): 99–100, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2004:IRA

[1323] Peter J. Denning and Dorothy E. Denning. Inside risks: Artificial stupidity. Communications of the ACM, 47 (5):112, May 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPf

[1324] Diane Crawford. Editorial pointers.

Communications of the ACM, 47(6):5, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTSb

[1325] CACM Staff. News track: Special skills draft; MBA outsourcing track; ancient Egypt online; Rx PlayStation; camera ready; blogging Bahgdad. Communications of the ACM, 47(6):9–10, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FTW

[1326] Diane Crawford, Mark Sadler, Anonymous, Michael Cusumano, Alan Karp, Lee A. Hollaar, Chuck Walbourn, Thomas W. Moran, Seung Kyoon Shin, Ram D. Gopal, G. Lawrence Sanders, and Andrew B. Whinston. Forum: The threat from within; no excuses for Microsoft; how about a manageable password?; when is copyright infringement legal? Don’t miss the lessons in games; inconsistent statistics unconvincing. Communications of the ACM, 47(6):11–13, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2004:BSR

[1327] Phillip G. Armour. The business of software: Real work, necessary friction, optional chaos. Communications of the ACM, 47(6):15–18, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2004:LSWb

[1328] Pamela Samuelson. Legally speaking: Why reform the U.S. patent system? Communications of the ACM, 47(6):19–23, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lau:2004:VSC

[1329] Stephen Lau. Viewpoint: The spinning cube of potential doom. Communications of the ACM, 47(6):25–26, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glaser:2004:SGP

[1330] Allan Glaser, Snow Fu, and Mark Tumelty. On site: Growing a participatory programming environment. Communications of the ACM, 47(6):27–29, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Culler:2004:IWS

[1331] David E. Culler and Wei Hong. Introduction: Wireless sensor networks. Communications of the ACM, 47(6): 30–33, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Szewczyk:2004:HMS

[1332] Robert Szewczyk, Eric Osterweil, Joseph Polastre, Michael Hamilton, Alan Mainwaring, and Deborah Estrin. Habitat monitoring with sensor networks. Communications of the ACM, 47(6):34–40, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:2004:PEW

[1333] Jason Hill, Mike Horton, Ralph Kling, and Lakshman Krishnamurthy. The platforms enabling wireless sensor networks. Communications of the ACM, 47(6):41–46, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woo:2004:NSQ

[1334] Alec Woo, Sam Madden, and Ramesh Govindan. Networking support for query processing in sensor networks.

Communications of the ACM, 47(6): 47–52, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Perrig:2004:SWS

[1335] Adrian Perrig, John Stankovic, and David Wagner. Security in wireless sensor networks. Communications of the ACM, 47(6):53–57, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sipior:2004:SSM

[1336] Janice C. Sipior, Burke T. Ward, and P. Gregory Bonner. Should spam be on the menu? Communications of the ACM, 47(6):59–63, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vatanasombut:2004:HRO

[1337] Banphot Vatanasombut, Antonis C. Stylianou, and Magid Igbaria. How to retain online customers. Communications of the ACM, 47(6):64–70, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Looney:2004:EBM

[1338] Clayton A. Looney, Leonard M. Jessup, and Joseph S. Valacich. Emerging business models for mobile brokerage services. Communications of the ACM, 47(6):71–77, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gorla:2004:WSW

[1339] Narasimhaiah Gorla and Yan Wah Lam. Who should work with whom?: Building effective software project teams. Communications of the ACM, 47(6):79–82, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2004:PSC

[1340] Hal Berghel and David L. Sallach. A paradigm shift in computing and IT education. Communications of the ACM, 47(6):83–88, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2004:ARC

[1341] Robert L. Glass, V. Ramesh, and Iris Vessey. An analysis of research in computing disciplines. Communications of the ACM, 47(6):89–94, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ke:2004:SGS

[1342] Weiling Ke and Kwok Kee Wei.

Successful e-government in Singapore. Communications of the ACM, 47(6): 95–99, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wedekind:2004:TOT

[1343] Hartmut Wedekind and Erich Ortner.

Technical opinion: Toward universal literacy: from computer science upward. Communications of the ACM, 47(6):101–104, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2004:IRO

[1344] Peter G. Neumann. Inside risks: Optimistic optimization. Communications of the ACM, 47(6):112, June 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPg

[1345] Diane Crawford. Editorial pointers. Communications of the ACM, 47(7):5, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTD

[1346] CACM Staff. News track: Dominance lost; vocal cards; flexible flying; fantastic voyage; uninvited consequences; snooze control. Communications of the ACM, 47(7):9–10, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FDN

[1347] Diane Crawford, Industry Usability Reporting Project Steering Committee, National Institute of Standards and Technology, Peter J. Denning, Lennart Meier, Jim Farrans, Li Deng, Xuedong Huang, and James D. Mooney. Forum: Don’t neglect usability in the total cost of ownership; where’s the steak?; discourage the sale of credentials in anonymous employment; for voice interfaces, hold the SALT; look deeper for markup roots. Communications of the ACM, 47(7):11–13, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2004:PIF

[1348] Peter J. Denning. The profession of IT: The field of programmers myth. Communications of the ACM, 47(7):15–20, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mbarika:2004:STP

[1349] Victor W. A. Mbarika. On site:

Is telemedicine the panacea for Sub-Saharan Africa’s medical nightmare? Communications of the ACM, 47(7): 21–24, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2004:SWH

[1350] Rebecca T. Mercuri. Security watch:

The HIPAA-potamus in health care data security. Communications of the ACM, 47(7):25–28, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2004:TSMb

[1351] Michael A. Cusumano. Technology strategy and management: More lawyers than programmers? Microsoft must alter its strategy to avoid an increasingly litigious future. Communications of the ACM, 47(7):29–31, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mann:2004:VWG

[1352] Catherine L. Mann. Viewpoint: What global sourcing means for U.S. IT workers and for the U.S. economy. Communications of the ACM, 47(7):33–35, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:2004:IBI

[1353] Donna L. Hoffman, Thomas P. Novak, and Alladi Venkatesh. Has the Internet become indispensable? Communications of the ACM, 47(7):37–42, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jackson:2004:IIU

[1354] Linda A. Jackson, Alexander von Eye, Gretchen Barbatsis, Frank Biocca, Hiram E. Fitzgerald, and Yong Zhao. The impact of Internet use on the other side of the digital divide. Communications of the ACM, 47(7):43–47, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harrison:2004:HTF

[1355] Warren Harrison, George Heuston, Sarah Mocas, Mark Morrissey, and John Richardson. High-tech forensics: An Oregon police department successfully enlists tech-savvy citizens to serve as police reserve specialists. Communications of the ACM, 47(7):48–52, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schell:2004:UMO

[1356] George P. Schell. Universities marginalize online courses. Communications of the ACM, 47(7):53–56, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chari:2004:DI

[1357] Kaushal Chari and Saravanan Seshadri. Demystifying integration. Communications of the ACM, 47(7):58–63, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martinsons:2004:ECO

[1358] Maris G. Martinsons. ERP in China: one package, two profiles. Communications of the ACM, 47(7):65–68, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liang:2004:WWV

[1359] Huigang Liang, Yajiong Xue, William R.

Boulton, and Terry Anthony Byrd. Why Western vendors don’t dominate China’s ERP market. Communications of the ACM, 47(7):69–72, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Strader:2004:IPT

[1360] Troy J. Strader and Sridhar N. Ramaswami. Investor perceptions of traditional and online channels. Communications of the ACM, 47(7):73–76, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shanks:2004:RCC

[1361] Graeme Shanks, Elizabeth Tansley, and Ron Weber. Representing composites in conceptual modeling. Communications of the ACM, 47(7):77–80, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bolloju:2004:IQB

[1362] Narasimha Bolloju. Improving the quality of business object models using collaboration patterns. Communications of the ACM, 47(7):81–86, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cavusoglu:2004:MEI

[1363] Huseyin Cavusoglu, Birendra Mishra, and Srinivasan Raghunathan. A model for evaluating IT security investments. Communications of the ACM, 47(7): 87–92, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kocher:2004:IRI

[1364] Paul Kocher and Bruce Schneier. Inside risks: Insider risks in elections. Communications of the ACM, 47(7):104, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPh

[1365] Diane Crawford. Editorial pointers. Communications of the ACM, 47(8):5, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTWa

[1366] CACM Staff. News track: Who gets the last word? pay up or scram; smart stepping; China follows own drummer; elusive E-government; in the groove. Communications of the ACM, 47(8):9–10, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FFA

[1367] Diane Crawford, Donn B. Parker, Dani Oderbolz, Kurt Gunteroth, Nicholas Ourosoff, Michael J. Lewchuk, and Ian Perry. Forum: The folk art of information security needs an upgrade; don’t let patents inhibit software innovation; hold the methodologies: Give me tools; measuring project friction and chaos; less is more code. Communications of the ACM, 47(8):11–13, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2004:PPMb

[1368] Robert L. Glass. Practical programmer: The mystery of formal methods disuse. Communications of the ACM, 47(8):15–17, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:2004:VSD

[1369] Peter Freeman and David Hart. Viewpoint: A science of design for softwareintensive systems. Communications of the ACM, 47(8):19–21, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bush:2004:SCS

[1370] Martin E. Bush. On site: Computer science education in Japan. Communications of the ACM, 47(8):23–26, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2004:III

[1371] Andrew Rosenbloom. Introduction:

Interactive immersion in 3D graphics. Communications of the ACM, 47(8):28–31, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scharver:2004:DCI

[1372] Chris Scharver, Ray Evenhouse, Andrew Johnson, and Jason Leigh. Designing cranial implants in a haptic augmented reality environment. Communications of the ACM, 47(8):32–38, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lin:2004:PBV

[1373] Ming Lin, William Baxter, Vincent Scheib, and Jeremy Wendt. Physically based virtual painting. Communications of the ACM, 47(8):40–47, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lok:2004:TMR

[1374] Benjamin C. Lok. Toward the merging of real and virtual spaces. Communications of the ACM, 47(8):48–53, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Houston:2004:VDA

[1375] Mike Houston, Chris Niederauer, Maneesh Agrawala, and Greg Humphreys. Visualizing dynamic architectural environments. Communications of the ACM, 47(8):54–59, August 2004. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jacobs:2004:ADL

[1376] Charles Jacobs, Wil Li, Evan Schrier, David Bargeron, and David Salesin. Adaptive document layout. Communications of the ACM, 47(8):60–66, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burmester:2004:HPK

[1377] Mike Burmester and Yvo G. Desmedt. Is hierarchical public-key certification the next target for hackers? Communications of the ACM, 47(8):68–74, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oppliger:2004:CMN

[1378] Rolf Oppliger. Certified mail: the next challenge for secure messaging. Communications of the ACM, 47(8):75–79, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gallaugher:2004:EFF

[1379] John Gallaugher and Nigel Melville.

Electronic frontiers in foreign exchange trading. Communications of the ACM, 47(8):81–87, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:AAB

[1380] CACM Staff. 2004 ACM awards banquet: a night to celebrate. Communications of the ACM, 47(8):89–92, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Westfall:2004:DTR

[1381] Ralph D. Westfall. Does telecommuting really increase productivity? Communications of the ACM, 47(8):93–96, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paper:2004:VSI

[1382] David J. Paper, James A. Rodger, and Steven J. Simon. Voice says it all in the Navy. Communications of the ACM, 47(8):97–101, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elfatatry:2004:NSO

[1383] Ahmed Elfatatry and Paul Layzell. Negotiating in service-oriented environments. Communications of the ACM, 47(8):103–108, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2004:IRC

[1384] Lauren Weinstein. Inside risks: Close exposures of the digital kind. Communications of the ACM, 47(8):120, August 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPi

[1385] Diane Crawford. Editorial pointers. Communications of the ACM, 47(9):5, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTPb

[1386] CACM Staff. News track: Political lines; virtual visions; E-earache; sensors on ice; Coke promo fizzles. Communications of the ACM, 47(9):9–10, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FNS

[1387] Diane Crawford, Peter Wegner, Kurt Gunteroth, Joshua Stern, David Marks, and Peter J. Denning. Forum: No scientific discipline was ever built on a myth. Communications of the ACM, 47(9):11–13, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2004:SCW

[1388] Meg McGinity. Staying connected: Weaving a wireless safety net. Communications of the ACM, 47(9):15–18, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2004:DVWb

[1389] Hal Berghel. Digital village: Wireless infidelity I: war driving. Communications of the ACM, 47(9):21–26, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jacobs:2004:VPR

[1390] James A. Jacobs and Charles Humphrey.

Viewpoint: Preserving research data. Communications of the ACM, 47(9): 27–29, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sutcliffe:2004:IEU

[1391] Alistair Sutcliffe and Nikolay Mehandjiev. Introduction: End-user development. Communications of the ACM, 47(9):31–32, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fischer:2004:MDM

[1392] G. Fischer, E. Giaccardi, Y. Ye, A. G.

Sutcliffe, and N. Mehandjiev. Metadesign: a manifesto for end-user development. Communications of the ACM, 47(9):33–37, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beringer:2004:RET

[1393] Joerg Beringer. Reducing expertise tension. Communications of the ACM, 47(9):39–40, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wulf:2004:EEU

[1394] Volker Wulf and Matthias Jarke. The economics of end-user development. Communications of the ACM, 47(9): 41–42, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Repenning:2004:ABE

[1395] Alexander Repenning and Andri Ioannidou. Agent-based end-user development. Communications of the ACM, 47(9):43–46, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Myers:2004:NPL

[1396] Brad A. Myers, John F. Pane, and Andy Ko. Natural programming languages and environments. Communications of the ACM, 47(9):47–52, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burnett:2004:EUS

[1397] Margaret Burnett, Curtis Cook, and Gregg Rothermel. End-user software engineering. Communications of the ACM, 47(9):53–58, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morch:2004:CBT

[1398] Anders I. Mørch, Gunnar Stevens, Markus Won, Markus Klann, Yvonne Dittrich, and Volker Wulf. Componentbased technologies for end-user development. Communications of the ACM, 47(9):59–62, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berti:2004:NDU

[1399] Silvia Berti, Fabio Paternò, and Carmen Santoro. Natural development of ubiquitous interfaces. Communications of the ACM, 47(9):63–64, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blackwell:2004:EUD

[1400] Alan F. Blackwell. End-user developers at home. Communications of the ACM, 47(9):65–66, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Resnick:2004:CER

[1401] Paul J. Resnick, Derek L. Hansen, and Caroline R. Richardson. Calculating error rates for filtering software. Communications of the ACM, 47 (9):67–71, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Turetken:2004:SSA

[1402] Ozgur Turetken, David Schuff, Ramesh Sharda, and Terence T. Ow. Supporting systems analysis and design through fisheye views. Communications of the ACM, 47(9):72–77, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhuge:2004:TEG

[1403] Hai Zhuge and Xiaoqing Shi. Toward the eco-grid: a harmoniously evolved interconnection environment. Communications of the ACM, 47(9): 78–83, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hine:2004:EPD

[1404] Michael J. Hine, Mark Hill, David Ruth, Bob Carlson, Dave Banks, and Jim Troxell. Empowering persons with disabilities with decision-support technology. Communications of the ACM, 47(9):85–89, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lippert:2004:AVP

[1405] Susan K. Lippert and Murugan Anandarajan. Academic vs. practitioner systems planning and analysis. Communications of the ACM, 47(9):91–94, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bailes:2004:TOM

[1406] Jason E. Bailes and Gary F. Templeton. Technical opinion: Managing P2P security. Communications of the ACM, 47(9):95–98, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2004:IRB

[1407] Peter G. Neumann. Inside risks: The big picture. Communications of the ACM, 47(9):112, September 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPj

[1408] Diane Crawford. Editorial pointers. Communications of the ACM, 47(10):5, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTE

[1409] CACM Staff. News track: Experts dispute passport plan; Parliament endorses open access; A shadow over the valley; paint by numbers; talk of the future; emotional ride. Communications of the ACM, 47(10):9–10, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FCL

[1410] Diane Crawford, Scott Ambler, Peter G. Neumann, Peter Turk, Hartmut Wedekind, John Stanley, Julian V. Noble, Hong-Lok Li, Michael Blaha, and Dean Wakerley. Forum: Consider long-term implications now and forever; the real philosophical roots of literacy; give me common carrier Internet service; for science and math, skip the online course; prepare CS graduates for industrial reality; elegant code produces maximal comprehension. Communications of the ACM, 47(10):11–13, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2004:BSD

[1411] Phillip G. Armour. The business of software: Not-defect: the mature discipline of testing. Communications of the ACM, 47(10):15–18, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2004:HLH

[1412] David A. Patterson. Hot links: Hot links. Communications of the ACM, 47(10):19–20, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vonLohmann:2004:VVC

[1413] Fred von Lohmann. Viewpoint: Voluntary collective licensing for music file sharing. Communications of the ACM, 47(10):21–24, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2004:TSMc

[1414] Michael A. Cusumano. Technology strategy and management: Reflections on free and open software. Communications of the ACM, 47(10):25–27, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2004:IPP

[1415] Peter G. Neumann. Introduction: The problems and potentials of voting systems. Communications of the ACM, 47(10):28–30, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Deutsch:2004:VSS

[1416] Herb Deutsch and Stephen Berger. Voting systems standards and certifications. Communications of the ACM, 47(10):31–33, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coggins:2004:ITV

[1417] Carolyn Coggins. Independent testing of voting systems. Communications of the ACM, 47(10):34–38, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williams:2004:IVS

[1418] Brit J. Williams and Merle S. King.

Implementing voting systems: the Georgia method. Communications of the ACM, 47(10):39–42, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Franco:2004:SVM

[1419] Anthony Di Franco, Andrew Petro, Emmett Shear, and Vladimir Vladimirov. Small vote manipulations can swing elections. Communications of the ACM, 47(10):43–45, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:2004:AE

[1420] Douglas W. Jones. Auditing elections. Communications of the ACM, 47(10):46–50, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2004:CE

[1421] Rebecca T. Mercuri and L. Jean Camp.

The code of elections. Communications of the ACM, 47(10):52–57, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jefferson:2004:AIV

[1422] David Jefferson, Aviel D. Rubin, Barbara Simons, and David Wagner. Analyzing Internet voting security. Communications of the ACM, 47(10):59– 64, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kitcat:2004:SAV

[1423] Jason Kitcat. Source availability and e-voting: an advocate recants. Communications of the ACM, 47(10):65– 67, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grove:2004:ASV

[1424] Jeff Grove. ACM statement on voting systems. Communications of the ACM, 47(10):69–70, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2004:LLB

[1425] David A. Patterson. Latency lags bandwith. Communications of the ACM, 47(10):71–75, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2004:CWS

[1426] Huaiqing Wang and Shuozhong Wang. Cyber warfare: steganography vs. steganalysis. Communications of the ACM, 47(10):76–82, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samoladas:2004:OSS

[1427] Ioannis Samoladas, Ioannis Stamelos, Lefteris Angelis, and Apostolos Oikonomou. Open source software development should strive for even greater code maintainability. Communications of the ACM, 47(10):83– 87, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lightner:2004:ECF

[1428] Nancy J. Lightner. Evaluating ecommerce functionality with a focus on customer service. Communications of the ACM, 47(10):88–92, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2004:PTA

[1429] Michael Alan Smith. Portals: toward an application framework for interoperability. Communications of the ACM, 47(10):93–97, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Southard:2004:SOB

[1430] Peter B. Southard and Keng Siau. A survey of online e-banking retail initiatives. Communications of the ACM, 47 (10):99–102, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cegielski:2004:TOW

[1431] Casey G. Cegielski. Technical opinion: Who values technology certification? Communications of the ACM, 47 (10):103–105, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneier:2004:IRN

[1432] Bruce Schneier. Inside risks: The nonsecurity of secrecy. Communications of the ACM, 47(10):120, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPk

[1433] Diane Crawford. Editorial pointers. Communications of the ACM, 47 (11):5–6, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTSc

[1434] CACM Staff. News track: Spam spotting the DNA way; majoring in homeland security; in the driver’s seat; campus security; Bay Area attractions; hello, I must be going. Communications of the ACM, 47(11):9–10, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FDP

[1435] Diane Crawford, Alex Simonelis, Mike Brenner, Spike McLeary, Jack Nilles, and Ralph D. Westfall. Forum: Design principles are where you find them; the computer didn’t make you do it; why doesn’t everyone telecommute? Communications of the ACM, 47(11):11–13, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2004:PIN

[1436] Peter J. Denning. The profession of IT: Network laws. Communications of the ACM, 47(11):15–20, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:HLTa

[1437] Diane Crawford. Hot links: Top 10 downloads from ACM’s Portal. Communications of the ACM, 47(11):21–22, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2004:PPR

[1438] Robert L. Glass. Practical programmer: Is this a revolutionary idea, or not? Communications of the ACM, 47 (11):23–25, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Matloff:2004:VGA

[1439] Norman Matloff. Viewpoint: Globalization and the American IT worker. Communications of the ACM, 47(11): 27–29, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stewart:2004:IBT

[1440] Craig A. Stewart. Introduction: Bioinformatics: transforming biomedical research and medical care. Communications of the ACM, 47(11):30–33, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bader:2004:CBH

[1441] David A. Bader. Computational biology and high-performance computing. Communications of the ACM, 47 (11):34–41, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ebisuzaki:2004:PC

[1442] Toshikazu Ebisuzaki, Robert Germain, and Makoto Taiji. PetaFLOPS computing. Communications of the ACM, 47(11):42–45, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Overbeek:2004:SPP

[1443] Ross Overbeek, Terry Disz, and Rick Stevens. The SEED: a peer-to-peer environment for genome annotation. Communications of the ACM, 47(11): 46–51, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ellisman:2004:ERB

[1444] Mark Ellisman, Michael Brady, David Hart, Fang-Pang Lin, Matthias Müller, and Larry Smarr. The emerging role of biogrids. Communications of the ACM, 47(11):52–57, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Javahery:2004:BPM

[1445] Homa Javahery, Ahmed Seffah, and Thiruvengadam Radhakrishnan. Beyond power: making bioinformatics tools user-centered. Communications of the ACM, 47(11):58–63, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2004:BCV

[1446] Chris R. Johnson, Rob MacLeod, Steven G. Parker, and David Weinstein. Biomedical computing and visualization software environments. Communications of the ACM, 47(11):64–71, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tiwana:2004:OMR

[1447] Amrit Tiwana and Mark Keil. The one-minute risk assessment tool. Communications of the ACM, 47(11):73–77, November 2004. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Thong:2004:WLA

[1448] James Y. L. Thong, Weiyin Hong, and Kar Yan Tam. What leads to acceptance of digital libraries? Communications of the ACM, 47(11):78–83, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Edgington:2004:AOF

[1449] Theresa Edgington, Beomjin Choi, Katherine Henson, T. S. Raghu, and Ajay Vinze. Adopting ontology to facilitate knowledge sharing. Communications of the ACM, 47(11):85–90, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ganapathy:2004:VST

[1450] Shobha Ganapathy, C. Ranganathan, and Balaji Sankaranarayanan. Visualization strategies and tools for enhancing customer relationship management. Communications of the ACM, 47 (11):92–99, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Biehl:2004:RRD

[1451] Markus Biehl, Edmund Prater, and John R. McIntyre. Remote repair, diagnostics, and maintenance. Communications of the ACM, 47(11):100–106, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stafford:2004:WAL

[1452] Thomas F. Stafford and Dennis Gonier. What Americans like about being online. Communications of the ACM, 47 (11):107–112, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Donaldson:2004:TON

[1453] Ken Donaldson and Vicki Stone. Technical opinion: Nanoscience fact versus fiction. Communications of the ACM, 47(11):113–115, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vora:2004:IRE

[1454] P. L. Vora, B. Adida, R. Bucholz, D. Chaum, D. L. Dill, D. Jefferson, D. W. Jones, W. Lattin, A. D. Rubin, M. I. Shamos, and M. Yung. Inside risks: Evaluation of voting systems. Communications of the ACM, 47(11):144, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:EPl

[1455] Diane Crawford. Editorial pointers. Communications of the ACM, 47(12):5, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:NTWb

[1456] CACM Staff. News track: A world connected; invisa-bills; starting over; bots wild; creative measures; by George! Communications of the ACM, 47(12):9–10, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:FFE

[1457] Diane Crawford, Ben Bederson, Harry Hochheiser, Jeff Johnson, Clare-Marie Karat, Jonathan Lazar, John David, Frances C. Bernstein, Andrew D. Wolfe, Jr., Hong-Lok Li, Mark Burgin, and Nicholas Roeder. Forum: For a fair election, ensure usability; don’t skip secrecy if security is the goal; let users program their own systems; put more science in computer science; what we don’t know. Communications of the ACM, 47(12):11–13, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2004:DVWc

[1458] Hal Berghel and Jacob Uecker. Digital village: Wireless infidelity II: airjacking. Communications of the ACM, 47 (12):15–20, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2004:HLTb

[1459] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 47(12):21–22, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2004:PLH

[1460] David A. Patterson. President’s letter: The health of research conferences and the dearth of big idea papers. Communications of the ACM, 47 (12):23–24, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2004:SWM

[1461] Rebecca T. Mercuri. Security watch: The many colors of multimedia security. Communications of the ACM, 47 (12):25–29, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2004:ISI

[1462] Andrew Rosenbloom. Introduction: Special issue: The blogosphere. Communications of the ACM, 47(12):30–33, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kumar:2004:SEB

[1463] Ravi Kumar, Jasmine Novak, Prabhakar Raghavan, and Andrew Tomkins. Structure and evolution of blogspace. Communications of the ACM, 47(12): 35–39, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nardi:2004:WWB

[1464] Bonnie A. Nardi, Diane J. Schiano, Michelle Gumbrecht, and Luke Swartz. Why we blog. Communications of the ACM, 47(12):41–46, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cayzer:2004:SBD

[1465] Steve Cayzer. Semantic blogging and decentralized knowledge management. Communications of the ACM, 47 (12):47–52, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blood:2004:HBS

[1466] Rebecca Blood. How blogging software reshapes the online community. Communications of the ACM, 47(12): 53–55, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sunstein:2004:DF

[1467] Cass R. Sunstein. Democracy and filtering. Communications of the ACM, 47(12):57–59, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Birman:2004:VLI

[1468] Kenneth P. Birman. Viewpoint: Like it or not, Web services are distributed objects. Communications of the ACM, 47 (12):60–62, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2004:AARb

[1469] CACM Staff. ACM annual report for FY04. Communications of the ACM, 47(12):63–69, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Seffah:2004:OMU

[1470] Ahmed Seffah and Eduard Metzker.

The obstacles and myths of usability and software engineering. Communications of the ACM, 47(12):71–76, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bharati:2004:UCC

[1471] Pratyush Bharati and Abhijit Chaudhury. Using choiceboards to create business value. Communications of the ACM, 47(12):77–81, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loiacono:2004:CWA

[1472] Eleanor T. Loiacono. Cyberaccess: Web accessibility and corporate America. Communications of the ACM, 47 (12):82–87, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malhotra:2004:BSU

[1473] Yogesh Malhotra and Dennis F. Galletta. Building systems that users want to use. Communications of the ACM, 47(12):88–94, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sawyer:2004:SDT

[1474] Steve Sawyer. Software development teams. Communications of the ACM, 47(12):95–99, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Werthner:2004:CT

[1475] Hannes Werthner and Francesco Ricci. E-commerce and tourism. Communications of the ACM, 47(12):101–105, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bellovin:2004:IRS

[1476] Steven M. Bellovin. Inside risks: Spamming, phishing, authentication, and privacy. Communications of the ACM, 47(12):144, December 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPa

[1477] Diane Crawford. Editorial pointers.

Communications of the ACM, 48(1): 5, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTSa

[1478] CACM Staff. News track: Sniffing out trouble; home improvements; patent number 977; on the nose; take a number; sick thoughts. Communications of the ACM, 48(1):9–10, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FBL

[1479] Diane Crawford, Brad Kidwell, Jeffrey P. Buzen, David A. Patterson, Bradley Dilger, Brit J. Williams, Merle S. King, Chris Morris, David H. Jameson, Ralph Johnson, I. Samoladas, I. Stamelos, and L. Angelis. Forum: Bandwidth, latency, and physical reality; share disturbing information on e-voting; ensure quality assurance for bioinformatics applications; follow the money; share open source sources. Communications of the ACM, 48(1):11–13, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2005:BSU

[1480] Phillip G. Armour. The business of software: The unconscious art of software testing. Communications of the ACM, 48(1):15–18, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTa

[1481] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(1):19–20, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGinity:2005:SCL

[1482] Meg McGinity. Staying connected: Let your fingers do the talking. Communications of the ACM, 48(1):21–23, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLM

[1483] David A. Patterson. President’s letter: Minority-minority and minoritymajority technology transfer. Communications of the ACM, 48(1):25–26, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2005:VGL

[1484] Deborah Anderson. Viewpoint: Global linguistic diversity for the Internet. Communications of the ACM, 48 (1):27–28, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stephenson:2005:VCN

[1485] Chris Stephenson. Viewpoint: Creating a national K–12 computer science community. Communications of the ACM, 48(1):29–31, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Druin:2005:IDC

[1486] Allison Druin and Juan Pablo Hourcade. Interaction design and children. Communications of the ACM, 48(1):32–34, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kestenbaum:2005:CIW

[1487] David Kestenbaum. The challenges of IDC: what have we learned from our past? Communications of the ACM, 48(1):35–38, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guha:2005:WYC

[1488] Mona Leigh Guha, Allison Druin, Gene Chipman, Jerry Alan Fails, Sante Simms, and Allison Farber. Working with young children as technology design partners. Communications of the ACM, 48(1):39–42, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoysniemi:2005:CIG

[1489] Johanna Höysniemi, Perttu Hämäläinen, Laura Turkki, and Teppo Rouvi. Children’s intuitive gestures in vision-based action games. Communications of the ACM, 48(1):44–50, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eisenberg:2005:MSE

[1490] Michael Eisenberg. The material side of educational technology. Communications of the ACM, 48(1):51–54, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rogers:2005:ULI

[1491] Yvonne Rogers, Sara Price, Cliff Randell, Danae Stanton Fraser, Mark Weal, and Geraldine Fitzpatrick. Ubilearning integrates indoor and outdoor experiences. Communications of the ACM, 48(1):55–59, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robertson:2005:SCV

[1492] Judy Robertson and Judith Good. Story creation in virtual game worlds. Communications of the ACM, 48 (1):61–65, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mahapatra:2005:EEU

[1493] RadhaKanta Mahapatra and Vincent S. Lai. Evaluating end-user training programs. Communications of the ACM, 48(1):66–70, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holzinger:2005:UEM

[1494] Andreas Holzinger. Usability engineering methods for software developers. Communications of the ACM, 48(1):71–74, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tillquist:2005:UAS

[1495] John Tillquist and Waymond Rodgers. Using asset specificity and asset scope to measure the value of IT. Communications of the ACM, 48(1):75–80, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fonseca:2005:LWS

[1496] Rodrigo Fonseca, Virgílio Almeida, and Mark Crovella. Locality in a web of streams. Communications of the ACM, 48(1):82–88, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vaidyanathan:2005:FET

[1497] Ganesh Vaidyanathan. A framework for evaluating third-party logistics. Communications of the ACM, 48(1):89–94, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ettredge:2005:TIR

[1498] Michael Ettredge and John Gerdes Jr.

Timeliness of investor relations data at corporate Web sites. Communications of the ACM, 48(1):95–100, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herrero:2005:UDM

[1499] Ines Herrero and Jose L. Salmeron. Using the DEA methodology to rank software technical efficiency. Communications of the ACM, 48(1):101–105, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barnes:2005:TOA

[1500] Stuart J. Barnes. Technical opinion: Assessing the value of IS journals. Communications of the ACM, 48 (1):110–112, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ledin:2005:IRT

[1501] George Ledin, Jr. Inside risks: Not teaching viruses and worms is harmful. Communications of the ACM, 48(1):144, January 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPb

[1502] Diane Crawford. Editorial pointers. Communications of the ACM, 48(2):5, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTAa

[1503] CACM Staff. News track: Not all fun and games; IT hiring up; what lies Ahead; IT budget ups and downs; phishing expeditions; the (really) big picture; number one “blog”. Communications of the ACM, 48(2):9–10, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.tpd.tno.
nl/smartsite600.html.

Crawford:2005:FGA

[1504] Diane Crawford, Serdar Tasiran, Norman Matloff, Dana Honeycutt, Cass R. Sunstein, Wolter Pieters, Joseph R. Kintry, Tyrone Grandison, and Peter G. Neumann. Forum: A global audience deserves a global perspective; blogs no threat to democracy; Internet voting not impossible; don’t trade privacy for security. Communications of the ACM, 48(2):11–13, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLW

[1505] David A. Patterson. President’s letter: Why join ACM? Communications of the ACM, 48(2):14, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2005:TSMa

[1506] Michael A. Cusumano. Technology strategy and management: Google: what it is and what it is not. Communications of the ACM, 48(2):15–17, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crowcroft:2005:VNC

[1507] Jon Crowcroft. Viewpoint: On the nature of computing. Communications of the ACM, 48(2):19–20, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:SACa

[1508] CACM Staff. SIGs announce candidates for election. Communications of the ACM, 48(2):21–22, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTb

[1509] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(2):23–24, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Metaxas:2005:MIM

[1510] Dimitris Metaxas. Medical image modeling tools and applications. Communications of the ACM, 48(2):26–29, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delingette:2005:HSS

[1511] Herve Delingette and Nicholas Ayache. Hepatic surgery simulation. Communications of the ACM, 48(2):31–36, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaufman:2005:VC

[1512] Arie E. Kaufman, Sarang Lakare, Kevin Kreeger, and Ingmar Bitter. Virtual colonoscopy. Communications of the ACM, 48(2):37–41, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Park:2005:VHM

[1513] Kyoungju Park, Albert Montillo, Dimitris Metaxas, and Leon Axel. Volumetric heart modeling and analysis. Communications of the ACM, 48 (2):43–48, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Imielinska:2005:IVA

[1514] Celina Imielinska and Pat Molholt. Incorporating 3D virtual anatomy into the medical curriculum. Communications of the ACM, 48(2):49–54, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yoo:2005:OSS

[1515] Terry S. Yoo and Michael J. Ackerman. Open source software for medical image processing and visualization. Communications of the ACM, 48(2):55–59, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NICa

[1516] CACM Staff. Notice of intent to decharter professional chapters. Communications of the ACM, 48(2):61–62, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NICb

[1517] CACM Staff. Notice of intent to decharter student chapters. Communications of the ACM, 48(2):63–65, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Evaristo:2005:CMP

[1518] J. Roberto Evaristo, Kevin C. Desouza, and Kevin Hollister. Centralization momentum: the pendulum swings back again. Communications of the ACM, 48(2):66–71, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Salam:2005:TC

[1519] A. F. Salam, Lakshmi Iyer, Prashant Palvia, and Rahul Singh. Trust in e-commerce. Communications of the ACM, 48(2):72–77, February 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bodin:2005:EIS

[1520] Lawrence D. Bodin, Lawrence A. Gordon, and Martin P. Loeb. Evaluating information security investments using the analytic hierarchy process. Communications of the ACM, 48(2):78–83, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nah:2005:VMA

[1521] Fiona Fui-Hoon Nah, Keng Siau, and Hong Sheng. The value of mobile applications: a utility company study. Communications of the ACM, 48(2):85–90, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rainer:2005:EDA

[1522] R. Kelly Rainer, Jr. and Mark D. Miller. Examining differences across journal rankings. Communications of the ACM, 48(2):91–94, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yuan:2005:IMC

[1523] Yufei Yuan and Brian Detlor. Intelligent mobile crisis response systems. Communications of the ACM, 48(2):95–98, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hu:2005:TOD

[1524] Chenglie Hu. Technical opinion: Dataless objects considered harmful. Communications of the ACM, 48(2):99–101, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Becker:2005:TOG

[1525] Shirley Ann Becker. Technical opinion: E-government usability for older adults. Communications of the ACM, 48(2):102–104, February 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2005:IRR

[1526] Peter G. Neumann. Inside risks: Responsibilities of technologists. Communications of the ACM, 48(2):128, February 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPc

[1527] Diane Crawford. Editorial pointers.

Communications of the ACM, 48(3):5, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTO

[1528] CACM Staff. News track: Ominous clouds; in the wrong hands; China’s netizens to soar; chess champs in check; painful plateau; war is no game. Communications of the ACM, 48(3):9–10, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FBS

[1529] Diane Crawford, Marius Matioc, Steven M. Bellovin, Richard Hubert, Andrew D. Wolfe, Jr., David Foulser, and Andrew R. Kilner. Forum: To block spam, demand sender authentication; not revolutionary (thank goodness); how to know the known from the unknowns; user first in user-centered design. Communications of the ACM, 48(3):11–13, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:AF

[1530] CACM Staff. ACM fellows. Communications of the ACM, 48(3):14, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLC

[1531] David A. Patterson. President’s letter: 20^{th} century vs. 21^{st} century C&C: the SPUR manifesto. Communications of the ACM, 48(3):15–16, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2005:BSPa

[1532] Phillip G. Armour. The business of software: Project portfolios: organizational management of risk. Communications of the ACM, 48(3):17–20, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTc

[1533] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(3):21–22, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conti:2005:VWC

[1534] Gregory Conti. Viewpoint: Why computer scientists should attend hacker conferences. Communications of the ACM, 48(3):23–24, March 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2005:PPF

[1535] Robert L. Glass. Practical programmer: The first business application: a significant milestone in software history. Communications of the ACM, 48(3):25–26, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2005:LSL

[1536] Pamela Samuelson. Legally speaking: Legislative challenges to the sony safe harbor rule. Communications of the ACM, 48(3):27–31, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Streitz:2005:DC

[1537] Norbert Streitz and Paddy Nixon. The disappearing computer. Communications of the ACM, 48(3):32–35, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borriello:2005:DRW

[1538] Gaetano Borriello, Matthew Chalmers, Anthony LaMarca, and Paddy Nixon. Delivering real-world ubiquitous location systems. Communications of the ACM, 48(3):36–41, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Russell:2005:BDC

[1539] Daniel M. Russell, Norbert A. Streitz, and Terry Winograd. Building disappearing computers. Communications of the ACM, 48(3):42–48, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coutaz:2005:CK

[1540] Joëlle Coutaz, James L. Crowley, Simon Dobson, and David Garlan. Context is key. Communications of the ACM, 48(3):49–53, March 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Benford:2005:BPD

[1541] Steve Benford, Carsten Magerkurth, and Peter Ljungstrand. Bridging the physical and digital in pervasive gaming. Communications of the ACM, 48(3):54–57, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lahlou:2005:PTI

[1542] Saadi Lahlou, Marc Langheinrich, and Carsten Röcker. Privacy and trust issues with invisible computers. Communications of the ACM, 48(3):59–60, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bannon:2005:HDC

[1543] Liam Bannon, Steve Benford, John Bowers, and Christian Heath. Hybrid design creates innovative museum experiences. Communications of the ACM, 48(3):62–65, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gellersen:2005:SCA

[1544] Hans Gellersen. Smart-its: computers for artifacts in the physical world. Communications of the ACM, 48(3):66, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lalis:2005:AHC

[1545] Spyros Lalis, Alexandros Karypidis, and Anthony Savidis. Ad-hoc composition in wearable and mobile computing. Communications of the ACM, 48(3):67–68, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kameas:2005:EG

[1546] Achilles Kameas and Irene Mavrommati. Extrovert gadgets. Communications of the ACM, 48(3):69, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Paradiso:2005:SM

[1547] Joseph A. Paradiso. Sensate media.

Communications of the ACM, 48(3):70, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gershman:2005:ECA

[1548] Anatole Gershman and Andrew Fano.

Examples of commercial applications of ubiquitous computing. Communications of the ACM, 48(3):71, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2005:CSB

[1549] Jacques Cohen. Computer science and bioinformatics. Communications of the ACM, 48(3):72–78, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sen:2005:CDW

[1550] Arun Sen and Atish P. Sinha. A comparison of data warehousing methodologies. Communications of the ACM, 48(3):79–84, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:SACb

[1551] CACM Staff. SIGs announce candidates for election. Communications of the ACM, 48(3):85, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moores:2005:DCU

[1552] Trevor Moores. Do consumers understand the role of privacy seals in e-commerce? Communications of the ACM, 48(3):86–91, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Montano:2005:MSU

[1553] Bonnie Rubenstein Montano, Victoria Yoon, Stuart Lowry, and Teresa Merlau. A multiagent system for U.S. defense research contracting. Communications of the ACM, 48(3):93–97, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hibbard:2005:JDC

[1554] William Hibbard, Curtis Rueden, Steve Emmerson, Tom Rink, David Glowacki, Tom Whittaker, Don Murray, David Fulker, and John Anderson. Java distributed components for numerical visualization in VisAD. Communications of the ACM, 48(3):98–104, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mao:2005:SUC

[1555] Ji-Ye Mao, Karel Vredenburg, Paul W. Smith, and Tom Carey. The state of user-centered design practice. Communications of the ACM, 48(3):105–109, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2005:HMC

[1556] Xiaoni Zhang and Victor R. Prybutok.

How the mobile communication markets differ in China, the U.S., and Europe. Communications of the ACM, 48(3):111–114, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gori:2005:TOB

[1557] Marco Gori and Ian Witten. Technical opinion: The bubble of Web visibility. Communications of the ACM, 48(3):115–117, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2005:IRA

[1558] Peter G. Neumann. Inside risks: Anticipating disasters. Communications of the ACM, 48(3):128, March 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPd

[1559] Diane Crawford. Editorial pointers.

Communications of the ACM, 48(4):5, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTSb

[1560] CACM Staff. News track: Setting the bar; China’s café count; portrait of excess; U.S. visa rules relax; knock on wood. Communications of the ACM, 48(4):9–10, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FPO

[1561] Diane Crawford, Rodney M. Bates, David Armstrong, Robin Evans, Phillip G. Armour, Scott Duncan, and Hong-Lok Li. Forum: For programmers: Objects are not the only tool; computing is mathematical engineering; blog on blog. Communications of the ACM, 48(4):11–13, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2005:DVT

[1562] Hal Berghel. Digital village: The two sides of ROI: return on investment vs. risk of incarceration. Communications of the ACM, 48(4):15–20, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLS

[1563] David A. Patterson. President’s letter: The state of funding for new initiatives in computer science and engineering. Communications of the ACM, 48(4):21–25, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2005:PIC

[1564] Peter J. Denning. The profession of IT: Is computer science science? Communications of the ACM, 48(4):27–31, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTd

[1565] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(4):33–34, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Suttmeier:2005:VNT

[1566] Richard P. Suttmeier. Viewpoint: A new technonationalism?: China and the development of technical standards. Communications of the ACM, 48(4):35–37, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cerf:2005:VSS

[1567] Vinton G. Cerf. Viewpoint: Spam, spim, and spit. Communications of the ACM, 48(4):39–43, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martinsons:2005:ITC

[1568] Maris G. Martinsons. Introduction:

Transforming China. Communications of the ACM, 48(4):44–48, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhu:2005:DUE

[1569] Jonathan J. H. Zhu and Enhai Wang.

Diffusion, use, and effect of the Internet in China. Communications of the ACM, 48(4):49–53, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guo:2005:IDC

[1570] Xunhua Guo and Guoqing Chen. Internet diffusion in Chinese companies. Communications of the ACM, 48(4):54–58, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martinsons:2005:IEE

[1571] Maris G. Martinsons. The Internet enlightens and empowers Chinese society. Communications of the ACM, 48(4):59–60, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davison:2005:TWC

[1572] Robert M. Davison, Douglas R. Vogel, and Roger W. Harris. The etransformation of western China. Communications of the ACM, 48(4):62–67, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Quan:2005:IEC

[1573] Jing Quan, Qing Hu, and Xinan Wang. IT is not for everyone in China. Communications of the ACM, 48(4):69–72, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burrows:2005:KMC

[1574] Glen R. Burrows, Damon L. Drummond, and Maris G. Martinsons. Knowledge management in China. Communications of the ACM, 48(4): 73–76, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tu:2005:CRT

[1575] Qiang Tu, Kanliang Wang, and Qin Shu. Computer-related technostress in China. Communications of the ACM, 48(4):77–81, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Iacovou:2005:SIP

[1576] Charalambos L. Iacovou and Albert S.

Dexter. Surviving IT project cancellations. Communications of the ACM, 48(4):83–86, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Collberg:2005:SPC

[1577] Christian Collberg and Stephen Kobourov. Self-plagiarism in computer science. Communications of the ACM, 48(4):88–94, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Christiaanse:2005:PBT

[1578] Ellen Christiaanse. Performance benefits through integration hubs. Communications of the ACM, 48(4):95–100, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berendt:2005:PCS

[1579] Bettina Berendt, Oliver Günther, and Sarah Spiekermann. Privacy in ecommerce: stated preferences vs. actual behavior. Communications of the ACM, 48(4):101–106, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delen:2005:IMK

[1580] Dursun Delen, Nikunj P. Dalal, and Perakath C. Benjamin. Integrated modeling: the key to holistic understanding of the enterprise. Communications of the ACM, 48(4):107–112, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tan:2005:PGA

[1581] K. C. Tan, M. L. Wang, and W. Peng. A P2P genetic algorithm environment for the Internet. Communications of the ACM, 48(4):113–116, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhuge:2005:SGS

[1582] Hai Zhuge. Semantic Grid: scientific issues, infrastructure, and methodology. Communications of the ACM, 48(4):117–119, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneier:2005:TFA

[1583] Bruce Schneier. Two-factor authentication: too little, too late. Communications of the ACM, 48(4):136, April 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPe

[1584] Diane Crawford. Editorial pointers. Communications of the ACM, 48(5):5, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTHa

[1585] CACM Staff. News track: Hackers sacked; back to the future; U.K. alerts netizens; schoolyard E-bullies; IT and work, work, work; clocking family moves à la Potter. Communications of the ACM, 48(5):9–10, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FDP

[1586] Diane Crawford, Richard Stallman, Peter Denning, Jon Crowcroft, Herbert Kanner, Edwin D. Reilly, Jr., Len Cohen, Larry Brunelle, Philip Burgess, and Jonathan Grudin. Forum: To defend privacy, don’t collect personal data; look beyond abstraction to define computing; voter anonymity and vote security still impossible; hold the LEO and celebrate tools; must reading for software engineers; when is it OK to republish? Communications of the ACM, 48(5):11–13, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2005:SWT

[1587] Rebecca T. Mercuri. Security watch:

Trusting in transparency. Communications of the ACM, 48(5):15–19, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTe

[1588] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(5):21–22, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2005:SCF

[1589] Meg McGinity Shannon. Staying connected: It’s the FCC on the line. Communications of the ACM, 48(5):23–26, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLRa

[1590] David A. Patterson. President’s letter: Recognizing individual excellence helps us all. Communications of the ACM, 48(5):27–28, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nambisan:2005:VHP

[1591] Satish Nambisan. Viewpoint: How to prepare tomorrow’s technologists for global networks of innovation. Communications of the ACM, 48(5):29–31, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desai:2005:ISI

[1592] Anand Desai. Introduction: Special issue: Adaptive complex enterprises. Communications of the ACM, 48(5):32–35, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tan:2005:HCS

[1593] Joseph Tan, H. Joseph Wen, and Neveen Awad. Health care and services delivery systems as complex adaptive systems. Communications of the ACM, 48(5):36–44, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:2005:TBC

[1594] Albert Jones and Abhijit Deshmukh. Test beds for complex systems. Communications of the ACM, 48(5):45–50, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramanathan:2005:FAA

[1595] Jay Ramanathan. Fractal architecture for the adaptive complex enterprise. Communications of the ACM, 48(5):51–57, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramnath:2005:IES

[1596] Rajiv Ramnath and David Landsbergen. IT-enabled sense-and-respond strategies in complex public organizations. Communications of the ACM, 48(5):58–64, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bush:2005:DSK

[1597] Ashley A. Bush and Amrit Tiwana.

Designing sticky knowledge networks. Communications of the ACM, 48(5): 66–71, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nerur:2005:CMA

[1598] Sridhar Nerur, RadhaKanta Mahapatra, and George Mangalaraj. Challenges of migrating to agile methodologies. Communications of the ACM, 48(5):72–78, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lai:2005:BTS

[1599] Vincent S. Lai and Bo K. Wong.

Business types, e-strategies, and performance. Communications of the ACM, 48(5):80–85, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pawlowski:2005:GCF

[1600] Suzanne D. Pawlowski, Pratim Datta, and Andrea L. Houston. The (gradually) changing face of state IT jobs. Communications of the ACM, 48(5):

87–91, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fleischmann:2005:CTO

[1601] Kenneth R. Fleischmann and William A. Wallace. A covenant with transparency: opening the black box of models. Communications of the ACM, 48(5):93–97, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Potanin:2005:SFG

[1602] Alex Potanin, James Noble, Marcus Frean, and Robert Biddle. Scale-free geometry in OO programs. Communications of the ACM, 48(5):99–103, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zheng:2005:IDW

[1603] Baihua Zheng and Dik Lun Lee. Information dissemination via wireless broadcast. Communications of the ACM, 48(5):105–110, May 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abiteboul:2005:LDR

[1604] Serge Abiteboul, Rakesh Agrawal, Phil Bernstein, Mike Carey, Stefano Ceri, Bruce Croft, David DeWitt, Mike Franklin, Hector Garcia Molina, Dieter Gawlick, Jim Gray, Laura Haas, Alon Halevy, Joe Hellerstein, Yannis Ioannidis, Martin Kersten, Michael Pazzani, Mike Lesk, David Maier, Jeff Naughton, Hans Schek, Timos Sellis, Avi Silberschatz, Mike Stonebraker, Rick Snodgrass, Jeff Ullman, Gerhard Weikum, Jennifer Widom, and Stan Zdonik. The Lowell database research self-assessment. Communications of the ACM, 48(5):111–118, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Light:2005:TOP

[1605] Ben Light. Technical opinion: Potential pitfalls in packaged software adoption. Communications of the ACM, 48(5):119–121, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneier:2005:IRR

[1606] Bruce Schneier. Inside risks: Risks of third-party data. Communications of the ACM, 48(5):136, May 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPf

[1607] Diane Crawford. Editorial pointers.

Communications of the ACM, 48(6):5, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTB

[1608] CACM Staff. News track: Betting the pharm; mindful of trust; Asian jam session; schools unplugged, FCC undone; data for a song; A fitting solution. Communications of the ACM, 48(6):9–10, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FIC

[1609] Diane Crawford, David Jameson, Vinton Cerf, Eugen Weinstein, Robert M. Davison, Douglas R. Vogel, Matt Wimble, Michael Cusumano, Charles W. Neville, and Philip Wandler. Forum:

The imperceptible cost of eliminating spam; Taiwan’s place in China; don’t discount capital structure; beware the consequences of citing selfplagiarism. Communications of the ACM, 48(6):11–13, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2005:BSS

[1610] Phillip G. Armour. The business of software: Sarbanes–Oxley and software projects. Communications of the ACM, 48(6):15–17, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTf

[1611] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(6):19–20, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2005:LSS

[1612] Pamela Samuelson. Legally speaking: The Supreme Court revisits the sony safe harbor. Communications of the ACM, 48(6):21–25, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLDa

[1613] David A. Patterson. President’s letter: Do you queue? Communications of the ACM, 48(6):27–28, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2005:VUC

[1614] Jacques Cohen. Viewpoint: Updating computer science education. Communications of the ACM, 48(6):29–31, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2005:IHC

[1615] Andrew Rosenbloom. Introduction: 3D hard copy. Communications of the ACM, 48(6):32–35, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2005:RPC

[1616] Paul K. Wright. Rapid prototyping in consumer product design. Communications of the ACM, 48(6):36–41, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bailey:2005:LMS

[1617] Mike Bailey. Layered manufacturing for scientific visualization. Communications of the ACM, 48(6):42–48, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McMains:2005:LMT

[1618] Sara McMains. Layered manufacturing technologies. Communications of the ACM, 48(6):50–56, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Funkhouser:2005:SBR

[1619] Thomas Funkhouser, Michael Kazhdan, Patrick Min, and Philip Shilane. Shape-based retrieval and analysis of 3D models. Communications of the ACM, 48(6):58–64, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sequin:2005:RPV

[1620] Carlo H. Séquin. Rapid prototyping: a 3D visualization tool takes on sculpture and mathematical forms. Communications of the ACM, 48(6):66–73, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koller:2005:PGC

[1621] David Koller and Marc Levoy. Protecting 3D graphics content. Communications of the ACM, 48(6):74–80, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Umesh:2005:CSE

[1622] U. N. Umesh, Minh Q. Huynh, and Len Jessup. Creating successful entrepreneurial ventures in IT. Communications of the ACM, 48(6):82–87, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moore:2005:IPO

[1623] Jo Ellen Moore and Mary Sue Love. IT professionals as organizational citizens. Communications of the ACM, 48(6):88–93, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McHugh:2005:ISR

[1624] James A. M. McHugh and Fadi P. Deek. An incentive system for reducing malware attacks. Communications of the ACM, 48(6):94–99, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Xu:2005:CNA

[1625] Jennifer Xu and Hsinchun Chen. Criminal network analysis and visualization. Communications of the ACM, 48(6):100–107, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2005:ASC

[1626] Rahul Singh, A. F. Salam, and Lakshmi Iyer. Agents in e-supply chains. Communications of the ACM, 48(6):108– 115, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nah:2005:KMM

[1627] Fiona Fui-Hoon Nah, Keng Siau, and Yuhong Tian. Knowledge management mechanisms of financial service sites. Communications of the ACM, 48(6):117–123, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kanungo:2005:BMO

[1628] Tapas Kanungo, Philip Resnik, Song Mao, Doe-Wan Kim, and Qigong Zheng. The Bible and multilingual optical character recognition. Communications of the ACM, 48(6):124–130, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:TOC

[1629] Younghwa Lee. Technical opinion: The CAN-SPAM act: a silver bullet solution? Communications of the ACM, 48(6):131–132, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landau:2005:WLW

[1630] Susan Landau. Inside risks: What lessons are we teaching? Communications of the ACM, 48(6):144, June 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPg

[1631] Diane Crawford. Editorial pointers. Communications of the ACM, 48(7):5, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTHb

[1632] CACM Staff. News track: Hacker posse; Moore at home; sniffing out cancer; retiring old friends; A rose by any other name?; robo-jockeys. Communications of the ACM, 48(7):9–10, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FDU

[1633] Diane Crawford, Rajesh Gupta, Ashley Braganza, Raymond L. Robert, Hal Berghel, Richard Stallman, Michael Cusumano, Ephraim McLean, and Ralph Westfall. Forum: DARPA (and U.S.) opportunities lost; don’t ignore the CIO’s legal burdens; free is not open software; LEO lives; resolving the dataless dilemma in OO programming. Communications of the ACM, 48(7):11–13, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLRb

[1634] David A. Patterson. President’s letter: Reflections on a programming Olympiad. Communications of the ACM, 48(7):15–16, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTg

[1635] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(7):17–18, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2005:PIL

[1636] Peter J. Denning. The profession of IT: The locality principle. Communications of the ACM, 48(7):19–24, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2005:TSMb

[1637] Michael A. Cusumano. Technology strategy and management: The puzzle of Japanese software. Communications of the ACM, 48(7):25–27, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baron:2005:VIM

[1638] Naomi S. Baron. Viewpoint: Instant Messaging and the future of language. Communications of the ACM, 48(7):29–31, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holtzblatt:2005:IDM

[1639] Karen Holtzblatt. Introduction: Designing for the mobile device: Experiences, challenges, and methods. Communications of the ACM, 48(7):32–35, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blom:2005:CCC

[1640] Jan Blom, Jan Chipchase, and Jaakko Lehikoinen. Contextual and cultural challenges for user mobility research. Communications of the ACM, 48(7): 37–41, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Page:2005:MRS

[1641] Colleen Page. Mobile research strategies for a global market. Communications of the ACM, 48(7):42–48, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Balaji:2005:CPC

[1642] T. S. Balaji, Brian Landers, Jesse Kates, and Bob Moritz. A carrier’s perspective on creating a mobile multimedia service. Communications of the ACM, 48(7):49–53, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kangas:2005:AUC

[1643] Eeva Kangas and Timo Kinnunen. Applying user-centered design to mobile application development. Communications of the ACM, 48(7):55–59, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rondeau:2005:MAB

[1644] David B. Rondeau. For mobile applications, branding is experience. Communications of the ACM, 48(7):61–66, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heath:2005:MLI

[1645] Barbara P. Heath, David J. McArthur, Marilyn K. McClelland, and Ronald J. Vetter. Metadata lessons from the iLumina digital library. Communications of the ACM, 48(7):68–74, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:IAW

[1646] Gun A. Lee, Gerard J. Kim, and Mark Billinghurst. Immersive authoring: What You eXperience Is What You Get (WYXIWYG). Communications of the ACM, 48(7):76–81, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Galletta:2005:DSC

[1647] Dennis F. Galletta, Alexandra Durcikova, Andrea Everard, and Brian M. Jones. Does spell-checking software need a warning label? Communications of the ACM, 48(7):82–86, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cappel:2005:CLA

[1648] James J. Cappel, Victor R. Prybutok, and Benny Varghese. A closer look at attention to detail. Communications of the ACM, 48(7):87–92, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thelwall:2005:SWI

[1649] Mike Thelwall. Scientific Web intelligence: finding relationships in university Webs. Communications of the ACM, 48(7):93–96, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschheim:2005:IBE

[1650] Rudy Hirschheim. The Internet-based education bandwagon: look before you leap. Communications of the ACM, 48(7):97–101, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davis:2005:TOU

[1651] Robert Davis and David Yung. Technical opinion: Understanding the interactivity between television and mobile commerce. Communications of the ACM, 48(7):103–105, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Felten:2005:DPP

[1652] Edward W. Felten. Inside risks: DRM and public policy. Communications of the ACM, 48(7):112, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPh

[1653] Diane Crawford. Editorial pointers. Communications of the ACM, 48(8):5, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTSc

[1654] CACM Staff. News track: A scattering flock? transistor treasure; unions lose legion; piracy pays; pacman unleashed. Communications of the ACM, 48(8):9–10, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FRW

[1655] Diane Crawford, Amir Ben-Amram, Christian Collbert, Stephen Kobourov, Alex Simonelis, Ashley A. Bush, Amrit Tiwana, Larry Bernstein, and Rich Altmaier. Forum: Rights and wrongs in scientific publications; give users more than propaganda; define accountability in software development; stick to public safety, not business software, in laws and regulations. Communications of the ACM, 48(8):11–13, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2005:PPS

[1656] Robert L. Glass. Practical programmer: “silver bullet” milestones in software history. Communications of the ACM, 48(8):15–18, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTh

[1657] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(8):19–20, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2005:DVW

[1658] Hal Berghel and Jacob Uecker. Digital village: WiFi attack vectors. Communications of the ACM, 48(8):21–28, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLDb

[1659] David A. Patterson. President’s letter:

Does ACM support matter to conferences or journals? Communications of the ACM, 48(8):29–30, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arora:2005:VTG

[1660] Sanjeev Arora and Bernard Chazelle. Viewpoint: Is the thrill gone? Communications of the ACM, 48(8):31–33, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stafford:2005:IS

[1661] Thomas F. Stafford. Introduction: Spyware. Communications of the ACM, 48(8):34–36, August 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gibson:2005:SWI

[1662] Steve Gibson. Spyware was inevitable. Communications of the ACM, 48(8):37–39, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thompson:2005:WSP

[1663] Roger Thompson. Why spyware poses multiple threats to security. Communications of the ACM, 48(8):41–43, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2005:WDC

[1664] Xiaoni Zhang. What do consumers really know about spyware? Communications of the ACM, 48(8):44–48, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:2005:WDP

[1665] Lee A. Freeman and Andrew Urbaczewski. Why do people hate spyware? Communications of the ACM, 48(8):50–53, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Awad:2005:DBO

[1666] Neveen Farag Awad and Kristina Fitzgerald. The deceptive behaviors that offend us most about spyware. Communications of the ACM, 48(8):55–60, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hu:2005:SIN

[1667] Qing Hu and Tamara Dinev. Is spyware an Internet nuisance or public menace? Communications of the ACM, 48(8):61–66, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schmidt:2005:SLK

[1668] Mark B. Schmidt and Kirk P. Arnett. Spyware: a little knowledge is a wonderful thing. Communications of the ACM, 48(8):67–70, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:IFA

[1669] Younghwa Lee and Kenneth A.

Kozar. Investigating factors affecting the adoption of anti-spyware systems. Communications of the ACM, 48(8):72–77, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Warkentin:2005:FSA

[1670] Merrill Warkentin, Xin Luo, and Gary F. Templeton. A framework for spyware assessment. Communications of the ACM, 48(8):79–84, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shukla:2005:WBS

[1671] Sudhindra Shukla and Fiona Fui-Hoon Nah. Web browsing and spyware intrusion. Communications of the ACM, 48(8):85–90, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arnett:2005:BGM

[1672] Kirk P. Arnett and Mark B. Schmidt. Busting the ghost in the machine. Communications of the ACM, 48 (8):92–95, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Poston:2005:SVO

[1673] Robin Poston, Thomas F. Stafford, and Amy Hennington. Spyware: a view from the (online) street. Communications of the ACM, 48(8):96–99, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bose:2005:AIM

[1674] Indranil Bose and Raktim Pal. Auto-ID: managing anything, anywhere, anytime in the supply chain. Communications of the ACM, 48(8):100–106, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sheng:2005:CAD

[1675] Yihua Philip Sheng, Peter P. Mykytyn, Jr., and Charles R. Litecky. Competitor analysis and its defenses in the e-marketplace. Communications of the ACM, 48(8):107–112, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cegielski:2005:EIT

[1676] Casey G. Cegielski, Brian J. Reithel, and Carl M. Rebman. Emerging information technologies: developing a timely IT strategy. Communications of the ACM, 48(8):113–117, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bond:2005:SA

[1677] Gregory W. Bond. Software as art.

Communications of the ACM, 48(8): 118–124, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hardaway:2005:SRC

[1678] Donald E. Hardaway. Sharing research in the 21st Century: borrowing a page from open source software. Communications of the ACM, 48(8):125–128, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:2005:WSI

[1679] Jeanette Nasem Morgan. Why the software industry needs a good ghostbuster. Communications of the ACM, 48(8):129–133, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reif:2005:TOH

[1680] Harry L. Reif and Michel Mitri. Technical opinion: How university professors teach project management for information systems. Communications of the ACM, 48(8):134–136, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2005:IRD

[1681] Peter G. Neumann and Michael D. Byrne. Inside risks: Disability-related risks. Communications of the ACM, 48(8):144, August 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPi

[1682] Diane Crawford. Editorial pointers. Communications of the ACM, 48(9):5, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTL

[1683] CACM Staff. News track: Low-tech security prevails; change for the meter; quicky wiki; kicking and screaming; career swings; A news item we couldn’t refuse. Communications of the ACM, 48(9):9–10, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FWU

[1684] Diane Crawford, Lee Hollaar, Edward W. Felten, Bradley Rhodes, Richard H. Veith, and Scott Reynolds. Forum: Which use is fair in digital rights management?; we all own David; less than meets the eye in online coursework; software lessons from the Dallas Cowboys. Communications of the ACM, 48(9):11–13, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2005:BSPb

[1685] Phillip G. Armour. The business of software: To plan, two plans. Communications of the ACM, 48(9): 15–19, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2005:SCN

[1686] Meg McGinity Shannon. Staying connected: Nanotechnology’s shadow. Communications of the ACM, 48(9): 21–23, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLRc

[1687] David A. Patterson. President’s letter:

Restoring the popularity of computer science. Communications of the ACM, 48(9):25–28, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTi

[1688] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(9):29–30, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stajano:2005:RXR

[1689] Frank Stajano. RFID is x-ray vision. Communications of the ACM, 48 (9):31–33, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Borriello:2005:ISI

[1690] Gaetano Borriello. Introduction: Special issue: RFID: Tagging the world. Communications of the ACM, 48(9): 34–37, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2005:RBT

[1691] Joshua R. Smith, Kenneth P. Fishkin, Bing Jiang, Alexander Mamishev, Matthai Philipose, Adam D. Rea, Sumit Roy, and Kishore Sundara-Rajan. RFID-based techniques for human-activity detection. Communications of the ACM, 48(9):39–44, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raskar:2005:PWT

[1692] Ramesh Raskar, Paul Beardsley, Paul Dietz, and Jeroen van Baar. Photosensing wireless tags for geometric procedures. Communications of the ACM, 48(9):46–51, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pering:2005:SMM

[1693] Trevor Pering, Rafael Ballagas, and Roy Want. Spontaneous marriages of mobile devices and interactive spaces. Communications of the ACM, 48(9): 53–59, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hsi:2005:REV

[1694] Sherry Hsi and Holly Fait. RFID enhances visitors’ museum experience at the exploratorium. Communications of the ACM, 48(9):60–65, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ohkubo:2005:RPI

[1695] Miyako Ohkubo, Koutarou Suzuki, and Shingo Kinoshita. RFID privacy issues and technical challenges. Communications of the ACM, 48(9): 66–71, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunther:2005:RPC

[1696] Oliver Günther and Sarah Spiekermann. RFID and the perception of control: the consumer’s view. Communications of the ACM, 48(9):73–76, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eckfeldt:2005:WDR

[1697] Bruce Eckfeldt. What does RFID do for the consumer? Communications of the ACM, 48(9):77–79, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shah:2005:NS

[1698] Rajiv C. Shah and Jay P. Kesan.

Nurturing software. Communications of the ACM, 48(9):80–85, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Backhouse:2005:QT

[1699] James Backhouse, Carol Hsu, Jimmy C.

Tseng, and John Baptista. A question of trust. Communications of the ACM, 48(9):87–91, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jiang:2005:MBI

[1700] Zhenhui Jiang, Weiquan Wang, and Izak Benbasat. Multimedia-based interactive advising technology for online consumer decision support. Communications of the ACM, 48(9):92–98, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wood:2005:WSE

[1701] Charles A. Wood and Terence T. Ow. WEBVIEW: an SQL extension for joining corporate data to data derived from the web. Communications of the ACM, 48(9):99–104, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2005:IAQ

[1702] Ping Zhang and Na Li. The importance of affective quality. Communications of the ACM, 48(9):105–108, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhuge:2005:EES

[1703] Hai Zhuge. Exploring an epidemic in an e-science environment. Communications of the ACM, 48(9):109–114, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2005:PYM

[1704] Deborah K. Smith, Trevor Moores, and Jerry Chang. Prepare your mind for learning. Communications of the ACM, 48(9):115–118, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gorry:2005:TOS

[1705] G. Anthony Gorry. Technical opinion: As simple as possible, but not simpler. Communications of the ACM, 48 (9):119–122, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simons:2005:IRR

[1706] Barbara Simons and Jim Horning.

Inside risks: Risks of technologyoblivious policy. Communications of the ACM, 48(9):136, September 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPj

[1707] Diane Crawford. Editorial pointers. Communications of the ACM, 48(10):5, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTM

[1708] CACM Staff. News track: Medicare to the E-rescue; virus motors running; border security tags; Japan sets ‘super’ sites; beautiful music; energysaving appliances. Communications of the ACM, 48(10):9–10, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FWC

[1709] Diane Crawford, Diomidis Spinellis, Peter J. Denning, Henry Baker, Peter J. Denning, Alex Simonelis, David Patterson, and Jeff Johnson. Forum: Why computing parallels brain function; the government’s response to tech and CS downturns; disband the hacker posse. Communications of the ACM, 48(10):11–13, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLN

[1710] David A. Patterson. President’s letter: The new Professional Development Centre boasts 1,000 courses, O’Reilly books, and CS classics. Communications of the ACM, 48(10):15–16, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTj

[1711] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(10):17–18, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2005:LSD

[1712] Pamela Samuelson. Legally speaking: Did MGM really win the grokster case? Communications of the ACM, 48(10):19–24, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2005:TSMc

[1713] Michael A. Cusumano. Technology strategy and management: Software in Ireland: a balance of entrepreneurship and… lifestyle management? Communications of the ACM, 48(10):25–27, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:2005:VAD

[1714] Kenneth A. Ross. Viewpoint: Academic dishonesty and the Internet.

Communications of the ACM, 48 (10):29–31, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mowshowitz:2005:IDS

[1715] Abbe Mowshowitz and Murray Turoff. Introduction: The digital society. Communications of the ACM, 48 (10):32–35, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wolff:2005:GIW

[1716] Edward N. Wolff. The growth of information workers in the U.S. economy. Communications of the ACM, 48(10):37–42, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Czaja:2005:DAA

[1717] Sara J. Czaja and Starr Roxanne Hiltz. Digital aids for an aging society. Communications of the ACM, 48(10):43–44, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zureik:2005:CPD

[1718] Elia Zureik and Abbe Mowshowitz.

Consumer power in the digital society. Communications of the ACM, 48 (10):46–51, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wellman:2005:CNN

[1719] Barry Wellman. Community: from neighborhood to network. Communications of the ACM, 48(10):53–55, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Noam:2005:WIB

[1720] Eli M. Noam. Why the Internet is bad for democracy. Communications of the ACM, 48(10):57–58, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hiltz:2005:EGD

[1721] Starr Roxanne Hiltz and Murray Turoff. Education goes digital: the evolution of online learning and the revolution in higher education. Communications of the ACM, 48(10):59–64, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varian:2005:UAI

[1722] Hal R. Varian. Universal access to information. Communications of the ACM, 48(10):65–66, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldschmidt:2005:HMI

[1723] Peter G. Goldschmidt. HIT and MIS: implications of health information technology and medical information systems. Communications of the ACM, 48(10):68–74, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2005:DEU

[1724] Yong Jin Kim, Rajiv Kishore, and G. Lawrence Sanders. From DQ to EQ: understanding data quality in the context of e-business systems. Communications of the ACM, 48(10):75–81, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adomavicius:2005:PTP

[1725] Gediminas Adomavicius and Alexander Tuzhilin. Personalization technologies: a process-oriented perspective. Communications of the ACM, 48(10):83–90, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Prabhakar:2005:IST

[1726] Bipin Prabhakar, Charles R. Litecky, and Kirk Arnett. IT skills in a tough job market. Communications of the ACM, 48(10):91–94, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Radev:2005:NSO

[1727] Dragomir Radev, Jahna Otterbacher, Adam Winkel, and Sasha Blair-Goldensohn. NewsInEssence: summarizing online news topics. Communications of the ACM, 48(10):95–98, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:CPL

[1728] Sang M. Lee, Xin Tan, and Silvana Trimi. Current practices of leading e-government countries. Communications of the ACM, 48(10):99–104, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malhotra:2005:IWW

[1729] Arvind Malhotra and Albert H. Segars.

Investigating wireless Web adoption patterns in the U.S. Communications of the ACM, 48(10):105–110, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peslak:2005:EPP

[1730] Alan R. Peslak. Technical opinion: The educational productivity paradox. Communications of the ACM, 48(10): 111–114, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2005:IRB

[1731] Lauren Weinstein. Inside risks: The best-laid plans: a cautionary tale for developers. Communications of the ACM, 48(10):128, October 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPk

[1732] Diane Crawford. Editorial pointers. Communications of the ACM, 48(11):5, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTAb

[1733] CACM Staff. News track: Attracting hackers; super competition; blog profiles; improvements in hand; tattooed fruit; love lost. Communications of the ACM, 48(11):9–10, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FWD

[1734] Diane Crawford, Lawrence C. Paulson, Michael A. Cusumano, Stephen Schmid, Fabio A. Schreiber, Hal Lowe, Günter Rote, William Adams, David Anderson, Thomas E. Stafford, and Arthur M. Rosenberg. Forum: Welcome defects as a sign of innovation?; homage to the relational database; the cultural roots of computer science; look to human nature for the roots of spyware; wireless at work. Communications of the ACM, 48(11): 11–13, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2005:PIR

[1735] Peter J. Denning and Andrew McGettrick. The profession of IT: Recentering computer science. Communications of the ACM, 48(11):15–19, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2005:PPP

[1736] Robert L. Glass. Practical programmer: The plot to deskill software engineering. Communications of the ACM, 48(11):21–24, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTk

[1737] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48 (11):25–26, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klawe:2005:VCO

[1738] Maria Klawe and Ben Shneiderman.

Viewpoint: Crisis and opportunity in computer science. Communications of the ACM, 48(11):27–28, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:PLRd

[1739] David A. Patterson. President’s letter: Rescuing our families, our neighbors, and ourselves. Communications of the ACM, 48(11):29–31, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kugel:2005:TTO

[1740] Peter Kugel. It’s time to think outside the computational box. Communications of the ACM, 48(11):32–37, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chan:2005:MCM

[1741] Alvin T. S. Chan. Mobile cookies management on a smart card. Communications of the ACM, 48(11):38–43, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frost:2005:CPD

[1742] Richard A. Frost. Call for a publicdomain SpeechWeb. Communications of the ACM, 48(11):45–49, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuperman:2005:DPS

[1743] Benjamin A. Kuperman, Carla E. Brodley, Hilmi Ozdoganoglu, T. N. Vijaykumar, and Ankit Jalote. Detection and prevention of stack buffer overflow attacks. Communications of the ACM, 48(11):50–56, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shih:2005:RLI

[1744] Chuan-Fong Shih, Jason Dedrick, and Kenneth L. Kraemer. Rule of law and the international diffusion of ecommerce. Communications of the ACM, 48(11):57–62, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tassabehji:2005:BEK

[1745] Rana Tassabehji and Maria Vakola.

Business email: the killer impact. Communications of the ACM, 48(11): 64–70, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nerur:2005:ARI

[1746] Sridhar Nerur, Riyaz Sikora, George Mangalaraj, and VenuGopal Balijepally. Assessing the relative influence of journals in a citation network. Communications of the ACM, 48 (11):71–74, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mookerjee:2005:MES

[1747] Radha Mookerjee. Maintaining enterprise software applications. Communications of the ACM, 48(11):75–79, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wild:2005:ADS

[1748] Rosemary H. Wild, Kenneth A. Griggs, and Eldon Y. Li. An architecture for distributed scenario building and evaluation. Communications of the ACM, 48(11):80–86, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ettredge:2005:UWB

[1749] Michael Ettredge, John Gerdes, and Gilbert Karuga. Using Web-based search data to predict macroeconomic statistics. Communications of the ACM, 48(11):87–92, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hardgrave:2005:TOS

[1750] Bill C. Hardgrave and Deborah J. Armstrong. Technical opinion: Software process improvement: it’s a journey, not a destination. Communications of the ACM, 48(11):93–96, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bellovin:2005:RNS

[1751] Steven M. Bellovin, Matt Blaze, and Susan Landau. Inside risks: The real national-security needs for VoIP. Communications of the ACM, 48(11):120, November 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:EPl

[1752] Diane Crawford. Editorial pointers. Communications of the ACM, 48 (12):5–6, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2005:NTC

[1753] CACM Staff. News track: Consumers strike back; addictive hackers; fish’s tale; A gamer’s life; rankings irk profs. Communications of the ACM, 48(12):9–10, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:FNR

[1754] Diane Crawford, Kevin Markey, Scott Peer, Philip Machanick, Jim Chorn, C. J. Fearnley, Melvin Perez, and Jeanette Nasem Morgan. Forum: No RFID choice, no thanks; booms, bust, and CS enrollment; who needs a certificate? Communications of the ACM, 48 (12):11–13, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2005:HLTl

[1755] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 48(12):15–16, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2005:SWC

[1756] Rebecca Mercuri. Security watch:

Challenges in forensic computing. Communications of the ACM, 48(12): 17–21, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2005:DVP

[1757] Hal Berghel and David Hoelzer. Digital village: Pernicious ports. Communications of the ACM, 48(12):23–30, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2005:RDR

[1758] David A. Patterson. Robots in the desert: a research parable for our times. Communications of the ACM, 48(12):31–33, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruckman:2005:VSR

[1759] Amy S. Bruckman. Viewpoint: Student research and the Internet. Communications of the ACM, 48(12):35–37, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2005:ISB

[1760] Rahul Singh, Lakshmi S. Iyer, and A. F. Salam. Introduction: The semantic E-business vision. Communications of the ACM, 48(12):38–41, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deBruijn:2005:UWS

[1761] Jos de Bruijn, Dieter Fensel, Uwe Keller, and Rubén Lara. Using the Web service modeling ontology to enable semantic e-business. Communications of the ACM, 48(12):43–47, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:SKM

[1762] JinKyu Lee, Shambhu J. Upadhyaya, H. Raghav Rao, and Raj Sharman. Secure knowledge management and the Semantic Web. Communications of the ACM, 48(12):48–54, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thomas:2005:SAM

[1763] Manoj Thomas, Richard Redmond, Victoria Yoon, and Rahul Singh. A semantic approach to monitor business process. Communications of the ACM, 48(12):55–59, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kajan:2005:TOD

[1764] Ejub Kajan and Leonid Stoimenov. Toward an ontology-driven architectural framework for B2B. Communications of the ACM, 48(12):60–66, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moran:2005:UAM

[1765] Thomas P. Moran, Alex Cozzi, and Stephen P. Farrell. Unified activity management: supporting people in ebusiness. Communications of the ACM, 48(12):67–70, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thuraisingham:2005:DSP

[1766] Bhavani Thuraisingham. Directions for security and privacy for semantic ebusiness applications. Communications of the ACM, 48(12):71–73, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:MDB

[1767] Juhnyoung Lee. Model-driven business transformation and the Semantic Web. Communications of the ACM, 48 (12):75–77, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ODonnell:2005:SHN

[1768] Michael J. O’Donnell. Separate handles from names on the internet. Communications of the ACM, 48(12):78–83, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Augustine:2005:APM

[1769] Sanjiv Augustine, Bob Payne, Fred Sencindiver, and Susan Woodcock. Agile project management: steering from the edges. Communications of the ACM, 48(12):85–89, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2005:DAE

[1770] Younghwa Lee, Kenneth A. Kozar, and Kai R. Larsen. Does avatar email improve communication? Communications of the ACM, 48(12):91–95, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dingsoyr:2005:PKM

[1771] Torgeir Dingsøyr, Hans Karim Djarraya, and Emil Røyrvik. Practical knowledge management tool use in a software consulting company. Communications of the ACM, 48(12):96–100, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chandra:2005:CCD

[1772] Akhilesh Chandra and Thomas Calderon. Challenges and constraints to the diffusion of biometrics in information systems. Communications of the ACM, 48 (12):101–106, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kudyba:2005:EEH

[1773] Stephan Kudyba, G. Brent Hamar, and William M. Gandy. Enhancing efficiency in the health care industry. Communications of the ACM, 48(12):

107–110, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Musa:2005:CPT

[1774] Philip F. Musa, Victor W. Mbarika, and Peter Meso. Calling for programmed technology transfer and adoption strategies for sustainable LDC growth. Communications of the ACM, 48(12):111–116, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2005:WR

[1775] Peter Denning, Jim Horning, David Parnas, and Lauren Weinstein. Inside risks: Wikipedia risks. Communications of the ACM, 48(12):152, December 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPa

[1776] Diane Crawford. Editorial pointers. Communications of the ACM, 49(1):5–7, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTL

[1777] CACM Staff. News track: Litter unleashed; China’s net preferences; brain drained; cell phones alter African life; U.S. IT salaries to rise; runway technochic. Communications of the ACM, 49(1):9–10, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FCI

[1778] Diane Crawford, George Dodd, Thomas Joos, Anthony Ralston, Peter J. Denning, Andrew McGettrick, Todd Hartle, Robert L. Glass, Ed Callaway, George Hacken, and David P. Duggan. Forum: Call it anything but programming, even science; put creativity back in programming; the Internet’s greatest contribution to democracy; Japanese rigor for everyone. Communications of the ACM, 49(1):11–13, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTa

[1779] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(1):15–16, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2006:BSC

[1780] Phillip G. Armour. The business of software: Counting boulders and measuring mountains. Communications of the ACM, 49(1):17–20, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:AAR

[1781] David A. Patterson and CACM Staff. ACM annual report for FY05. Communications of the ACM, 49(1):21–30, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2006:SCN

[1782] Meg McGinity Shannon. Staying connected: The network of life. Communications of the ACM, 49(1):29–32, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLN

[1783] David A. Patterson. President’s letter:

New directions for CACM ? Communications of the ACM, 49(1):33–35, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell:2006:VHT

[1784] Scott Campbell. Viewpoint: How to think about security failures. Communications of the ACM, 49(1):37–39, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teevan:2006:IPI

[1785] Jaime Teevan, William Jones, and Benjamin B. Bederson. Introduction: Personal information management. Communications of the ACM, 49(1):40–43, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Czerwinski:2006:DME

[1786] Mary Czerwinski, Douglas W. Gage, Jim Gemmell, Catherine C. Marshall, Manuel A. Pérez-Quiñonesis, Meredith M. Skeels, and Tiziana Catarci. Digital memories in an era of ubiquitous computing and abundant storage. Communications of the ACM, 49(1):44–50, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pratt:2006:PHI

[1787] Wanda Pratt, Kenton Unruh, Andrea Civan, and Meredith Skeels. Personal health information management. Communications of the ACM, 49(1):51–55, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karat:2006:UPS

[1788] Clare-Marie Karat, Carolyn Brodie, and John Karat. Usable privacy and security for personal information management. Communications of the ACM, 49(1):56–57, January 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cutrell:2006:SEP

[1789] Edward Cutrell, Susan T. Dumais, and Jaime Teevan. Searching to eliminate personal information management. Communications of the ACM, 49(1):58–64, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marshall:2006:KEI

[1790] Catherine C. Marshall and William Jones. Keeping encountered information. Communications of the ACM, 49(1):66–67, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whittaker:2006:EPI

[1791] Steve Whittaker, Victoria Bellotti, and Jacek Gwizdka. Email in personal information management. Communications of the ACM, 49(1):68–73, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2006:PGP

[1792] Thomas Erickson. From PIM to GIM: personal information management in group contexts. Communications of the ACM, 49(1):74–75, January 2006.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Karger:2006:DUP

[1793] David R. Karger and William Jones.

Data unification in personal information management. Communications of the ACM, 49(1):77–82, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kelly:2006:EPI

[1794] Diane Kelly. Evaluating personal information management behaviors and tools. Communications of the ACM, 49(1):84–86, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gemmell:2006:MPD

[1795] Jim Gemmell, Gordon Bell, and Roger Lueder. MyLifeBits: a personal database for everything. Communications of the ACM, 49(1):88–95, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Desouza:2006:FDB

[1796] Kevin C. Desouza, Yukika Awazu, and Amrit Tiwana. Four dynamics for bringing use back into software reuse. Communications of the ACM, 49(1):96–100, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pallis:2006:IPC

[1797] George Pallis and Athena Vakali. Insight and perspectives for content delivery networks. Communications of the ACM, 49(1):101–106, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2006:NTP

[1798] Grace J. Johnson and Paul J. Ambrose. Neo-tribes: the power and potential of online communities in health care. Communications of the ACM, 49 (1):107–113, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chenoweth:2006:SKI

[1799] Tim Chenoweth, Karen Corral, and Haluk Demirkan. Seven key interventions for data warehouse success. Communications of the ACM, 49(1): 114–119, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gordon:2006:BPI

[1800] Lawrence A. Gordon and Martin P. Loeb. Budgeting process for information security expenditures. Communications of the ACM, 49(1):121– 125, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tedre:2006:EIC

[1801] Matti Tedre, Erkki Sutinen, Esko Kähkönen, and Piet Kommers. Ethnocomputing: ICT in cultural and social context. Communications of the ACM, 49(1):126–130, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meinel:2006:TOF

[1802] Christoph Meinel and Volker Klotz. Technical opinion: The first 10 years of the ECCC digital library. Communications of the ACM, 49(1):131–134, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Knight:2006:IRS

[1803] John C. Knight and Nancy G. Leveson.

Inside risks: Software and higher education. Communications of the ACM, 49(1):160, January 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPb

[1804] Diane Crawford. Editorial pointers. Communications of the ACM, 49(2):5, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTM

[1805] CACM Staff. News track: Moon match; billboard ratings; StAIRway to heaven; Google eyes; owning to nano unknowns; what’s the buzz? Communications of the ACM, 49(2):9–10, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FOM

[1806] Diane Crawford, Kurt Guntheroth, Andy K. Y. Poon, John Leary, Mike Brenner, Robert J. DuWors, Gediminas Adomavicius, Alexander Tuzhilin, and Alex Simonelis. Forum: Only more original research can save computer science; testing Wiki credibility; no one programs from scratch; emphasize positive feedback in personalization; why the Internet is good for democracy. Communications of the ACM, 49(2):11–13, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Templeton:2006:TOV

[1807] Gary F. Templeton. Technical opinion: Video image stabilization and registration technology. Communications of the ACM, 49(2):15–18, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Uldal:2006:IPC

[1808] Siri Birgitte Uldal and Muhammad Aimal Marjan. International perspectives: Computing in post-war Afghanistan. Communications of the ACM, 49(2):19–24, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2006:LSR

[1809] Pamela Samuelson. Legally speaking: Regulating technical design. Communications of the ACM, 49(2):25–30, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2006:TSMa

[1810] Michael A. Cusumano. Technology strategy and management: Where does Russia fit into the global software industry? Communications of the ACM, 49(2):31–34, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kumar:2006:VWS

[1811] Nanda Kumar and Abbe Mowshowitz. Viewpoint: Who should govern the Internet? Communications of the ACM, 49(2):35–37, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTb

[1812] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(2):39–40, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLO

[1813] David A. Patterson. President’s letter: Offshoring: Finally facts vs. folklore. Communications of the ACM, 49(2):41–42, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Panda:2006:I

[1814] Brajendra Panda, Joseph Giordano, and Daniel Kalil. Introduction: Nextgeneration cyber forensics. Communications of the ACM, 49(2):44–47, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Casey:2006:ISS

[1815] Eoghan Casey. Investigating sophisticated security breaches. Communications of the ACM, 49(2):48–55, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carrier:2006:RLD

[1816] Brian D. Carrier. Risks of live digital forensic analysis. Communications of the ACM, 49(2):56–61, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adelstein:2006:LFD

[1817] Frank Adelstein. Live forensics: Diagnosing your system without killing it first. Communications of the ACM, 49(2):63–66, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

CDESFWG:2006:SDE

[1818] The Common Digital Evidence Storage Format Working Group. Standardizing digital evidence storage. Communications of the ACM, 49(2):67–68, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hosmer:2006:DEB

[1819] Chet Hosmer. Digital evidence bag. Communications of the ACM, 49(2): 69–70, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teelink:2006:ICF

[1820] Sheldon Teelink and Robert F. Erbacher. Improving the computer forensic analysis process through visualization. Communications of the ACM, 49(2):71–75, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Richard:2006:NGD

[1821] Golden G. Richard, III and Vassil Roussev. Next-generation digital forensics. Communications of the ACM, 49(2):76–80, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bhaskar:2006:SLL

[1822] Rahul Bhaskar. State and local law enforcement is not ready for a cyber Katrina. Communications of the ACM, 49(2):81–83, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garfinkel:2006:ANF

[1823] Simson L. Garfinkel. AFF: a new format for storing hard drive images. Communications of the ACM, 49(2):85–87, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shankaranarayanan:2006:ME

[1824] Ganesan Shankaranarayanan and Adir Even. The metadata enigma. Communications of the ACM, 49(2):88–94, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Poon:2006:PBI

[1825] Pak-Lok Poon and Amy H. L. Lau. The present B2C implementation framework. Communications of the ACM, 49(2):96–103, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Damsgaard:2006:WCP

[1826] Jan Damsgaard, Mihir A. Parikh, and Bharat Rao. Wireless commons perils in the common good. Communications of the ACM, 49(2):104–109, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:SCN

[1827] CACM Staff. Student de-charter notice. Communications of the ACM, 49(2):110, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gladney:2006:PDP

[1828] H. M. Gladney. Principles for digital preservation. Communications of the ACM, 49(2):111–116, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Owen:2006:PAN

[1829] Stephen Owen, Pearl Brereton, and David Budgen. Protocol analysis: a neglected practice. Communications of the ACM, 49(2):117–122, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armstrong:2006:QOO

[1830] Deborah J. Armstrong. The quarks of object-oriented development. Communications of the ACM, 49(2):123–128, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varma:2006:MCS

[1831] Roli Varma. Making computer science minority-friendly. Communications of the ACM, 49(2):129–134, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2006:IRT

[1832] Peter G. Neumann. Inside risks: Trustworthy systems revisited. Communications of the ACM, 49(2):152, February 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPc

[1833] Diane Crawford. Editorial pointers.

Communications of the ACM, 49(3):5, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTWa

[1834] CACM Staff. News track: Wars blogged; snap judgments; CMU’s roboreceptionist; have we got mail! Communications of the ACM, 49(3):9–10, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FIC

[1835] Diane Crawford, Peter Wegner, Dina Goldin, Jean-Pierre Rosen, Curtis Rhodes, C. J. Fearnley, Peter J. Denning, Andrew McGettrick, Mike Segel, and Michael J. O’Donnell. Forum: Interactive computing is already outside the box; lack of Ada reflects software immaturity; be skeptical of rhetorical slight of hand; more to innovation than innovation alone; handles not a naming solution. Communications of the ACM, 49(3):11–13, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2006:BSO

[1836] Phillip G. Armour. The business of software: The operational executive sponsor. Communications of the ACM, 49(3):15–18, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2006:PPC

[1837] Robert L. Glass. Practical programmer: Is the crouching tiger a threat? Communications of the ACM, 49(3):19–20, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:2006:ECS

[1838] Bertrand Meyer and Willy Zwaenepoel. The European report: Europe’s computer scientists take fate into their own hands. Communications of the ACM, 49(3):21–24, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:CNA

[1839] CACM Staff. Call for nominations for advanced member grades in the ACM. Communications of the ACM, 49(3):25–26, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLC

[1840] David A. Patterson. President’s letter: Computer science education in the 21^{st} century. Communications of the ACM, 49(3):27–30, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:RYF

[1841] David A. Patterson. Reviving your favorite CS books. Communications of the ACM, 49(3):31, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wing:2006:VCT

[1842] Jeannette M. Wing. Viewpoint: Computational thinking. Communications of the ACM, 49(3):33–35, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martin-Flatin:2006:ISM

[1843] Jean-Philippe Martin-Flatin, Joe Sventek, and Kurt Geihs. Introduction: Self-managed systems and services. Communications of the ACM, 49(3):36–39, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robertson:2006:ARS

[1844] Paul Robertson and Brian Williams.

Automatic recovery from software failure. Communications of the ACM, 49(3):41–47, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Porter:2006:EWS

[1845] George Porter and Randy H. Katz.

Effective Web service load balancing through statistical monitoring. Communications of the ACM, 49(3):48–54, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rolia:2006:SAQ

[1846] Jerry Rolia, Ludmila Cherkasova, Martin Arlitt, and Vijay Machiraju. Supporting application quality of service in shared resource pools. Communications of the ACM, 49(3):55–60, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leibnitz:2006:BIS

[1847] Kenji Leibnitz, Naoki Wakamiya, and Masayuki Murata. Biologically inspired self-adaptive multi-path routing in overlay networks. Communications of the ACM, 49(3):62–67, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2006:CAN

[1848] Atul Singh and Mads Haahr. Creating an adaptive network of hubs using Schelling’s model. Communications of the ACM, 49(3):69–73, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:AF

[1849] CACM Staff. ACM fellows. Communications of the ACM, 49(3):74–75, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Halvey:2006:MWS

[1850] Martin Halvey, Mark T. Keane, and Barry Smyth. Mobile Web surfing is the same as Web surfing. Communications of the ACM, 49(3):76–81, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKinney:2006:MDR

[1851] Vicki R. McKinney and Mary M. Whiteside. Maintaining distributed relationships. Communications of the ACM, 49(3):82–86, March 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Walter:2006:MII

[1852] Zhiping Walter and George Scott.

Management issues of Internet/Web systems. Communications of the ACM, 49(3):87–91, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deMoor:2006:AST

[1853] Aldo de Moor and Mark Aakhus. Argumentation support: From technologies to tools. Communications of the ACM, 49(3):93–98, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gilbert:2006:AQC

[1854] Juan E. Gilbert. Applications quest:

Computing diversity. Communications of the ACM, 49(3):99–104, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beatty:2006:EIB

[1855] Robert C. Beatty and Craig D.

Williams. ERP II: Best practices for successfully implementing an ERP upgrade. Communications of the ACM, 49(3):105–109, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Othman:2006:WCS

[1856] Mazliza Othman and Rodziah Latih.

Women in computer science: No shortage here! Communications of the ACM, 49(3):111–114, March 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Santos:2006:TOU

[1857] Brian L. Dos Santos and Andrew L. Wright. Technical opinion: Using bulletin boards in an educational setting. Communications of the ACM, 49(3):115–118, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTc

[1858] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(3):126–127, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rotenberg:2006:IRR

[1859] Marc Rotenberg. Inside risks: Real ID, real trouble? Communications of the ACM, 49(3):128, March 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPd

[1860] Diane Crawford. Editorial pointers. Communications of the ACM, 49(4):5, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTBa

[1861] CACM Staff. News track: The big sweep; BMW blasted; Americans turn to the Net; high-performance help; ACM issues offshoring report; the boss is watching; fast(er) food to go. Communications of the ACM, 49(4):9–10, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FPG

[1862] Diane Crawford, William Adams, Phillip G. Armour, Peter Farrel-Vinay, Sanjiv Augustine, Bob Payne, Susan Woodcock, David Lorge Parnas, Amy S. Bruckman, Adrian Bowyer, Alessandro Berni, and Jack Ganssle. Forum: To produce a good plan, start with a good process; who defined the requirements?; proff depends on evidence; end the tragedy of the wireless commons; look further for domain name persistence; before coding, learn disciplined engineering. Communications of the ACM, 49(4):11–13, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2006:PIH

[1863] Peter J. Denning. The profession of IT:

Hastily formed networks. Communications of the ACM, 49(4):15–20, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2006:DVP

[1864] Hal Berghel. Digital village: Phishing mongers and posers. Communications of the ACM, 49(4):21–25, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:SDR

[1865] CACM Staff. Statewide databases of registered voters. Communications of the ACM, 49(4):26–28, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTd

[1866] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(4):29–30, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLS

[1867] David A. Patterson. President’s letter: Seven reasons to shave your head and three reasons not to: the bald truth. Communications of the ACM, 49(4):31–32, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Argamon:2006:VTM

[1868] Shlomo Argamon and Mark Olsen. Viewpoint: Toward meaningful computing. Communications of the ACM, 49(4):33–35, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:2006:ISE

[1869] Ryen W. White, Bill Kules, Steven M. Drucker, and m. c. schraefel. Introduction: Supporting exploratory search. Communications of the ACM, 49(4):36–39, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marchionini:2006:ESF

[1870] Gary Marchionini. Exploratory search: From finding to understanding. Communications of the ACM, 49(4):41–46, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

schraefel:2006:MII

[1871] m. c. schraefel, Max Wilson, Alistair Russell, and Daniel A. Smith. mSpace: Improving information access to multimedia domains with multimodal exploratory search. Communications of the ACM, 49(4):47–49, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cutrell:2006:EPI

[1872] Edward Cutrell and Susan T. Dumais. Exploring personal information. Communications of the ACM, 49(4):50–51, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fox:2006:ECL

[1873] Edward A. Fox, Fernando Das Neves, Xiaoyan Yu, Rao Shen, Seonho Kim, and Weiguo Fan. Exploring the computing literature with visualization and stepping stones & pathways. Communications of the ACM, 49(4):52–58, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hearst:2006:CVF

[1874] Marti A. Hearst. Clustering versus faceted categories for information exploration. Communications of the ACM, 49(4):59–61, April 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gersh:2006:SIB

[1875] John Gersh, Bessie Lewis, Jaime Montemayor, Christine Piatko, and Russell Turner. Supporting insight-based information exploration in intelligence analysis. Communications of the ACM, 49(4):63–68, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:2006:FPI

[1876] Ben Shneiderman, Benjamin B. Bederson, and Steven M. Drucker. Find that photo!: Interface strategies to annotate, browse, and share. Communications of the ACM, 49(4):69–71, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jansen:2006:UTP

[1877] Bernard J. Jansen. Using temporal patterns of interactions to design effective automated searching assistance. Communications of the ACM, 49(4):72–74, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Li:2006:FW

[1878] Jiexun Li, Rong Zheng, and Hsinchun Chen. From fingerprint to writeprint. Communications of the ACM, 49(4): 76–82, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jukic:2006:MSA

[1879] Nenad Jukic. Modeling strategies and alternatives for data warehousing projects. Communications of the ACM, 49(4):83–88, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ordanini:2006:WDM

[1880] Andrea Ordanini. What drives market transactions in B2B exchanges? Communications of the ACM, 49(4):89–93, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zviran:2006:DCE

[1881] Moshe Zviran, Dov Te’eni, and Yuval Gross. Does color in email make a difference? Communications of the ACM, 49(4):94–99, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vidal:2006:TAC

[1882] Rosario Vidal and Elena Mulet. Thinking about computer systems to support design synthesis. Communications of the ACM, 49(4):100–104, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Enns:2006:BSI

[1883] Harvey G. Enns, Thomas W. Ferratt, and Jayesh Prasad. Beyond stereotypes of IT professionals: Implications for IT HR practices. Communications of the ACM, 49(4):105–109, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2006:IMR

[1884] Sang M. Lee and Choong Kwon Lee.

Technical opinion: IT managers’ requisite skills. Communications of the ACM, 49(4):111–114, April 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2006:FIB

[1885] Lauren Weinstein. Inside risks: Fake ID: Batteries not included. Communications of the ACM, 49(4):120, April 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPe

[1886] Diane Crawford. Editorial pointers. Communications of the ACM, 49(5):5, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTE

[1887] CACM Staff. News track: Embedded risks; coat of silence; buggy concept; net oversight; virtual visitation; campus campaigns. Communications of the ACM, 49(5):9–10, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FDF

[1888] Diane Crawford, Bob Elias, Claud D. Price, Bruce Ballengee, Jack Yang, Heman Robinson, Charles Betz, Ganesan Shankaranarayanan, Adrian Bowyer, Peter Denning, James Horning, David Parnas, and Lauren Weinstein. Forum: Don’t fear the tiger; give metadata its due; who do you trust? Communications of the ACM, 49(5):11–13, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTe

[1889] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(5):15–16, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2006:SWS

[1890] Rebecca T. Mercuri. Security watch: Scoping identity theft. Communications of the ACM, 49(5):17–21, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2006:SCP

[1891] Meg McGinity Shannon. Staying connected: Private lives. Communications of the ACM, 49(5):23–26, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLR

[1892] David A. Patterson. President’s letter: Revisiting my campaign statement. Communications of the ACM, 49(5):27–30, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Easton:2006:VBA

[1893] Thomas A. Easton. Viewpoint: Beyond the algorithmization of the sciences. Communications of the ACM, 49(5):31–33, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weigand:2006:I

[1894] Hans Weigand. Introduction. Communications of the ACM, 49(5):44–46, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2006:ILA

[1895] Peter J. Denning and Robert Dunham.

Innovation as language action. Communications of the ACM, 49(5):47–52, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldkuhl:2006:AMI

[1896] Göran Goldkuhl. Action and media in interorganizational interaction. Communications of the ACM, 49(5):53–57, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dietz:2006:DSB

[1897] Jan L. G. Dietz. The deep structure of business processes. Communications of the ACM, 49(5):58–64, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teeni:2006:LAP

[1898] Dov Te’eni. The language-action perspective as a basis for communication support systems. Communications of the ACM, 49(5):65–70, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winograd:2006:DNF

[1899] Terry Winograd. Designing a new foundation for design. Communications of the ACM, 49(5):71–74, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schoop:2006:PWM

[1900] Mareike Schoop, Aldo de Moor, and Jan L. G. Dietz. The pragmatic Web: a manifesto. Communications of the ACM, 49(5):75–76, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gregg:2006:AWI

[1901] Dawn G. Gregg and Steven Walczak. Adaptive Web information extraction. Communications of the ACM, 49(5):78–84, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andriole:2006:CIB

[1902] Stephen J. Andriole. The collaborate/ integrate business technology strategy. Communications of the ACM, 49(5):85–90, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitworth:2006:WSP

[1903] Brian Whitworth, Jerry Fjermestad, and Edward Mahinda. The web of system performance. Communications of the ACM, 49(5):92–99, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhuge:2006:DKF

[1904] Hai Zhuge. Discovery of knowledge flow in science. Communications of the ACM, 49(5):101–107, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dobing:2006:HUU

[1905] Brian Dobing and Jeffrey Parsons. How UML is used. Communications of the ACM, 49(5):109–113, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katerattanakul:2006:ISR

[1906] Pairin Katerattanakul, Bernard Han, and Alan Rea. Is information systems a reference discipline? Communications of the ACM, 49(5):114–118, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spangler:2006:EPI

[1907] William E. Spangler, Kathleen S.

Hartzel, and Mordechai Gal-Or. Exploring the privacy implications of addressable advertising and viewer profiling. Communications of the ACM, 49(5):119–123, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jenkins:2006:TOM

[1908] Stephen B. Jenkins. Technical opinion: Musings of an “old-school” programmer. Communications of the ACM, 49(5):124–126, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2006:RR

[1909] Peter G. Neumann and Lauren Weinstein. Inside risks: Risks of RFID. Communications of the ACM, 49(5): 136, May 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPf

[1910] Diane Crawford. Editorial pointers. Communications of the ACM, 49(6):5, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTBb

[1911] CACM Staff. News track: Britannica bristles; snooze headlines; the geek factor; sharper image; riding the brain train. Communications of the ACM, 49(6):9–10, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FMC

[1912] Diane Crawford, Richard L. Lozes, Andy Belk, Wayne Huang, Umesh Panchaksharaiah, Zu Kim, Collin Park, David Evans, Moshe Zviran, Dov Te’eni, Yuval Gross, David Randolph, and Amy S. Bruckman. Forum: To make CS relevant, give it an industrial focus; the tiger means healthy competition; shocked by email color; pursue reality as a competitive advantage. Communications of the ACM, 49(6):11–13, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2006:PLF

[1913] David A. Patterson. President’s letter: Farewell address: The growing and graying of ACM. Communications of the ACM, 49(6):15–18, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2006:BSL

[1914] Phillip G. Armour. The business of software: The learning edge. Communications of the ACM, 49(6):19–22, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boisvert:2006:APP

[1915] Ronald F. Boisvert and Mary Jane Irwin. ACM plagiarism policy: Plagiarism on the rise. Communications of the ACM, 49(6):23–24, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTf

[1916] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(6):25–26, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2006:LSC

[1917] Pamela Samuelson. Legally speaking: Copyrighting standards. Communications of the ACM, 49(6):27–31, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conti:2006:IHI

[1918] Gregory Conti. Introduction: Hacking and innovation. Communications of the ACM, 49(6):32–36, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cross:2006:AFH

[1919] Tom Cross. Academic freedom and the hacker ethic. Communications of the ACM, 49(6):37–40, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bono:2006:STL

[1920] Stephen Bono, Aviel Rubin, Adam Stubblefield, and Matthew Green. Security through legality. Communications of the ACM, 49(6):41–43, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grand:2006:RLH

[1921] Joe Grand. Research lessons from hardware hacking. Communications of the ACM, 49(6):44–49, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Potter:2006:WHP

[1922] Bruce Potter. Wireless hotspots: Petri dish of wireless security. Communications of the ACM, 49(6):50–56, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindner:2006:SSS

[1923] Felix F. X. Lindner. Software security is software reliability. Communications of the ACM, 49(6):57–61, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaminsky:2006:ENW

[1924] Dan Kaminsky. Explorations in namespace: White-hat hacking across the domain name system. Communications of the ACM, 49(6):62–69, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bagchi:2006:GSP

[1925] Kallol Bagchi, Peeter Kirs, and Robert Cerveny. Global software piracy: Can economic factors alone explain the trend? Communications of the ACM, 49(6):70–76, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carr:2006:RGR

[1926] Christopher L. Carr. Reciprocity: The golden rule of IS-user service relationship quality and cooperation. Communications of the ACM, 49(6):77–83, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cronan:2006:PCC

[1927] Timothy Paul Cronan, C. Bryan Foltz, and Thomas W. Jones. Piracy, computer crime, and IS misuse at the university. Communications of the ACM, 49(6):84–90, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kang:2006:SAM

[1928] David Kang and Roger Chiang. A systematic approach in managing postdeployment system changes. Communications of the ACM, 49(6):91–96, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2006:AMC

[1929] Bonnie Brinton Anderson, James V. Hansen, Paul Benjamin Lowry, and Scott L. Summers. The application of model checking for securing ecommerce transactions. Communications of the ACM, 49(6):97–101, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shore:2006:EIA

[1930] Barry Shore. Enterprise integration across the globally disbursed service organization. Communications of the ACM, 49(6):102–106, June 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jackson:2006:TOS

[1931] Thomas W. Jackson, Anthony Burgess, and Janet Edwards. Technical opinion: A simple approach to improving email communication. Communications of the ACM, 49(6):107–109, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Charette:2006:IRE

[1932] Robert Charette. Inside risks: EHRs: Electronic health records or exceptional hidden risks? Communications of the ACM, 49(6):120, June 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPg

[1933] Diane Crawford. Editorial pointers.

Communications of the ACM, 49(7):5, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTWb

[1934] CACM Staff. News track: Weather cells; urban (ground) legends; coralling bot rustlers; who do you trust?; China’s “pedia” product; the peak of devotion. Communications of the ACM, 49(7):9–10, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FTP

[1935] Diane Crawford, Richard Veith, Gregory Aicklen, John Knight, Rebecca Mercuri, D. J. Howorth, Mazliza Othman, Rodziah Latih, and Chin-Teng Lin. Forum: Think of programming as an elegant essay; nurture the joy of engineering; don’t equate bankruptcy and identity theft; author’s response; what Malaysia looks for when funding CS study; author’s response; we’re from Taiwan, not from China. Communications of the ACM, 49(7):11–13, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:AER

[1936] Diane Crawford. ACM election results:

ACM’s 2006 general elections. Communications of the ACM, 49(7):13, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2006:PII

[1937] Peter J. Denning. The profession of IT: Infoglut. Communications of the ACM, 49(7):15–19, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2006:TSMb

[1938] Michael A. Cusumano. Technology strategy and management: What road ahead for Microsoft and Windows? Communications of the ACM, 49(7): 21–23, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTg

[1939] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(7):25–26, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wegner:2006:VPP

[1940] Peter Wegner and Dina Goldin. Viewpoint: Principles of problem solving. Communications of the ACM, 49(7):27–29, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spohrer:2006:ISI

[1941] Jim Spohrer and Doug Riecken. Introduction: Special issues on services science. Communications of the ACM, 49(7):30–32, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chesbrough:2006:RMS

[1942] Henry Chesbrough and Jim Spohrer. A research manifesto for services science. Communications of the ACM, 49(7):35–40, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sheehan:2006:USS

[1943] Jerry Sheehan. Understanding service sector innovation. Communications of the ACM, 49(7):42–47, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zysman:2006:ARF

[1944] John Zysman. The algorithmic revolution—the fourth service transformation. Communications of the ACM, 49(7):48, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rust:2006:WAR

[1945] Roland T. Rust and Carol Miu. What academic research tells us about service. Communications of the ACM, 49(7):49–54, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sheth:2006:SEF

[1946] Amit Sheth, Kunal Verma, and Karthik Gomadam. Semantics to energize the full services spectrum. Communications of the ACM, 49(7):55–61, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dietrich:2006:RPB

[1947] Brenda Dietrich. Resource planning for business services. Communications of the ACM, 49(7):62–64, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rouse:2006:ET

[1948] William B. Rouse and Marietta L. Baba. Enterprise transformation. Communications of the ACM, 49(7): 66–72, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bitner:2006:EDS

[1949] Mary Jo Bitner and Stephen W. Brown. The evolution and discovery of services science in business schools. Communications of the ACM, 49(7):73–78, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ganz:2006:GSE

[1950] Walter Ganz. Germany: service engineering. Communications of the ACM, 49(7):79, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinhardt:2006:GCA

[1951] Christof Weinhardt, Dirk Neumann, and Carsten Holtmann. Germany: computer-aided market engineering. Communications of the ACM, 49(7):79, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maglio:2006:SSS

[1952] Paul P. Maglio, Savitha Srinivasan, Jeffrey T. Kreulen, and Jim Spohrer. Service systems, service scientists, SSME, and innovation. Communications of the ACM, 49(7):81–85, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2006:CCM

[1953] Stuart I. Feldman, Krishna S. Nathan, Thomas Li, Kazuyoshi Hidaka, and Corinna Schulze. The Clarion Call for modern services: China, Japan, Europe, and the US. Communications of the ACM, 49(7):86–87, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Avison:2006:MIU

[1954] David Avison, Shirley Gregor, and David Wilson. Managerial IT unconsciousness. Communications of the ACM, 49(7):88–93, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:2006:TBS

[1955] Lance J. Hoffman, Kim Lawson-Jenkins, and Jeremy Blum. Trust beyond security: an expanded trust model. Communications of the ACM, 49(7):94–101, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bekkering:2006:TV

[1956] Ernst Bekkering and J. P. Shim. Trust in videoconferencing. Communications of the ACM, 49(7):103–107, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bolloju:2006:ANA

[1957] Narasimha Bolloju and Felix S. K. Leung. Assisting novice analysts in developing quality conceptual models with UML. Communications of the ACM, 49(7):108–112, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martin:2006:TOT

[1958] Fred Martin. Technical opinion: Toy projects considered harmful. Communications of the ACM, 49(7):113–116, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2006:IRR

[1959] Peter G. Neumann. Inside risks: Risks relating to system compositions. Communications of the ACM, 49(7):120, July 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPh

[1960] Diane Crawford. Editorial pointers. Communications of the ACM, 49(8):5, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTA

[1961] CACM Staff. News track: Airport security afoot; air fare; the human touch; Luddite politicians seek help; are we there yet?; tunes to go. Communications of the ACM, 49(8):9–10, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FSF

[1962] Diane Crawford, Umesk Panchaksharaiah, L. Peter Deutsch, Kallol Bagchi, Peeter Kirs, Robert Cerveny, Edward J. Ferrara, Frederick G. Volpicelli, Diomidis Spinellis, and C. Augusto Casas. Forum: Stress fundamentals in CS education, simplicity in production; no room for bias when covering the copyright system; don’t misuse or misteach UML; cover to cover; future CS course already here; to attract new members, increase ACM scope and mandate. Communications of the ACM, 49(8):11–13, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2006:PPS

[1963] Robert L. Glass. Practical programmer: The Standish report: does it really describe a software crisis? Communications of the ACM, 49(8):15–16, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2006:DVD

[1964] Hal Berghel and David Hoelzer. Digital village: Disk wiping by any other name. Communications of the ACM, 49(8):17–21, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTh

[1965] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(8):23–24, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldman:2006:VBB

[1966] Dan B. Goldman. Viewpoint: Blockbusters and blackboards. Communications of the ACM, 49(8):25–27, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pardo:2006:IMI

[1967] Bryan Pardo. Introduction: Music information retrieval. Communications of the ACM, 49(8):28–31, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ellis:2006:EIM

[1968] Daniel P. W. Ellis. Extracting information from music audio. Communications of the ACM, 49(8):32–37, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dannenberg:2006:MSA

[1969] Roger B. Dannenberg and Christopher Raphael. Music score alignment and computer accompaniment. Communications of the ACM, 49(8):38–43, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2006:SMR

[1970] Avery Wang. The Shazam music recognition service. Communications of the ACM, 49(8):44–48, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Birmingham:2006:QHV

[1971] William Birmingham, Roger Dannenberg, and Bryan Pardo. Query by humming with the VocalSearch system. Communications of the ACM, 49(8):49–52, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dunn:2006:VRU

[1972] Jon W. Dunn, Donald Byrd, Mark Notess, Jenn Riley, and Ryan Scherle. Variations2: retrieving and using music in an academic setting. Communications of the ACM, 49(8):53–58, August 2006.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pons:2006:BMT

[1973] Alexander P. Pons. Biometric marketing: targeting the online consumer. Communications of the ACM, 49(8):60–66, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rajlich:2006:CPS

[1974] Vaclav Rajlich. Changing the paradigm of software engineering. Communications of the ACM, 49(8):67–70, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nord:2006:MWP

[1975] G. Daryl Nord, Tipton F. McCubbins, and Jeretta Horn Nord. E-monitoring in the workplace: privacy, legislation, and surveillance software. Communications of the ACM, 49(8):72–77, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Procaccino:2006:DCS

[1976] J. Drew Procaccino, June M. Verner, and Steven J. Lorenzet. Defining and contributing to software development success. Communications of the ACM, 49(8):79–83, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:2006:TAP

[1977] Eugenia Y. Huang. Tightening access policy: the hidden trade-off. Communications of the ACM, 49(8):85–89, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McDowell:2006:PPI

[1978] Charlie McDowell, Linda Werner, Heather E. Bullock, and Julian Fernald. Pair programming improves student retention, confidence, and program quality. Communications of the ACM, 49(8):90–95, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siponen:2006:TOI

[1979] Mikko Siponen. Technical opinion: Information security standards focus on the existence of process, not its content. Communications of the ACM, 49(8):97–100, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mulligan:2006:IRR

[1980] Deirdre K. Mulligan, Ari Schwartz, and Indrani Mondal. Inside risks: Risks of online storage. Communications of the ACM, 49(8):112, August 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPi

[1981] Diane Crawford. Editorial pointers. Communications of the ACM, 49(9):5, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTH

[1982] CACM Staff. News track: HP’s hot spot [memory spot]; spare cycles against malaria; the caveman’s disconnection [neanderthal DNA analysis]; privacy policy advice; take me home, country roads. Communications of the ACM, 49(9):9–10, September 2006.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FWIa

[1983] Diane Crawford, William Adams, Peter G. Neumann, John C. Nash, Dru Lavigne, Russell McOrmond, Charles McDonald, Michael Richardson, Raymond Wood, Kallol Bagchi, Peeter Kirs, Robert Cerveny, Berhoooz Parhami, Peter Denning, Rodney M. Bates, and Robert L. Glass. Forum: When it comes to software, don’t just do it; author’s response; A single vested source is not data; author’s response; how to welcome uninvited guests; author’s response; skip toy-program thinking in real-world projects; call it problem solving, not computational thinking. Communications of the ACM, 49(9):11–13, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2006:SHD

[1984] Phillip G. Armour. The business of software: Software: hard data. Communications of the ACM, 49(9): 15–17, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2006:SCH

[1985] Meg McGinity Shannon. Staying connected: Home turf. Communications of the ACM, 49(9):19–21, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTi

[1986] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(9):23–24, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adams:2006:VWL

[1987] Dennis Adams. Viewpoint: Wireless laptops in the classroom (and the Sesame Street syndrome). Communications of the ACM, 49(9):25–27, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Muller:2006:IPS

[1988] Günter Müller. Introduction: Privacy and security in highly dynamic systems. Communications of the ACM, 49 (9):28–31, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sackmann:2006:PPA

[1989] Stefan Sackmann, Jens Strüker, and Rafael Accorsi. Personalization in privacy-aware highly dynamic systems. Communications of the ACM, 49(9): 32–38, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pretschner:2006:DUC

[1990] Alexander Pretschner, Manuel Hilty, and David Basin. Distributed usage control. Communications of the ACM, 49(9):39–44, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Seshadri:2006:EVC

[1991] Arvind Seshadri, Mark Luk, Adrian Perrig, Leendert van Doorn, and Pradeep Khosla. Externally verifiable code execution. Communications of the ACM, 49(9):45–49, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buchmann:2006:PCL

[1992] Johannes Buchmann, Alexander May, and Ulrich Vollmer. Perspectives for cryptographic long-term security. Communications of the ACM, 49(9): 50–55, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Subirana:2006:LP

[1993] Brian Subirana and Malcolm Bain. Legal programming. Communications of the ACM, 49(9):57–62, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:2006:WDP

[1994] Mark D. Hill, Jean-Luc Gaudiot, Mary Hall, Joe Marks, Paolo Prinetto, and Donna Baglio. A Wiki for discussing and promoting best practices in research. Communications of the ACM, 49(9):63–64, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wiederhold:2006:WYS

[1995] Gio Wiederhold. What is your software worth? Communications of the ACM, 49(9):65–75, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fan:2006:TPT

[1996] Weiguo Fan, Linda Wallace, Stephanie Rich, and Zhongju Zhang. Tapping the power of text mining. Communications of the ACM, 49(9):76–82, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stiegler:2006:PVS

[1997] Marc Stiegler, Alan H. Karp, Ka-Ping Yee, Tyler Close, and Mark S. Miller. Polaris: virus-safe computing for Windows XP. Communications of the ACM, 49(9):83–88, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hong:2006:MDS

[1998] Se-Joon Hong, Kar Yan Tam, and Jinwoo Kim. Mobile data service fuels the desire for uniqueness. Communications of the ACM, 49(9):89–94, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chae:2006:ATA

[1999] Bongsug (Kevin) Chae and Roger McHaney. Asian trio’s adoption of Linux-based open source development. Communications of the ACM, 49(9): 95–99, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2006:IRF

[2000] Peter G. Neumann. Inside risks: The foresight saga. Communications of the ACM, 49(9):112, September 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPj

[2001] Diane Crawford. Editorial pointers. Communications of the ACM, 49(10):5, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTO

[2002] CACM Staff. News track: Outsourcing twist; Iraq’s cool factor; Segway’s segue; A picture is worth even more spam; blog-free CEOs; sponge count. Communications of the ACM, 49(10):9–10, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FWIb

[2003] Diane Crawford, Donn B. Parker, Phillip Armour, and Bhupinder Singh Anand. Forum: Why information security is still a folk art;; in programming, as in life, know the problem domain; measure software success as delivered value to the customer; how definitive is the classical Church–Turing thesis? Communications of the ACM, 49(10):11–13, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2006:TSMc

[2004] Michael A. Cusumano. Technology strategy and management: Envisioning the future of India’s software services business. Communications of the ACM, 49(10):15–17, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTj

[2005] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49(10):19–20, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2006:LSI

[2006] Pamela Samuelson. Legally speaking: IBM’s pragmatic embrace of open source. Communications of the ACM, 49(10):21–25, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aagerfalk:2006:ISI

[2007] Pär J. Ågerfalk and Brian Fitzgerald. Introduction: Special issue: Flexible and distributed software processes: old petunias in new bowls? Communications of the ACM, 49(10):26–34, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2006:ACS

[2008] Gwanhoo Lee, William DeLone, and J. Alberto Espinosa. Ambidextrous coping strategies in globally distributed software development projects. Communications of the ACM, 49(10):35– 40, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramesh:2006:CDS

[2009] Balasubramaniam Ramesh, Lan Cao, Kannan Mohan, and Peng Xu. Can distributed software development be agile? Communications of the ACM, 49(10):41–46, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2006:AIC

[2010] One-Ki (Daniel) Lee, Probir Banerjee, Kai H. Lim, Kuldeep Kumar, Jos van Hillegersberg, and Kwok Kee Wei. Aligning IT components to achieve agility in globally distributed system development. Communications of the ACM, 49(10):48–54, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagstrom:2006:DFD

[2011] Patrick Wagstrom and James Herbsleb. Dependency forecasting in the distributed agile organization. Communications of the ACM, 49(10):55–56, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Flor:2006:GDS

[2012] Nick V. Flor. Globally distributed software development and pair programming. Communications of the ACM, 49(10):57–58, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liu:2006:TNI

[2013] Jiming Liu and K. C. Tsui. Toward nature-inspired computing. Communications of the ACM, 49(10):59–64, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Giguette:2006:BOP

[2014] Ray Giguette. Building objects out of Plato: applying philosophy, symbolism, and analogy to software design. Communications of the ACM, 49(10):66–71, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cegielski:2006:WMG

[2015] Casey G. Cegielski and Dianne J.

Hall. What makes a good programmer? Communications of the ACM, 49(10):73–75, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dinev:2006:WSS

[2016] Tamara Dinev. Why spoofing is serious Internet fraud. Communications of the ACM, 49(10):76–82, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hann:2006:WGS

[2017] Il-Horn Hann, Kai-Lung Hui, Yee-Lin Lai, S. Y. T. Lee, and I. P. L. Png. Who gets spammed? Communications of the ACM, 49(10):83–87, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bharati:2006:SCS

[2018] Pratyush Bharati and Abhijit Chaudhury. Studying the current status of technology adoption. Communications of the ACM, 49(10):88–93, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bellovin:2006:IRV

[2019] Steven M. Bellovin. Inside risks: Virtual machines, virtual security? Communications of the ACM, 49(10):104, October 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPk

[2020] Diane Crawford. Editorial pointers. Communications of the ACM, 49(11):5, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTS

[2021] CACM Staff. News track: Super game plan; E-learning roots disputed; Chinese history; U.K. seeks popular science input; encryption commitment; improve-ment news; vid kid. Communications of the ACM, 49(11):9–10, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FDT

[2022] Diane Crawford, Karl Wiegers, Marc Stiegler, Alan H. Karp, Ka-Ping Yee, Tyler Close, Mark S. Miller, Jeffrey A. Rosenwald, Jim Chorn, and David Dreyfus. Forum: Don’t target me, unless I say you can; avoid the trap of unrealistic expectations; fewer degrees of separation means more precision in software projects. Communications of the ACM, 49(11):11–13, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2006:PPL

[2023] Robert L. Glass. Practical programmer: Looking into the challenges of complex IT projects. Communications of the ACM, 49(11):15–17, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2006:PID

[2024] Peter J. Denning and Rick Hayes-Roth. The profession of IT: Decision making in very large networks. Communications of the ACM, 49(11): 19–23, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:HLTk

[2025] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 49 (11):25–26, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Teuscher:2006:VBU

[2026] Christof Teuscher. Viewpoint: Biologically uninspired computer science. Communications of the ACM, 49(11): 27–29, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sreenan:2006:ISI

[2027] Cormac J. Sreenan and Kobus van der Merwe. Introduction to special issue on entertainment networking. Communications of the ACM, 49(11):30–33, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2006:HSO

[2028] Kuan-Ta Chen, Polly Huang, and Chin-Laung Lei. How sensitive are online gamers to network quality? Communications of the ACM, 49(11):34–38, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Claypool:2006:LPA

[2029] Mark Claypool and Kajal Claypool. Latency and player actions in online games. Communications of the ACM, 49(11):40–45, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brun:2006:MLF

[2030] Jeremy Brun, Farzad Safaei, and Paul Boustead. Managing latency and fairness in networked games. Communications of the ACM, 49(11):46–51, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kurtisi:2006:ENC

[2031] Zefir Kurtisi, Xiaoyuan Gu, and Lars Wolf. Enabling network-centric music performance in wide-area networks. Communications of the ACM, 49(11): 52–54, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rejaie:2006:ACB

[2032] Reza Rejaie. Anyone can broadcast video over the internet. Communications of the ACM, 49(11):55–57, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahmann:2006:CSH

[2033] Verena Kahmann, Jens Brandt, and Lars Wolf. Collaborative streaming in heterogeneous and dynamic scenarios. Communications of the ACM, 49 (11):58–63, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:2006:NAI

[2034] Dijiang Huang, Qing Cao, Amit Sinha, Marc J. Schniederjans, Cory Beard, Lein Harn, and Deep Medhi. New architecture for intra-domain network security issues. Communications of the ACM, 49(11):64–72, November 2006.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taylor:2006:CRO

[2035] Hazel Taylor. Critical risks in outsourced IT projects: the intractable and the unforeseen. Communications of the ACM, 49(11):74–79, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Payne:2006:UCE

[2036] Dinah Payne and Brett J. L. Landry. A uniform code of ethics: business and IT professional ethics. Communications of the ACM, 49(11):81–84, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sen:2006:CTW

[2037] Arun Sen, Peter A. Dacin, and Christos Pattichis. Current trends in web data analysis. Communications of the ACM, 49(11):85–91, November 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mercuri:2006:IRC

[2038] Rebecca T. Mercuri, Vincent J. Lipsio, and Beth Feehan. Inside risks: COTS and other electronic voting backdoors. Communications of the ACM, 49(11):112, November 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:EPl

[2039] Diane Crawford. Editorial pointers. Communications of the ACM, 49(12):5, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:NTT

[2040] CACM Staff. News track: Tech 20/20 on 2020; patents not pending; A matter of style; nano products, giga concerns; the tiny engine that could; “houston, we have a problem”. Communications of the ACM, 49(12):9–10, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:FGB

[2041] Diane Crawford, Girish Seshagiri, Pär Ågerfalk, Brian Fitzgerald, Trevor R. H. Clarke, Niels J. Bjergstrom, Alex Simonelis, Charlie McDowell, Linda Werner, Heather E. Bullock, Julian Fernald, Melvin Fitting, Bhupinder Singh Anand, and Thorkil Naur. Forum: GSD benefits still not worth the trouble; let interactive genetic algorithms provide music accompaniment; insist on risk management in software engineering; could one head be better than two?; would Church or Turing make every function computable?; still no proof of the unsolvability of Turing’s halting problem. Communications of the ACM, 49(12):11–13, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2006:DVF

[2042] Hal Berghel. Digital village: Fungible credentials and next-generation fraud. Communications of the ACM, 49(12):15–19, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2006:AHD

[2043] CACM Staff. ACM honors distinguished members for contributions to computing. Communications of the ACM, 49(12):20–21, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2006:TDA

[2044] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 49(12):23–24, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blank:2006:VRM

[2045] Douglas Blank. Viewpoint: Robots make computer science personal. Communications of the ACM, 49(12):25–27, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sugumaran:2006:ISP

[2046] Vijayan Sugumaran, Sooyong Park, and Kyo C. Kang. Introduction: Software product line engineering. Communications of the ACM, 49(12):28–32, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clements:2006:GTH

[2047] Paul C. Clements, Lawrence G. Jones, John D. McGregor, and Linda M. Northrop. Getting there from here: a roadmap for software product line adoption. Communications of the ACM, 49(12):33–36, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Krueger:2006:NMS

[2048] Charles W. Krueger. New methods in software product line practice. Communications of the ACM, 49 (12):37–40, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bosch:2006:CBS

[2049] Jan Bosch. The challenges of broadening the scope of software product families. Communications of the ACM, 49 (12):41–44, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Batory:2006:AAF

[2050] Don Batory, David Benavides, and Antonio Ruiz-Cortes. Automated analysis of feature models: challenges ahead. Communications of the ACM, 49(12):45–47, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eriksson:2006:SPL

[2051] Magnus Eriksson, Jürgen Börstler, and Kjell Borg. Software product line modeling made practical. Communications of the ACM, 49(12):49–54, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2006:FOV

[2052] Jaejoon Lee and Dirk Muthig. Featureoriented variability management in product line engineering. Communications of the ACM, 49(12):55–59, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Czarnecki:2006:MLC

[2053] Krzysztof Czarnecki, Michal Antkiewicz, and Chang Hwan Peter Kim.

Multi-level customization in application engineering. Communications of the ACM, 49(12):60–65, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Helferich:2006:PMS

[2054] Andreas Helferich, Klaus Schmid, and Georg Herzwurm. Product management for software product lines: an unsolved problem? Communications of the ACM, 49(12):66–67, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mohan:2006:CMP

[2055] Kannan Mohan and Balasubramaniam Ramesh. Change management patterns in software product lines. Communications of the ACM, 49(12):68–72, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kishi:2006:FVS

[2056] Tomoji Kishi and Natsuko Noda. Formal verification and software product lines. Communications of the ACM, 49 (12):73–77, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pohl:2006:SPL

[2057] Klaus Pohl and Andreas Metzger. Software product line testing. Communications of the ACM, 49(12):78–81, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pena:2006:MAS

[2058] Joaquin Peña, Michael G. Hinchey, and Antonio Ruiz-Cortés. Multi-agent system product lines: challenges and benefits. Communications of the ACM, 49 (12):82–84, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

In:2006:QBC

[2059] Hoh Peter In, Jongmoon Baik, Sangsoo Kim, Ye Yang, and Barry Boehm. A quality-based cost estimation model for the product line life cycle. Communications of the ACM, 49(12):85–88, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Angell:2006:REC

[2060] Ian Angell and Jan Kietzmann. RFID and the end of cash? Communications of the ACM, 49(12):90–96, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Maamar:2006:WCC

[2061] Zakaria Maamar, Djamal Benslimane, and Nanjangud C. Narendra. What can context do for Web services? Communications of the ACM, 49(12): 98–103, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bichler:2006:SFA

[2062] Martin Bichler and Jayant R. Kalagnanam.

Software frameworks for advanced procurement auction markets. Communications of the ACM, 49(12):104–108, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Montazemi:2006:HTM

[2063] Ali Reza Montazemi. How they manage IT: SMEs in Canada and the U.S. Communications of the ACM, 49(12): 109–112, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Belanger:2006:TOW

[2064] France Belanger, Weiguo Fan, L. Christian Schaupp, Anjala Krishen, Jeannine Everhart, David Poteet, and Kent Nakamoto. Technical opinion: Web site success metrics: addressing the duality of goals. Communications of the ACM, 49(12):114–116, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hasselbring:2006:LRR

[2065] Wilhelm Hasselbring, Matthias Rohr, Jurgen Taeger, and Daniel Winteler. Inside risks: Liability risks in reusing third-party software. Communications of the ACM, 49(12):144, December 2006. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPa

[2066] Diane Crawford. Editorial pointers. Communications of the ACM, 50(1): 5, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTa

[2067] CACM Staff. News track. Communications of the ACM, 50(1):9–10, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fa

[2068] Diane Crawford. Forum. Communications of the ACM, 50(1):11–12, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2007:BSA

[2069] Phillip G. Armour. The business of software: Agile … and offshore. Communications of the ACM, 50(1):13–16, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2007:SCSa

[2070] Meg McGinity Shannon. Staying connected: Shannon’s eleven. Communications of the ACM, 50(1):17–20, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beale:2007:VSD

[2071] Russell Beale. Viewpoint: Slanty design. Communications of the ACM, 50(1):21–24, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:AAR

[2072] CACM Staff. ACM annual report for FY06. Communications of the ACM, 50(1):25–32, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abril:2007:PHD

[2073] Patricia S. Abril and Robert Plant. The patent holder’s dilemma: buy, sell, or troll? Communications of the ACM, 50(1):36–44, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhao:2007:SMA

[2074] Huimin Zhao. Semantic matching across heterogeneous data sources. Communications of the ACM, 50 (1):45–50, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Biehl:2007:SFI

[2075] Markus Biehl. Success factors for implementing global information systems. Communications of the ACM, 50(1):52–58, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Singh:2007:IOP

[2076] Nitin Singh, Kee hung Lai, and T. C. E. Cheng. Intra-organizational perspectives on IT-enabled supply chains. Communications of the ACM, 50 (1):59–65, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McDonald:2007:WST

[2077] Sharon McDonald and Helen M. Edwards. Who should test whom? Communications of the ACM, 50(1):66–71, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Surakka:2007:WSS

[2078] Sami Surakka. What subjects and skills are important for software developers? Communications of the ACM, 50(1):73–78, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoepman:2007:IST

[2079] Jaap-Henk Hoepman and Bart Jacobs. Increased security through open source. Communications of the ACM, 50(1):79–83, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naur:2007:CVH

[2080] Peter Naur. Computing versus human thinking. Communications of the ACM, 50(1):85–94, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:HLT

[2081] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(1):95–96, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2007:IRM

[2082] Lauren Weinstein. Inside risks: Ma Bell’s revenge: the battle for network neutrality. Communications of the ACM, 50(1):128, January 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPb

[2083] Diane Crawford. Editorial pointers. Communications of the ACM, 50(2):5, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTb

[2084] CACM Staff. News track. Communications of the ACM, 50(2):9–10, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fb

[2085] Diane Crawford. Forum. Communications of the ACM, 50(2):11–13, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:SCN

[2086] CACM Staff. Student de-charter notice. Communications of the ACM, 50(2):14, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2007:TSMa

[2087] Michael A. Cusumano. Technology strategy and management: What road ahead for Microsoft the company? Communications of the ACM, 50(2):15–18, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2007:LSP

[2088] Pamela Samuelson. Legally speaking: Patent reform through the courts. Communications of the ACM, 50(2): 19–23, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goodman:2007:SOB

[2089] Joshua Goodman, Gordon V. Cormack, and David Heckerman. Spam and the ongoing battle for the inbox. Communications of the ACM, 50(2):24–33, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agrawal:2007:ECH

[2090] Rakesh Agrawal, Tyrone Grandison, Christopher Johnson, and Jerry Kiernan. Enabling the 21st century health care information technology revolution. Communications of the ACM, 50(2):34–42, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jacoby:2007:IMM

[2091] Grant A. Jacoby and Luqi. Intranet model and metrics. Communications of the ACM, 50(2):43–50, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glassberg:2007:EMH

[2092] Bonnie C. Glassberg. Electronic markets hypothesis redux: where are we now? Communications of the ACM, 50(2):51–55, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grimes:2007:CCS

[2093] Galen A. Grimes. Compliance with the CAN-SPAM Act of 2003. Communications of the ACM, 50(2):56–62, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baraglia:2007:DPW

[2094] Ranieri Baraglia and Fabrizio Silvestri.

Dynamic personalization of web sites without user intervention. Communications of the ACM, 50(2):63–67, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Koh:2007:EPV

[2095] Joon Koh, Young-Gul Kim, Brian Butler, and Gee-Woo Bock. Encouraging participation in virtual communities. Communications of the ACM, 50(2):68–73, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2007:WCA

[2096] Dongsong Zhang. Web content adaptation for mobile handheld devices. Communications of the ACM, 50(2):75–79, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:SACa

[2097] CACM Staff. SIGs announce candidates for election. Communications of the ACM, 50(2):80, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bhatti:2007:IAF

[2098] Rafae Bhatti, Elisa Bertino, and Arif Ghafoor. An integrated approach to federated identity and privilege management in open systems. Communications of the ACM, 50(2):81–87, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shao:2007:IOO

[2099] Benjamin B. Shao and Julie Smith David. The impact of offshore outsourcing on IT workers in developed countries. Communications of the ACM, 50(2):89–94, February 2007. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oppliger:2007:TOI

[2100] Rolf Oppliger. Technical opinion:

IT security: in search of the Holy Grail. Communications of the ACM, 50(2):96–98, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAb

[2101] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(2):99–100, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2007:IRW

[2102] Peter G. Neumann. Inside risks:

Widespread network failures. Communications of the ACM, 50(2):112, February 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPc

[2103] Diane Crawford. Editorial pointers.

Communications of the ACM, 50(3):5, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTc

[2104] CACM Staff. News track. Communications of the ACM, 50(3):9–10, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fc

[2105] Diane Crawford. Forum. Communications of the ACM, 50(3):11–13, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:AMT

[2106] CACM Staff. ACM A. M. Turing Award nominations solicited. Communications of the ACM, 50(3):14, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2007:BSM

[2107] Phillip G. Armour. The business of software: Mortality play. Communications of the ACM, 50(3):15–18, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barr:2007:VFF

[2108] Earl Barr, Matt Bishop, and Mark Gondree. Viewpoint: Fixing federal e-voting standards. Communications of the ACM, 50(3):19–24, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:SACb

[2109] CACM Staff. SIGs announce candidates for election. Communications of the ACM, 50(3):25, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:AF

[2110] CACM Staff. ACM fellows. Communications of the ACM, 50(3):26–27, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

VandeWalle:2007:ERI

[2111] Bartel Van de Walle and Murray Turoff. Emergency response information systems: Emerging trends and technologies: Introduction. Communications of the ACM, 50(3):29–31, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carver:2007:HCI

[2112] Liz Carver and Murray Turoff. Humancomputer interaction: the human and computer as a team in emergency management information systems. Communications of the ACM, 50(3):33–38, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

French:2007:DSS

[2113] Simon French and Murray Turoff. Decision support systems. Communications of the ACM, 50(3):39–40, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fiedrich:2007:ABS

[2114] Frank Fiedrich and Paul Burghardt.

Agent-based systems for disaster management. Communications of the ACM, 50(3):41–42, March 2007. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mendonca:2007:CAM

[2115] David Mendonça, Theresa Jefferson, and John Harrald. Collaborative adhocracies and mix-and-match technologies in emergency management. Communications of the ACM, 50(3):44–49, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Manoj:2007:CCE

[2116] B. S. Manoj and Alexandra Hubenko Baker. Communication challenges in emergency response. Communications of the ACM, 50(3):51–53, March 2007.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Palen:2007:OFS

[2117] Leysia Palen, Starr Roxanne Hiltz, and Sophia B. Liu. Online forums supporting grassroots participation in emergency preparedness and response. Communications of the ACM, 50(3):54–58, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Botterell:2007:PWN

[2118] Art Botterell and Ronja Addams-Moring. Public warning in the networked age: open standards to the rescue? Communications of the ACM, 50(3):59–60, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Currion:2007:OSS

[2119] Paul Currion, Chamindra de Silva, and Bartel Van de Walle. Open source software for disaster management. Communications of the ACM, 50(3):61–65, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andriole:2007:HHE

[2120] Stephen J. Andriole. The 7 habits of highly effective technology leaders. Communications of the ACM, 50(3):66–72, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horan:2007:TCI

[2121] Thomas A. Horan and Benjamin L. Schooley. Time-critical information services. Communications of the ACM, 50(3):73–78, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nerur:2007:TRA

[2122] Sridhar Nerur and VenuGopal Balijepally. Theoretical reflections on agile development methodologies. Communications of the ACM, 50(3):79–83, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCoy:2007:EOA

[2123] Scott McCoy, Andrea Everard, Peter Polak, and Dennis F. Galletta. The effects of online advertising. Communications of the ACM, 50(3):84–88, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhong:2007:EII

[2124] Ning Zhong, Jiming Liu, and Yiyu Yao. Envisioning intelligent information technologies through the prism of web intelligence. Communications of the ACM, 50(3):89–94, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:2007:SOA

[2125] James A. Hall and Stephen L. Liedtka.

The Sarbanes–Oxley Act: implications for large-scale IT outsourcing. Communications of the ACM, 50(3):95–100, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAc

[2126] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(3):101–102, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ward:2007:TOP

[2127] Nigel Ward. Technical opinion: The (un)predictability of computer science graduate school admissions. Communications of the ACM, 50(3):104–106, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parker:2007:IRR

[2128] Donn B. Parker. Inside risks: Risks of risk-based security. Communications of the ACM, 50(3):120, March 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPd

[2129] Diane Crawford. Editorial pointers. Communications of the ACM, 50(4):5, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:SCA

[2130] Diane Crawford. Forum: Stop chasing the AI illusion. Communications of the ACM, 50(4):7–9, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTd

[2131] CACM Staff. News track. Communications of the ACM, 50(4):13–14, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2007:DVH

[2132] Hal Berghel. Digital village: Hiding data, forensics, and anti-forensics.

Communications of the ACM, 50(4): 15–20, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2007:MM

[2133] Peter J. Denning. The profession of IT:

Mastering the mess. Communications of the ACM, 50(4):21–25, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:AHD

[2134] CACM Staff. ACM honors distinguished members. Communications of the ACM, 50(4):26–27, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:CNA

[2135] CACM Staff. Call for nominations for advanced member grades in ACM. Communications of the ACM, 50(4): 28–29, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:PDS

[2136] CACM Staff. Proposal to dissolve SIGAPL. Communications of the ACM, 50(4):29, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2007:VFG

[2137] Jenova Chen. Viewpoint: Flow in games (and everything else). Communications of the ACM, 50(4):31–34, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kramer:2007:AKC

[2138] Jeff Kramer. Is abstraction the key to computing? Communications of the ACM, 50(4):36–42, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nevo:2007:HAD

[2139] Dorit Nevo and Michael R. Wade. How to avoid disappointment by design. Communications of the ACM, 50(4):43–48, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Orman:2007:CSS

[2140] Levent V. Orman. Consumer support systems. Communications of the ACM, 50(4):49–54, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ceri:2007:DDI

[2141] Stefano Ceri, Maristella Matera, Francesca Rizzo, and Vera Demaldé. Designing data-intensive web applications for content accessibility using web marts. Communications of the ACM, 50(4):55–61, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Green:2007:AAI

[2142] Ido Green, Tzvi Raz, and Moshe Zviran. Analysis of active intrusion prevention data for predicting hostile activity in computer networks. Communications of the ACM, 50(4):63–68, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sakthivel:2007:MRO

[2143] S. Sakthivel. Managing risk in offshore systems development. Communications of the ACM, 50(4):69–75, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stone:2007:ID

[2144] Jeffrey A. Stone and Elinor Madigan.

Inconsistencies and disconnects. Communications of the ACM, 50(4):76–79, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:2007:IVB

[2145] Dianne J. Hall and David Paradice.

Investigating value-based decision bias and mediation: do you do as you think? Communications of the ACM, 50(4):81–85, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chau:2007:WPS

[2146] Michael Chau, Xiao Fang, and Olivia R. Liu Sheng. What are people searching on government web sites? Communications of the ACM, 50(4):87–92, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAd

[2147] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(4):93–94, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Butler:2007:DTI

[2148] E. Sonny Butler, Cheryl Aasheim, and Susan Williams. Technical opinion: Does telecommuting improve productivity? Communications of the ACM, 50(4):101–103, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horning:2007:IRR

[2149] Jim Horning. Inside risks: Risks of virtual professionalism. Communications of the ACM, 50(4):104, April 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPe

[2150] Diane Crawford. Editorial pointers. Communications of the ACM, 50(5):5– 6, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fd

[2151] Diane Crawford. Forum. Communications of the ACM, 50(5):7, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2007:APH

[2152] Stuart I. Feldman. President’s letter: ACM’s past helps steer its future. Communications of the ACM, 50(5):10, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTe

[2153] CACM Staff. News track. Communications of the ACM, 50(5):11–12, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2007:SCP

[2154] Meg McGinity Shannon. Staying connected: The pearly gates of Internet TV. Communications of the ACM, 50(5):13–15, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jenkins:2007:VRA

[2155] Stephen B. Jenkins. Viewpoint: The rime of the ancyent programmer. Communications of the ACM, 50(5):17–20, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2007:OMQ

[2156] Robert L. Glass. Practical programmer: One man’s quest for the state of software engineering’s practice. Communications of the ACM, 50(5):21–23, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wise:2007:ALB

[2157] David S. Wise. ACM at 60: a look back in time: Introduction. Communications of the ACM, 50(5):24–26, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Longo:2007:AED

[2158] Bernadette Longo. ACM established to develop communication about computing. Communications of the ACM, 50(5):27–29, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Akera:2007:EBO

[2159] Atsushi Akera. Edmund Berkeley and the origins of ACM. Communications of the ACM, 50(5):30–35, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Haigh:2007:SAH

[2160] Thomas Haigh, Elisabeth Kaplan, and Carrie Seib. Sources for ACM history: what, where, why. Communications of the ACM, 50(5):36–41, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Care:2007:ODR

[2161] Charles Care. Not only digital: a review of ACM’s early involvement with analog computing technology. Communications of the ACM, 50(5):42–45, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hemmendinger:2007:AIC

[2162] David Hemmendinger. The ACM and IEEE-CS guidelines for undergraduate CS education. Communications of the ACM, 50(5):46–53, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:2007:HAS

[2163] Judy Brown and Steve Cunningham. A history of ACM SIGGRAPH. Communications of the ACM, 50(5):54–61, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Edmondson-Yurkanan:2007:SAJ

[2164] Chris Edmondson-Yurkanan. SIG-COMM’s archaeological journey into networking’s past. Communications of the ACM, 50(5):63–68, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bergin:2007:HHP

[2165] Thomas J. (Tim) Bergin. A history of the history of programming languages. Communications of the ACM, 50(5):69–74, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McBride:2007:SA

[2166] Matthew R. McBride. The software architect. Communications of the ACM, 50(5):75–81, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Obrenovic:2007:UAM

[2167] Zeljko Obrenovic, Julio Abascal, and Dusan Starcevic. Universal accessibility as a multimodal design issue. Communications of the ACM, 50(5):83–88, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harris:2007:WPS

[2168] Michael Harris, Kris Aebischer, and Tim Klaus. The whitewater process: software product development in small IT businesses. Communications of the ACM, 50(5):89–93, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

He:2007:ADW

[2169] Bin He, Mitesh Patel, Zhen Zhang, and Kevin Chen-Chuan Chang. Accessing the deep web. Communications of the ACM, 50(5):94–101, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

TorresdaSilva:2007:MMA

[2170] Viviane Torres da Silva and Carlos J. P. de Lucena. Modeling multi-agent systems. Communications of the ACM, 50(5):103–108, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DevitoDaCunha:2007:DPM

[2171] Alessandra Devito Da Cunha and David Greathead. Does personality matter?: an analysis of code-review ability. Communications of the ACM, 50(5):109–112, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goodell:2007:TOR

[2172] Geoffrey Goodell and Paul Syverson. Technical opinion: The right place at the right time. Communications of the ACM, 50(5):113–117, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAe

[2173] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(5):124–125, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schneier:2007:IRP

[2174] Bruce Schneier. Inside risks: The psychology of security. Communications of the ACM, 50(5):128, May 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPf

[2175] Diane Crawford. Editorial pointers.

Communications of the ACM, 50(6):5, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fe

[2176] Diane Crawford. Forum. Communications of the ACM, 50(6):7–9, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTf

[2177] CACM Staff. News track. Communications of the ACM, 50(6):13–14, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2007:LSS

[2178] Pamela Samuelson. Legally speaking:

Software patents and the metaphysics of Section 271(f). Communications of the ACM, 50(6):15–19, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2007:BST

[2179] Phillip G. Armour. The business of software: Twenty percent. Communications of the ACM, 50(6):21–23, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:2007:VWS

[2180] Ben Shneiderman. Viewpoint: Web science: a provocative invitation to computer science. Communications of the ACM, 50(6):25–27, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanHeck:2007:SBN

[2181] Eric van Heck and Peter Vervest. Smart business networks: how the network wins. Communications of the ACM, 50(6):28–37, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dekleva:2007:EEI

[2182] Sasha Dekleva, J. P. Shim, Upkar Varshney, and Geoffrey Knoerzer. Evolution and emerging issues in mobile wireless networks. Communications of the ACM, 50(6):38–43, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Friedlander:2007:DPT

[2183] Amy Friedlander, Allison Mankin, W. Douglas Maughan, and Stephen D. Crocker. DNSSEC: a protocol toward securing the Internet infrastructure. Communications of the ACM, 50(6):44–50, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Constantiou:2007:FIS

[2184] Ioanna D. Constantiou, Jan Damsgaard, and Lars Knutsen. The four incremental steps toward advanced mobile service adoption. Communications of the ACM, 50(6):51–55, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cotteleer:2007:CCC

[2185] Mark J. Cotteleer, Christopher A. Cotteleer, and Andrew Prochnow. Cutting checks: challenges and choices in B2B e-payments. Communications of the ACM, 50(6):56–61, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jeyaraj:2007:EIE

[2186] Anand Jeyaraj and Vicki L. Sauter. An empirical investigation of the effectiveness of systems modeling and verification tools. Communications of the ACM, 50(6):62–67, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Avison:2007:TRH

[2187] David Avison and Terry Young. Time to rethink health care and ICT? Communications of the ACM, 50(6):69–74, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Framling:2007:DPM

[2188] Kary Främling, Timo Ala-Risku, Mikko Kärkkäinen, and Jan Holmström. Design patterns for managing product life cycle information. Communications of the ACM, 50(6):75–79, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ren:2007:AVP

[2189] Jie Ren and Richard N. Taylor. Automatic and versatile publications ranking for research institutions and scholars. Communications of the ACM, 50(6):81–85, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAf

[2190] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(6):86–87, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cocosila:2007:TOP

[2191] Mihail Cocosila, Ofir Turel, Norm Archer, and Yufei Yuan. Technical opinion: Perceived health risks of 3G cell phones: do users care? Communications of the ACM, 50(6):89–92, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:2007:IRR

[2192] Peter A. Freeman. Inside risks: Risks are your responsibility. Communications of the ACM, 50(6):104, June 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPg

[2193] Diane Crawford. Editorial pointers. Communications of the ACM, 50(7):5, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTg

[2194] CACM Staff. News track. Communications of the ACM, 50(7):7–8, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2007:PICa

[2195] Peter J. Denning. The profession of IT: Computing is a natural science. Communications of the ACM, 50(7):13–18, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2007:TSMb

[2196] Michael A. Cusumano. Technology strategy and management: The changing labyrinth of software pricing. Communications of the ACM, 50(7):19–22, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mowshowitz:2007:VPV

[2197] Abbe Mowshowitz and Nanda Kumar.

Viewpoint: Public vs. private interest on the Internet. Communications of the ACM, 50(7):23–25, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zyda:2007:CSG

[2198] Michael Zyda. Creating a science of games: Introduction. Communications of the ACM, 50(7):26–29, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mayo:2007:GSE

[2199] Merrilea J. Mayo. Games for science and engineering education. Communications of the ACM, 50(7):30–35, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chatham:2007:GT

[2200] Ralph E. Chatham. Games for training. Communications of the ACM, 50(7):36–43, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kelly:2007:HBS

[2201] Henry Kelly, Kay Howell, Eitan Glinert, Loring Holding, Chris Swain, Adam Burrowbridge, and Michelle Roper. How to build serious games. Communications of the ACM, 50(7): 44–49, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pausch:2007:CME

[2202] Randy Pausch and Don Marinelli. Carnegie Mellon’s entertainment technology center: combining the left and right brain. Communications of the ACM, 50(7):50–57, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kelleher:2007:USM

[2203] Caitlin Kelleher and Randy Pausch. Using storytelling to motivate programming. Communications of the ACM, 50(7):58–64, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raghuvanshi:2007:RTS

[2204] Nikunj Raghuvanshi, Christian Lauterbach, Anish Chandak, Dinesh Manocha, and Ming C. Lin. Real-time sound synthesis and propagation for games. Communications of the ACM, 50(7): 66–73, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schumaker:2007:EET

[2205] Robert P. Schumaker, Ying Liu, Mark Ginsburg, and Hsinchun Chen. Evaluating the efficacy of a terrorism question/answer system. Communications of the ACM, 50(7):74–80, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2007:ANG

[2206] In Lee. An architecture for a nextgeneration holistic e-recruiting system. Communications of the ACM, 50(7): 81–85, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wu:2007:CSO

[2207] Harris Wu and Michael Gordon. Collaborative structuring: organizing document repositories effectively and efficiently. Communications of the ACM, 50(7):86–91, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Niederman:2007:ERA

[2208] Fred Niederman, Richard G. Mathieu, Roger Morley, and Ik-Whan Kwon. Examining RFID applications in supply chain management. Communications of the ACM, 50(7):92–101, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kesh:2007:KAI

[2209] Someswar Kesh and Pauline Ratnasingam. A knowledge architecture for IT security. Communications of the ACM, 50(7):103–108, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anton:2007:RCO

[2210] Annie I. Antón, Elisa Bertino, Ninghui Li, and Ting Yu. A roadmap for comprehensive online privacy policy management. Communications of the ACM, 50(7):109–116, July 2007. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rodgers:2007:EWB

[2211] Waymond Rodgers and Solomon Negash. The effects of Web-based technologies on knowledge transfer. Communications of the ACM, 50(7):117– 122, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Al-Fedaghi:2007:TOC

[2212] Sabah Al-Fedaghi. Technical opinion: Conferences under scrutiny. Communications of the ACM, 50(7):123–126, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAg

[2213] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(7): 135, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Perrow:2007:IRD

[2214] Charles Perrow. Inside risks: Disasters evermore? Communications of the ACM, 50(7):136, July 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPh

[2215] Diane Crawford. Editorial pointers. Communications of the ACM, 50(8):5, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTh

[2216] CACM Staff. News track. Communications of the ACM, 50(8):7–8, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Ff

[2217] Diane Crawford. Forum. Communications of the ACM, 50(8):13–14, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:CNA

[2218] Diane Crawford. Call for nominations for ACM general election. Communications of the ACM, 50(8):14, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2007:DVB

[2219] Hal Berghel. Digital village: Betterthan-nothing security practices. Communications of the ACM, 50(8):15–18, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruckman:2007:VSC

[2220] Amy S. Bruckman. Viewpoint: Software copyright and role models. Communications of the ACM, 50(8):19–20, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2007:PPD

[2221] Robert L. Glass. Practical programmer: A deja-vu look at software engineering researchers who care about practice. Communications of the ACM, 50(8):21–23, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kobsa:2007:PEP

[2222] Alfred Kobsa. Privacy-enhanced personalization. Communications of the ACM, 50(8):24–33, August 2007. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elfatatry:2007:DCC

[2223] Ahmed Elfatatry. Dealing with change: components versus services. Communications of the ACM, 50(8):35–39, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thatcher:2007:MIV

[2224] Matt E. Thatcher and David E. Pingry. Modeling the IT value paradox. Communications of the ACM, 50(8):41–45, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Erickson:2007:TPC

[2225] John Erickson and Keng Siau. Theoretical and practical complexity of modeling methods. Communications of the ACM, 50(8):46–51, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simonsen:2007:ITM

[2226] Jesper Simonsen. Involving top management in IT projects. Communications of the ACM, 50(8):52–58, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

LeBlanc:2007:DLS

[2227] Larry J. LeBlanc and Michael R.

Galbreth. Designing large-scale supply chain linear programs in spreadsheets. Communications of the ACM, 50(8):59–64, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wan:2007:CPC

[2228] Yun Wan, Satya Menon, and Arkalgud Ramaprasad. A classification of product comparison agents. Communications of the ACM, 50(8):65–71, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sipior:2007:UIO

[2229] Janice C. Sipior and Burke T. Ward.

Unintended invitation: organizational wi-fi use by external roaming users. Communications of the ACM, 50 (8):72–77, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tse:2007:TOO

[2230] T. H. Tse, Francis C. M. Lau, W. K.

Chan, Peter C. K. Liu, and Colin K. F. Luk. Testing object-oriented industrial software without precise oracles or results. Communications of the ACM, 50(8):78–85, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sen:2007:DDS

[2231] Sagnika Sen, Haluk Demirkan, and Michael Goul. Dynamic decision support through instantiation of UEML representations. Communications of the ACM, 50(8):87–93, August 2007.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raghupathi:2007:CGI

[2232] W. “RP” Raghupathi. Corporate governance of IT: a framework for development. Communications of the ACM, 50(8):94–99, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAh

[2233] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(8):100, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramingwong:2007:TOO

[2234] Sakgasit Ramingwong and A. S. M. Sajeev. Technical opinion: Offshore outsourcing: the risk of keeping mum. Communications of the ACM, 50(8):101–103, August 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:2007:IRW

[2235] David Lorge Parnas. Inside risks: Which is riskier: OS diversity or OS monopoly? Communications of the ACM, 50(8):112, August 2007. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPi

[2236] Diane Crawford. Editorial pointers. Communications of the ACM, 50(9):5, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTi

[2237] CACM Staff. News track. Communications of the ACM, 50(9):7–8, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fg

[2238] Diane Crawford. Forum. Communications of the ACM, 50(9):13–14, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waters:2007:WYC

[2239] Neil L. Waters. Viewpoint: Why you can’t cite Wikipedia in my class. Communications of the ACM, 50(9): 15–17, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boisvert:2007:OFE

[2240] Ronald F. Boisvert, Mary Jane Irwin, and Holly Rushmeier. Online first: Evolving the ACM journal distribution program. Communications of the ACM, 50(9):19–20, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2007:SCSb

[2241] Meg McGinity Shannon. Staying connected: Shaking hands, kissing babies, and … blogging? Communications of the ACM, 50(9):21–24, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2007:BSC

[2242] Phillip G. Armour. The business of software: The conservation of uncertainty. Communications of the ACM, 50(9):25–28, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Munakata:2007:BSN

[2243] Toshinori Munakata. Beyond silicon: new computing paradigms: Introduction. Communications of the ACM, 50 (9):30–34, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robinett:2007:CTC

[2244] Warren Robinett, Gregory S. Snider, Philip J. Kuekes, and R. Stanley Williams. Computing with a trillion crummy components. Communications of the ACM, 50(9):35–39, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kong:2007:CCN

[2245] Jing Kong. Computation with carbon nanotube devices. Communications of the ACM, 50(9):40–42, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stadler:2007:MCO

[2246] Robert Stadler. Molecular, chemical, and organic computing. Communications of the ACM, 50(9):43–45, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reif:2007:APB

[2247] John H. Reif and Thomas H. LaBean. Autonomous programmable biomolecular devices using self-assembled DNA nanostructures. Communications of the ACM, 50(9):46–53, September 2007.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bacon:2007:TWQ

[2248] Dave Bacon and Debbie Leung. Toward a world with quantum computers. Communications of the ACM, 50 (9):55–59, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abdeldayem:2007:OCN

[2249] Hossin Abdeldayem and Donald O.

Frazier. Optical computing: need and challenge. Communications of the ACM, 50(9):60–62, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marr:2007:MNC

[2250] David W. M. Marr and Toshinori Munakata. Micro/nanofluidic computing. Communications of the ACM, 50 (9):64–68, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aono:2007:ABN

[2251] Masashi Aono, Masahiko Hara, and Kazuyuki Aihara. Amoeba-based neurocomputing with chaotic dynamics. Communications of the ACM, 50(9): 69–72, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marowka:2007:PCD

[2252] Ami Marowka. Parallel computing on any desktop. Communications of the ACM, 50(9):74–78, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lepouras:2007:DEU

[2253] George Lepouras, Costas Vassilakis, Constantin Halatsis, and Panagiotis Georgiadis. Domain expert user development: the SmartGov approach. Communications of the ACM, 50(9): 79–83, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Valacich:2007:OCH

[2254] Joseph S. Valacich, D. Veena Parboteeah, and John D. Wells. The online consumer’s hierarchy of needs. Communications of the ACM, 50(9): 84–90, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brown:2007:WMW

[2255] Susan A. Brown, Norman L. Chervany, and Bryan A. Reinicke. What matters when introducing new information technology. Communications of the ACM, 50(9):91–96, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Braganza:2007:SCP

[2256] Ashley Braganza and Arnoud Franken. SOX, compliance, and power relationships. Communications of the ACM, 50 (9):97–102, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pollach:2007:WWO

[2257] Irene Pollach. What’s wrong with online privacy policies? Communications of the ACM, 50(9):103–108, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stafford:2007:TOO

[2258] Thomas F. Stafford and Dennis Gonier. Technical opinion: The online research ’bubble’. Communications of the ACM, 50(9):109–112, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAi

[2259] Diane Crawford. Hot links: Top 10 downloads from ACM’s digital library. Communications of the ACM, 50(9):118, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2007:IRM

[2260] Peter G. Neumann. Inside risks: Emigrating risks? Communications of the ACM, 50(9):120, September 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPj

[2261] Diane Crawford. Editorial pointers. Communications of the ACM, 50(10):5, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTj

[2262] CACM Staff. News track. Communications of the ACM, 50(10):7–8, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fh

[2263] Diane Crawford. Forum. Communications of the ACM, 50(10):13–14, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2007:TSMc

[2264] Michael A. Cusumano. Technology strategy and management: Extreme programming compared with Microsoft-style iterative development. Communications of the ACM, 50 (10):15–18, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2007:LSD

[2265] Pamela Samuelson. Legally speaking: Does copyright law need to be reformed? Communications of the ACM, 50(10):19–23, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Passerini:2007:TOR

[2266] Katia Passerini, Karen Patten, Michael R. Bartolacci, and Jerry Fjermestad. Technical opinion: Reflections and trends in the expansion of cellular wireless services in the U.S. and China. Communications of the ACM, 50(10):25–28, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thrun:2007:VWW

[2267] Sebastian Thrun. Viewpoint: Why we compete in DARPA’s Urban Challenge autonomous robot race. Communications of the ACM, 50(10):29–31, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Best:2007:IPP

[2268] Michael L. Best, Kipp Jones, Illenin Kondo, Dhanaraj Thakur, Edem Wornyo, and Calvin Yu. International perspectives: Post-conflict communications: the case of Liberia. Communications of the ACM, 50(10):33–39, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carmel:2007:WNM

[2269] Erran Carmel and Pamela Abbott.

Why ’nearshore’ means that distance matters. Communications of the ACM, 50(10):40–46, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carmel:2007:NLT

[2270] Erran Carmel and Pamela Abbott.

Nearshore locations in three principal global clusters and their respective attributes. Communications of the ACM, 50(10):45, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thatcher:2007:SPG

[2271] Matt E. Thatcher and David E. Pingry.

[software patents] the good, the bad, and the messy. Communications of the ACM, 50(10):47–52, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mentzas:2007:KSS

[2272] Gregoris Mentzas, Kostas Kafentzis, and Panos Georgolios. Knowledge services on the semantic web. Communications of the ACM, 50(10):53–58, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Umesh:2007:GIM

[2273] U. N. Umesh, Len Jessup, and Minh Q. Huynh. [getting ideas to market] current issues faced by technology entrepreneurs. Communications of the ACM, 50(10):60–66, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Akbulut:2007:ISP

[2274] Asli Yagmur Akbulut and Clayton Arlen Looney. Inspiring students to pursue computing degrees. Communications of the ACM, 50(10):67–71, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Konana:2007:PPR

[2275] Prabhudev Konana and Gautam Ray. Physical product reengineering with embedded information technology. Communications of the ACM, 50 (10):72–78, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Papazoglou:2007:BPD

[2276] Michael P. Papazoglou and Willem-Jan van den Heuvel. Business process development life cycle methodology. Communications of the ACM, 50 (10):79–85, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuechler:2007:BAU

[2277] William L. Kuechler. Business applications of unstructured text. Communications of the ACM, 50(10):86–93, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jagatic:2007:SP

[2278] Tom N. Jagatic, Nathaniel A. Johnson, Markus Jakobsson, and Filippo Menczer. Social phishing. Communications of the ACM, 50(10):94–100, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baker:2007:NMM

[2279] Wade H. Baker, Loren Paul Rees, and Peter S. Tippett. Necessary measures: metric-driven information security risk assessment and decision making. Communications of the ACM, 50 (10):101–106, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goeke:2007:LFY

[2280] Richard J. Goeke and Robert H. Faley.

Leveraging the flexibility of your data warehouse. Communications of the ACM, 50(10):107–111, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DArcy:2007:DII

[2281] John D’Arcy and Anat Hovav. Deterring internal information systems misuse. Communications of the ACM, 50 (10):113–117, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAj

[2282] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(10):118, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lin:2007:IRT

[2283] Herbert S. Lin, Alfred Z. Spector, Peter G. Neumann, and Seymour E. Goodman. Inside risks: Toward a safer and more secure cyberspace. Communications of the ACM, 50(10):128, October 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPk

[2284] Diane Crawford. Editorial pointers. Communications of the ACM, 50(11):5, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTk

[2285] CACM Staff. News track. Communications of the ACM, 50(11):7–8, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2007:PICb

[2286] Peter J. Denning. The profession of IT: The choice uncertainty principle. Communications of the ACM, 50 (11):9–14, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2007:PPR

[2287] Robert L. Glass. Practical programmer: A research project with important practitioner-oriented findings. Communications of the ACM, 50(11): 15–16, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2007:RAA

[2288] Stuart I. Feldman. President’s letter: The rewards of ACM’s awards. Communications of the ACM, 50(11): 17–18, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parnas:2007:VSN

[2289] David Lorge Parnas. Viewpoint: Stop the numbers game. Communications of the ACM, 50(11):19–21, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feitelson:2007:ECS

[2290] Dror G. Feitelson. Experimental computer science: Introduction. Communications of the ACM, 50(11):24–26, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGeoch:2007:EA

[2291] Catherine C. McGeoch. Experimental algorithmics. Communications of the ACM, 50(11):27–31, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Basili:2007:ESB

[2292] Victor R. Basili and Marvin V.

Zelkowitz. Empirical studies to build a science of computer science. Communications of the ACM, 50(11):33–37, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Peterson:2007:EDE

[2293] Larry Peterson and Vivek S. Pai.

Experience-driven experimental systems research. Communications of the ACM, 50(11):38–44, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2007:WCU

[2294] Alan Jay Smith. Workloads (creation and use). Communications of the ACM, 50(11):45–50, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Voorhees:2007:TCI

[2295] Ellen M. Voorhees. TREC: Continuing information retrieval’s tradition of experimentation. Communications of the ACM, 50(11):51–54, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jackel:2007:HDS

[2296] L. D. Jackel, Douglas Hackett, Eric Krotkov, Michael Perschbacher, James Pippine, and Charles Sullivan. How DARPA structures its robotics programs to improve locomotion and navigation. Communications of the ACM, 50(11):55–59, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nov:2007:WMW

[2297] Oded Nov. What motivates wikipedians? Communications of the ACM, 50 (11):60–64, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tallon:2007:ILC

[2298] Paul P. Tallon and Richard Scannell. Information life cycle management. Communications of the ACM, 50 (11):65–69, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Foshay:2007:DDW

[2299] Neil Foshay, Avinandan Mukherjee, and Andrew Taylor. Does data warehouse end-user metadata add value? Communications of the ACM, 50(11): 70–77, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sauer:2007:ISV

[2300] Chris Sauer, Andrew Gemino, and Blaize Horner Reich. The impact of size and volatility on IT project performance. Communications of the ACM, 50(11):79–84, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thomas:2007:MKW

[2301] Dominic M. Thomas, Robert P.

Bostrom, and Marianne Gouge. Making knowledge work in virtual teams. Communications of the ACM, 50(11): 85–90, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2007:DPP

[2302] Hee-Woong Kim and Yunjie Xu.

Drivers of price premium in e-markets. Communications of the ACM, 50(11): 91–95, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAk

[2303] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(11):97, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bishop:2007:IRR

[2304] Matt Bishop and David Wagner. Inside risks: Risks of e-voting. Communications of the ACM, 50(11):120, November 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:EPl

[2305] Diane Crawford. Editorial pointers. Communications of the ACM, 50(12):5, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NTl

[2306] CACM Staff. News track. Communications of the ACM, 50(12):7–8, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:Fi

[2307] Diane Crawford. Forum. Communications of the ACM, 50(12):9–10, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2007:DVC

[2308] Hal Berghel. Digital village: Credit card forensics. Communications of the ACM, 50(12):11–14, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gaskins:2007:VPB

[2309] Robert Gaskins. Viewpoint: Power-Point at 20: back to basics. Communications of the ACM, 50(12):15–17, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shneiderman:2007:CST

[2310] Ben Shneiderman. Creativity support tools: accelerating discovery and innovation. Communications of the ACM, 50(12):20–32, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2007:YCC

[2311] Dongwon Lee, Jaewoo Kang, Prasenjit Mitra, C. Lee Giles, and Byung-Won On. Are your citations clean? Communications of the ACM, 50(12): 33–38, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2007:NER

[2312] CACM Staff. Nominees for elections and report of the ACM nominating committee. Communications of the ACM, 50(12):39, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aalberts:2007:TNS

[2313] Robert J. Aalberts, Percy Poon, and Paul Thistle. Trespass, nuisance, and spam: 11th century common law meets the internet. Communications of the ACM, 50(12):40–45, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lopez:2007:MRD

[2314] Antonio M. Lopez, Jr., Frederick G. Lopez, Robert W. Lent, and Madonna G. Constantine. Multidisciplinary research on the datapath of the computing disciplines. Communications of the ACM, 50(12):46–50, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:2007:MBU

[2315] Erica L. Wagner and Gabriele Piccoli. Moving beyond user participation to achieve successful IS design. Communications of the ACM, 50 (12):51–55, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boh:2007:SDD

[2316] Wai Fong Boh, Christina Soh, and Steven Yeo. Standards development and diffusion: a case study of RosettaNet. Communications of the ACM, 50(12):57–62, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oshri:2007:ICR

[2317] Ilan Oshri, Sue Newell, and Shan L. Pan. Implementing component reuse strategy in complex products environments. Communications of the ACM, 50(12):63–67, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mohan:2007:TVS

[2318] Kannan Mohan and Balasubramaniam Ramesh. Tracing variations in software product families. Communications of the ACM, 50(12):68–73, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schild:2007:SOM

[2319] Klaus Schild and Stefan Bussmann.

Self-organization in manufacturing operations. Communications of the ACM, 50(12):74–79, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grover:2007:FIB

[2320] Varun Grover, Raymond M. Henry, and Jason B. Thatcher. Fix IT-business relationships through better decision rights. Communications of the ACM, 50(12):80–86, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ray:2007:CIR

[2321] Gautam Ray, Waleed A. Muhanna, and Jay B. Barney. Competing with IT: the role of shared IT-business understanding. Communications of the ACM, 50 (12):87–91, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2007:TDAl

[2322] Diane Crawford. Hot links: Top 10 downloads from ACM’s Digital Library. Communications of the ACM, 50(12):92, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

dePillis:2007:TOC

[2323] Emmeline de Pillis and Kimberly Furumo. Technical opinion: Counting the cost of virtual teams. Communications of the ACM, 50(12):93–95, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tan:2007:TOT

[2324] Chuan-Hoo Tan and Hock-Hai Teo. Technical opinion: Training future software developers to acquire agile development skills. Communications of the ACM, 50(12):97–98, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bellovin:2007:ISE

[2325] Steven M. Bellovin, Matt Blaze, Whitfield Diffie, Susan Landau, Jennifer Rexford, and Peter G. Neumann. Inside risks: Internal surveillance, external risks. Communications of the ACM, 50(12):128, December 2007. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPa

[2326] Diane Crawford. Editorial pointers. Communications of the ACM, 51(1): 5, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:Fa

[2327] Diane Crawford. Forum. Communications of the ACM, 51(1):7–9, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2008:BSD

[2328] Phillip G. Armour. The business of software: Digging CACM. Communications of the ACM, 51(1):13–15, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2008:SCH

[2329] Meg McGinity Shannon. Staying connected: Happy anniversary, CACM. Communications of the ACM, 51 (1):17–19, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Freeman:2008:VBE

[2330] Peter A. Freeman. Viewpoint: Back to experimentation. Communications of the ACM, 51(1):21–22, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2008:FYS

[2331] Stuart I. Feldman. President’s letter:

Fifty years and still growing. Communications of the ACM, 51(1):23, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EDI

[2332] Diane Crawford. From the editor’s desk: Introduction. Communications of the ACM, 51(1):24–26, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gotlieb:2008:TRP

[2333] Calvin C. Gotlieb. A time to retrospect and prospect. Communications of the ACM, 51(1):26–29, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lynn:2008:BC

[2334] M. Stuart Lynn. The battle of the covers. Communications of the ACM, 51(1):30–32, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ashenhurst:2008:BBS

[2335] Robert L. Ashenhurst. The battle behind the scenes. Communications of the ACM, 51(1):33–34, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2008:DVA

[2336] Peter J. Denning. Déjà vu all over again. Communications of the ACM, 51(1):35–39, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2008:AE

[2337] Jacques Cohen. From academia to the editorship. Communications of the ACM, 51(1):41–43, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2008:CPP

[2338] Moshe Y. Vardi. CACM: past, present, and future. Communications of the ACM, 51(1):44–48, January 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rosenbloom:2008:VI

[2339] Andrew Rosenbloom. Voices: Introduction. Communications of the ACM, 51(1):50–51, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:2008:RIC

[2340] Jon Bentley. In the realm of insight and creativity. Communications of the ACM, 51(1):52–54, January 2008.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Diffie:2008:ISY

[2341] Whitfield Diffie. Information security:

50 years behind, 50 years ahead. Communications of the ACM, 51(1):55–57, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wing:2008:FDQ

[2342] Jeannette M. Wing. Five deep questions in computing. Communications of the ACM, 51(1):58–60, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spafford:2008:IT

[2343] Eugene H. Spafford. Inspiration and trust. Communications of the ACM, 51(1):61–62, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brooks:2008:NY

[2344] Rodney Brooks. The next 50 years.

Communications of the ACM, 51 (1):63–64, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2008:HIP

[2345] Pamela Samuelson. Hacking intellectual property law. Communications of the ACM, 51(1):65–67, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jenkins:2008:OC

[2346] Stephen B. Jenkins. Ode to code.

Communications of the ACM, 51(1):

67, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agha:2008:CPC

[2347] Gul Agha. Computing in pervasive cyberspace. Communications of the ACM, 51(1):68–70, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Conti:2008:CGT

[2348] Gregory Conti. Could googling take down a president? Communications of the ACM, 51(1):71–73, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crowcroft:2008:TNA

[2349] Jon Crowcroft. Toward a network architecture that does everything. Communications of the ACM, 51(1):74–77, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2008:RCR

[2350] Peter G. Neumann. Reflections on computer-related risks. Communications of the ACM, 51(1):78–80, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leigh:2008:CCM

[2351] Jason Leigh and Maxine D. Brown. Cyber-commons: merging real and virtual worlds. Communications of the ACM, 51(1):82–85, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2008:BLB

[2352] Gordon Bell. Bell’s law for the birth and death of computer classes. Communications of the ACM, 51(1):86–94, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Polites:2008:CPC

[2353] Greta L. Polites and Richard T. Watson. The centrality and prestige of CACM. Communications of the ACM, 51(1):95–100, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:BRP

[2354] Diane Crawford. Breakthrough research: a preview of things to come: Introduction. Communications of the ACM, 51(1):104, January 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2008:TPD

[2355] David A. Patterson. Technical perspective: the data center is the computer. Communications of the ACM, 51(1):105, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dean:2008:MSD

[2356] Jeffrey Dean and Sanjay Ghemawat.

MapReduce: simplified data processing on large clusters. Communications of the ACM, 51(1):107–113, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chazelle:2008:TPF

[2357] Bernard Chazelle. Technical perspective: finding a good neighbor, near and fast. Communications of the ACM, 51(1):115, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andoni:2008:NOH

[2358] Alexandr Andoni and Piotr Indyk. Near-optimal hashing algorithms for approximate nearest neighbor in high dimensions. Communications of the ACM, 51(1):117–122, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zegans:2008:PR

[2359] Leonard S. Zegans. Inside risks: The psychology of risks. Communications of the ACM, 51(1):152, January 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPb

[2360] Diane Crawford. Editorial pointers. Communications of the ACM, 51(2):5–6, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:NTa

[2361] CACM Staff. News track. Communications of the ACM, 51(2):7–8, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:Fb

[2362] Diane Crawford. Forum. Communications of the ACM, 51(2):9–11, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:SAC

[2363] Diane Crawford. SIGs announce candidates for election. Communications of the ACM, 51(2):14, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2008:TSMa

[2364] Michael A. Cusumano. Technology strategy and management: Managing software development in globally distributed teams. Communications of the ACM, 51(2):15–17, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oram:2008:VPP

[2365] Andy Oram. Viewpoint: Peer to patent needs your expertise. Communications of the ACM, 51(2):19–20, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2008:LSH

[2366] Pamela Samuelson, Jerome H. Reichman, and Graeme Dinwoodie. Legally speaking: How to achieve (some) balance in anti-circumvention laws. Communications of the ACM, 51(2):21–25, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2008:AAR

[2367] Stuart I. Feldman. ACM annual report for FY07. Communications of the ACM, 51(2):27–34, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2008:ARG

[2368] Jeffrey Y. Kim, Jonathan P. Allen, and Elan Lee. Alternate reality gaming. Communications of the ACM, 51 (2):36–42, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shih:2008:IDD

[2369] Eric Shih, Kenneth L. Kraemer, and Jason Dedrick. IT diffusion in developing countries. Communications of the ACM, 51(2):43–48, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Keane:2008:PBT

[2370] Mark T. Keane, Maeve O’Brien, and Barry Smyth. Are people biased in their use of search engines? Communications of the ACM, 51(2):49–52, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taylor:2008:ASB

[2371] Nolan J. Taylor, Eleanor T. Loiacono, and Richard T. Watson. Alternative scenarios to the ’banner’ years. Communications of the ACM, 51(2):53–58, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marks:2008:SK

[2372] Peter Marks, Peter Polak, Scott Mc-Coy, and Dennis Galletta. Sharing knowledge. Communications of the ACM, 51(2):60–65, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loo:2008:MTW

[2373] Alfred Loo. The myths and truths of wireless security. Communications of the ACM, 51(2):66–71, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mandviwalla:2008:MBW

[2374] Munir Mandviwalla, Abhijit Jain, Julie Fesenmaier, Jeff Smith, Paul Weinberg, and Greg Meyers. Municipal broadband wireless networks. Communications of the ACM, 51(2):72–80, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKinney:2008:WMI

[2375] Vicki R. McKinney, Darryl D. Wilson, Nita Brooks, Anne O’Leary-Kelly, and Bill Hardgrave. Women and men in the IT profession. Communications of the ACM, 51(2):81–84, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liu:2008:MLC

[2376] Bing Liu and Alexander Tuzhilin. Managing large collections of data mining models. Communications of the ACM, 51(2):85–89, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kan:2008:TOR

[2377] Min-Yen Kan and Yee Fan Tan. Technical opinion: Record matching in digital library metadata. Communications of the ACM, 51(2):91–94, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meunier:2008:IRS

[2378] Pascal Meunier. Inside risks: Software transparency and purity. Communications of the ACM, 51(2):104, February 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPc

[2379] Diane Crawford. Editorial pointers. Communications of the ACM, 51(3):5– 6, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:NTb

[2380] CACM Staff. News track. Communications of the ACM, 51(3):7–8, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2008:BSI

[2381] Phillip G. Armour. The business of software: The inaccurate conception. Communications of the ACM, 51(3):13–16, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2008:PPM

[2382] Robert L. Glass. Practical programmer: Managing for innovation. Communications of the ACM, 51(3):17–18, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hu:2008:JSC

[2383] Chenglie Hu. Viewpoint: Just say ’A class defines a data type’. Communications of the ACM, 51(3):19–21, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:AF

[2384] CACM Staff. ACM fellows. Communications of the ACM, 51(3):22–23, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cuff:2008:USW

[2385] Dana Cuff, Mark Hansen, and Jerry Kang. Urban sensing: out of the woods. Communications of the ACM, 51(3):24–33, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Soh:2008:MG

[2386] Jason O. B. Soh and Bernard C. Y.

Tan. Mobile gaming. Communications of the ACM, 51(3):35–39, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhao:2008:PSS

[2387] Liping Zhao. Patterns, symmetry, and symmetry breaking. Communications of the ACM, 51(3):40–46, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pinsker:2008:CBX

[2388] Robert Pinsker and Shaomin Li. Costs and benefits of XBRL adoption: early evidence. Communications of the ACM, 51(3):47–50, March 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stahl:2008:CCR

[2389] Bernd Carsten Stahl and Carole Brooke. The contribution of critical IS research. Communications of the ACM, 51(3):51–55, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2008:IS

[2390] David Wright, Michael Friedewald, Wim Schreurs, Michiel Verlinden, Serge Gutwirth, Yves Punie, Ioannis Maghiros, Elena Vildjiounaite, and Petteri Alahuhta. The illusion of security. Communications of the ACM, 51(3):56–63, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Junglas:2008:LBS

[2391] Iris A. Junglas and Richard T. Watson. Location-based services. Communications of the ACM, 51(3):65–69, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

LaRose:2008:PPR

[2392] Robert LaRose, Nora J. Rifon, and Richard Enbody. Promoting personal responsibility for internet safety. Communications of the ACM, 51(3):71–76, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Safayeni:2008:REN

[2393] Frank Safayeni, P. Robert Duimering, Kimberley Zheng, Natalia Derbentseva, Christopher Poile, and Bing Ran. Requirements engineering in new product development. Communications of the ACM, 51(3):77–82, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cho:2008:PBC

[2394] Kwangsu Cho, Tingting Rachel Chung, William R. King, and Christian Schunn. Peer-based computersupported knowledge refinement: an empirical investigation. Communications of the ACM, 51(3):83–88, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gauri:2008:TOR

[2395] Dinesh K. Gauri, Amit Bhatnagar, and Raghav Rao. Technical opinion: Role of word of mouth in online store loyalty. Communications of the ACM, 51(3):89–91, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lu:2008:WSN

[2396] Xiaoming Lu and George Ledin, Jr.

Inside risks: Wireless sensor networks and the risks of vigilance. Communications of the ACM, 51(3):104, March 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPd

[2397] Diane Crawford. Editorial pointers.

Communications of the ACM, 51(4): 5–6, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:Fc

[2398] Diane Crawford. Forum. Communications of the ACM, 51(4):10, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:NTc

[2399] CACM Staff. News track. Communications of the ACM, 51(4):11–12, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2008:DVF

[2400] Hal Berghel. Digital village: Faithbased security. Communications of the ACM, 51(4):13–17, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2008:PIG

[2401] Peter J. Denning and Peter Yaholkovsky. The profession of IT: Getting to ’we’. Communications of the ACM, 51(4):19–24, April 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:CNA

[2402] CACM Staff. Call for nominations for advanced member grades in the ACM. Communications of the ACM, 51(4):25–26, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:AHD

[2403] CACM Staff. ACM honors distinguished and senior members. Communications of the ACM, 51(4):27–30, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kunkle:2008:SRC

[2404] Daniel Kunkle and Gene Cooperman. Viewpoint: Solving Rubik’s Cube: disk is the new RAM. Communications of the ACM, 51(4):31–33, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

West:2008:PS

[2405] Ryan West. The psychology of security. Communications of the ACM, 51(4):34–40, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Watson:2008:BOS

[2406] Richard T. Watson, Marie-Claude Boudreau, Paul T. York, Martina E. Greiner, and Donald Wynn, Jr. The business of open source. Communications of the ACM, 51(4):41–46, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liao:2008:MCS

[2407] Ziqi Liao and Michael Tow Cheung. Measuring consumer satisfaction in internet banking: a core framework. Communications of the ACM, 51(4):47–51, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moreau:2008:PED

[2408] Luc Moreau, Paul Groth, Simon Miles, Javier Vazquez-Salceda, John Ibbotson, Sheng Jiang, Steve Munroe, Omer Rana, Andreas Schreiber, Victor Tan, and Laszlo Varga. The provenance of electronic data. Communications of the ACM, 51(4):52–58, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yueh:2008:DLM

[2409] Hsiu-Ping Yueh and Shihkuan Hsu. Designing a learning management system to support instruction. Communications of the ACM, 51(4):59–63, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bodin:2008:ISR

[2410] Lawrence D. Bodin, Lawrence A. Gordon, and Martin P. Loeb. Information security and risk management. Communications of the ACM, 51(4):64–68, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gregg:2008:TCA

[2411] Dawn G. Gregg and Judy E. Scott. A typology of complaints about eBay sellers. Communications of the ACM, 51(4):69–74, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oshri:2008:MLB

[2412] Ilan Oshri, Julia Kotlarsky, and Leslie Willcocks. Missing links: building critical social ties for global collaborative teamwork. Communications of the ACM, 51(4):76–81, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chan:2008:KMS

[2413] Ivy Chan and Chee-Kwong Chao. Knowledge management in small and medium-sized enterprises. Communications of the ACM, 51(4):83–88, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khalifa:2008:DCR

[2414] Mohamed Khalifa and Kathy Ning. Demographic changes in IS research productivity and impact. Communications of the ACM, 51(4):89–94, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitten:2008:MWH

[2415] Pamela Whitten, Deirdre Mylod, Goran Gavran, and Howard Sypher. ’most wired hospitals’ rate patient satisfaction. Communications of the ACM, 51(4):96–102, April 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Litecky:2008:TOS

[2416] Chuck Litecky, Bipin Prabhakar, and Kirk Arnett. Technical opinion: The size of the IT job market. Communications of the ACM, 51(4):107–109, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weinstein:2008:CA

[2417] Lauren Weinstein. Inside risks: A current affair. Communications of the ACM, 51(4):112, April 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPe

[2418] Diane Crawford. Editorial pointers.

Communications of the ACM, 51(5):5– 6, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:NTd

[2419] CACM Staff. News track. Communications of the ACM, 51(5):7–8, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:Fd

[2420] Diane Crawford. Forum. Communications of the ACM, 51(5):9–11, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:AMT

[2421] CACM Staff. ACM A. M. Turing Award nominations solicited. Communications of the ACM, 51(5):14, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cohen:2008:VCR

[2422] Jacques Cohen. Viewpoint: The crucial role of CS in systems and synthetic biology. Communications of the ACM, 51(5):15–18, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shannon:2008:SCM

[2423] Meg McGinity Shannon. Staying connected: This menu has changed. Communications of the ACM, 51(5):19–21, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:AE

[2424] CACM Staff. 2008 ACM elections. Communications of the ACM, 51(5): 22–31, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chung:2008:WSM

[2425] Wingyan Chung. Web searching in a multilingual world. Communications of the ACM, 51(5):32–40, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hadar:2008:HIO

[2426] Irit Hadar and Uri Leron. How intuitive is object-oriented design? Communications of the ACM, 51(5):41–46, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dori:2008:WPD

[2427] Dov Dori. Words from pictures for dual-channel processing. Communications of the ACM, 51(5):47–52, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Trimi:2008:ETM

[2428] Silvana Trimi and Hong Sheng. Emerging trends in M-government. Communications of the ACM, 51(5):53–58, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mohan:2008:ICM

[2429] Kannan Mohan, Peng Xu, and Balasubramaniam Ramesh. Improving the change-management process. Communications of the ACM, 51(5):59–64, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chen:2008:CER

[2430] Rui Chen, Raj Sharman, H. Raghav Rao, and Shambhu J. Upadhyaya. Coordination in emergency response management. Communications of the ACM, 51(5):66–73, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garcia:2008:THA

[2431] Alessandro Garcia and Carlos Lucena.

Taming heterogeneous agent architectures. Communications of the ACM, 51(5):75–81, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nov:2008:TFG

[2432] Oded Nov and Bharat Rao. Technologyfacilitated ’give according to your abilities, receive according to your needs’. Communications of the ACM, 51(5): 83–87, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gavish:2008:RIA

[2433] Bezalel Gavish and Christopher L. Tucci. Reducing Internet auction fraud. Communications of the ACM, 51(5):89–97, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Whitley:2008:TOD

[2434] Edgar A. Whitley and Ian R. Hosein. Technical opinion: Departmental influences on policy design. Communications of the ACM, 51(5):98–100, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bellovin:2008:IRP

[2435] Steven M. Bellovin. Inside risks: The physical world and the real world. Communications of the ACM, 51(5): 104, May 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EPf

[2436] Diane Crawford. Editorial pointers. Communications of the ACM, 51(6):5– 6, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:NTe

[2437] CACM Staff. News track. Communications of the ACM, 51(6):7–8, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:Fe

[2438] Diane Crawford. Forum. Communications of the ACM, 51(6):9–11, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:2008:DVB

[2439] Hal Berghel. Digital village: BRAP forensics. Communications of the ACM, 51(6):15–20, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Glass:2008:PPS

[2440] Robert L. Glass. Practical programmer: Software design and the monkey’s brain. Communications of the ACM, 51(6):21–22, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ragowsky:2008:VGM

[2441] Arik Ragowsky, Paul S. Licker, and David Gefen. Viewpoint: Give me information, not technology. Communications of the ACM, 51(6):23–25, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vertegaal:2008:OUI

[2442] Roel Vertegaal and Ivan Poupyrev.

Organic user interfaces: Introduction. Communications of the ACM, 51(6):26–30, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ishii:2008:TUI

[2443] Hiroshi Ishii. The tangible user interface and its evolution. Communications of the ACM, 51(6):32–36, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rekimoto:2008:OIT

[2444] Jun Rekimoto. Organic interaction technologies: from stone to skin. Communications of the ACM, 51(6):38–44, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Co:2008:EDT

[2445] Elise Co and Nikita Pashenkov. Emerging display technologies for organic user interfaces. Communications of the ACM, 51(6):45–47, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holman:2008:OUI

[2446] David Holman and Roel Vertegaal. Organic user interfaces: designing computers in any way, shape, or form. Communications of the ACM, 51(6): 48–55, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blevis:2008:SIO

[2447] Eli Blevis. Sustainability implications of organic user interface technologies: an inky problem. Communications of the ACM, 51(6):56–57, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parkes:2008:DKI

[2448] Amanda Parkes, Ivan Poupyrev, and Hiroshi Ishii. Designing kinetic interactions for organic user interfaces. Communications of the ACM, 51(6):58–65, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwesig:2008:WMI

[2449] Carsten Schwesig. What makes an interface feel organic? Communications of the ACM, 51(6):67–69, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oosterhuis:2008:IPA

[2450] Kas Oosterhuis and Nimish Biloria. Interactions with proactive architectural spaces: the muscle projects. Communications of the ACM, 51(6):70–78, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kodama:2008:DFS

[2451] Sachiko Kodama. Dynamic ferrofluid sculpture: organic shape-changing art forms. Communications of the ACM, 51(6):79–81, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weitzner:2008:IA

[2452] Daniel J. Weitzner, Harold Abelson, Tim Berners-Lee, Joan Feigenbaum, James Hendler, and Gerald Jay Sussman. Information accountability. Communications of the ACM, 51(6):82–87, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Iacovou:2008:RPO

[2453] Charalambos L. Iacovou and Robbie Nakatsu. A risk profile of offshoreoutsourced development projects. Communications of the ACM, 51(6): 89–94, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Abraham:2008:UAF

[2454] Chon Abraham, Richard T. Watson, and Marie-Claude Boudreau. Ubiquitous access: on the front lines of patient care and safety. Communications of the ACM, 51(6):95–99, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:2008:ODP

[2455] Wayne Wei Huang, Jonathon Greene, and John Day. Outsourcing and the decrease of IS program enrollment. Communications of the ACM, 51(6):101–104, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horning:2008:IRR

[2456] Jim Horning and Peter G. Neumann. Inside risks: Risks of neglecting infrastructure. Communications of the ACM, 51(6):112, June 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2008:ELW

[2457] Moshe Y. Vardi. Editor’s letter: ’where do you come from? and where are you going?’. Communications of the ACM, 51(7):5, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2008:PCA

[2458] Scott E. Delman. Publisher’s corner: The art and business of revitalizing a 50-year-old science and technology magazine. Communications of the ACM, 51(7):7, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:COY

[2459] David Roman. CACM online: Your attention, please. Communications of the ACM, 51(7):8, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hayes:2008:CC

[2460] Brian Hayes. Cloud computing. Communications of the ACM, 51(7):9–11, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:2008:QC

[2461] Michael Ross and Mark Oskin. Quantum computing. Communications of the ACM, 51(7):12–13, July 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:2008:SDD

[2462] Leah Hoffman. In search of dependable design. Communications of the ACM, 51(7):14–16, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aggarwal:2008:EMI

[2463] Alok Aggarwal. Emerging markets:

India’s role in the globalization of IT. Communications of the ACM, 51(7):17–19, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2008:LSR

[2464] Pamela Samuelson. Legally speaking: Revisiting patentable subject matter. Communications of the ACM, 51(7):20–22, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2008:KVB

[2465] George V. Neville-Neil. Kode vicious: Beautiful code exists, if you know where to look. Communications of the ACM, 51(7):23–25, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andriole:2008:PCT

[2466] Stephen J. Andriole and Eric Roberts. Point/counterpoint: Technology curriculum for the early 21st Century. Communications of the ACM, 51(7): 27–32, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rashid:2008:ICI

[2467] Rick Rashid. Image crisis: Inspiring a new generation of computer scientists. Communications of the ACM, 51(7):33–34, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shustek:2008:IAB

[2468] Len Shustek. Interview: The ’art’ of being Donald Knuth. Communications of the ACM, 51(7):35–39, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilde:2008:XF

[2469] Erik Wilde and Robert J. Glushko.

XML fever. Communications of the ACM, 51(7):40–46, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leventhal:2008:FSM

[2470] Adam Leventhal. Flash storage memory. Communications of the ACM, 51(7):47–51, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Seltzer:2008:BRD

[2471] Margo Seltzer. Beyond relational databases. Communications of the ACM, 51(7):52–58, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hendler:2008:WSI

[2472] James Hendler, Nigel Shadbolt, Wendy Hall, Tim Berners-Lee, and Daniel Weitzner. Web science: an interdisciplinary approach to understanding the Web. Communications of the ACM, 51(7):60–69, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oskin:2008:RIB

[2473] Mark Oskin. The revolution inside the box. Communications of the ACM, 51(7):70–78, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Larus:2008:TM

[2474] James Larus and Christos Kozyrakis.

Transactional memory. Communications of the ACM, 51(7):80–88, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Colwell:2008:TPC

[2475] Bob Colwell. Technical perspective: Computer science takes on molecular dynamics. Communications of the ACM, 51(7):90, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shaw:2008:ASP

[2476] David E. Shaw, Martin M. Deneroff, Ron O. Dror, Jeffrey S. Kuskin, Richard H. Larson, John K. Salmon, Cliff Young, Brannon Batson, Kevin J. Bowers, Jack C. Chao, Michael P. Eastwood, Joseph Gagliardo, J. P. Grossman, C. Richard Ho, Douglas J. Lerardi, István Kolossváry, John L. Klepeis, Timothy Layman, Christine McLeavey, Mark A. Moraes, Rolf Mueller, Edward C. Priest, Yibing Shan, Jochen Spengler, Michael Theobald, Brian Towles, and Stanley C. Wang. Anton, a specialpurpose machine for molecular dynamics simulation. Communications of the ACM, 51(7):91–97, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhao:2008:TPP

[2477] Feng Zhao. Technical perspective: The physical side of computing. Communications of the ACM, 51(7):98, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levis:2008:ENP

[2478] Philip Levis, Eric Brewer, David Culler, David Gay, Samuel Madden, Neil Patel, Joe Polastre, Scott Shenker, Robert Szewczyk, and Alec Woo. The emergence of a networking primitive in wireless sensor networks. Communications of the ACM, 51(7):99–106, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffman:2008:LBQ

[2479] Leah Hoffman. Last byte: Q&A talking model-checking technology. Communications of the ACM, 51(7):112, 1??– 1??, July 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2008:PLN

[2480] Stuart I. Feldman. President’s letter: A new beginning, a fond farewell. Communications of the ACM, 51(8):5, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:WBK

[2481] CACM Staff. Words both kind and contrary. Communications of the ACM, 51(8):7, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:SCH

[2482] David Roman. Small changes hint at bigger things. Communications of the ACM, 51(8):8, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varian:2008:DPA

[2483] Hal R. Varian. Designing the perfect auction. Communications of the ACM, 51(8):9–11, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aycinena:2008:AA

[2484] Peggy Aycinena. Access for all. Communications of the ACM, 51(8):12–14, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Underwood:2008:CP

[2485] Sarah Underwood. Challenging poverty. Communications of the ACM, 51(8):15–17, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:2008:RJ

[2486] Michael Ross. Remembering Jim. Communications of the ACM, 51(8):18, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2008:PIV

[2487] Peter J. Denning. The profession of IT: Voices of computing. Communications of the ACM, 51(8):19–21, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2008:FLS

[2488] Alex E. Bell. From the front lines: Software development amidst the whiz of silver bullets. Communications of the ACM, 51(8):22–24, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guzdial:2008:EPW

[2489] Mark Guzdial. Education: Paving the way for computational thinking. Communications of the ACM, 51(8):25–27, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lazowska:2008:VEF

[2490] Ed Lazowska. Viewpoint: Envisioning the future of computing research. Communications of the ACM, 51(8):28–30, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shustek:2008:IDK

[2491] Len Shustek. Interview: Donald Knuth: A life’s work interrupted. Communications of the ACM, 51(8):31–35, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waldo:2008:SGV

[2492] Jim Waldo. Scaling in games and virtual worlds. Communications of the ACM, 51(8):38–44, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2008:CSR

[2493] Mache Creeger. CTO storage roundtable. Communications of the ACM, 51(8):45–51, August 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henning:2008:RFC

[2494] Michi Henning. The rise and fall of CORBA. Communications of the ACM, 51(8):52–57, August 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vonAhn:2008:DGP

[2495] Luis von Ahn and Laura Dabbish. Designing games with a purpose. Communications of the ACM, 51(8):58–67, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spinellis:2008:COK

[2496] Diomidis Spinellis and Panagiotis Louridas. The collaborative organization of knowledge. Communications of the ACM, 51(8):68–73, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shoham:2008:CSG

[2497] Yoav Shoham. Computer science and game theory. Communications of the ACM, 51(8):74–79, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pugh:2008:TPM

[2498] William Pugh. Technical perspective: A methodology for evaluating computer system performance. Communications of the ACM, 51(8):82, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Blackburn:2008:WSC

[2499] Stephen M. Blackburn, Kathryn S. McKinley, Robin Garner, Chris Hoffmann, Asjad M. Khan, Rotem Bentzur, Amer Diwan, Daniel Feinberg, Daniel Frampton, Samuel Z. Guyer, Martin Hirzel, Antony Hosking, Maria Jump, Han Lee, J. Eliot B. Moss, Aashish Phansalkar, Darko Stefanovik, Thomas VanDrunen, Daniel von Dincklage, and Ben Wiedermann. Wake up and smell the coffee: evaluation methodology for the 21st Century. Communications of the ACM, 51(8):83–89, August 2008.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shavit:2008:TPT

[2500] Nir Shavit. Technical perspective:

Transactions are tomorrow’s loads and stores. Communications of the ACM, 51(8):90, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harris:2008:CMT

[2501] Tim Harris, Simon Marlow, Simon Peyton Jones, and Maurice Herlihy. Composable memory transactions. Communications of the ACM, 51(8):91–100, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2008:LBPa

[2502] Peter Winkler. Last byte: Puzzled delightful graph theory. Communications of the ACM, 51(8):104, August 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:2008:PLA

[2503] Wendy Hall. President’s letter: ACM’s place in the global picture. Communications of the ACM, 51(9):5, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2008:PCY

[2504] Scott E. Delman. Publisher’s corner:

50 years young. Communications of the ACM, 51(9):7, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:KAR

[2505] CACM Staff. Letters to the editor: Knuth’s art of recovering from errors. Communications of the ACM, 51(9):9, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:COCa

[2506] David Roman. CACM online: Current features. Communications of the ACM, 51(9):10, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2008:NFD

[2507] Kirk L. Kroeker. News: Finding diamonds in the rough. Communications of the ACM, 51(9):11–13, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kugler:2008:NUV

[2508] Logan Kugler. News: Ubiquitous video. Communications of the ACM, 51(9):14–16, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2008:NPM

[2509] Samuel Greengard. News: Privacy matters. Communications of the ACM, 51(9):17–18, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2008:NWR

[2510] Leah Hoffmann. News: Wisdom from Randy Pausch. Communications of the ACM, 51(9):19, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2008:TSMb

[2511] Michael Cusumano. Technology strategy and management: The puzzle of Apple. Communications of the ACM, 51(9):22–24, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2008:KVPa

[2512] George Neville-Neil. Kode vicious:

Pride and prejudice (the vasa). Communications of the ACM, 51(9):25–26, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:2008:IPS

[2513] Cameron Wilson and Peter Harsha.

IT policy: Science policy isn’t always about science. Communications of the ACM, 51(9):27–29, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baraniuk:2008:VGW

[2514] Richard G. Baraniuk and C. Sidney Burrus. Viewpoint: Global warming toward open educational resources. Communications of the ACM, 51(9): 30–32, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Foster:2008:PHD

[2515] Ian Foster, Savas Parastatidis, Paul Watson, and Mark Mckeown. Practice: How do I model state?: Let me count the ways. Communications of the ACM, 51(9):34–41, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garrett:2008:PP

[2516] Matthew Garrett. Practice: Powering down. Communications of the ACM, 51(9):42–46, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2008:PCS

[2517] Mache Creeger. Practice: CTO storage roundtable, part two. Communications of the ACM, 51(9):47–52, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hinchey:2008:PSE

[2518] Mike Hinchey, Michael Jackson, Patrick Cousot, Byron Cook, Jonathan P. Bowen, and Tiziana Margaria. Practice: Software engineering and formal methods. Communications of the ACM, 51(9):54–59, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roussinov:2008:BKA

[2519] Dmitri Roussinov, Weiguo Fan, and José Robles-Flores. Beyond keywords: Automated question answering on the Web. Communications of the ACM, 51 (9):60–65, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:2008:DCR

[2520] Bertrand Meyer. Design and code reviews in the age of the Internet. Communications of the ACM, 51(9): 66–71, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bernstein:2008:IIE

[2521] Philip A. Bernstein and Laura M.

Haas. Information integration in the enterprise. Communications of the ACM, 51(9):72–79, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moir:2008:TPT

[2522] Mark Moir. Technical perspective:

Transactional memory in the operating system. Communications of the ACM, 51(9):82, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rossbach:2008:TMT

[2523] Christopher J. Rossbach, Hany E. Ramadan, Owen S. Hofmann, Donald E. Porter, Aditya Bhandari, and Emmett Witchel. TxLinux and MetaTM: transactional memory and the operating system. Communications of the ACM, 51 (9):83–91, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Attiya:2008:TPD

[2524] Hagit Attiya. Technical perspective:

Distributing your data and having it, too. Communications of the ACM, 51(9):92, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuhn:2008:DSM

[2525] Fabian Kuhn, Thomas Locher, and Roger Wattenhofer. Distributed selection: a missing piece of data aggregation. Communications of the ACM, 51 (9):93–99, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2008:LBPb

[2526] Peter Winkler. Last byte: Puzzled: Solutions and sources. Communications of the ACM, 51(9):103, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bainbridge:2008:LBF

[2527] William Sims Bainbridge. Last byte: Future tense: Will. Communications of the ACM, 51(9):104, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Poston:2008:UFB

[2528] Robin S. Poston. Using and fixing biased rating schemes. Communications of the ACM, 51(9):105–109, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mohan:2008:UTM

[2529] Kannan Mohan and Radhika Jain. Using traceability to mitigate cognitive biases in software development. Communications of the ACM, 51(9): 110–114, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pons:2008:UUP

[2530] Alexander P. Pons and Peter Polak. Understanding user perspectives on biometric technology. Communications of the ACM, 51(9):115–118, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhou:2008:FLF

[2531] Lina Zhou and Dongsong Zhang. Following linguistic footprints: automatic deception detection in online communication. Communications of the ACM, 51(9):119–122, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kotlarsky:2008:TAD

[2532] Julia Kotlarsky, Ilan Oshri, Kuldeep Kumar, and Jos van Hillegersberg. Towards agility in design in global component-based development. Communications of the ACM, 51(9):123– 127, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hikmet:2008:SPP

[2533] Neset Hikmet, Eileen Z. Taylor, and Christopher J. Davis. The student productivity paradox: technology mediated learning in schools. Communications of the ACM, 51(9):128–131, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chau:2008:WFD

[2534] Patrick Y. K. Chau, Fujun Lai, and Dahui Li. What factors drive the assimilation of Internet technologies in China? Communications of the ACM, 51(9):132–135, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2008:ELL

[2535] Moshe Y. Vardi. Editor’s letter:

Let us—together—make CACM exciting. Communications of the ACM, 51(10):5, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2008:PCS

[2536] Scott E. Delman. Publisher’s corner: The softer side of communications. Communications of the ACM, 51(10):7, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:LEP

[2537] CACM Staff. Letters to the editor: Prep students for irreversible software trends. Communications of the ACM, 51(10):8–9, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:COCb

[2538] David Roman. CACM online: Communications site to launch in January. Communications of the ACM, 51(10):10, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kurp:2008:NGC

[2539] Patrick Kurp. News: Green computing. Communications of the ACM, 51 (10):11–13, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2008:NSD

[2540] Alex Wright. News: Searching the Deep Web. Communications of the ACM, 51(10):14–15, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farivar:2008:NCE

[2541] Cyrus Farivar. News: Clean elections. Communications of the ACM, 51(10):16–18, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2008:NIL

[2542] Leah Hoffmann. News: An inspiring legacy. Communications of the ACM, 51(10):19, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell-Kelly:2008:VHR

[2543] Martin Campbell-Kelly. Viewpoints:

Historical reflections: Will the future of software be open source? Communications of the ACM, 51(10):21–23, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2008:VCE

[2544] Deborah G. Johnson. Viewpoints:

Computing ethics: Computer experts: guns-for-hire or professionals? Communications of the ACM, 51(10):24– 26, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2008:VFL

[2545] Alex E. Bell. Viewpoints: From the front lines DOA with SOA. Communications of the ACM, 51(10):27–28, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dill:2008:VPC

[2546] David L. Dill and Daniel Castro. Viewpoints: Point/counterpoint: The U.S. should ban paperless electronic voting machines. Communications of the ACM, 51(10):29–33, October 2008.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2008:PCS

[2547] George V. Neville-Neil. Practice: Code spelunking redux. Communications of the ACM, 51(10):36–41, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilde:2008:PDD

[2548] Erik Wilde and Robert J. Glushko. Practice: Document design matters. Communications of the ACM, 51 (10):43–49, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fatahalian:2008:PCL

[2549] Kayvon Fatahalian and Mike Houston. Practice: A closer look at GPUs. Communications of the ACM, 51(10):50–57, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Xu:2008:TDN

[2550] Jennifer Xu and Hsinchun Chen. The topology of dark networks. Communications of the ACM, 51(10):58–65, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Slonim:2008:CCC

[2551] Jacob Slonim, Sam Scully, and Michael McAllister. Crossroads for Canadian CS enrollment. Communications of the ACM, 51(10):66–70, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kari:2008:MFN

[2552] Lila Kari and Grzegorz Rozenberg.

The many facets of natural computing. Communications of the ACM, 51 (10):72–83, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Levoy:2008:TPC

[2553] Marc Levoy. Technical perspective:

Computational photography on large collections of images. Communications of the ACM, 51(10):86, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hays:2008:SCU

[2554] James Hays and Alexei A. Efros.

Scene completion using millions of photographs. Communications of the ACM, 51(10):87–94, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tardos:2008:TPN

[2555] Éva Tardos. Technical perspective:

New developments in graph partitioning. Communications of the ACM, 51(10):95, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arora:2008:GFG

[2556] Sanjeev Arora, Satish Rao, and Umesh Vazirani. Geometry, flows, and graphpartitioning algorithms. Communications of the ACM, 51(10):96–105, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2008:LBQ

[2557] Leah Hoffmann. Last byte: Q&A: A complex thinker. Communications of the ACM, 51(10):112–ff, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:2008:LSP

[2558] Mary C. Jones. Large scale project team building: beyond the basics. Communications of the ACM, 51(10): 113–116, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adomavicius:2008:UET

[2559] Gediminas Adomavicius, Jesse Bockstedt, Alok Gupta, and Robert J. Kauffman. Understanding evolution in technology ecosystems. Communications of the ACM, 51(10):117–122, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Song:2008:UIN

[2560] Seokwoo Song, Sridhar Nerur, and James T. C. Teng. Understanding the influence of network positions and knowledge processing styles. Communications of the ACM, 51(10):123–126, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Santos:2008:RSC

[2561] Brian L. Dos Santos and Lars S. Smith. RFID in the supply chain: panacea or Pandora’s box? Communications of the ACM, 51(10):127–131, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ye:2008:SBC

[2562] Chen Ye, Kevin C. Desouza, Sridhar R. Papagari Sangareddy, and Sanjeev Jha. Switching between consumer technologies. Communications of the ACM, 51(10):132–136, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Darking:2008:GDD

[2563] Mary Darking, Edgar A. Whitley, and Paolo Dini. Governing diversity in the digital ecosystem. Communications of the ACM, 51(10):137–140, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tiwana:2008:MPJ

[2564] Amrit Tiwana, Ashley A. Bush, Hiroshi Tsuji, Kenichi Yoshida, and Akito Sakurai. Myths and paradoxes in Japanese IT offshoring. Communications of the ACM, 51(10):141–145, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ariyachandra:2008:TOW

[2565] Thilini Ariyachandra and Hugh J.

Watson. Technical opinion: Which data warehouse architecture is best? Communications of the ACM, 51(10): 146–147, October 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:2008:CLA

[2566] John R. White. CEO’s letter: On the 10th anniversary of ACM’s Digital Library. Communications of the ACM, 51(11):5, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crawford:2008:EEC

[2567] Diane Crawford. Executive editor’s corner: Jim Gray: humble visionary. Communications of the ACM, 51(11):7, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:LEE

[2568] CACM Staff. Letters to the editor: Even science would benefit from auctions. Communications of the ACM, 51(11):8–9, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:COF

[2569] David Roman. CACM online: A first look at the redesigned site. Communications of the ACM, 51(11):10, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2008:NDC

[2570] Leah Hoffmann. News: Damage control. Communications of the ACM, 51 (11):11–13, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Howard:2008:NAO

[2571] Bill Howard. News: Analyzing online social networks. Communications of the ACM, 51(11):14–16, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindley:2008:NLC

[2572] David Lindley. News: The limits of computability. Communications of the ACM, 51(11):17–19, November 2008.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldfarb:2008:VEB

[2573] Avi Goldfarb and Catherine Tucker.

Viewpoints: Economic and business dimensions: Search engine advertising. Communications of the ACM, 51 (11):22–24, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landau:2008:VPS

[2574] Susan Landau. Viewpoints: Privacy and security: A multidimensional problem. Communications of the ACM, 51 (11):25–26, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2008:VLS

[2575] Pamela Samuelson. Viewpoints:

Legally speaking: Quantafying the value of patent exhaustion. Communications of the ACM, 51(11):27–29, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goode:2008:VER

[2576] Joanna Goode. Viewpoints: Education: Reprogramming: college preparatory computer science. Communications of the ACM, 51(11):31–33, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cantrill:2008:PRW

[2577] Bryan Cantrill and Jeff Bonwick. Practice: Real-world concurrency. Communications of the ACM, 51(11):34–39, November 2008. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Cascaval:2008:PST

[2578] Calin Cascaval, Colin Blundell, Maged Michael, Harold W. Cain, Peng Wu, Stefanie Chiras, and Siddhartha Chatterjee. Practice: Software transactional memory: why is it only a research toy? Communications of the ACM, 51 (11):40–46, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2008:PCR

[2579] Mache Creeger. Practice: CTO roundtable on virtualization. Communications of the ACM, 51(11):47–53, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stonebraker:2008:TJG

[2580] Michael Stonebraker and David J. De-Witt. A tribute to Jim Gray. Communications of the ACM, 51(11):54–57, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Szalay:2008:JGA

[2581] Alexander S. Szalay. Jim Gray, astronomer. Communications of the ACM, 51(11):58–65, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kleinberg:2008:CST

[2582] Jon Kleinberg. The convergence of social and technological networks. Communications of the ACM, 51(11):66–72, November 2008. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:2008:TPP

[2583] Jim Gray. Technical perspective: The Polaris Tableau system. Communications of the ACM, 51(11):74, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stolte:2008:PSQ

[2584] Chris Stolte, Diane Tang, and Pat Hanrahan. Polaris: a system for query, analysis, and visualization of multidimensional databases. Communications of the ACM, 51(11):75–84, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liskov:2008:TPS

[2585] Barbara Liskov. Technical perspective: Safeguarding online information against failures and attacks. Communications of the ACM, 51(11):85, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kotla:2008:ZSB

[2586] Ramakrishna Kotla, Allen Clement, Edmund Wong, Lorenzo Alvisi, and Mike Dahlin. Zyzzyva: speculative Byzantine fault tolerance. Communications of the ACM, 51(11):86–95, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2008:LBPc

[2587] Peter Winkler. Last byte: Puzzled:

Circular food. Communications of the ACM, 51(11):112, November 2008.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2008:SSA

[2588] DaeSoo Kim, Terence T. Ow, and Minjoon Junc. SME strategies: an assessment of high vs. low performers. Communications of the ACM, 51(11): 113–117, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feijoo:2008:EPV

[2589] Claudio Feijóo, José Luis Gómez-Barroso, and David Rojo-Alonso. A European perspective of VoIP in market competition. Communications of the ACM, 51(11):118–120, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scott:2008:TAE

[2590] Judy E. Scott. Technology acceptance and ERP documentation usability. Communications of the ACM, 51 (11):121–124, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Currim:2008:CMW

[2591] Faiz Currim and Sudha Ram. Conceptually modeling windows and bounds for space and time in database constraints. Communications of the ACM, 51(11):125–129, November 2008. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sheng:2008:UCS

[2592] Quan Z. Sheng, Boualem Benatallah, and Zakaria Maamar. User-centric services provisioning in wireless environments. Communications of the ACM, 51(11):130–135, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Seo:2008:EDS

[2593] DongBack Seo and Ariel I. La Paz. Exploring the dark side of IS in achieving organizational agility. Communications of the ACM, 51(11):136–139, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Venkatraman:2008:SSE

[2594] Srinivasan Venkatraman, Hillol Bala, Viswanath Venkatesh, and Jack Bates. Six strategies for electronic medical records systems. Communications of the ACM, 51(11):140–144, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhang:2008:TOM

[2595] Ping Zhang. Technical opinion: Motivational affordances: reasons for ICT design and use. Communications of the ACM, 51(11):145–147, November 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boisvert:2008:APA

[2596] Ronald F. Boisvert and Holly Rushmeier. ACM publications: access and sustainability. Communications of the ACM, 51(12):5, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2008:LBF

[2597] Scott E. Delman. Looking back and forward. Communications of the ACM, 51(12):7, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2008:ABG

[2598] CACM Staff. Apple builds great platforms, too, not just products. Communications of the ACM, 51(12):8, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2008:PNC

[2599] David Roman. Previewing the new CACM web site. Communications of the ACM, 51(12):10, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2008:LM

[2600] Kirk L. Kroeker. Living machines. Communications of the ACM, 51(12): 11–13, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Selker:2008:TF

[2601] Ted Selker. Touching the future. Communications of the ACM, 51(12): 14–16, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2008:UM

[2602] Samuel Greengard. Upwardly mobile. Communications of the ACM, 51 (12):17–19, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2008:MD

[2603] Leah Hoffmann. Making a difference. Communications of the ACM, 51(12):20, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Aggarwal:2008:EML

[2604] Alok Aggarwal. Emerging markets:

Labor supply in the Indian IT industry. Communications of the ACM, 51 (12):21–23, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:2008:IPA

[2605] Cameron Wilson and Peter Harsha. IT policy: Advising policymakers is more than just providing advice. Communications of the ACM, 51(12):24–26, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2008:KVPb

[2606] George V. Neville-Neil. Kode vicious:

Permanence and change. Communications of the ACM, 51(12):27–28, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2008:PIE

[2607] Peter J. Denning, Chris Gunderson, and Rick Hayes-Roth. The profession of IT: Evolutionary system development. Communications of the ACM, 51 (12):29–31, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hamilton:2008:IDD

[2608] James Hamilton. Interview: Database dialogue with Pat Selinger. Communications of the ACM, 51(12):32–35, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Souders:2008:HPW

[2609] Steve Souders. High-performance web sites. Communications of the ACM, 51 (12):36–41, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2008:CVR

[2610] Mache Creeger. CTO virtualization roundtable, part II. Communications of the ACM, 51(12):43–49, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berman:2008:GDG

[2611] Francine Berman. Got data?: a guide to data preservation in the information age. Communications of the ACM, 51 (12):50–56, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Horrocks:2008:OSW

[2612] Ian Horrocks. Ontologies and the Semantic Web. Communications of the ACM, 51(12):58–67, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Etzioni:2008:OIE

[2613] Oren Etzioni, Michele Banko, Stephen Soderland, and Daniel S. Weld. Open information extraction from the web. Communications of the ACM, 51(12): 68–74, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stonebraker:2008:TPO

[2614] Michael Stonebraker. Technical perspective: One size fits all: an idea whose time has come and gone. Communications of the ACM, 51(12):76, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boncz:2008:BMW

[2615] Peter A. Boncz, Martin L. Kersten, and Stefan Manegold. Breaking the memory wall in MonetDB. Communications of the ACM, 51(12):77–85, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rinard:2008:TPP

[2616] Martin C. Rinard. Technical perspective: Patching program errors. Communications of the ACM, 51(12):86, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Novark:2008:EAC

[2617] Gene Novark, Emery D. Berger, and Benjamin G. Zorn. Exterminator: Automatically correcting memory errors with high probability. Communications of the ACM, 51(12):87–95, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2008:LBPd

[2618] Peter Winkler. Last byte: Puzzled: Solutions and sources. Communications of the ACM, 51(12):118, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Doctorow:2008:FTP

[2619] Cory Doctorow. Future tense: Pester power. Communications of the ACM, 51(12):120–ff, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Duchessi:2008:EBP

[2620] Peter Duchessi and Indushobha Chengalur-Smith. Enhancing business performance via vendor managed inventory applications. Communications of the ACM, 51(12):121–127, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phang:2008:FIE

[2621] Chee Wei Phang and Atreyi Kankanhalli. A framework of ICT exploitation for e-participation initiatives. Communications of the ACM, 51(12):128–132, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davern:2008:EIT

[2622] Michael J. Davern and Carla L. Wilkin. Evolving innovations through design and use. Communications of the ACM, 51(12):133–137, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Naf:2008:TAI

[2623] Michael Näf and David Basin. Two approaches to an information security laboratory. Communications of the ACM, 51(12):138–142, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stacey:2008:OAG

[2624] Patrick Stacey and Joe Nandhakumar. Opening up to agile games development. Communications of the ACM, 51 (12):143–146, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malaga:2008:WPS

[2625] Ross A. Malaga. Worst practices in search engine optimization. Communications of the ACM, 51(12):147–150, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berthon:2008:GRT

[2626] Pierre Berthon, Leyland F. Pitt, and Richard T. Watson. From genesis to revelations: the technology diaspora. Communications of the ACM, 51(12): 151–154, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lim:2008:TOB

[2627] Vivien K. G. Lim, Thompson S. H.

Teo, and Jen Yuin Chin. Technical opinion: Bosses and their e-manners. Communications of the ACM, 51(12): 155–157, December 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:ELH

[2628] Moshe Y. Vardi. Editor’s letter: How are we doing? Communications of the ACM, 52(1):5, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LET

[2629] CACM Staff. Letters to the editor:

True seeds of open source software. Communications of the ACM, 52(1): 6, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COCa

[2630] David Roman. CACM online: Communications preps Web-enhanced articles. Communications of the ACM, 52(1): 8, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindley:2009:NCF

[2631] David Lindley. News: Calculating the future. Communications of the ACM, 52(1):9–11, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kanipe:2009:NUY

[2632] Jeff Kanipe. News: The universe in your computer. Communications of the ACM, 52(1):12–14, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2009:NGS

[2633] Alex Wright. News: Get smart. Communications of the ACM, 52(1):15–16, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ross:2009:NPW

[2634] Michael Ross. News: A pioneer woman. Communications of the ACM, 52(1):17, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zittrain:2009:VLT

[2635] Jonathan Zittrain. Viewpoints: Law and technology: The end of the generative Internet. Communications of the ACM, 52(1):18–20, January 2009.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bell:2009:VFL

[2636] Alex E. Bell. Viewpoints: From the front lines: Ground control to Architect Tom. … Communications of the ACM, 52(1):21–22, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2009:VBS

[2637] Phillip G. Armour. Viewpoints: The business of software: The ontology of paper. Communications of the ACM, 52(1):23–24, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2009:VTS

[2638] Michael Cusumano. Viewpoints: Technology strategy and management: The legacy of Bill Gates. Communications of the ACM, 52(1):25–26, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Crowcroft:2009:VSA

[2639] Jon Crowcroft, S. Keshav, and Nick McKeown. Viewpoint: Scaling the academic publication process to Internet scale. Communications of the ACM, 52(1):27–30, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feldman:2009:AAR

[2640] Stuart I. Feldman. ACM’s annual report. Communications of the ACM, 52(1):33–37, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vogels:2009:EC

[2641] Werner Vogels. Eventually consistent. Communications of the ACM, 52(1):40–44, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mashey:2009:LRB

[2642] John Mashey. The long road to 64 bits. Communications of the ACM, 52(1):45–53, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parikh:2009:ERD

[2643] Tapan S. Parikh. Engineering rural development. Communications of the ACM, 52(1):54–63, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jutla:2009:WPS

[2644] Dawn N. Jutla and Dimitri Kanevsky. wisepad services for vision-, hearing-, and speech-impaired users. Communications of the ACM, 52(1):64–69, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Feigenbaum:2009:CCC

[2645] Joan Feigenbaum, David C. Parkes, and David M. Pennock. Computational challenges in e-commerce. Communications of the ACM, 52(1):70–74, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shum:2009:TPC

[2646] Harry Shum. Technical perspective: Customizing media to displays. Communications of the ACM, 52(1):76, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shamir:2009:SCM

[2647] Ariel Shamir and Shai Avidan. Seam carving for media retargeting. Communications of the ACM, 52(1):77–85, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mackinlay:2009:TPF

[2648] Jock D. Mackinlay. Technical perspective: Finding and telling stories with data. Communications of the ACM, 52(1):86, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Heer:2009:VVS

[2649] Jeffrey Heer, Fernanda B. Viégas, and Martin Wattenberg. Voyagers and voyeurs: Supporting asynchronous collaborative visualization. Communications of the ACM, 52(1):87–97, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:LBQa

[2650] Leah Hoffmann. Last byte: Q&A:

The upper limit. Communications of the ACM, 52(1):112ff, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mueller:2009:ACR

[2651] Erik T. Mueller. Automating commonsense reasoning using the event calculus. Communications of the ACM, 52 (1):113–117, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kumar:2009:ACP

[2652] Nanda Kumar and Roumen Vragov. Active citizen participation using ICT tools. Communications of the ACM, 52 (1):118–121, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dwivedi:2009:UAN

[2653] Yogesh Dwivedi and Zahir Irani. Understanding the adopters and nonadopters of broadband. Communications of the ACM, 52(1):122–125, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnston:2009:IST

[2654] Allen C. Johnston and Ron Hale. Improved security through information security governance. Communications of the ACM, 52(1):126–129, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gebauer:2009:EBB

[2655] Judith Gebauer and Mark Ginsburg. Exploring the black box of tasktechnology fit. Communications of the ACM, 52(1):130–135, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bala:2009:DRH

[2656] Hillol Bala, Viswanath Venkatesh, Srinivasan Venkatraman, Jack Bates, and Steven H. Brown. Disaster response in health care: A design extension for enterprise data warehouse. Communications of the ACM, 52(1): 136–140, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pendharkar:2009:RBS

[2657] Parag C. Pendharkar and James A.

Rodger. The relationship between software development team size and software development cost. Communications of the ACM, 52(1):141–144, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ou:2009:TOW

[2658] Carol Xiaojuan Ou and Robert M.

Davison. Technical opinion: Why eBay lost to TaoBao in China: the global advantage. Communications of the ACM, 52(1):145–148, January 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spafford:2009:UPR

[2659] Eugene H. Spafford. USACM’s policy role. Communications of the ACM, 52(2):5, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:SPS

[2660] CACM Staff. Seven principles for secure e-voting. Communications of the ACM, 52(2):8–9, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:DOD

[2661] David Roman. The dot-org difference. Communications of the ACM, 52(2):10, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:PBF

[2662] Kirk L. Kroeker. Photography’s bright future. Communications of the ACM, 52(2):11–13, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2009:MSS

[2663] Alex Wright. Making sense of sensors. Communications of the ACM, 52(2):14–15, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2009:FIP

[2664] Samuel Greengard. The first Internet President. Communications of the ACM, 52(2):16–18, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:SDA

[2665] Kirk L. Kroeker. SIGGRAPH debuts in Asia. Communications of the ACM, 52(2):19, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Arora:2009:EBD

[2666] Ashish Arora, Matej Drev, and Chris Forman. Economic and business dimensions: The extent of globalization of software innovation. Communications of the ACM, 52(2):20–22, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fletcher:2009:EHC

[2667] George H. L. Fletcher and James J. Lu. Education human: computing skills: rethinking the K-12 experience. Communications of the ACM, 52(2):23–25, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Irion:2009:PSI

[2668] Kristina Irion. Privacy and security:

International communications surveillance. Communications of the ACM, 52(2):26–28, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neumann:2009:IRU

[2669] Peter G. Neumann. Inside risks: U.S.

election after-math. Communications of the ACM, 52(2):29, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

vanSchewick:2009:PCN

[2670] Barbara van Schewick and David Farber. Point/counterpoint: Network neutrality nuances. Communications of the ACM, 52(2):31–37, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Drepper:2009:PPT

[2671] Ulrich Drepper. Parallel programming with transactional memory. Communications of the ACM, 52(2):38–43, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leighton:2009:IPI

[2672] Tom Leighton. Improving performance on the Internet. Communications of the ACM, 52(2):44–51, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Raman:2009:TBW

[2673] T. V. Raman. Toward 2^{W}, beyond Web 2.0. Communications of the ACM, 52(2):52–59, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:2009:CRN

[2674] Mary Hall, David Padua, and Keshav Pingali. Compiler research: the next 50 years. Communications of the ACM, 52(2):60–67, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Klawe:2009:WCT

[2675] Maria Klawe, Telle Whitney, and Caroline Simard. Women in computing— take 2. Communications of the ACM, 52(2):68–76, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wallach:2009:TPT

[2676] Dan Wallach. Technical perspective: Tools for information to flow securely and swift-ly. Communications of the ACM, 52(2):78, February 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chong:2009:BSW

[2677] Stephen Chong, Jed Liu, Andrew C. Myers, Xin Qi, K. Vikram, Lantian Zheng, and Xin Zheng. Building secure Web applications with automatic partitioning. Communications of the ACM, 52(2):79–87, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kalai:2009:TPC

[2678] Ehud Kalai. Technical perspective:

The complexity of computing Nash equilibrium. Communications of the ACM, 52(2):88, February 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Daskalakis:2009:CCN

[2679] Constantinos Daskalakis, Paul W.

Goldberg, and Christos H. Papadimitriou. The complexity of computing a Nash equilibrium. Communications of the ACM, 52(2):89–97, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:LBPa

[2680] Peter Winkler. Last byte: Puzzled:

Will my algorithm terminate? Communications of the ACM, 52(2):104, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carter:2009:NCD

[2681] Pamela E. Carter and Gina Green.

Networks of contextualized data: a framework for cyberinfrastructure data management. Communications of the ACM, 52(2):105–109, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Li:2009:SOO

[2682] Xiao-Bai Li and Luvai Motiwalla. For sale by owner online: who gets the saved commission? Communications of the ACM, 52(2):110–114, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elmacioglu:2009:OWS

[2683] Ergin Elmacioglu and Dongwon Lee. Oracle, where shall I submit my papers? Communications of the ACM, 52 (2):115–118, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Argamon:2009:APA

[2684] Shlomo Argamon, Moshe Koppel, James W. Pennebaker, and Jonathan Schler. Automatically profiling the author of an anonymous text. Communications of the ACM, 52(2):119–123, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Behrens:2009:SSG

[2685] Sandy Behrens. Shadow systems: the good, the bad and the ugly. Communications of the ACM, 52(2): 124–129, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ragowsky:2009:WMT

[2686] Arik Ragowsky and David Gefen. Why IS management is in trouble and how to save it: lessons learned in the automotive industry. Communications of the ACM, 52(2):130–133, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2009:WFS

[2687] M. Eric Johnson, Dan McGuire, and Nicholas D. Willey. Why file sharing networks are dangerous? Communications of the ACM, 52(2):134–138, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Benbunan-Fich:2009:TOM

[2688] Raquel Benbunan-Fich and Gregory E.

Truman. Technical opinion: Multitasking with laptops during meetings. Communications of the ACM, 52 (2):139–141, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:ELY

[2689] Moshe Y. Vardi. Editor’s letter: ’yes, it can be done’. Communications of the ACM, 52(3):5, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2009:PCCa

[2690] Scott E. Delman. Publisher’s corner:

Communications’ Web site to launch in March. Communications of the ACM, 52(3):7, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:CMW

[2691] CACM Staff. Children’s magic won’t deliver the Semantic Web. Communications of the ACM, 52(3):8–9, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:PL

[2692] David Roman. Prepare for launch.

Communications of the ACM, 52(3):12, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goth:2009:BI

[2693] Gregory Goth. Betting on ideas. Communications of the ACM, 52(3):13–15, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:CC

[2694] Leah Hoffmann. Crowd control. Communications of the ACM, 52(3):16–17, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:EV

[2695] Kirk L. Kroeker. The evolution of virtualization. Communications of the ACM, 52(3):18–20, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frenkel:2009:DUI

[2696] Karen A. Frenkel. A difficult, unforgettable idea. Communications of the ACM, 52(3):21, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:AFH

[2697] CACM Staff. ACM Fellows honored. Communications of the ACM, 52(3):22, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2009:PIS

[2698] Peter J. Denning and Richard D. Riehle. The profession of IT: Is software engineering engineering? Communications of the ACM, 52(3):24–26, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2009:LSW

[2699] Pamela Samuelson. Legally speaking: When is a ’license’ really a sale? Communications of the ACM, 52(3):27–29, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patterson:2009:VYS

[2700] David A. Patterson. Viewpoint: Your students are your legacy. Communications of the ACM, 52(3):30–33, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ullman:2009:VAS

[2701] Jeffrey D. Ullman. Viewpoint: Advising students for success. Communications of the ACM, 52(3):34–37, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shustek:2009:IIC

[2702] Len Shustek. Interview: An interview with C. A. R. Hoare. Communications of the ACM, 52(3):38–41, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

White:2009:BSB

[2703] Walker White, Christoph Koch, Johannes Gehrke, and Alan Demers. Better scripts, better games. Communications of the ACM, 52(3):42–47, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Larson:2009:ECP

[2704] Jim Larson. Erlang for concurrent programming. Communications of the ACM, 52(3):48–56, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sellen:2009:RHV

[2705] Abigail Sellen, Yvonne Rogers, Richard Harper, and Tom Rodden. Reflecting human values in the digital age. Communications of the ACM, 52(3):58–66, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Harel:2009:SMP

[2706] David Harel. Statecharts in the making: a personal account. Communications of the ACM, 52(3):67–75, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sudan:2009:PCP

[2707] Madhu Sudan. Probabilistically checkable proofs. Communications of the ACM, 52(3):76–84, March 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spielman:2009:TPB

[2708] Daniel A. Spielman. Technical perspective: The beauty of error-correcting codes. Communications of the ACM, 52(3):86, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guruswami:2009:ECI

[2709] Venkatesan Guruswami and Atri Rudra. Error correction up to the information-theoretic limit. Communications of the ACM, 52(3):87–95, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mishra:2009:TPW

[2710] Bud Mishra. Technical perspective:

Where biology meets computing. Communications of the ACM, 52(3):96, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grosu:2009:LDE

[2711] Radu Grosu, Scott A. Smolka, Flavio Corradini, Anita Wasilewska, Emilia Entcheva, and Ezio Bartocci. Learning and detecting emergent behavior in networks of cardiac myocytes. Communications of the ACM, 52(3):97–105, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:LBPb

[2712] Peter Winkler. Last byte: Puzzled: Solutions and sources. Communications of the ACM, 52(3):111, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Garreau:2009:LBF

[2713] Joel Garreau. Last byte: Future Tense:

Radical evolution. Communications of the ACM, 52(3):112–ff, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sacchi:2009:DIM

[2714] Alexandre Sacchi, Emerson Giannini, Regiane Bochic, Nicolau Reinhard, and Alexandre B. Lopes. Digital inclusion with the McInternet: would you like fries with that? Communications of the ACM, 52(3):113–116, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holsapple:2009:NMK

[2715] Clyde W. Holsapple. A new map for knowledge dissemination channels. Communications of the ACM, 52(3): 117–125, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Culnan:2009:OPP

[2716] Mary J. Culnan and Thomas J. Carlin. Online privacy practices in higher education: making the grade? Communications of the ACM, 52(3):126–130, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fleischmann:2009:ETC

[2717] Kenneth R. Fleischmann and William A. Wallace. Ensuring transparency in computational modeling. Communications of the ACM, 52(3):131–134, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mann:2009:OAP

[2718] Florian Mann, Benedikt von Walter, Thomas Hess, and Rolf T. Wigand. Open access publishing in science. Communications of the ACM, 52(3): 135–139, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linden:2009:WCV

[2719] Greg Linden, Kenneth L. Kraemer, and Jason Dedrick. Who captures value in a global innovation network?: the case of Apple’s iPod. Communications of the ACM, 52(3):140–144, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Formica:2009:CSE

[2720] Anna Formica. Concept similarity by evaluating information contents and feature vectors: a combined approach. Communications of the ACM, 52(3):145–149, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loo:2009:TOS

[2721] Alfred Loo. Technical opinion: Security threats of smart phones and Bluetooth. Communications of the ACM, 52(3):150–152, March 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McGettrick:2009:ELC

[2722] Andrew McGettrick. Education letter:

Computing education matters. Communications of the ACM, 52(4):5, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LEW

[2723] CACM Staff. Letters to the editor:

What role for computer science in the war on terror? Communications of the ACM, 52(4):9, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COO

[2724] David Roman. CACM online: An ongoing study in usability. Communications of the ACM, 52(4):10, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stemp-Morlock:2009:NLM

[2725] Graeme Stemp-Morlock. News: Learning more about active learning. Communications of the ACM, 52(4):11–13, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2009:NOS

[2726] Alex Wright. News: Our sentiments, exactly. Communications of the ACM, 52(4):14–15, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Essex:2009:NDS

[2727] David Essex. News: Did somebody say virtual colonoscopy? Communications of the ACM, 52(4):16–18, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Violino:2009:NTR

[2728] Bob Violino. News: Time to reboot. Communications of the ACM, 52(4):19, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Joch:2009:NIE

[2729] Alan Joch. News: IT ecosystem in peril. Communications of the ACM, 52(4):20, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:2009:EMI

[2730] Anonymous. Emerging markets IT and the world’s ’bottom billion’. Communications of the ACM, 52(4):22–24, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

changes:2009:KVS

[2731] System changes and side effects. Kode vicious: System changes and side effects. Communications of the ACM, 52(4):25–26, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2009:TSMa

[2732] Michael Cusumano. Technology strategy and management: Strategies for difficult (and Darwinian) economic times. Communications of the ACM, 52(4):27–28, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buckley:2009:VCS

[2733] Michael Buckley. Viewpoint: Computing as social science. Communications of the ACM, 52(4):29–30, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:2009:VRE

[2734] Bertrand Meyer, Christine Choppy, Jørgen Staunstrup, and Jan van Leeuwen. Viewpoint: Research evaluation for computer science. Communications of the ACM, 52(4):31–34, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shapiro:2009:PBL

[2735] Mike Shapiro. Purpose-built languages. Communications of the ACM, 52(4):36–41, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Provos:2009:CWC

[2736] Niels Provos, Moheeb Abu Rajab, and Panayiotis Mavrommatis. Cybercrime 2.0: when the cloud turns dark. Communications of the ACM, 52(4):42–47, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Richardson:2009:ODL

[2737] Chris Richardson. ORM in dynamic languages. Communications of the ACM, 52(4):48–55, April 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weikum:2009:DIR

[2738] Gerhard Weikum, Gjergji Kasneci, Maya Ramanath, and Fabian Suchanek. Database and information-retrieval methods for knowledge discovery. Communications of the ACM, 52(4):56–64, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Williams:2009:RIV

[2739] Samuel Williams, Andrew Waterman, and David Patterson. Roofline: an insightful visual performance model for multicore architectures. Communications of the ACM, 52(4):65–76, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jackson:2009:DPD

[2740] Daniel Jackson. A direct path to dependable software. Communications of the ACM, 52(4):78–88, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Merchan:2009:TPD

[2741] Arif Merchan. Technical perspective: Disk array models for automating storage management. Communications of the ACM, 52(4):90, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mesnier:2009:RFM

[2742] Michael P. Mesnier, Matthew Wachs, Raja R. Sambasivan, Alice X. Zheng, and Gregory R. Ganger. Relative fitness modeling. Communications of the ACM, 52(4):91–96, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graefe:2009:TPI

[2743] Goetz Graefe. Technical perspective:

Integrating flash devices. Communications of the ACM, 52(4):97, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roberts:2009:INF

[2744] David Roberts, Taeho Kgil, and Trevor Mudge. Integrating NAND flash devices onto servers. Communications of the ACM, 52(4):98–103, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hofmann:2009:LBQ

[2745] Leah Hofmann. Last byte: Q&A: Our dame commander. Communications of the ACM, 52(4):112–ff, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nunamaker:2009:PEV

[2746] Jay F. Nunamaker, Jr., Bruce A.

Reinig, and Robert O. Briggs. Principles for effective virtual teamwork. Communications of the ACM, 52 (4):113–117, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Martinsons:2009:HCI

[2747] Maris G. Martinsons, Robert M. Davison, and Valdis Martinsons. How culture influences IT-enabled organizational change and information systems. Communications of the ACM, 52(4):118–123, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bock:2009:NWR

[2748] Gee-Woo Bock and Swee Ling Ho. Non-work related computing (NWRC). Communications of the ACM, 52 (4):124–128, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scheibe:2009:OLI

[2749] Kevin P. Scheibe, James C. McElroy, and Paula C. Morrow. Object language and impression management. Communications of the ACM, 52(4):129–131, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Belanger:2009:IDD

[2750] France Bélanger and Lemuria Carter. The impact of the digital divide on egovernment use. Communications of the ACM, 52(4):132–135, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shalhoub:2009:AIS

[2751] Zeinab Karake Shalhoub. Analysis of industry-specific concentration of CPOs in Fortune 500 companies. Communications of the ACM, 52(4):136– 141, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scornavacca:2009:MPC

[2752] Eusebio Scornavacca, Sid Huff, and Stephen Marshall. Mobile phones in the classroom: if you can’t beat them, join them. Communications of the ACM, 52(4):142–146, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goes:2009:TOO

[2753] Paulo Goes, Yanbin Tu, and Y. Alex Tung. Technical opinion: Online auctions hidden metrics. Communications of the ACM, 52(4):147–149, April 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:ELC

[2754] Moshe Y. Vardi. Editor’s letter:

Conferences vs. journals in computing research. Communications of the ACM, 52(5):5, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LEL

[2755] CACM Staff. Letters to the Editor: Logic of lemmings in compiler innovation. Communications of the ACM, 52(5):7–9, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:BRA

[2756] CACM Staff. blog@CACM: Recommendation algorithms, online privacy, and more. Communications of the ACM, 52(5):10–11, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COP

[2757] David Roman. CACM online: The print-Web partnership turns the page. Communications of the ACM, 52(5):12, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:NRS

[2758] Kirk L. Kroeker. News: Rethinking signal processing. Communications of the ACM, 52(5):13–15, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Essex:2009:NMM

[2759] David Essex. News: Matchmaker, matchmaker. Communications of the ACM, 52(5):16–17, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2009:NLG

[2760] Samuel Greengard. News: Learning goes global. Communications of the ACM, 52(5):18–19, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:NLW

[2761] CACM Staff. News: Liskov wins Turing Award. Communications of the ACM, 52(5):21, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Larouche:2009:LTN

[2762] Pierre Larouche. Law and technology: The network neutrality debate hits Europe. Communications of the ACM, 52(5):22–24, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coder:2009:EBD

[2763] LeAnne Coder, Joshua L. Rosenbloom, Ronald A. Ash, and Brandon R. Dupont. Economic and business dimensions: Increasing gender diversity in the IT work force. Communications of the ACM, 52(5):25–27, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell-Kelly:2009:HRR

[2764] Martin Campbell-Kelly. Historical reflections: The rise, fall, and resurrection of software as a service. Communications of the ACM, 52(5):28–30, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guzdial:2009:ETC

[2765] Mark Guzdial. Education: Teaching computing to everyone. Communications of the ACM, 52(5):31–33, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Birman:2009:VPC

[2766] Ken Birman and Fred B. Schneider.

Viewpoint: Program committee overload in systems. Communications of the ACM, 52(5):34–37, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wadlow:2009:SB

[2767] Thomas Wadlow and Vlad Gorelik. Security in the browser. Communications of the ACM, 52(5):40–45, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Henning:2009:ADM

[2768] Michi Henning. API design matters. Communications of the ACM, 52(5):46–56, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schrock:2009:DAP

[2769] Eric Schrock. Debugging AJAX in production. Communications of the ACM, 52(5):57–60, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Larus:2009:SMD

[2770] James Larus. Spending Moore’s dividend. Communications of the ACM, 52(5):62–69, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lee:2009:CNT

[2771] Edward A. Lee. Computing needs time. Communications of the ACM, 52(5):70–79, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Priami:2009:ASB

[2772] Corrado Priami. Algorithmic systems biology. Communications of the ACM, 52(5):80–88, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2009:TPC

[2773] Ross Anderson. Technical perspective: A chilly sense of security. Communications of the ACM, 52(5):90, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Halderman:2009:LWR

[2774] J. Alex Halderman, Seth D. Schoen, Nadia Heninger, William Clarkson, William Paul, Joseph A. Calandrino, Ariel J. Feldman, Jacob Appelbaum, and Edward W. Felten. Lest we remember: cold-boot attacks on encryption keys. Communications of the ACM, 52(5):91–98, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herlihy:2009:TPH

[2775] Maurice Herlihy. Technical perspective: Highly concurrent data structures. Communications of the ACM, 52 (5):99, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Scherer:2009:SSQ

[2776] William N. Scherer, III, Doug Lea, and Michael L. Scott. Scalable synchronous queues. Communications of the ACM, 52(5):100–111, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:LBPc

[2777] Peter Winkler. Last byte: Puzzled:

Understanding relationships among numbers. Communications of the ACM, 52(5):112, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Procaccino:2009:SDV

[2778] J. Drew Procaccino and June M.

Verner. Software developers’ views of end-users and project success. Communications of the ACM, 52(5):113–116, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cousins:2009:DUC

[2779] Karlene C. Cousins and Upkar Varshney. Designing ubiquitous computing environments to support work life balance. Communications of the ACM, 52(5):117–123, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Galup:2009:OIS

[2780] Stuart D. Galup, Ronald Dattero, Jim J. Quan, and Sue Conger. An overview of IT service management. Communications of the ACM, 52 (5):124–127, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Plice:2009:TIC

[2781] Robert K. Plice, Nigel P. Melville, and Oleg V. Pavlov. Toward an information-compatible anti-spam strategy. Communications of the ACM, 52(5):128–130, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCart:2009:CBS

[2782] James A. McCart, Varol O. Kayhan, and Anol Bhattacherjee. Crossbidding in simultaneous online auctions. Communications of the ACM, 52(5):131–134, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ou:2009:TDQ

[2783] Carol Xiaojuan Ou and Choon Ling Sia. To trust or to distrust, that is the question: investigating the trustdistrust paradox. Communications of the ACM, 52(5):135–139, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brock:2009:RTP

[2784] Gary W. Brock, Denise J. McManus, and Joanne E. Hale. Reflections today prevent failures tomorrow. Communications of the ACM, 52(5):140–144, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rebstock:2009:TOS

[2785] Michael Rebstock. Technical opinion: Semantic ambiguity: Babylon, Rosetta or beyond? Communications of the ACM, 52(5):145–146, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weyuker:2009:AWL

[2786] Elaine Weyuker. ACM-W letter:

ACM-W celebrates women in computing. Communications of the ACM, 52 (6):5, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LES

[2787] CACM Staff. Letters to the Editor:

Share the threats. Communications of the ACM, 52(6):9, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lau:2009:BSA

[2788] Tessa Lau and Daniel Reed. blog@CACM:

Speech-activated user interfaces and climbing Mt. Exascale. Communications of the ACM, 52(6):10–11, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COMa

[2789] David Roman. CACM online: Making that connection. Communications of the ACM, 52(6):12, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Monroe:2009:NMR

[2790] Don Monroe. News: Micromedicine to the rescue. Communications of the ACM, 52(6):13–15, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:NCC

[2791] Leah Hoffmann. News: Content control. Communications of the ACM, 52(6):16–17, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goth:2009:NAH

[2792] Gregory Goth. News: Autonomous helicopters. Communications of the ACM, 52(6):18–20, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Violino:2009:NLB

[2793] Bob Violino. News: Looking backward and forward. Communications of the ACM, 52(6):21, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spafford:2009:PSA

[2794] Eugene H. Spafford. Privacy and security: Answering the wrong questions is no answer. Communications of the ACM, 52(6):22–24, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fu:2009:IRR

[2795] Kevin Fu. Inside risks: Reducing risks of implantable medical devices. Communications of the ACM, 52(6):25–27, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2009:PIB

[2796] Peter J. Denning. The profession of IT: Beyond computational thinking. Communications of the ACM, 52(6):28–30, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallman:2009:VWO

[2797] Richard Stallman. Viewpoint: Why ‘open source’ misses the point of free software. Communications of the ACM, 52(6):31–33, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2009:KVO

[2798] George V. Neville-Neil. Kode vicious:

Obvious truths. Communications of the ACM, 52(6):34–35, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Elerath:2009:HDD

[2799] Jon Elerath. Hard-disk drives: the good, the bad, and the ugly. Communications of the ACM, 52(6):38–45, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ODell:2009:NFE

[2800] Mike O’Dell. Network front-end processors, yet again. Communications of the ACM, 52(6):46–50, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2009:WS

[2801] George V. Neville-Neil. Whither sockets? Communications of the ACM, 52(6):51–55, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Agrawal:2009:CRD

[2802] Rakesh Agrawal, Anastasia Ailamaki, Philip A. Bernstein, Eric A. Brewer, Michael J. Carey, Surajit Chaudhuri, Anhai Doan, Daniela Florescu, Michael J. Franklin, Hector Garcia-Molina, Johannes Gehrke, Le Gruenwald, Laura M. Haas, Alon Y. Halevy, Joseph M. Hellerstein, Yannis E. Ioannidis, Hank F. Korth, Donald Kossmann, Samuel Madden, Roger Magoulas, Beng Chin Ooi, Tim O’Reilly, Raghu Ramakrishnan, Sunita Sarawagi, Michael Stonebraker, Alexander S. Szalay, and Gerhard Weikum. The Claremont Report on Database Research. Communications of the ACM, 52(6):56–65, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kraemer:2009:OLP

[2803] Kenneth L. Kraemer, Jason Dedrick, and Prakul Sharma. One laptop per child: vision vs. reality. Communications of the ACM, 52(6):66–73, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dias:2009:HCS

[2804] M. Bernardine Dias and Eric Brewer.

How computer science serves the developing world. Communications of the ACM, 52(6):74–80, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Myers:2009:TPR

[2805] Andrew Myers. Technical perspective: Reframing security for the Web. Communications of the ACM, 52(6):82, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Barth:2009:SFC

[2806] Adam Barth, Collin Jackson, and John C. Mitchell. Securing frame communication in browsers. Communications of the ACM, 52(6):83–91, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jouppi:2009:TPS

[2807] Norman P. Jouppi. Technical perspective: Software and hardware support for deterministic replay of parallel programs. Communications of the ACM, 52(6):92, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hower:2009:THB

[2808] Derek R. Hower, Pablo Montesinos, Luis Ceze, Mark D. Hill, and Josep Torrellas. Two hardware-based approaches for deterministic multiprocessor replay. Communications of the ACM, 52(6):93–100, June 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:PSSa

[2809] Peter Winkler. Puzzled: Solutions and sources. Communications of the ACM, 52(6):103, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sawyer:2009:FTW

[2810] Robert J. Sawyer. Future tense: Webmind says hello. Communications of the ACM, 52(6):104, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Park:2009:HWS

[2811] Jonghun Park and Kwanho Kim. Hyperlinking the work for selfmanagement of flexible workflows. Communications of the ACM, 52 (6):113–117, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thompson:2009:AIT

[2812] Steven M. Thompson and Matthew D. Dean. Advancing information technology in health care. Communications of the ACM, 52(6):118–121, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gupta:2009:DMB

[2813] Amar Gupta. Deriving mutual benefits from offshore outsourcing. Communications of the ACM, 52(6):122–126, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lycett:2009:CED

[2814] Mark Lycett and Chris Partridge. The challenge of epistemic divergence in IS development. Communications of the ACM, 52(6):127–131, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lim:2009:FCH

[2815] Nena Lim and Anne Khoo. Forensics of computers and handheld devices: identical or fraternal twins? Communications of the ACM, 52(6):132–135, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bhattacharjee:2009:RTM

[2816] Sudip Bhattacharjee, Ram D. Gopal, James R. Marsden, and Ramesh Sankaranarayanan. Re-tuning the music industry: can they re-attain business resonance? Communications of the ACM, 52(6):136–140, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delen:2009:HFK

[2817] Dursun Delen and Suliman Al-Hawamdeh. A holistic framework for knowledge discovery and management. Communications of the ACM, 52(6):141–145, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Liang:2009:TOL

[2818] T. P. Liang, Andrew J. Czaplewski, Gary Klein, and James J. Jiang. Technical opinion: Leveraging first-mover advantages in Internet-based consumer services. Communications of the ACM, 52(6):146–148, June 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:ELO

[2819] Moshe Y. Vardi. Editor’s letter: Open, closed, or clopen access? Communications of the ACM, 52(7):5, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2009:PCCb

[2820] Scott E. Delman. Publisher’s corner: Communications’ annual report card. Communications of the ACM, 52(7):6– 7, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LEI

[2821] CACM Staff. Letters to the editor: Inspire with introductory computer science. Communications of the ACM, 52(7):8–9, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Guzdial:2009:BSI

[2822] Mark Guzdial, Greg Linden, and Tessa Lau. blog@CACM: Sharing ideas, writing apps, and creating a professional Web presence. Communications of the ACM, 52(7):10–11, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COMb

[2823] David Roman. CACM online: Moving forward and backward. Communications of the ACM, 52(7):12, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2009:NCA

[2824] Alex Wright. News: Contemporary approaches to fault tolerance. Communications of the ACM, 52(7):13–15, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:NTN

[2825] Kirk L. Kroeker. News: Toward native Web execution. Communications of the ACM, 52(7):16–17, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2009:NWL

[2826] Samuel Greengard. News: Are we losing our ability to think critically? Communications of the ACM, 52(7): 18–19, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frenkel:2009:NLC

[2827] Karen A. Frenkel. News: Liskov’s creative joy. Communications of the ACM, 52(7):20–22, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Joch:2009:NMC

[2828] Alan Joch. News: Master of connections. Communications of the ACM, 52(7):23, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:NAA

[2829] CACM Staff. News: ACM award winners. Communications of the ACM, 52 (7):25, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2009:LSD

[2830] Pamela Samuelson. Legally speaking:

The dead souls of the Google book search settlement. Communications of the ACM, 52(7):28–30, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sako:2009:TSM

[2831] Mari Sako. Technology strategy and management: Globalization of knowledge-intensive professional services. Communications of the ACM, 52(7):31–33, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2009:BSCa

[2832] Phillip G. Armour. The business of software: The cliché defense. Communications of the ACM, 52(7):34–36, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Felleisen:2009:VWC

[2833] Matthias Felleisen and Shriram Krishnamurthi. Viewpoint: Why computer science doesn’t matter. Communications of the ACM, 52(7):37–40, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dewar:2009:PCC

[2834] Robert Dewar and Owen Astrachan. Point/counterpoint: CS education in the U.S.: heading in the wrong direction? Communications of the ACM, 52(7):41–45, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Graefe:2009:FMR

[2835] Goetz Graefe. The five-minute rule 20 years later (and how flash memory changes the rules). Communications of the ACM, 52(7):48–59, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2009:FPT

[2836] Jonathan M. Smith. Fighting physics:

a tough battle. Communications of the ACM, 52(7):60–65, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gelenbe:2009:STS

[2837] Erol Gelenbe. Steps toward self-aware networks. Communications of the ACM, 52(7):66–75, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kazman:2009:MMN

[2838] Rick Kazman and Hong-Mei Chen.

The Metropolis Model: a new logic for development of crowdsourced systems. Communications of the ACM, 52(7):76–84, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dalvi:2009:PDD

[2839] Nilesh Dalvi, Christopher Ré, and Dan Suciu. Probabilistic databases: diamonds in the dirt. Communications of the ACM, 52(7):86–94, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Russell:2009:TPU

[2840] Stuart Russell and Lawrence Saul.

Technical perspective: The ultimate pilot program. Communications of the ACM, 52(7):96, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coates:2009:ALH

[2841] Adam Coates, Pieter Abbeel, and Andrew Y. Ng. Apprenticeship learning for helicopter control. Communications of the ACM, 52(7):97–105, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morrisett:2009:TPC

[2842] Greg Morrisett. Technical perspective: A compiler’s story. Communications of the ACM, 52(7):106, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leroy:2009:FVR

[2843] Xavier Leroy. Formal verification of a realistic compiler. Communications of the ACM, 52(7):107–115, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:LBQb

[2844] Leah Hoffmann. Last byte: Q&A: Liskov on Liskov. Communications of the ACM, 52(7):120–ff, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fabian:2009:VES

[2845] Benjamin Fabian and Oliver Günther. Virtual extension: Security challenges of the EPCglobal network. Communications of the ACM, 52(7):121–125, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parrish:2009:VEC

[2846] James L. Parrish, Jr. and James F. Courtney, Jr. Virtual extension:

Churchman’s inquirers as design templates for knowledge management systems. Communications of the ACM, 52(7):126–129, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malhotra:2009:VER

[2847] Arvind Malhotra and Claudia Kubowicz Malhotra. Virtual extension: A relevancy-based services view for driving adoption of wireless Web services in the U.S. Communications of the ACM, 52(7):130–134, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rost:2009:VEI

[2848] Johann Rost and Robert L. Glass.

Virtual extension: The impact of subversive stakeholders on software projects. Communications of the ACM, 52(7):135–138, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Patel:2009:VEI

[2849] Sandip C. Patel, Ganesh D. Bhatt, and James H. Graham. Virtual extension: Improving the cyber security of SCADA communication networks. Communications of the ACM, 52(7):139–142, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hovav:2009:VEA

[2850] Anat Hovav and Ciprian Popoviciu.

Virtual extension: Adoption leadership and early planners: Comcast’s IP upgrade strategy. Communications of the ACM, 52(7):143–146, July 2009.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Woolridge:2009:VES

[2851] Richard W. Woolridge, David P. Hale, Joanne E. Hale, and R. Shane Sharpe. Virtual extension: Software project scope alignment: an outcome-based approach. Communications of the ACM, 52(7):147–152, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davison:2009:VET

[2852] Robert M. Davison, Maris G. Martinsons, Henry W. H. Lo, and Yuan Li. Virtual extension: Technical opinion: The ethics of IT professionals in China. Communications of the ACM, 52(7):153–155, July 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bourne:2009:CP

[2853] Steve Bourne and Bryan Cantrill. Communications and the practitioner. Communications of the ACM, 52(8):5, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Delman:2009:RB

[2854] Scott E. Delman. Responding to the blogosphere. Communications of the ACM, 52(8):7, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:WII

[2855] CACM Staff. Why invention and innovation diverge. Communications of the ACM, 52(8):8–9, August 2009.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:VEa

[2856] CACM Staff. In the virtual extension.

Communications of the ACM, 52(8):10, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wing:2009:IRA

[2857] Jeannette M. Wing, Daniel Reed, and Mark Guzdial. An ICT research agenda, HPC and innovation, and why only the developed world lacks women in computing. Communications of the ACM, 52(8):12–13, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:NS

[2858] David Roman. The new searchers.

Communications of the ACM, 52(8):14, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Monroe:2009:JY

[2859] Don Monroe. Just for you. Communications of the ACM, 52(8):15–17, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:FRB

[2860] Kirk L. Kroeker. Face recognition breakthrough. Communications of the ACM, 52(8):18–19, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geller:2009:IDP

[2861] Tom Geller. IT drives policy—and vice versa. Communications of the ACM, 52(8):20, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:LTG

[2862] Leah Hoffmann. Learning through games. Communications of the ACM, 52(8):21–22, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goth:2009:UUC

[2863] Gregory Goth. U.S. unveils cybersecurity plan. Communications of the ACM, 52(8):23, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Draper:2009:EBD

[2864] Tim Draper. Economic and business dimensions: Entrepreneurship during a slump. Communications of the ACM, 52(8):24–26, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dann:2009:EAC

[2865] Wanda Dann and Stephen Cooper. Education Alice 3: concrete to abstract. Communications of the ACM, 52(8):27–29, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Snow:2009:PSE

[2866] Brian Snow and Clinton Brooks. Privacy and security: An ethics code for U.S. intelligence officers. Communications of the ACM, 52(8):30–32, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fortnow:2009:VTC

[2867] Lance Fortnow. Viewpoint: Time for computer science to grow up. Communications of the ACM, 52(8):33–35, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jacobs:2009:PBD

[2868] Adam Jacobs. The pathologies of big data. Communications of the ACM, 52(8):36–44, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reis:2009:BSL

[2869] Charles Reis, Adam Barth, and Carlos Pizano. Browser security: lessons from Google Chrome. Communications of the ACM, 52(8):45–49, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Creeger:2009:CRC

[2870] Mache Creeger. CTO roundtable:

cloud computing. Communications of the ACM, 52(8):50–56, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shinohara:2009:BPI

[2871] Kristen Shinohara and Josh Tenenberg.

A blind person’s interactions with technology. Communications of the ACM, 52(8):58–66, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morelli:2009:RCE

[2872] Ralph Morelli, Allen Tucker, Norman Danner, Trishan R. De Lanerolle, Heidi J. C. Ellis, Ozgur Izmirli, Danny Krizanc, and Gary Parker. Revitalizing computing education through free and open source software for humanity. Communications of the ACM, 52(8):67–75, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Malik:2009:BST

[2873] Sharad Malik and Lintao Zhang. Boolean satisfiability from theoretical hardness to practical success. Communications of the ACM, 52(8):76–82, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herbsleb:2009:TPM

[2874] James Herbsleb. Technical perspective: Maintaining quality in the face of distributed development. Communications of the ACM, 52(8):84, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bird:2009:DDD

[2875] Christian Bird, Nachiappan Nagappan, Premkumar Devanbu, Harald Gall, and Brendan Murphy. Does distributed development affect software quality?: an empirical case study of Windows Vista. Communications of the ACM, 52(8):85–93, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sapatnekar:2009:TPW

[2876] Sachin S. Sapatnekar. Technical perspective: Where the chips may fall.

Communications of the ACM, 52(8):94, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Orshansky:2009:SAC

[2877] Michael Orshansky and Wei-Shen Wang. Statistical analysis of circuit timing using majorization. Communications of the ACM, 52(8):95–100, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:PPI

[2878] Peter Winkler. Puzzled: Probability and intuition. Communications of the ACM, 52(8):104, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cha:2009:WDI

[2879] Hoon S. Cha, David E. Pingry, and Matt E. Thatcher. What determines IT spending priorities? Communications of the ACM, 52(8):105–110, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lim:2009:DCQ

[2880] Andrew Lim, Hong Ma, Qi Wen, Zhou Xu, and Brenda Cheang. Distinguishing citation quality for journal impact assessment. Communications of the ACM, 52(8):111–116, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gerace:2009:CEP

[2881] Thomas Gerace and Huseyin Cavusoglu. The critical elements of the patch management process. Communications of the ACM, 52(8):117–121, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kohli:2009:LBI

[2882] Rajiv Kohli and Nigel P. Melville. Learning to build an IT innovation platform. Communications of the ACM, 52(8):122–126, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

OConchuir:2009:GSD

[2883] Eoin Ó Conchúir, Pär J. Ågerfalk, Helena H. Olsson, and Brian Fitzgerald. Global software development: where are the benefits? Communications of the ACM, 52(8):127–131, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kaul:2009:RBW

[2884] Rohit Kaul, Yeogirl Yun, and Seong-Gon Kim. Ranking billions of Web pages using diodes. Communications of the ACM, 52(8):132–136, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Varma:2009:ANA

[2885] Roli Varma. Attracting Native Americans to computing. Communications of the ACM, 52(8):137–140, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bose:2009:TOW

[2886] Indranil Bose and Alvin Chung Man Leung. Technical opinion: What drives the adoption of antiphishing measures by Hong Kong banks? Communications of the ACM, 52(8):141–143, August 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:FMC

[2887] Moshe Y. Vardi. The financial meltdown and computing. Communications of the ACM, 52(9):5, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:CSD

[2888] CACM Staff. Computer science does matter. Communications of the ACM, 52(9):8–9, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:VEb

[2889] CACM Staff. In the virtual extension. Communications of the ACM, 52(9):10, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stonebraker:2009:SGB

[2890] Michael Stonebraker and Jason Hong.

Saying good-bye to DBMSs, designing effective interfaces. Communications of the ACM, 52(9):12–13, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:WYR

[2891] David Roman. What you read on your summer vacation. Communications of the ACM, 52(9):14, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goth:2009:EPU

[2892] Gregory Goth. Entering a parallel universe. Communications of the ACM, 52(9):15–17, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:MN

[2893] Kirk L. Kroeker. Medical nanobots. Communications of the ACM, 52(9): 18–19, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2009:FAO

[2894] Samuel Greengard. Facing an age-old problem. Communications of the ACM, 52(9):20–22, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frenkel:2009:CSM

[2895] Karen A. Frenkel. Computer Science meets environmental science. Communications of the ACM, 52(9):23, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Schwartz:2009:LTK

[2896] Paul M. Schwartz. Law and technology: Keeping track of telecommunications surveillance. Communications of the ACM, 52(9):24–26, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2009:PIC

[2897] Peter J. Denning and Paul S. Rosenbloom. The profession of IT: Computing: the fourth great domain of science. Communications of the ACM, 52 (9):27–29, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cleverley:2009:EMH

[2898] Mark Cleverley. Emerging markets:

How ICT advances might help developing nations. Communications of the ACM, 52(9):30–32, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wilson:2009:IPL

[2899] Cameron Wilson and Peter Harsha.

IT policy: The long road to Computer Science education reform. Communications of the ACM, 52(9):33–35, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ghuloum:2009:VFI

[2900] Anwar Ghuloum. Viewpoint: Face the inevitable, embrace parallelism. Communications of the ACM, 52(9): 36–38, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anderson:2009:IIM

[2901] David P. Anderson. Interview: An interview with Maurice Wilkes. Communications of the ACM, 52(9):39–42, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2009:RC

[2902] Bruce Johnson. Reveling in constraints. Communications of the ACM, 52(9):44–48, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Legrand:2009:MCL

[2903] Iosif Legrand, Ramiro Voicu, Catalin Cirstoiu, Costin Grigoras, Latchezar Betev, and Alexandru Costan. Monitoring and control of large systems with MonALISA. Communications of the ACM, 52(9):49–55, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

OSullivan:2009:MSR

[2904] Bryan O’Sullivan. Making sense of revision-control systems. Communications of the ACM, 52(9):56–62, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bhagwan:2009:SIC

[2905] Varun Bhagwan, Tyrone Grandison, and Daniel Gruhl. Sound index: charts for the people, by the people. Communications of the ACM, 52 (9):64–70, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyle:2009:WIP

[2906] James Boyle. What intellectual property law should learn from software. Communications of the ACM, 52(9): 71–76, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fortnow:2009:SPV

[2907] Lance Fortnow. The status of the P versus NP problem. Communications of the ACM, 52(9):78–86, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yelick:2009:TPA

[2908] Katherine Yelick. Technical perspective: Abstraction for parallelism. Communications of the ACM, 52(9):88, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kulkarni:2009:OPR

[2909] Milind Kulkarni, Keshav Pingali, Bruce Walter, Ganesh Ramanarayanan, Kavita Bala, and L. Paul Chew. Optimistic parallelism requires abstractions. Communications of the ACM, 52(9):89–97, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dacier:2009:TPT

[2910] Marc Dacier. Technical perspective:

They do click, don’t they? Communications of the ACM, 52(9):98, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kanich:2009:SEA

[2911] Chris Kanich, Christian Kreibich, Kirill Levchenko, Brandon Enright, Geoffrey M. Voelker, Vern Paxson, and Stefan Savage. Spamalytics: an empirical analysis of spam marketing conversion. Communications of the ACM, 52 (9):99–107, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:PSSb

[2912] Peter Winkler. Puzzled: Solutions and sources. Communications of the ACM, 52(9):110, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lanier:2009:FTC

[2913] Jaron Lanier. Future tense: Confusions of the hive mind. Communications of the ACM, 52(9):112–ff, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Muthitacharoen:2009:EUI

[2914] Achita (Mi) Muthitacharoen and Khawaja A. Saeed. Examining user involvement in continuous software development: (a case of error reporting system). Communications of the ACM, 52(9):113–117, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pasko:2009:CFB

[2915] Alexander Pasko and Valery Adzhiev. Constructive function-based modeling in multilevel education. Communications of the ACM, 52(9):118–122, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zhu:2009:OSD

[2916] Hongwei Zhu and Stuart E. Madnick. One size does not fit all: legal protection for non-copyrightable data. Communications of the ACM, 52(9): 123–128, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loiacono:2009:SCW

[2917] Eleanor T. Loiacono, Nicholas C. Romano, Jr., and Scott McCoy. The state of corporate website accessibility. Communications of the ACM, 52 (9):128–132, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Willison:2009:OIR

[2918] Robert Willison and Mikko Siponen.

Overcoming the insider: reducing employee computer crime through Situational Crime Prevention. Communications of the ACM, 52(9):133–137, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Xia:2009:BBC

[2919] Mu Xia, Yun Huang, Wenjing Duan, and Andrew B. Whinston. Ballot box communication in online communities. Communications of the ACM, 52(9): 138–142, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Batra:2009:MAP

[2920] Dinesh Batra. Modified agile practices for outsourced software projects. Communications of the ACM, 52(9): 143–148, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katz:2009:TOF

[2921] James Katz and Ronald E. Rice. Technical opinion: Falling into the net: Main Street America playing games and making friends online. Communications of the ACM, 52(9):149–150, September 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hall:2009:PLA

[2922] Wendy Hall. President’s letter: ACM Europe. Communications of the ACM, 52(10):5, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:TC

[2923] CACM Staff. Letters to the editor: Time and computing. Communications of the ACM, 52(10):6, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:VEc

[2924] CACM Staff. In the Virtual Extension. Communications of the ACM, 52(10):7, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Linden:2009:BNP

[2925] Greg Linden, Michael Conover, and Judy Robertson. blog@CACM: The Netflix prize, computer science outreach, and Japanese mobile phones. Communications of the ACM, 52(10): 8–9, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:COF

[2926] CACM Staff. CACM online: Following the leaders. Communications of the ACM, 52(10):10, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lindley:2009:NMD

[2927] David Lindley. News: Managing data. Communications of the ACM, 52(10):11–13, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Joch:2009:NDN

[2928] Alan Joch. News: Debating net neutrality. Communications of the ACM, 52(10):14–15, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Geller:2009:NSF

[2929] Tom Geller. News: Shaping the future. Communications of the ACM, 52(10):16–18, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Armour:2009:BSCb

[2930] Phillip G. Armour. The business of software: Contagious craziness, spreading sanity. Communications of the ACM, 52(10):19–20, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Campbell-Kelly:2009:HRC

[2931] Martin Campbell-Kelly. Historical reflections: Computing in the depression era. Communications of the ACM, 52 (10):21–22, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Porras:2009:IRR

[2932] Phillip Porras. Inside risks: Reflections on Conficker. Communications of the ACM, 52(10):23–24, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cusumano:2009:TSMb

[2933] Michael Cusumano. Technology strategy and management: Dealing with the venture capital crisis. Communications of the ACM, 52(10):25–27, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2009:KVK

[2934] George V. Neville-Neil. Kode Vicious: Kode reviews 101. Communications of the ACM, 52(10):28–29, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoare:2009:VRA

[2935] C. A. R. Hoare. Viewpoint: Retrospective: an axiomatic basis for computer programming. Communications of the ACM, 52(10):30–32, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Phillips:2009:PBM

[2936] James C. Phillips and John E. Stone. Probing biomolecular machines with graphics processors. Communications of the ACM, 52(10):34–41, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dougherty:2009:UBI

[2937] Matthew T. Dougherty, Michael J. Folk, Erez Zadok, Herbert J. Bernstein, Frances C. Bernstein, Kevin W. Eliceiri, Werner Benger, and Christoph Best. Unifying biological image formats with HDF5. Communications of the ACM, 52(10):42–47, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:CDS

[2938] CACM Staff. A conversation with David E. Shaw. Communications of the ACM, 52(10):48–54, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Asanovic:2009:VPC

[2939] Krste Asanovic, Rastislav Bodik, James Demmel, Tony Keaveny, Kurt Keutzer, John Kubiatowicz, Nelson Morgan, David Patterson, Koushik Sen, John Wawrzynek, David Wessel, and Katherine Yelick. A view of the parallel computing landscape. Communications of the ACM, 52(10):56– 67, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cleland-Huang:2009:ASM

[2940] Jane Cleland-Huang, Horatiu Dumitru, Chuan Duan, and Carlos Castro-Herrera. Automated support for managing feature requests in open forums. Communications of the ACM, 52(10):68–74, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Spielman:2009:SAA

[2941] Daniel A. Spielman and Shang-Hua Teng. Smoothed analysis: an attempt to explain the behavior of algorithms in practice. Communications of the ACM, 52(10):76–84, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chaudhuri:2009:TPR

[2942] Surajit Chaudhuri. Technical perspective: Relational query optimization: data management meets statistical estimation. Communications of the ACM, 52(10):86, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beyer:2009:DVS

[2943] Kevin Beyer, Rainer Gemulla, Peter J. Haas, Berthold Reinwald, and Yannis Sismanis. Distinct-value synopses for multiset operations. Communications of the ACM, 52(10):87–95, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gehrke:2009:TPD

[2944] Johannes Gehrke. Technical perspective: Data stream processing: when you only get one look. Communications of the ACM, 52(10):96, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cormode:2009:FFI

[2945] Graham Cormode and Marios Hadjieleftheriou. Finding the frequent items in streams of data. Communications of the ACM, 52(10):97–105, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:QN

[2946] Leah Hoffmann. Q&A: The networker. Communications of the ACM, 52(10):112–ff, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goeke:2009:TOD

[2947] Richard J. Goeke and Robert H. Faley. Technical opinion: Do SAP successes outperform themselves and their competitors? Communications of the ACM, 52(10):113–117, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Subramanian:2009:BFF

[2948] Girish H. Subramanian, Gary Klein, James J. Jiang, and Chien-Lung Chan. Balancing four factors in system development projects. Communications of the ACM, 52(10):118–121, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sangareddy:2009:ASC

[2949] Sridhar R. Papagari Sangareddy, Sanjeev Jha, Chen Ye, and Kevin C. Desouza. Attaining superior complaint resolution. Communications of the ACM, 52(10):122–126, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Waller:2009:MUC

[2950] Vivienne Waller and Robert B. Johnston. Making ubiquitous computing available. Communications of the ACM, 52(10):127–130, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Flynn:2009:EIP

[2951] Donal Flynn, Gary Pan, Mark Keil, and Magnus Mähring. De-escalating IT projects: the DMM model. Communications of the ACM, 52(10):131– 134, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Casacuberta:2009:HIH

[2952] Francisco Casacuberta, Jorge Civera, Elsa Cubel, Antonio L. Lagarda, Guy Lapalme, Elliott Macklovitch, and Enrique Vidal. Human interaction for high-quality machine translation. Communications of the ACM, 52(10):

135–138, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Savoy:2009:HEG

[2953] Jacques Savoy and Ljiljana Dolamic. How effective is Google’s translation service in search? Communications of the ACM, 52(10):139–143, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hu:2009:OJS

[2954] Nan Hu, Jie Zhang, and Paul A. Pavlou. Overcoming the J-shaped distribution of product reviews. Communications of the ACM, 52(10):144– 147, October 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vardi:2009:ELI

[2955] Moshe Y. Vardi. Editor’s letter: Is the image crisis over? Communications of the ACM, 52(11):5, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LEP

[2956] CACM Staff. Letters to the editor: Pay for editorial independence. Communications of the ACM, 52(11):6–7, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:VEd

[2957] CACM Staff. In the virtual extension. Communications of the ACM, 52(11):9, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rao:2009:BCS

[2958] Ramana Rao and Greg Linden.

blog@CACM: Computer science curriculum, deceptive advertising. Communications of the ACM, 52(11):10–11, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COI

[2959] David Roman. CACM online: Internet addiction: it’s spreading, but is it real? Communications of the ACM, 52(11):12, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anthes:2009:NDD

[2960] Gary Anthes. News: Deep data dives discover natural laws. Communications of the ACM, 52(11):13–14, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kroeker:2009:NEP

[2961] Kirk L. Kroeker. News: Electronic paper’s next chapter. Communications of the ACM, 52(11):15–17, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:NIE

[2962] Leah Hoffmann. News: Implementing electronic medical records. Communications of the ACM, 52(11):18–20, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goth:2009:NEN

[2963] Gregory Goth. News: Exploring new frontiers. Communications of the ACM, 52(11):21–23, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lampson:2009:PSU

[2964] Butler Lampson. Privacy and security: Usable security: how to get it. Communications of the ACM, 52 (11):25–27, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Samuelson:2009:LSB

[2965] Pamela Samuelson. Legally speaking: Are business methods patentable? Communications of the ACM, 52(11): 28–30, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greenstein:2009:EBD

[2966] Shane Greenstein. Economic and business dimensions: The broadband price is not right. Communications of the ACM, 52(11):31–33, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Smith:2009:VPS

[2967] Jonathan M. Smith. Viewpoint: On public service and computer science. Communications of the ACM, 52(11): 34–35, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cramblitt:2009:IIP

[2968] Bob Cramblitt. Interview: An interview with Ping Fu. Communications of the ACM, 52(11):36–39, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Diffie:2009:CSP

[2969] Whitfield Diffie and Susan Landau.

Communications surveillance: privacy and security at risk. Communications of the ACM, 52(11):42–47, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shilton:2009:FBL

[2970] Katie Shilton. Four billion little brothers?: privacy, mobile phones, and ubiquitous data collection. Communications of the ACM, 52(11):48–53, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stachour:2009:YDK

[2971] Paul Stachour and David Collier-Brown. You don’t know jack about software maintenance. Communications of the ACM, 52(11):54–58, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Resnick:2009:SPA

[2972] Mitchel Resnick, John Maloney, Andrés Monroy-Hernández, Natalie Rusk, Evelyn Eastmond, Karen Brennan, Amon Millner, Eric Rosenbaum, Jay Silver, Brian Silverman, and Yasmin Kafai. Scratch: programming for all. Communications of the ACM, 52 (11):60–67, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bandyopadhyay:2009:WIM

[2973] Tridib Bandyopadhyay, Vijay S. Mookerjee, and Ram C. Rao. Why IT managers don’t go for cyber-insurance products. Communications of the ACM, 52(11):68–73, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clarke:2009:MCA

[2974] Edmund M. Clarke, E. Allen Emerson, and Joseph Sifakis. Model checking: algorithmic verification and debugging. Communications of the ACM, 52(11):74–84, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Druschel:2009:TPN

[2975] Peter Druschel. Technical perspective: Narrowing the semantic gap in distributed programming. Communications of the ACM, 52(11):86, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Loo:2009:DN

[2976] Boon Thau Loo, Tyson Condie, Minos Garofalakis, David E. Gay, Joseph M. Hellerstein, Petros Maniatis, Raghu Ramakrishnan, Timothy Roscoe, and Ion Stoica. Declarative networking. Communications of the ACM, 52(11): 87–95, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Shawe-Taylor:2009:TPM

[2977] John Shawe-Taylor. Technical perspective: Machine learning for complex predictions. Communications of the ACM, 52(11):96, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Joachims:2009:PSO

[2978] Thorsten Joachims, Thomas Hofmann, Yisong Yue, and Chun-Nam Yu. Predicting structured objects with support vector machines. Communications of the ACM, 52(11):97–104, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:LBPd

[2979] Peter Winkler. Last byte: Puzzled:

Covering the plane. Communications of the ACM, 52(11):112, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khansa:2009:QBI

[2980] Lara Khansa and Divakaran Liginlal.

Quantifying the benefits of investing in information security. Communications of the ACM, 52(11):113–117, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chang:2009:IHP

[2981] Wei-Lun Chang, Soe-Tsyr Yuan, and Eldon Y. Li. iCare home portal: an extended model of quality aging e-services. Communications of the ACM, 52(11):118–124, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Taneja:2009:CJT

[2982] Aakash Taneja, Anil Singh, and M. K.

Raja. Computing journals and their emerging roles in knowledge exchange. Communications of the ACM, 52(11): 125–131, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschheim:2009:ONW

[2983] Rudy Hirschheim. Offshoring and the new world order. Communications of the ACM, 52(11):132–135, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bolchini:2009:WCC

[2984] C. Bolchini, C. A. Curino, G. Orsi, E. Quintarelli, R. Rossato, F. A. Schreiber, and L. Tanca. And what can context do for data? Communications of the ACM, 52(11):136–140, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McCown:2009:WWS

[2985] Frank McCown, Catherine C. Marshall, and Michael L. Nelson. Why Web sites are lost (and how they’re sometimes found). Communications of the ACM, 52(11):141–145, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Westfall:2009:IYP

[2986] Ralph Westfall. If your pearls of wisdom fall in a forest. … Communications of the ACM, 52(11):146–149, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nielsen:2009:TOS

[2987] Frank Nielsen. Technical opinion: Steering self-learning distance algorithms. Communications of the ACM, 52(11):150–152, November 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stephenson:2009:CLI

[2988] Chris Stephenson. CSTA letter: It is a pivotal time for K–12 computer science. Communications of the ACM, 52(12):5, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:LEC

[2989] CACM Staff. Letters to the editor: In CS education, educate the educators first. Communications of the ACM, 52(12):6–7, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wing:2009:BCW

[2990] Jeannette M. Wing and Mark Guzdial. blog@CACM: CS woes: deadlinedriven research, academic inequality. Communications of the ACM, 52 (12):8–9, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Staff:2009:VE

[2991] CACM Staff. In the virtual extension. Communications of the ACM, 52(12):11, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Roman:2009:COCb

[2992] David Roman. CACM online: Crowdsourcing and the question of expertise. Communications of the ACM, 52(12):12, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anthes:2009:NBS

[2993] Gary Anthes. News: Blueprints for self-assembly. Communications of the ACM, 52(12):13–15, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wright:2009:NRW

[2994] Alex Wright. News: Ready for a Web OS? Communications of the ACM, 52 (12):16–17, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Greengard:2009:NMA

[2995] Samuel Greengard. News: Making automation work. Communications of the ACM, 52(12):18–19, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hoffmann:2009:NPS

[2996] Leah Hoffmann. News: Problem solvers. Communications of the ACM, 52(12):20, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ladner:2009:VBP

[2997] Richard E. Ladner. Viewpoints: Broadening participation: Opening remarks. Communications of the ACM, 52(12):22–24, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berry:2009:VEM

[2998] Orna Berry and Yigal Grayeff. Viewpoints: Emerging markets Israel’s technology industry as an economic growth engine. Communications of the ACM, 52(12):25–27, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Denning:2009:VPI

[2999] Peter J. Denning and Peter A. Freeman. Viewpoints: The profession of IT: Computing’s paradigm. Communications of the ACM, 52(12):28–30, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Neville-Neil:2009:VKV

[3000] George V. Neville-Neil. Viewpoints:

Kode vicious: Broken builds. Communications of the ACM, 52(12):31–32, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Parastatidis:2009:VSC

[3001] Savas Parastatidis, Evelyne Viegas, and Tony Hey. Viewpoint: A ’smart’ cyberinfrastructure for research. Communications of the ACM, 52(12):33–37, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramos:2009:PTA

[3002] Alan Ramos, Weina Scott, William Scott, Doug Lloyd, Katherine O’Leary, and Jim Waldo. Practice: A threat analysis of RFID passports. Communications of the ACM, 52(12):38–42, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vixie:2009:PWD

[3003] Paul Vixie. Practice: What DNS is not. Communications of the ACM, 52 (12):43–47, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fedorova:2009:PMP

[3004] Alexandra Fedorova, Juan Carlos Saez, Daniel Shelepov, and Manuel Prieto. Practice: Maximizing power efficiency with asymmetric multicore systems. Communications of the ACM, 52 (12):48–57, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Torrellas:2009:BMA

[3005] Josep Torrellas, Luis Ceze, James Tuck, Calin Cascaval, Pablo Montesinos, Wonsun Ahn, and Milos Prvulovic. The Bulk Multicore architecture for improved programmability. Communications of the ACM, 52 (12):58–65, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zyda:2009:CSC

[3006] Michael Zyda. Computer science in the conceptual age. Communications of the ACM, 52(12):66–72, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sung:2009:CGT

[3007] Kelvin Sung. Computer games and traditional CS courses. Communications of the ACM, 52(12):74–78, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Landay:2009:TPD

[3008] James A. Landay. Technical perspective: Design tools for the rest of us. Communications of the ACM, 52(12):80, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Igarashi:2009:DPT

[3009] Yuki Igarashi and Takeo Igarashi. Designing plush toys with a computer. Communications of the ACM, 52(12): 81–88, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanrahan:2009:TPG

[3010] Pat Hanrahan. Technical perspective:

A graphical sense of touch. Communications of the ACM, 52(12):89, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Izadi:2009:TTF

[3011] Shahram Izadi, Steve Hodges, Alex Butler, Darren West, Alban Rrustemi, Mike Molloy, and William Buxton. ThinSight: a thin form-factor interactive surface technology. Communications of the ACM, 52(12):90–98, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Winkler:2009:LBPe

[3012] Peter Winkler. Last byte: Puzzled: Solutions and sources. Communications of the ACM, 52(12):111, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Haldeman:2009:LBF

[3013] Joe Haldeman. Last byte: Future tense: Mightier than the pen. Communications of the ACM, 52(12):112, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Panteli:2009:PTG

[3014] Niki Panteli and Robert Tucker. Power and trust in global virtual teams. Communications of the ACM, 52(12): 113–115, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lim:2009:OPG

[3015] Sun Sun Lim, Hichang Cho, and Milagros Rivera Sanchez. Online privacy, government surveillance and national ID cards. Communications of the ACM, 52(12):116–120, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vaidyanathan:2009:SDW

[3016] Ganesh Vaidyanathan and Steven Mautone. Security in dynamic Web Content Management Systems applications. Communications of the ACM, 52 (12):121–125, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hafer:2009:AOS

[3017] Lou Hafer and Arthur E. Kirkpatrick. Assessing open source software as a scholarly contribution. Communications of the ACM, 52(12):126–129, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cerpa:2009:WDY

[3018] Narciso Cerpa and June M. Verner. Why did your project fail? Communications of the ACM, 52(12):130–134, December 2009. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic).

Renaud:2009:VPC

[3019] Karen Renaud and Antonella De Angeli. Visual passwords: cure-all or snake-oil? Communications of the ACM, 52(12):135–140, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kshetri:2009:PEI

[3020] Nir Kshetri. Positive externality, increasing returns, and the rise in cybercrimes. Communications of the ACM, 52(12):141–144, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Siponen:2009:TOE

[3021] Mikko Siponen, M. Adam Mahmood, and Seppo Pahnila. Technical opinion: Are employees putting your company at risk by not following information security policies? Communications of the ACM, 52(12):145–147, December 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).